### v1

# Computational Complexity of Sufficiency in Decision Problems

Abstract
We characterize the computational complexity of coordinate sufficiency in decision problems within the formal model. Given action set $A$, state space $S=X_{1}\times\cdots\times X_{n}$, and utility $u:A\times S\to\mathbb{R}$, a coordinate set $I$ is sufficient if $s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$.
The landscape in the formal model:
  • General case: SUFFICIENCY-CHECK is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
  • Tractable cases: Polynomial-time for bounded action sets under the explicit-state encoding; separable utilities ($u=f+g$) under any encoding; and tree-structured utility with explicit local factors.
  • Encoding-regime separation: Polynomial-time under the explicit-state encoding (polynomial in $|S|$). Under ETH, there exist succinctly encoded worst-case instances witnessed by a strengthened gadget construction (mechanized in Lean; see Appendix A) with $k^{*}=n$ for which SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
The tractable cases are stated with explicit encoding assumptions (Section 2.4). Together, these results answer the question “when is decision-relevant information identifiable efficiently?” within the stated regimes.
[…]
All results are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems).
Keywords: computational complexity, decision theory, polynomial hierarchy, tractability dichotomy, Lean 4
footnotetext: © 2026 Tristan Simas. This work is licensed under $C$ BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
Zenodo DOI: 10.5281/zenodo.18140965
1 Introduction
Consider a decision problem with actions $A$ and states $S=X_{1}\times\cdots\times X_{n}$. A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
This paper characterizes the efficient cases of coordinate sufficiency within the formal model:
Section 2.4 fixes the computational model and input encodings used for all complexity claims.
Problem Complexity When Tractable
SUFFICIENCY-CHECK coNP-complete Bounded actions (explicit-state), separable utility, tree-structured utility
MINIMUM-SUFFICIENT-SET coNP-complete Same conditions
ANCHOR-SUFFICIENCY $\Sigma_{2}^{\textsf{P}}$-complete Open
The tractable cases are stated with explicit encoding assumptions (Section 2.4). Outside those regimes, the succinct model yields hardness.
1.1 Landscape Summary
When is sufficiency checking tractable? We identify three sufficient conditions:
  1. 1.
    Bounded actions ($|A|\leq k$) under explicit-state encoding: with constantly many actions, we enumerate action pairs over the explicit utility table.
  2. 2.
    Separable utility ($u(a,s)=f(a)+g(s)$): The optimal action depends only on $f$, making all coordinates irrelevant to the decision.
  3. 3.
    Tree-structured utility: With explicit local factors over a tree, dynamic programming yields polynomial algorithms in the input length.
Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 3.6).
When is it intractable? The general problem is coNP-complete (Theorem 3.6), with a separation between explicit-state tractability and succinct worst-case hardness:
  • In the explicit-state model: SUFFICIENCY-CHECK is solvable in polynomial time in $|S|$ by explicitly computing $\text{Opt}(s)$ for all $s\in S$ and checking all pairs $(s,s^{\prime})$ with equal $I$-projection. In particular, instances with $k^{*}=O(\log|S|)$ are tractable in this model.
  • In the succinct model: under ETH there exist worst-case instances produced by our strengthened gadget in which the minimal sufficient set has size $\Omega(n)$ (indeed $n$) and SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
The lower-bound statement does not address intermediate regimes.
1.2 Main Theorems
  1. 1.
    Theorem 3.6: SUFFICIENCY-CHECK is coNP-complete via reduction from TAUTOLOGY.
  2. 2.
    Theorem 3.7: MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{\textsf{P}}$ structure collapses).
  3. 3.
    Theorem 3.9: ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete via reduction from $\exists\forall$-SAT.
  4. 4.
    Theorem 4.1: Encoding-regime separation: explicit-state polynomial-time (polynomial in $|S|$), and under ETH a succinct worst-case lower bound witnessed by a hard family with $k^{*}=n$.
  5. 5.
    Theorem 5.1: Polynomial algorithms for bounded actions, separable utility, tree structure.
1.3 Machine-Checked Proofs
All results are formalized in Lean 4 [7] ($\sim$5,000 lines, 200+ theorems). The formalization verifies the reduction mappings and combinatorial lemmas; complexity class membership follows by composition with TAUTOLOGY and $\exists\forall$-SAT.
What is new.
We contribute (i) a reusable Lean 4 framework for polynomial-time reductions with explicit polynomial bounds; (ii) the first machine-checked coNP-completeness proof for a decision-theoretic sufficiency problem; and (iii) a complete complexity landscape for coordinate sufficiency under explicit encoding assumptions. Prior work studies decision complexity in general or feature selection hardness, but does not formalize these reductions or establish this landscape in Lean.
1.4 Paper Structure
[…]
Section 2: foundations. Section 3: hardness proofs. Section 4: dichotomy. Section 5: tractable cases. Sections 7 and 8: corollaries and implications for practice. Section 9: related work. Appendix A: Lean listings.
2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [14, 13].
2.1 Decision Problems with Coordinate Structure
Definition 2.1 (Decision Problem).
A decision problem with coordinate structure is a tuple $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ where:
  • AA is a finite set of actions (alternatives)
  • $X_{1},\ldots,X_{n}$ are finite coordinate spaces
  • $S=X_{1}\times\cdots\times X_{n}$ is the state space
  • $U:A\times S\to\mathbb{Q}$ is the utility function
Definition 2.2 (Projection).
For state $s=(s_{1},\ldots,s_{n})\in S$ and coordinate set $I\subseteq\{1,\ldots,n\}$:
$s_{I}:=(s_{i})_{i\in I}$
is the projection of $s$ onto coordinates in $I$.
Definition 2.3 (Optimizer Map).
For state $s\in S$, the optimal action set is:
$\text{Opt}(s):=\arg\max_{a\in A}U(a,s)=\{a\in A:U(a,s)=\max_{a^{\prime}\in A}U(a^{\prime},s)\}$
2.2 Sufficiency and Relevance
Definition 2.4 (Sufficient Coordinate Set).
A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient for decision problem $\mathcal{D}$ if:
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
Equivalently, the optimal action depends only on coordinates in $I$.
Definition 2.5 (Minimal Sufficient Set).
A sufficient set $I$ is minimal if no proper subset $I^{\prime}\subsetneq I$ is sufficient.
Definition 2.6 (Relevant Coordinate).
Coordinate $i$ is relevant if it belongs to some minimal sufficient set.
Example 2.7 (Weather Decision).
Consider deciding whether to carry an umbrella:
  • Actions: $A=\{\text{carry},\text{don't carry}\}$
  • Coordinates: $X_{1}=\{\text{rain},\text{no rain}\}$, $X_{2}=\{\text{hot},\text{cold}\}$, $X_{3}=\{\text{Monday},\ldots,\text{Sunday}\}$
[…]
The minimal sufficient set is $I=\{1\}$ (only rain forecast matters). Coordinates 2 and 3 (temperature, day of week) are irrelevant.
2.3 The Decision Quotient
Definition 2.8 (Decision Equivalence).
For coordinate set $I$, states $s,s^{\prime}$ are $I$-equivalent (written $s\sim_{I}s^{\prime}$) if $s_{I}=s^{\prime}_{I}$.
Definition 2.9 (Decision Quotient).
The decision quotient for state $s$ under coordinate set $I$ is:
$\text{DQ}_{I}(s)=\frac{|\{a\in A:a\in\text{Opt}(s^{\prime})\text{ for some }s^{\prime}\sim_{I}s\}|}{|A|}$
This measures the fraction of actions that are optimal for at least one state consistent with $I$.
Proposition 2.10 (Sufficiency Characterization).
Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$.
Proof.
If $I$ is sufficient, then $s\sim_{I}s^{\prime}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$, so the set of actions optimal for some $s^{\prime}\sim_{I}s$ is exactly $\text{Opt}(s)$.
Conversely, if the condition holds, then for any $s\sim_{I}s^{\prime}$, the optimal actions form the same set (since $\text{DQ}_{I}(s)=\text{DQ}_{I}(s^{\prime})$ and both equal the relative size of the common optimal set).  $\blacksquare$
2.4 Computational Model and Input Encoding
We fix the computational model used by the complexity claims.
Succinct encoding (primary for hardness).
This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [2]. An instance is encoded as:
  • a finite action set $A$ given explicitly,
  • coordinate domains $X_{1},\ldots,X_{n}$ given by their sizes in binary,
  • a Boolean or arithmetic circuit $C_{U}$ that on input $(a,s)$ outputs $U(a,s)$.
The input length is $L=|A|+\sum_{i}\log|X_{i}|+|C_{U}|$. Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH, W[2]) are measured in $L$. All hardness results in Section 3 use this encoding.
Explicit-state encoding (used for enumeration algorithms and experiments).
The utility is given as a full table over $A\times S$. The input length is $L_{\text{exp}}=\Theta(|A||S|)$ (up to the bitlength of utilities). Polynomial time is measured in $L_{\text{exp}}$. Results stated in terms of $|S|$ use this encoding.
Unless explicitly stated otherwise, “polynomial time” refers to the succinct encoding.
3 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of determining which state coordinates are decision-relevant. We prove three main results:
  1. 1.
    SUFFICIENCY-CHECK is coNP-complete
  2. 2.
    MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{P}$ structure collapses)
  3. 3.
    ANCHOR-SUFFICIENCY (fixed coordinates) is $\Sigma_{2}^{P}$-complete
Within the formal model (Section 2) and the succinct encoding used for hardness (Section 2.4), these results sit beyond NP-completeness and provide a rigorous explanation for why over-modeling can be a rational strategy: in the worst case, finding (or even certifying) a minimal set of decision-relevant factors is computationally intractable.
3.1 Problem Definitions
Definition 3.1 (Decision Problem Encoding).
A decision problem instance is a tuple $(A,X_{1},\ldots,X_{n},U)$ where:
  • AA is a finite set of alternatives
  • $X_{1},\ldots,X_{n}$ are the coordinate domains, with state space $S=X_{1}\times\cdots\times X_{n}$
  • $U:A\times S\to\mathbb{Q}$ is the utility function (in the succinct encoding, $U$ is given as a Boolean circuit)
Definition 3.2 (Optimizer Map).
For state $s\in S$, define:
$\text{Opt}(s):=\arg\max_{a\in A}U(a,s)$
Definition 3.3 (Sufficient Coordinate Set).
A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if:
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
where $s_{I}$ denotes the projection of $s$ onto coordinates in $I$.
Problem 3.4 (SUFFICIENCY-CHECK).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I\subseteq\{1,\ldots,n\}$
Question: Is $I$ sufficient?
Problem 3.5 (MINIMUM-SUFFICIENT-SET).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and integer $k$
Question: Does there exist a sufficient set $I$ with $|I|\leq k$?
3.2 Hardness of SUFFICIENCY-CHECK
Theorem 3.6 (coNP-completeness of SUFFICIENCY-CHECK).
SUFFICIENCY-CHECK is coNP-complete [5, 10]. (Machine-verified in Lean 4; see Appendix A.)
Source Target Key property preserved
TAUTOLOGY SUFFICIENCY-CHECK Tautology iff $\emptyset$ sufficient
$\exists\forall$-SAT ANCHOR-SUFFICIENCY Witness anchors iff formula true
SET-COVER MINIMUM-SUFFICIENT-SET Set size maps to coordinate size
Proof.
Membership in coNP: The complementary problem INSUFFICIENCY is in NP. Given a decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I$, a witness for insufficiency is a pair $(s,s^{\prime})$ such that:
  1. 1.
    $s_{I}=s^{\prime}_{I}$ (verifiable in polynomial time)
  2. 2.
    $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$ (verifiable by evaluating $U$ on all alternatives)
coNP-hardness: We reduce from TAUTOLOGY.
Given Boolean formula $\varphi(x_{1},\ldots,x_{n})$, construct a decision problem with:
  • Alternatives: $A=\{\text{accept},\text{reject}\}$
  • State space: $S=\{\text{reference}\}\cup\{0,1\}^{n}$ (equivalently, encode this as a product space with one extra coordinate $r\in\{0,1\}$ indicating whether the state is the reference state)
  • Utility:
    $U(\text{accept},\text{reference})$ $=1$
    $U(\text{reject},\text{reference})$ $=0$
    $U(\text{accept},a)$ $=\varphi(a)$
    $U(\text{reject},a)$ $=0\quad\text{for assignments }a\in\{0,1\}^{n}$
  • Query set: $I=\emptyset$
Claim: $I=\emptyset$ is sufficient $\iff$ $\varphi$ is a tautology.
($\Rightarrow$) Suppose $I$ is sufficient. Then $\text{Opt}(s)$ is constant over all states. Since $U(\text{accept},a)=\varphi(a)$ and $U(\text{reject},a)=0$:
  • $\text{Opt}(a)=\text{accept}$ when $\varphi(a)=1$
  • $\text{Opt}(a)=\{\text{accept},\text{reject}\}$ when $\varphi(a)=0$
For Opt to be constant, $\varphi(a)$ must be true for all assignments $a$; hence $\varphi$ is a tautology.
($\Leftarrow$) If $\varphi$ is a tautology, then $U(\text{accept},a)=1>0=U(\text{reject},a)$ for all assignments $a$. Thus $\text{Opt}(s)=\{\text{accept}\}$ for all states $s$, making $I=\emptyset$ sufficient.  $\blacksquare$
Mechanized strengthening (all coordinates relevant).
The reduction above establishes coNP-hardness using a single witness set $I=\emptyset$. For the ETH-based lower bound in Theorem 4.1, we additionally need worst-case instances where the minimal sufficient set has linear size.
We formalized a strengthened reduction in Lean 4: given a Boolean formula $\varphi$ over $n$ variables, construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (so $k^{*}=n$). Intuitively, the construction places a copy of the base gadget at each coordinate and makes the global “accept” condition hold only when every coordinate’s local test succeeds; a single falsifying assignment at one coordinate therefore changes the global optimal set, witnessing that coordinate’s relevance. This strengthening is mechanized in Lean; see Appendix A.
3.3 Complexity of MINIMUM-SUFFICIENT-SET
Theorem 3.7 (MINIMUM-SUFFICIENT-SET is coNP-complete).
MINIMUM-SUFFICIENT-SET is coNP-complete.
Proof.
Structural observation: The $\exists\forall$ quantifier pattern suggests $\Sigma_{2}^{P}$:
$\exists I\,(|I|\leq k)\;\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
However, this collapses because sufficiency has a simple characterization.
Key lemma: A coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_contains_relevant in Lean):
$\text{sufficient}(I)\iff\text{Relevant}\subseteq I$
[…]
Consequence: The minimum sufficient set is exactly the set of relevant coordinates. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
coNP membership: A witness that the answer is NO is a set of $k+1$ coordinates, each proven relevant (by exhibiting $s,s^{\prime}$ pairs). Verification is polynomial.
coNP-hardness: The $k=0$ case asks whether no coordinates are relevant, i.e., whether $\emptyset$ is sufficient. This is exactly SUFFICIENCY-CHECK, which is coNP-complete by Theorem 3.6.  $\blacksquare$
3.4 Anchor Sufficiency (Fixed Coordinates)
We also formalize a strengthened variant that fixes the coordinate set and asks whether there exists an assignment to those coordinates that makes the optimal action constant on the induced subcube.
Problem 3.8 (ANCHOR-SUFFICIENCY).
Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and fixed coordinate set $I\subseteq\{1,\ldots,n\}$
Question: Does there exist an assignment $\alpha$ to $I$ such that $\text{Opt}(s)$ is constant for all states $s$ with $s_{I}=\alpha$?
Theorem 3.9 (ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete).
ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete [16] (already for Boolean coordinate spaces).
Proof.
Membership in $\Sigma_{2}^{P}$: The problem has the form
$\exists\alpha\;\forall s\in S:\;(s_{I}=\alpha)\implies\text{Opt}(s)=\text{Opt}(s_{\alpha}),$
which is an $\exists\forall$ pattern.
$\Sigma_{2}^{P}$-hardness: Reduce from $\exists\forall$-SAT. Given $\exists x\,\forall y\,\varphi(x,y)$ with $x\in\{0,1\}^{k}$ and $y\in\{0,1\}^{m}$, if $m=0$ we first pad with a dummy universal variable (satisfiability is preserved), construct a decision problem with:
  • Actions $A=\{\text{YES},\text{NO}\}$
  • State space $S=\{0,1\}^{k+m}$ representing $(x,y)$
  • Utility
[…]
  • Fixed coordinate set $I$ = the $x$-coordinates.
If $\exists x^{\star}\,\forall y\,\varphi(x^{\star},y)=1$, then for any $y$ we have $U(\text{YES})=2$ and $U(\text{NO})\leq 1$, so $\text{Opt}(x^{\star},y)=\{\text{YES}\}$ is constant. Conversely, if $\varphi(x,y)$ is false for some $y$, then either $y=0^{m}$ (where NO is optimal) or $y\neq 0^{m}$ (where YES and NO tie), so the optimal set varies across $y$ and the subcube is not constant. Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.  $\blacksquare$
3.5 Tractable Subcases
Despite the general hardness, several natural subcases admit efficient algorithms:
Proposition 3.10 (Small State Space).
When $|S|$ is polynomial in the input size (i.e., explicitly enumerable), MINIMUM-SUFFICIENT-SET is solvable in polynomial time.
Proof.
Compute $\text{Opt}(s)$ for all $s\in S$. The minimum sufficient set is exactly the set of coordinates that “matter” for the resulting function, computable by standard techniques.  $\blacksquare$
Proposition 3.11 (Linear Utility).
When $U(a,s)=w_{a}\cdot s$ for weight vectors $w_{a}\in\mathbb{Q}^{n}$, MINIMUM-SUFFICIENT-SET reduces to identifying coordinates where weight vectors differ.
3.6 Implications
Corollary 3.12 (Why Over-Modeling Is Rational).
Finding the minimal set of decision-relevant factors is coNP-complete. Even verifying that a proposed set is sufficient is coNP-complete.
This formally explains the engineering phenomenon:
  1. 1.
    It’s computationally easier to model everything than to find the minimum
  2. 2.
    “Which unknowns matter?” is a hard question, not a lazy one to avoid
  3. 3.
    Bounded scenario analysis (small $\hat{S}$) makes the problem tractable
This connects to the pentalogy’s leverage framework: the modeling budget for deciding what to model is itself a computationally constrained resource.
3.7 Remark: The Collapse to coNP
Early analysis of MINIMUM-SUFFICIENT-SET focused on the apparent $\exists\forall$ quantifier structure, which suggested a $\Sigma_{2}^{P}$-complete result. We initially explored certificate-size lower bounds for the complement, attempting to show MINIMUM-SUFFICIENT-SET was unlikely to be in coNP.
However, the key insight—formalized as sufficient_contains_relevant—is that sufficiency has a simple characterization: a set is sufficient iff it contains all relevant coordinates. This collapses the $\exists\forall$ structure because:
  • The minimum sufficient set is exactly the relevant coordinate set
  • Checking relevance is in coNP (witness: two states differing only at that coordinate with different optimal sets)
  • Counting relevant coordinates is also in coNP
This collapse explains why ANCHOR-SUFFICIENCY retains its $\Sigma_{2}^{P}$-completeness: fixing coordinates and asking for an assignment that works is a genuinely different question. The “for all suffixes” quantifier cannot be collapsed when the anchor assignment is part of the existential choice.
4 Encoding-Regime Separation
The hardness results of Section 3 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound.
Model note.
Part 1 is an explicit-state upper bound (time polynomial in $|S|$). Part 2 is a succinct-encoding lower bound under ETH (time exponential in $n$). The encodings are defined in Section 2.4. These two parts are stated in different encodings and are not directly comparable as functions of a single input length.
Theorem 4.1 (Explicit–Succinct Regime Separation).
Let $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ be a decision problem with $|S|=N$ states. Let $k^{*}$ be the size of the minimal sufficient set.
  1. 1.
    Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
  2. 2.
    Succinct lower bound under ETH (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
Proof.
Part 1 (Explicit-state upper bound): Under the explicit-state encoding, SUFFICIENCY-CHECK is decidable in time polynomial in $N$ by direct enumeration: compute $\text{Opt}(s)$ for all $s\in S$ and then check all pairs $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$.
Equivalently, for any fixed $I$, the projection map $s\mapsto s_{I}$ has image size $|S_{I}|\leq|S|=N$, so any algorithm that iterates over projection classes (or over all state pairs) runs in polynomial time in $N$. Thus, in particular, when $k^{*}=O(\log N)$, SUFFICIENCY-CHECK is solvable in polynomial time under the explicit-state encoding.
Remark (bounded coordinate domains).
In the general model $S=\prod_{i}X_{i}$, for a fixed $I$ one always has $|S_{I}|\leq\prod_{i\in I}|X_{i}|$ (and $|S_{I}|\leq N$). If the coordinate domains are uniformly bounded, $|X_{i}|\leq d$ for all $i$, then $|S_{I}|\leq d^{|I|}$.
Part 2 (Succinct ETH lower bound, worst case): A strengthened version of the TAUTOLOGY reduction used in Theorem 3.6 produces a family of instances in which the minimal sufficient set has size $k^{*}=n$: given a Boolean formula $\varphi$ over $n$ variables, we construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (thus $k^{*}=n$). This strengthening is mechanized in Lean (see Appendix A). Under the Exponential Time Hypothesis (ETH) [9], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.  $\blacksquare$
Corollary 4.2 (Regime Separation (by Encoding)).
There is a clean separation between explicit-state tractability and succinct worst-case hardness (with respect to the encodings in Section 2.4):
  • Under the explicit-state encoding, SUFFICIENCY-CHECK is polynomial in $N=|S|$.
  • Under ETH, there exist succinctly encoded instances with $k^{*}=\Omega(n)$ (indeed $k^{*}=n$) for which SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
For Boolean coordinate spaces ($N=2^{n}$), the explicit-state bound is polynomial in $2^{n}$ (exponential in $n$), while under ETH the succinct lower bound yields $2^{\Omega(n)}$ time for the hard family in which $k^{*}=\Omega(n)$.
This encoding-regime separation explains why some domains admit tractable model selection under explicit-state assumptions, while other domains (or encodings) exhibit worst-case hardness that forces heuristic approaches.
5 Tractable Special Cases: When You Can Solve It
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly.
Theorem 5.1 (Tractable Subcases).
SUFFICIENCY-CHECK is polynomial-time solvable in the following cases:
  1. 1.
    Explicit-state encoding: The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
  2. 2.
    Separable utility (any encoding): $U(a,s)=f(a)+g(s)$.
  3. 3.
    Tree-structured utility with explicit local factors (succinct structured encoding): There exists a rooted tree on coordinates and local functions $u_{i}$ such that
    $U(a,s)=\sum_{i}u_{i}\bigl(a,s_{i},s_{\mathrm{parent}(i)}\bigr),$
    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly as part of the input.
5.1 Explicit-State Encoding
Proof of Part 1.
Given the full table of $U(a,s)$, compute $\text{Opt}(s)$ for all $s\in S$ in $O(|S||A|)$ time. For SUFFICIENCY-CHECK on a given $I$, verify that for all pairs $(s,s^{\prime})$ with $s_{I}=s^{\prime}_{I}$, we have $\text{Opt}(s)=\text{Opt}(s^{\prime})$. This takes $O(|S|^{2}|A|)$ time by direct enumeration and is polynomial in the explicit input length. If $|A|$ is constant, the runtime is $O(|S|^{2})$.  $\blacksquare$
5.2 Separable Utility
Proof of Part 2.
If $U(a,s)=f(a)+g(s)$, then:
$\text{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
The optimal action is independent of the state. Thus $I=\emptyset$ is always sufficient.  $\blacksquare$
5.3 Tree-Structured Utility
Proof of Part 3.
Assume the tree decomposition and explicit local tables as stated. For each node $i$ and each value of its parent coordinate, compute the set of actions that are optimal for some assignment of the subtree rooted at $i$. This is a bottom-up dynamic program that combines local tables with child summaries; each table lookup is explicit in the input. A coordinate is relevant if and only if varying its value changes the resulting optimal action set. The total runtime is polynomial in $n$, $|A|$, and the size of the local tables.  $\blacksquare$
5.4 Practical Implications
Condition Examples
Explicit-state encoding Small or fully enumerated state spaces
Separable utility Additive costs, linear models
Tree-structured utility Hierarchies, causal trees
6 Implications for Practice: Why Over-Modeling Is Optimal
This section states corollaries for engineering practice. Within the formal model, the complexity results of Sections 3 and 4 shift parts of this workflow from informal judgment toward explicit, checkable criteria. The observed behaviors—configuration over-specification, absence of automated minimization tools, heuristic model selection—are not failures of discipline but provably optimal responses to computational constraints under the stated cost model.
Some common prescriptions implicitly require exact minimization of sufficient parameter sets. In the worst case, that task is coNP-complete in our model, so we should calibrate critiques and expectations accordingly. A rational response is to include everything and pay linear maintenance costs, rather than attempt exponential minimization costs.
6.1 Configuration Simplification is SUFFICIENCY-CHECK
Real engineering problems reduce directly to the decision problems studied in this paper.
Theorem 6.1 (Configuration Simplification Reduces to SUFFICIENCY-CHECK).
Given a software system with configuration parameters $P=\{p_{1},\ldots,p_{n}\}$ and observed behaviors $B=\{b_{1},\ldots,b_{m}\}$, the problem of determining whether parameter subset $I\subseteq P$ preserves all behaviors is equivalent to SUFFICIENCY-CHECK.
Proof.
Construct decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ where:
  • Actions $A=B$ (each behavior is an action)
  • Coordinates $X_{i}$ = domain of parameter $p_{i}$
  • State space $S=X_{1}\times\cdots\times X_{n}$
  • Utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else $U(b,s)=0$
Then $\text{Opt}(s)=\{b\in B:b\text{ occurs under configuration }s\}$.
Coordinate set $I$ is sufficient iff:
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
This holds iff configurations agreeing on parameters in $I$ exhibit identical behaviors.
Therefore, “does parameter subset $I$ preserve all behaviors?” is exactly SUFFICIENCY-CHECK for the constructed decision problem.  $\blacksquare$
Remark 6.2.
This reduction is parsimonious: configuration simplification can be cast as an instance of SUFFICIENCY-CHECK under a direct encoding.
6.2 Computational Rationality of Over-Modeling
We now prove that over-specification is an optimal engineering strategy given the stated cost model and complexity constraints.
Theorem 6.3 (Rational Over-Modeling).
Consider an engineer specifying a system configuration with $n$ parameters. Let:
  • $C_{\text{over}}(k)$ = cost of maintaining $k$ extra parameters beyond minimal
  • $C_{\text{find}}(n)$ = cost of finding minimal sufficient parameter set
  • $C_{\text{under}}$ = expected cost of production failures from underspecification
When SUFFICIENCY-CHECK is coNP-complete (Theorem 3.6):
  1. 1.
    Naive exhaustive search inspects $2^{n}$ candidate subsets, so the brute-force finding cost grows exponentially in $n$; more formally, exhaustive search yields an upper bound $C_{\text{find}}(n)=O(2^{n})$. A matching *unconditional* lower bound of $2^{\Omega(n)}$ does not follow from coNP-completeness alone. Instead, coNP-completeness implies there is no polynomial‑time algorithm for arbitrary inputs unless $\textsf{P}=\textsf{coNP}$. Stronger assumptions (e.g., the Exponential Time Hypothesis) yield explicit worst-case lower bounds: under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ worst-case time lower bound in the succinct encoding (witnessed by the strengthened TAUTOLOGY gadget family).
  2. 2.
    Maintenance cost is linear: $C_{\text{over}}(k)=O(k)$.
  3. 3.
    Under the ETH (or when considering naive exhaustive search), exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, over-modeling minimizes total expected cost:
$C_{\text{over}}(k)<C_{\text{find}}(n)+C_{\text{under}}$
Over-modeling is economically optimal under the stated model and complexity constraints.
Proof.
By Theorem 3.6, SUFFICIENCY-CHECK is coNP-complete, so under the assumption $\textsf{P}\neq\textsf{coNP}$ there is no polynomial‑time algorithm that decides sufficiency for arbitrary inputs.
Finding the minimal sufficient set by brute force requires checking many candidate sets. Exhaustive search examines
$\sum_{i=0}^{n}\binom{n}{i}=2^{n}$
candidate subsets, so naive search has cost $O(2^{n})$. This shows only that brute-force is exponential. To assert a matching lower bound of $2^{\Omega(n)}$ for *all* algorithms requires additional assumptions. For example, under the Exponential Time Hypothesis (ETH), TAUTOLOGY (and hence SUFFICIENCY-CHECK in the succinct encoding) requires $2^{\Omega(n)}$ time, and SUFFICIENCY-CHECK inherits this conditional lower bound via our reduction.
Maintaining $k$ extra parameters incurs:
  • Documentation cost: $O(k)$ entries
  • Testing cost: $O(k)$ test cases
  • Migration cost: $O(k)$ parameters to update
Total maintenance cost is $C_{\text{over}}(k)=O(k)$.
For concrete threshold: when $n=20$ parameters, exhaustive search requires $2^{20}\approx 10^{6}$ checks. Including $k=5$ extra parameters costs $O(5)$ maintenance overhead but avoids $10^{6}$ computational work.
Since $2^{n}$ grows faster than any polynomial in $k$ or $n$, there exists $n_{0}$ such that for all $n>n_{0}$:
$C_{\text{over}}(k)\ll C_{\text{find}}(n)$
Adding underspecification risk $C_{\text{under}}$ (production failures from missing parameters), which is unbounded in the model, makes over-specification strictly dominant.  $\blacksquare$
Corollary 6.4 (Impossibility of Automated Configuration Minimization).
Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
  1. 1.
    Takes an arbitrary configuration file with $n$ parameters
  2. 2.
    Identifies the minimal sufficient parameter subset
  3. 3.
    Guarantees correctness (no false negatives)
Proof.
Such an algorithm would solve MINIMUM-SUFFICIENT-SET in polynomial time, contradicting Theorem 3.7 (assuming $\textsf{P}\neq\textsf{coNP}$).  $\blacksquare$
Remark 6.5.
Corollary 6.4 explains the observed absence of “config cleanup” tools in software engineering practice. Engineers who include extra parameters are not exhibiting poor discipline—they are adapting to computational constraints. The problem is not lack of tooling effort; it is mathematical intractability.
6.3 Connection to Observed Practice
These theorems provide mathematical grounding for three widespread engineering behaviors:
1. Configuration files grow over time. Removing parameters requires solving coNP-complete problems. Engineers rationally choose linear maintenance cost over exponential minimization cost.
2. Heuristic model selection dominates. ML practitioners use AIC, BIC, cross-validation instead of optimal feature selection because optimal selection is intractable (Theorem 6.3).
3. “Include everything” is a legitimate strategy. When determining relevance/minimal sufficiency is exponential in the worst case (e.g., $O(2^{n})$ by naive subset enumeration, and under ETH a $2^{\Omega(n)}$ worst-case lower bound in the succinct encoding; see Theorem 4.1), including all $n$ parameters costs $O(n)$. For large $n$, this is the rational choice.
These behaviors are not ad hoc workarounds; under the stated computational model they are rational responses to worst-case intractability. The complexity results provide a mathematical lens: over-modeling is not a failure—it is the rational strategy under the model.
7 Implications for Software Architecture
This section states corollaries for software architecture. The complexity results have direct implications for software engineering practice.
7.1 Why Over-Specification Is Rational
Software architects routinely specify more configuration parameters than strictly necessary. Our results show this is computationally rational:
Corollary 7.1 (Rational Over-Specification).
Given a software system with $n$ configuration parameters, checking whether a proposed subset suffices is coNP-complete. Finding the minimum such set is also coNP-complete.
This explains why configuration files grow over time: removing “unnecessary” parameters requires solving a hard problem.
7.2 Architectural Decision Quotient
The sufficiency framework suggests a measure for architectural decisions:
Definition 7.2 (Architectural Decision Quotient).
For a software system with configuration space $S$ and behavior space $B$:
$\text{ADQ}(I)=\frac{|\{b\in B:b\text{ achievable with some }s\text{ where }s_{I}\text{ fixed}\}|}{|B|}$
High ADQ means the configuration subset $I$ leaves many behaviors achievable—it doesn’t constrain the system much. Low ADQ means $I$ strongly constrains behavior.
Proposition 7.3.
Decisions with low ADQ (strongly constraining) require fewer additional decisions to fully specify system behavior.
7.3 Practical Recommendations
Based on our theoretical results:
  1. 1.
    Accept over-modeling: Don’t penalize engineers for including “extra” parameters. The alternative (minimal modeling) is computationally hard.
  2. 2.
    Use bounded scenarios: When the scenario space is small (Proposition 2.10), minimal modeling becomes tractable.
  3. 3.
    Exploit structure: Tree-structured dependencies, bounded alternatives, and separable utilities admit efficient algorithms.
  4. 4.
    Invest in heuristics: For general problems, develop domain-specific heuristics rather than seeking optimal solutions.
7.4 Hardness Distribution: Right Place vs Wrong Place
A general principle emerges from the complexity results: problem hardness is conserved and is distributed across a system in qualitatively different ways.
Definition 7.4 (Hardness Distribution).
Let $P$ be a problem with intrinsic hardness $H(P)$ (measured in computational steps, implementation effort, or error probability). A solution architecture $S$ partitions this hardness into:
  • $H_{\text{central}}(S)$: hardness paid once, at design time or in a shared component
  • $H_{\text{distributed}}(S)$: hardness paid per use site
For $n$ use sites, total realized hardness is:
$H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)$
Definition 7.5 (Hardness Conservation Principle).
For any problem $P$ with intrinsic hardness $H(P)$, any solution $S$ satisfies:
$H_{\text{central}}(S)+H_{\text{distributed}}(S)\geq H(P).$
This statement is presented as a definitional principle grounded in the prior definition of intrinsic hardness: any correct solution must account for at least $H(P)$ units of work, which the architecture may allocate centrally or distribute across use sites. The substantive, theorem-like consequences (for example, Centralization Dominance) follow from this grounding.
Definition 7.6 (Hardness Efficiency).
The hardness efficiency of solution $S$ with $n$ use sites is:
$\eta(S,n)=\frac{H_{\text{central}}(S)}{H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)}$
High $\eta$ indicates centralized hardness (paid once); low $\eta$ indicates distributed hardness (paid repeatedly).
Theorem 7.7 (Centralization Dominance).
For $n>1$ use sites, solutions with higher $H_{\text{central}}$ and lower $H_{\text{distributed}}$ yield:
  1. 1.
    Lower total realized hardness: $H_{\text{total}}(S_{1})<H_{\text{total}}(S_{2})$ when $H_{\text{distributed}}(S_{1})<H_{\text{distributed}}(S_{2})$
  2. 2.
    Fewer error sites: errors in centralized components affect 1 location; errors in distributed components affect $n$ locations
  3. 3.
    Higher leverage: one unit of central effort affects $n$ sites
Proof.
(1) follows from the total hardness formula. (2) follows from error site counting. (3) follows from the definition of leverage as $L=\Delta\text{Effect}/\Delta\text{Effort}$.  $\blacksquare$
Corollary 7.8 (Right Hardness vs Wrong Hardness).
A solution exhibits hardness in the right place when:
  • Hardness is centralized (high $H_{\text{central}}$, low $H_{\text{distributed}}$)
  • Hardness is paid at design/compile time rather than runtime
  • Hardness is enforced by tooling (type checker, compiler) rather than convention
A solution exhibits hardness in the wrong place when:
  • Hardness is distributed (low $H_{\text{central}}$, high $H_{\text{distributed}}$)
  • Hardness is paid repeatedly at each use site
  • Hardness relies on human discipline rather than mechanical enforcement
Example: Type System Instantiation. Consider a capability $C$ (e.g., provenance tracking) that requires hardness $H(C)$:
Approach $H_{\text{central}}$ $H_{\text{distributed}}$
Native type system support High (learning cost) Low (type checker enforces)
Manual implementation Low (no new concepts) High (reimplement per site)
For $n$ use sites, manual implementation costs $n\cdot H_{\text{distributed}}$, growing without bound. Native support costs $H_{\text{central}}$ once, amortized across all uses. The “simpler” approach (manual) is only simpler at $n=1$; for $n>H_{\text{central}}/H_{\text{distributed}}$, native support dominates.
Remark 7.9 (Connection to Decision Quotient).
The decision quotient (Section 2) measures which coordinates are decision-relevant. Hardness distribution measures where the cost of handling those coordinates is paid. A high-axis system makes relevance explicit (central hardness); a low-axis system requires users to track relevance themselves (distributed hardness).
Mechanized strengthening (sketch).
We formalized a strengthened reduction in the mechanization: given a Boolean formula $\varphi$ over $n$ variables we construct a decision problem with $n$ coordinates so that if $\varphi$ is not a tautology then every coordinate is decision-relevant. Intuitively, the construction places a copy of the base gadget at each coordinate and makes the global “accept” condition hold only when every coordinate’s local test succeeds; a single falsifying assignment at one coordinate therefore changes the global optimal set, witnessing that coordinate’s relevance. The mechanized statement and proof appear in the development as Reduction_AllCoords.lean (the lemma all_coords_relevant_of_not_tautology).
The next section develops the major practical consequence of this framework: the Simplicity Tax Theorem.
8 Corollary: Complexity Conservation
A quantitative consequence of the hardness results: when a model handles fewer dimensions than required, the gap must be paid at each use site.
Definition 8.1.
Let $R(P)$ be the required dimensions (those affecting Opt) and $A(M)$ the dimensions model $M$ handles natively. The expressive gap is $\text{Gap}(M,P)=R(P)\setminus A(M)$.
Theorem 8.2 (Conservation).
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced—only redistributed between “handled natively” and “handled externally.”
Theorem 8.3 (Linear Growth).
For $n$ decision sites: $\text{TotalExternalWork}=n\times|\text{Gap}(M,P)|$.
Theorem 8.4 (Amortization).
Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists $n^{*}=H_{\text{central}}/|\text{Gap}|$ such that for $n>n^{*}$, the complete model has lower total cost.
Corollary 8.5.
Since identifying $R(P)$ is coNP-complete (Theorem 3.6), minimizing the expressive gap is also intractable.
These results are machine-checked in Lean 4 (HardnessDistribution.lean).
9 Related Work
9.1 Computational Decision Theory
The complexity of decision-making has been studied extensively. Papadimitriou [12] established foundational results on the complexity of game-theoretic solution concepts. Our work extends this to the meta-question of identifying relevant information. For a modern treatment of complexity classes, see Arora and Barak [2].
Closest prior work and novelty.
Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [3, amaldi1998complexity]. Those results analyze predictive relevance or compression objectives. We study decision relevance and show coNP-completeness for sufficiency checking, a different quantifier structure with distinct proof techniques and a full hardness/tractability landscape under explicit encoding assumptions, mechanized in Lean 4. The formalization aspect is also novel: prior work establishes hardness on paper, while we provide machine-checked reductions with explicit polynomial bounds.
9.2 Feature Selection
In machine learning, feature selection asks which input features are relevant for prediction. This is known to be NP-hard in general [3]. Our results show the decision-theoretic analog is coNP-complete for both checking and minimization.
9.3 Value of Information
The value of information (VOI) framework [8] quantifies the maximum rational payment for information. Our work addresses a different question: not the value of information, but the complexity of identifying which information has value.
9.4 Model Selection
Statistical model selection (AIC [1], BIC [15], cross-validation [17]) provides practical heuristics for choosing among models. Our results provide theoretical justification: optimal model selection is intractable, so heuristics are necessary.
10 Conclusion
Methodology and Disclosure
Role of LLMs in this work. This paper was developed through human-AI collaboration. The author provided the core intuitions—the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET. Large language models (Claude, GPT-4) served as implementation partners for proof drafting, Lean formalization, and generation.
The Lean 4 proofs were iteratively refined: the author specified the target statements, the LLM proposed proof strategies, and the Lean compiler served as the arbiter of correctness. The complexity-theoretic reductions required careful human oversight to ensure the polynomial bounds were correctly established.
What the author contributed: The problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the hardness conjectures, the tractability conditions, and the connection to over-modeling in engineering practice.
What LLMs contributed: drafting, Lean tactic exploration, reduction construction assistance, and prose refinement.
The proofs are machine-checked; their validity is independent of generation method. We disclose this methodology in the interest of academic transparency.
Summary of Results
This paper establishes the computational complexity of coordinate sufficiency problems:
  • SUFFICIENCY-CHECK is coNP-complete (Theorem 3.6)
  • MINIMUM-SUFFICIENT-SET is coNP-complete (Theorem 3.7)
  • ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorem 3.9)
  • An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ (Theorem 4.1)
  • Tractable subcases exist for explicit-state encoding, separable utility, and tree-structured utility with explicit local factors (Theorem 5.1)
These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
Proofs are machine-checked in Lean 4 (see Appendix A for proof listings). The formalization verifies the reduction mappings and equivalence theorems; complexity classifications follow from standard results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete) under the encoding model of Section 2.4. In particular, the strengthened gadget showing that non-tautologies yield instances with all coordinates relevant (hence $k^{*}=\Omega(n)$) is also mechanized (see Appendix A).
Complexity Characterization
The results provide precise complexity characterizations within the formal model:
  1. 1.
    Exact bounds. SUFFICIENCY-CHECK is coNP-complete—both coNP-hard and in coNP.
  2. 2.
    Constructive reductions. The reductions from TAUTOLOGY and $\exists\forall$-SAT are explicit and machine-checked.
  3. 3.
    Encoding-regime separation. Under the explicit-state encoding, SUFFICIENCY-CHECK is polynomial in $|S|$. Under ETH, there exist succinctly encoded worst-case instances (witnessed by a strengthened gadget family with $k^{*}=n$) requiring $2^{\Omega(n)}$ time. Intermediate regimes are not ruled out by the lower-bound statement.
The Complexity Conservation Law
Section 8 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site. For $n$ sites, total external work is $(k-j)\times n$.
This conservation law is formalized in Lean 4 (HardnessDistribution.lean), proving:
  • Conservation: complexity cannot be eliminated, only redistributed
  • Dominance: complete models have lower total work than incomplete models
  • Amortization: there exists a threshold $n^{*}$ beyond which higher-dimensional models have lower total cost
Open Questions
Several questions remain for future work:
  • Fixed-parameter tractability: Is SUFFICIENCY-CHECK FPT when parameterized by the size of the minimal sufficient set?
  • Average-case complexity: What is the complexity under natural distributions on decision problems?
  • Quantum complexity: Does quantum computation provide speedups for sufficiency checking?
  • Learning cost formalization: Can central cost $H_{\text{central}}$ be formalized as the rank of a concept matroid, making the amortization threshold precisely computable?
Practical interpretation
While this work is theoretical, the complexity bounds have practical meaning. Under ETH, worst-case instances of SUFFICIENCY-CHECK and related problems grow exponentially with problem size; practitioners should therefore not expect polynomial‑time worst‑case solvers for arbitrary inputs. In practice, real datasets often have structure (sparsity, bounded dependency depth, separable utilities) that the tractable subcases capture. Thus, engineering efforts should focus on (a) detecting and exploiting structural restrictions in inputs, and (b) designing heuristic or approximate methods for large unstructured instances. Our experiments illustrate performance on representative inputs, but the theoretical bounds underline the need for domain constraints to achieve scalable, worst‑case guarantees.
Appendix A Lean 4 Proof Listings
The complete Lean 4 formalization is included in the appendix and companion artifact (Zenodo DOI is listed on the paper title page).
The formalization is not a transcription; it exposes a reusable reduction library with compositional polynomial bounds, enabling later mechanized hardness proofs to reuse these components directly.
A.1 On the Nature of Definitional Proofs
The Lean proofs follow directly from the formal definitions and standard complexity-theoretic constructions. The mechanization provides machine-checked verification and precision.
Definitional vs. derivational proofs. The core theorems establish definitional properties and reduction constructions. For example, the polynomial reduction composition theorem (Theorem A.1) proves that composing two polynomial-time reductions yields a polynomial-time reduction. The proof follows from the definition of polynomial time: composing two polynomials yields a polynomial.
Precedent in complexity theory. This pattern occurs throughout classical complexity theory:
  • Cook-Levin Theorem (1971): SAT is NP-complete. The proof constructs a reduction from an arbitrary NP problem to SAT. The construction itself is straightforward (encode Turing machine computation as boolean formula), but the insight is recognizing that SAT captures all of NP.
  • Ladner’s Theorem (1975): If P $\neq$ NP, then NP-intermediate problems exist. The proof is a diagonal construction—conceptually simple once the right framework is identified.
  • Toda’s Theorem (1991): The polynomial hierarchy is contained in P#P. The proof uses counting arguments that are elegant but not technically complex. The profundity is in the connection between counting and the hierarchy.
Why simplicity indicates strength. A definitional theorem derived from precise formalization is stronger than an informal argument. When we prove that sufficiency checking is coNP-complete (Theorem 3.6), we are not saying “we tried many algorithms and they all failed.” We are saying something general: any algorithm solving sufficiency checking solves TAUTOLOGY, and vice versa. The proof is a reduction construction that follows from the problem definitions.
Where the insight lies. The semantic contribution of our formalization is:
  1. 1.
    Precision forcing. Formalizing “coordinate sufficiency” in Lean requires stating exactly what it means for a coordinate subset to contain all decision-relevant information. This precision eliminates ambiguity about edge cases (what if projections differ only on irrelevant coordinates?).
  2. 2.
    Reduction correctness. The TAUTOLOGY reduction (Section 3) is machine-checked to preserve the decision structure. Complexity reductions are subtle and easy to mis-specify. We therefore mechanize the definitions and key reductions in Lean to rule out common formalization mistakes.
  3. 3.
    Encoding-regime separation. Theorem 4.1 contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ requiring $2^{\Omega(n)}$ time. Intermediate regimes are not ruled out by the lower-bound statement.
What machine-checking guarantees. The Lean compiler verifies that each proof step is valid and each definition is consistent; the mechanization does not introduce additional axioms beyond the chosen foundations. The mechanization contains no sorry placeholders. Reviewers can independently verify the proofs by running lake build in the mechanization directory and inspecting the companion artifact.
Comparison to informal complexity arguments. Prior work on model selection complexity (Chickering et al. [4], Teyssier & Koller [18]) presents compelling informal arguments but lacks machine-checked proofs. Our contribution is not new wisdom—the insight that model selection is hard is old. Our contribution is formalization: making “coordinate sufficiency” precise enough to mechanize, constructing verified reductions, and proving the complexity results hold for the formalized definitions.
This follows the tradition of verified complexity theory: just as Nipkow & Klein [11] formalized automata theory and Cook [6] formalized NP-completeness in proof assistants, we formalize decision-theoretic complexity. The proofs are simple because the formalization makes the structure clear. Simple proofs from precise definitions are the goal, not a limitation.
A.2 Module Structure
The formalization consists of 33 files organized as follows:
  • Basic.lean – Core definitions (DecisionProblem, CoordinateSet, Projection)
  • AlgorithmComplexity.lean – Complexity definitions (polynomial time, reductions)
  • PolynomialReduction.lean – Polynomial reduction composition (Theorem A.1)
  • Reduction.lean – TAUTOLOGY reduction for sufficiency checking
  • Reduction_AllCoords.lean – Strengthened gadget: if $\varphi$ is not a tautology then every coordinate is relevant
  • Hardness/ – Counting complexity and approximation barriers
  • Tractability/ – Bounded actions, separable utilities, tree structure, FPT
  • Economics/ – Value of information and elicitation connections
  • Dichotomy.lean and ComplexityMain.lean – Summary results
  • HardnessDistribution.lean – Simplicity Tax Theorem (Section 8)
A.3 Key Theorems
Theorem A.1 (Polynomial Composition, Lean).
Polynomial-time reductions compose to polynomial-time reductions.
theorem PolyReduction.comp_exists
    (f : PolyReduction A B) (g : PolyReduction B C) :
    exists h : PolyReduction A C,
      forall a, h.reduce a = g.reduce (f.reduce a)
Theorem A.2 (Simplicity Tax Conservation, Lean).
The simplicity tax plus covered axes equals required axes (partition).
theorem simplicityTax_conservation :
    simplicityTax P T + (P.requiredAxes inter T.nativeAxes).card
      = P.requiredAxes.card
Theorem A.3 (Simplicity Preference Fallacy, Lean).
Incomplete tools always cost more than complete tools for $n>0$ use sites.
theorem simplicity_preference_fallacy (T_simple T_complex : Tool)
    (h_simple_incomplete : isIncomplete P T_simple)
    (h_complex_complete : isComplete P T_complex)
    (n : Nat) (hn : n > 0) :
    totalExternalWork P T_complex n < totalExternalWork P T_simple n
A.4 Verification Status
The mechanization and its verification status are documented in the companion artifact; reviewers may consult the artifact for exact counts and files. The proofs compile and can be verified via the provided build instructions.

### v2

@@
 […]
-All results are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems).
+The reduction constructions and key equivalence theorems are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems); complexity classifications follow by composition with standard results (see Appendix A).
 Keywords: computational complexity, decision theory, polynomial hierarchy, tractability dichotomy, Lean 4
@@
 1.3 Machine-Checked Proofs
-All results are formalized in Lean 4 [7] ($\sim$5,000 lines, 200+ theorems).
-The formalization verifies the reduction mappings and combinatorial lemmas; complexity class membership follows by composition with TAUTOLOGY and $\exists\forall$-SAT.
+The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [6] ($\sim$5,000 lines, 200+ theorems).
+The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets.
+Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
 What is new.
-We contribute (i) a reusable Lean 4 framework for polynomial-time reductions with explicit polynomial bounds; (ii) the first machine-checked coNP-completeness proof for a decision-theoretic sufficiency problem; and (iii) a complete complexity landscape for coordinate sufficiency under explicit encoding assumptions.
-Prior work studies decision complexity in general or feature selection hardness, but does not formalize these reductions or establish this landscape in Lean.
+We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; and (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions.
+Prior work establishes hardness informally; we formalize the constructions.
 1.4 Paper Structure
@@
 Closest prior work and novelty.
-Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [3, amaldi1998complexity].
+Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [4, 2].
 Those results analyze predictive relevance or compression objectives.
@@
 These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
-Proofs are machine-checked in Lean 4 (see Appendix A for proof listings).
-The formalization verifies the reduction mappings and equivalence theorems; complexity classifications follow from standard results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete) under the encoding model of Section 2.4.
-In particular, the strengthened gadget showing that non-tautologies yield instances with all coordinates relevant (hence $k^{*}=\Omega(n)$) is also mechanized (see Appendix A).
+The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings).
+The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete).
+The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized.
 Complexity Characterization
@@
 Appendix A Lean 4 Proof Listings
-The complete Lean 4 formalization is included in the appendix and companion artifact (Zenodo DOI is listed on the paper title page).
-The formalization is not a transcription; it exposes a reusable reduction library with compositional polynomial bounds, enabling later mechanized hardness proofs to reuse these components directly.
-A.1 On the Nature of Definitional Proofs
-The Lean proofs follow directly from the formal definitions and standard complexity-theoretic constructions. The mechanization provides machine-checked verification and precision.
-Definitional vs.
-derivational proofs.
-The core theorems establish definitional properties and reduction constructions.
-For example, the polynomial reduction composition theorem (Theorem A.1) proves that composing two polynomial-time reductions yields a polynomial-time reduction.
-The proof follows from the definition of polynomial time: composing two polynomials yields a polynomial.
-Precedent in complexity theory. This pattern occurs throughout classical complexity theory:
-  • Cook-Levin Theorem (1971): SAT is NP-complete.
-The proof constructs a reduction from an arbitrary NP problem to SAT.
-The construction itself is straightforward (encode Turing machine computation as boolean formula), but the insight is recognizing that SAT captures all of NP.
-  • Ladner’s Theorem (1975): If P $\neq$ NP, then NP-intermediate problems exist. The proof is a diagonal construction—conceptually simple once the right framework is identified.
-  • Toda’s Theorem (1991): The polynomial hierarchy is contained in P#P.
-The proof uses counting arguments that are elegant but not technically complex.
-The profundity is in the connection between counting and the hierarchy.
-Why simplicity indicates strength.
-A definitional theorem derived from precise formalization is stronger than an informal argument.
-When we prove that sufficiency checking is coNP-complete (Theorem 3.6), we are not saying “we tried many algorithms and they all failed.” We are saying something general: any algorithm solving sufficiency checking solves TAUTOLOGY, and vice versa.
-The proof is a reduction construction that follows from the problem definitions.
-Where the insight lies. The semantic contribution of our formalization is:
+The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of approximately 5,000 lines across 33 files.
+A.1 What Is Machine-Checked
+The Lean formalization establishes:
   1. 1.
-    Precision forcing.
-Formalizing “coordinate sufficiency” in Lean requires stating exactly what it means for a coordinate subset to contain all decision-relevant information.
-This precision eliminates ambiguity about edge cases (what if projections differ only on irrelevant coordinates?).
+    Correctness of the TAUTOLOGY reduction: The theorem tautology_iff_sufficient proves that the mapping from Boolean formulas to decision problems preserves the decision structure (accept iff the formula is a tautology).
   2. 2.
-    Reduction correctness.
-The TAUTOLOGY reduction (Section 3) is machine-checked to preserve the decision structure.
-Complexity reductions are subtle and easy to mis-specify.
-We therefore mechanize the definitions and key reductions in Lean to rule out common formalization mistakes.
+    Decision problem definitions: Formal definitions of sufficiency, optimality, and the decision quotient.
   3. 3.
-    Encoding-regime separation.
-Theorem 4.1 contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ requiring $2^{\Omega(n)}$ time.
-Intermediate regimes are not ruled out by the lower-bound statement.
-What machine-checking guarantees.
-The Lean compiler verifies that each proof step is valid and each definition is consistent; the mechanization does not introduce additional axioms beyond the chosen foundations.
-The mechanization contains no sorry placeholders.
-Reviewers can independently verify the proofs by running lake build in the mechanization directory and inspecting the companion artifact.
-Comparison to informal complexity arguments.
-Prior work on model selection complexity (Chickering et al.
-[4], Teyssier & Koller [18]) presents compelling informal arguments but lacks machine-checked proofs.
-Our contribution is not new wisdom—the insight that model selection is hard is old.
-Our contribution is formalization: making “coordinate sufficiency” precise enough to mechanize, constructing verified reductions, and proving the complexity results hold for the formalized definitions.
-This follows the tradition of verified complexity theory: just as Nipkow & Klein [11] formalized automata theory and Cook [6] formalized NP-completeness in proof assistants, we formalize decision-theoretic complexity.
-The proofs are simple because the formalization makes the structure clear.
-Simple proofs from precise definitions are the goal, not a limitation.
-A.2 Module Structure
-The formalization consists of 33 files organized as follows:
-  • Basic.lean – Core definitions (DecisionProblem, CoordinateSet, Projection)
-  • AlgorithmComplexity.lean – Complexity definitions (polynomial time, reductions)
-  • PolynomialReduction.lean – Polynomial reduction composition (Theorem A.1)
-  • Reduction.lean – TAUTOLOGY reduction for sufficiency checking
-  • Reduction_AllCoords.lean – Strengthened gadget: if $\varphi$ is not a tautology then every coordinate is relevant
-  • Hardness/ – Counting complexity and approximation barriers
-  • Tractability/ – Bounded actions, separable utilities, tree structure, FPT
-  • Economics/ – Value of information and elicitation connections
-  • Dichotomy.lean and ComplexityMain.lean – Summary results
-  • HardnessDistribution.lean – Simplicity Tax Theorem (Section 8)
-A.3 Key Theorems
-Theorem A.1 (Polynomial Composition, Lean).
-Polynomial-time reductions compose to polynomial-time reductions.
-theorem PolyReduction.comp_exists
-    (f : PolyReduction A B) (g : PolyReduction B C) :
-    exists h : PolyReduction A C,
-      forall a, h.reduce a = g.reduce (f.reduce a)
-Theorem A.2 (Simplicity Tax Conservation, Lean).
-The simplicity tax plus covered axes equals required axes (partition).
+    Economic theorems: The Simplicity Tax conservation laws and hardness distribution results.
+Complexity classifications (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follow from informal composition with standard results (TAUTOLOGY is coNP-complete, etc.).
+The Lean proofs verify the reduction constructions; the complexity class membership is derived by combining these with established theorems from complexity theory.
+A.2 Polynomial-Time Reduction Definition
+We use Mathlib’s Turing machine framework to define polynomial-time computability:
+/-- Polynomial-time computable function using Turing machines -/
+def PolyTime {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
+    (f :   ) : Prop :=
+  Nonempty (Turing.TM2ComputableInPolyTime ea eb f)
+/-- Polynomial-time many-one (Karp) reduction -/
+def ManyOneReducesPoly {  : Type} (ea : FinEncoding ) (eb : FinEncoding )
+    (A : Set ) (B : Set ) : Prop :=
+   f :   , PolyTime ea eb f   x, x  A  f x  B
+This uses the standard definition: a reduction is polynomial-time computable via Turing machines and preserves membership.
+A.3 The Main Reduction Theorem
+Theorem A.1 (TAUTOLOGY Reduction Correctness, Lean).
+The reduction from TAUTOLOGY to SUFFICIENCY-CHECK is correct:
+theorem tautology_iff_sufficient ( : Formula n) :
+    .isTautology  (reductionProblem ).isSufficient Finset.empty
+This theorem is proven by showing both directions:
+  • If $\varphi$ is a tautology, then the empty coordinate set is sufficient
+  • If the empty coordinate set is sufficient, then $\varphi$ is a tautology
+The proof verifies that the utility construction in reductionProblem creates the appropriate decision structure where:
+  • At reference states, accept is optimal with utility 1
+  • At assignment states, accept is optimal iff $\varphi(a)=\text{true}$
+A.4 Economic Results
+The hardness distribution theorems (Section 8) are fully formalized:
 theorem simplicityTax_conservation :
@@
       = P.requiredAxes.card
-Theorem A.3 (Simplicity Preference Fallacy, Lean).
-Incomplete tools always cost more than complete tools for $n>0$ use sites.
 theorem simplicity_preference_fallacy (T_simple T_complex : Tool)
@@
     totalExternalWork P T_complex n < totalExternalWork P T_simple n
-A.4 Verification Status
-The mechanization and its verification status are documented in the companion artifact; reviewers may consult the artifact for exact counts and files.
-The proofs compile and can be verified via the provided build instructions.
+A.5 Module Structure
+  • Basic.lean – Core definitions (DecisionProblem, sufficiency, optimality)
+  • Sufficiency.lean – Sufficiency checking algorithms and properties
+  • Reduction.lean – TAUTOLOGY reduction construction and correctness
+  • Complexity.lean – Polynomial-time reduction definitions using mathlib
+  • HardnessDistribution.lean – Simplicity Tax theorems
+  • Tractability/ – Bounded actions, separable utilities, tree structure
+A.6 Verification
+The proofs compile with Lean 4 and contain no sorry placeholders. Run lake build in the proof directory to verify.

### v3

@@
-Computational Complexity of Sufficiency in Decision Problems
+Computational Complexity of Physical Information Sufficiency
 Abstract
-We characterize the computational complexity of coordinate sufficiency in decision problems within the formal model.
-Given action set $A$, state space $S=X_{1}\times\cdots\times X_{n}$, and utility $u:A\times S\to\mathbb{R}$, a coordinate set $I$ is sufficient if $s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$.
-The landscape in the formal model:
-  • General case: SUFFICIENCY-CHECK is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
-  • Tractable cases: Polynomial-time for bounded action sets under the explicit-state encoding; separable utilities ($u=f+g$) under any encoding; and tree-structured utility with explicit local factors.
-  • Encoding-regime separation: Polynomial-time under the explicit-state encoding (polynomial in $|S|$).
-Under ETH, there exist succinctly encoded worst-case instances witnessed by a strengthened gadget construction (mechanized in Lean; see Appendix A) with $k^{*}=n$ for which SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
-The tractable cases are stated with explicit encoding assumptions (Section 2.4).
-Together, these results answer the question “when is decision-relevant information identifiable efficiently?” within the stated regimes.
-[…]
-The reduction constructions and key equivalence theorems are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems); complexity classifications follow by composition with standard results (see Appendix A).
-Keywords: computational complexity, decision theory, polynomial hierarchy, tractability dichotomy, Lean 4
-footnotetext: © 2026 Tristan Simas. This work is licensed under $C$ BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
-Zenodo DOI: 10.5281/zenodo.18140965
+We study physical information sufficiency as a decision-theoretic meta-problem.
+For $\mathcal{D}=(A,S,U)$ with factored state space $S=X_{1}\times\cdots\times X_{n}$, a coordinate set $I$ is sufficient when
+$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
+This asks whether projected information preserves the full optimal-action correspondence.
+Complexity landscape.
+  • SUFFICIENCY-CHECK is coNP-complete; MINIMUM-SUFFICIENT-SET is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
+  • Under explicit-state encoding, polynomial-time algorithms hold for bounded actions, separable utility, and tree-structured utility.
+  • Under succinct encoding, hardness is regime-dependent: with ETH, there are worst-case families with $k^{*}=n$ requiring $2^{\Omega(n)}$ time.
+  • Under query access, the finite-state core has worst-case Opt-oracle complexity $\Omega(|S|)$, with Boolean value-entry and state-batch refinements preserving the obstruction.
+Physical grounding.
+The paper formalizes a physical-to-core encoding $E:\mathcal{P}\to\mathcal{D}$ and a transport rule: declared physical assumptions transfer to core assumptions, and core claims lift back to encoded physical instances.
+Encoded physical counterexamples induce core failures on the encoded slice.
+Discrete-time interface semantics (decision event = one tick) and budgeted thermodynamic lifts (bit lower bounds to energy/carbon lower bounds under declared constants) are formalized in the same assumption-typed framework.
+[C:P2.34,2.36,2.28,2.29,2.30,2.31,5.18,5.28;C2.35,5.22;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541,PBC1,CC65,DQ569.)
+All theorem-level claims are machine-checked in Lean 4 (13391 lines, 597 theorem/lemma statements).
+Complexity-class completeness follows by composition with standard complexity results; regime-dependent and physical-transport consequences are proved as assumption-explicit closures.
+Keywords: computational complexity, physical information, decision theory, polynomial hierarchy, formal verification
+© 2026 Tristan Simas. This work is licensed under CC BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
 1 Introduction
-Consider a decision problem with actions $A$ and states $S=X_{1}\times\cdots\times X_{n}$.
+Consider any decision problem $\mathcal{D}$ with actions $A$ and factored state space $S=X_{1}\times\cdots\times X_{n}$.
 A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
 $s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
-This paper characterizes the efficient cases of coordinate sufficiency within the formal model:
+This is coordinate sufficiency, the formalization of information sufficiency (whether available information determines the optimal action) over a factored state space.
+In information-theoretic terms, this is a task-conditional coding problem: the object preserved by projection is the decision boundary $s\mapsto\text{Opt}(s)$, not the full source state.
+The same object appears in classical source-coding and fidelity viewpoints [57, 59, 14, 64], and in zero-error/confusability structure where exact recoverability depends on graph constraints rather than raw description length [58, 31, 38, 15].
+The same separation also appears in model selection: short descriptions and certifiable adequacy are distinct [54, 24, 35, 8].
+For physical instantiations, the same tuple has a direct substrate reading: coordinates are measurable state degrees of freedom, actions are physically realizable interventions or transitions, and the optimizer map is the decision boundary induced by the declared objective contract.
+The physical-to-core encoding discipline in Section 2.7 makes this bridge theorem-level rather than metaphorical.[C:P2.34,2.36,2.28,2.29,2.30,2.31;C2.35;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541.)
+Compact claim-stamp notation.
+We use compact provenance stamps of the form $C:\cdots;R:\cdots]$: claim references are grouped by type ($D/P/C/T/L/R$ for Definition/Proposition/Corollary/Theorem/Lemma/Remark), and regime tags are abbreviated (for example ARAR=active declared regime, DCDC=declared contract, CRCR=declared class/regime, QfQf=query-finite, QbQb=query-Boolean).
+Theorem-strength narrative statements are stamped and, where available, paired with compact Lean handle IDs; unstamped narrative remains explanatory or citation-only context.
+SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY are meta-problems: each is parameterized by an arbitrary decision problem $\mathcal{D}$ and asks a structural question about $\mathcal{D}$’s coordinate space.
+The complexity results therefore apply to every problem with factored state space, not to a specific problem domain.[C:T4.6,4.7,4.9;R:S] (L: CC61,CC36,CC5.)
+Remark 1.1 (Universality of the Decision-Problem Tuple).
+The formal object of this paper is the decision-problem tuple $\mathcal{D}=(A,S,U)$: actions, factored states, utility.
+Within this model family, any problem requiring a choice among alternatives given structured information instantiates the tuple: the alternatives are $A$, the structured information is $S=X_{1}\times\cdots\times X_{n}$, and the objective is $U$.
+Thus the qualifier “decision problem” does not narrow scope inside the declared coordinate model.
+The meta-problem results below are universal over problems representable with this coordinate structure; the $(A,S,U)$ formalism supplies the language for stating the sufficiency predicate.
+This paper characterizes the complexity landscape of these meta-problems within the formal model:
+The paper keeps two layers explicit throughout: a substrate-neutral decision core (for complexity classification) and an assumption-typed physical instantiation layer (for transport to concrete systems).
+This separation prevents conflating representational hardness with physical assumptions while preserving a formal bridge between them.[C:P2.34;T2.37;R:AR] (L: CT6,CT4.)
 Section 2.4 fixes the computational model and input encodings used for all complexity claims.
+Section 2.5 gives the model contract and regime tags used to type every strong claim.
+Completeness claims are class-internal: complete for the static sufficiency class fixed by C1–C4 and the declared access regimes.
+Adjacent objective classes that violate the bridge conditions are typed separately (Section 2.5).
 Problem Complexity When Tractable
@@
 The tractable cases are stated with explicit encoding assumptions (Section 2.4). Outside those regimes, the succinct model yields hardness.
+1.1 Hardness–Recovery Reading Map
+Theorems in this paper are intended to be read as hardness/recovery pairs for the same fixed decision question:
+  1. 1.
+    Sufficiency checking: [S] coNP-complete in the general succinct regime (Theorem 4.6), paired with polynomial-time recovery under explicit-state and structured-access regimes (Theorem 6.1).
+  2. 2.
+    Minimal sufficient sets: [S] coNP-complete in general (Theorem 4.7), paired with the collapse criterion and relevance-computation recovery route (Theorems 4.14, 6.1).
+  3. 3.
+    Anchor sufficiency: $\Sigma_{2}^{\textsf{P}}$-complete in the general anchored formulation (Theorem 4.9); no general polynomial-time recovery condition is established in this model.
+(L: CC60, CC36, CC34, CC4, CC71, S2P2.)
+An operational criterion follows later from the same chain: once decision-site count exceeds the amortization threshold $n^{*}$, avoiding structural identification is more expensive than paying the one-time centralization cost (Theorem 9.6).[C:T9.6;R:LC] (L: HD3, HD21.)
 1.2 Landscape Summary
@@
     Tree-structured utility: With explicit local factors over a tree, dynamic programming yields polynomial algorithms in the input length.
-Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 3.6).
+Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 4.6). [C:T4.6,6.1;R:E,S] (L: CC61,CC69,CC70,CC71.)
 When is it intractable? The general problem is coNP-complete (Theorem 4.6), with a separation between explicit-state tractability and succinct worst-case hardness:
@@
 In particular, instances with $k^{*}=O(\log|S|)$ are tractable in this model.
+  • In the intermediate query-access model: the finite-state core yields an Opt-oracle lower bound $\Omega(|S|)$ (Boolean instantiation: $\Omega(2^{n})$), and the mechanized Boolean interface refinements preserve the obstruction for value-entry and state-batch interfaces, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness theorems (Propositions 5.35.16).
   • In the succinct model: under ETH there exist worst-case instances produced by our strengthened gadget in which the minimal sufficient set has size $\Omega(n)$ (indeed $n$) and SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
-The lower-bound statement does not address intermediate regimes.
+The explicit ETH lower bound is still a succinct worst-case statement; Proposition 5.3 provides the finite-state query core and Propositions 5.75.16 provide theorem-level Boolean-interface refinements.
+[C:P5.3,5.7,5.8,5.11,5.15,5.16;R:Qf,Qb] (L: CC50,L9,L11,L8,L21,L20.)
 1.3 Main Theorems
@@
     Theorem 6.1: Polynomial algorithms for bounded actions, separable utility, tree structure.
+  6. 6.
+    Theorem 2.37: physical-system bridge composition (law-objective semantics, behavior-preserving configuration reduction, and required-work lower bound) is proved as one bundled mechanized statement under declared encoding assumptions.
+  7. 7.
+    Propositions 2.282.31: the interface-time model is discrete, each decision event advances time by one unit, run-length equals elapsed interface time, and the unit-time law is substrate-invariant under the declared interface-consistency condition.
+[C:T4.6,4.7,4.9,5.1,6.1,2.37;P2.28,2.29,2.30,2.31;R:E,S,S+ETH,AR] (L: CC61,CC36,CC5,CC16,CC69,CC70,CC71,CT4,DQ548,DQ537,DQ539,DQ541.)
+1.4 Claim-Typing Theorems
+  1. 1.
+    Proposition 3.16: report admissibility is typed by certificate class ($\mathsf{ABSTAIN}$, $\mathsf{EXACT}$, $\mathsf{EPSILON}(\varepsilon)$).
+  2. 2.
+    Proposition 3.18: first-class evidence-object existence is equivalent to typed report admissibility.
+  3. 3.
+    Propositions 2.5 and 2.6: the $I=\emptyset$ query is exactly universal constancy of the decision boundary, and non-sufficiency is exactly existence of a counterexample pair.
+  4. 4.
+    Proposition 3.21: exact admissibility is equivalent to exact evidence existence.
+  5. 5.
+    Proposition 3.23 and Corollary 3.25: certified confidence is evidence-gated; for exact reports in no-competence regimes, certified confidence must be zero.
+  6. 6.
+    Proposition 3.24: under signal consistency, no evidence implies zero certified confidence.
+  7. 7.
+    Proposition 3.29 and Proposition 3.30: abstain may carry tentative guesses and self-reflected confidence, but self-confidence alone does not certify exactness.
+  8. 8.
+    Propositions 3.26, 3.27, 3.28: scalar step-count witnesses are always exact/falsifiable, while completion percentages are meaningful only under explicit declared bounds.
+  9. 9.
+    Theorem 3.56: outside declared carve-outs, admissible exact-claim policy requires explicit assumptions.
+  10. 10.
+    Proposition 3.61: under non-collapse, integrity and exact competence cannot both hold universally on hardness-blocked declared regimes.
+  11. 11.
+    Proposition 3.62 and Corollary 3.63: schema-level consequence for declared class/regime objects and exact-report inadmissibility.
+[C:T3.56;P3.16,3.18,2.5,2.6,3.21,3.23,3.24,3.29,3.30,3.26,3.27,3.28,3.61,3.62;C3.25,3.63;R:DC,AR] (L: CC75,IC2,DP6,DP7,IC35,IC38,IC40,IC46,IC47,IC42,IC44,IC45,CC30,CC15,CC38.)
+1.5 Regime Separation Theorems
+  1. 1.
+    Theorem 5.1: same-question separation between explicit-state and succinct regimes (with [S+ETH] hard-family lower bound).
+  2. 2.
+    Proposition 5.3: finite-state Opt-oracle lower bound in [Q_fin].
+  3. 3.
+    Propositions 5.75.16: Boolean-interface refinements and transfer/strictness in [Q_bool].
+[C:T5.1;P5.3,5.7,5.11,5.8,5.15,5.16;R:E,S+ETH,Qf,Qb] (L: CC16,CC50,L9,L8,L11,L21,L20,L23.)
 1.6 Machine-Checked Proofs
-The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [6] ($\sim$5,000 lines, 200+ theorems).
+The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [16, 17, 67, 21, 32] (13391 lines, 597 theorem/lemma statements).
 The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets.
@@
 What is new.
-We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; and (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions.
-Prior work establishes hardness informally; we formalize the constructions.
+We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions; and (iv) a formal separation between structural complexity and representational hardness used to derive theorem-indexed engineering corollaries, including theorem-level typed coverage/completeness closures for the declared static class (Theorems 2.38, 3.4).
+Prior work establishes hardness informally; we formalize the constructions and their regime-typed consequences.
+[C:T2.38,3.4,4.6,4.9;R:CR,S] (L: CC14,CC76,CC61,CC5.)
 1.7 Paper Structure
 […]
-Section 2: foundations.
+Section 2: decision-problem foundations and encoding model.
+Section 3: structural vs representational hardness; integrity vs competence.
 Section 4: hardness proofs.
-Section 4: dichotomy.
+Section 5: regime separation.
 Section 6: tractable cases.
-Sections 7 and 8: corollaries and implications for practice.
+Section 7: regime-conditional corollaries.
+Section 8: dominance theorems for hardness placement.
+Section 9: conservation law for decision-relevant coordinates.
 Section 10: related work.
 Appendix A: Lean listings.
+Sections 79 are not applied commentary.
+Every claim in those sections is a formal corollary, regime-typed by the framework of Section 3 and indexed to a machine-checked theorem handle.
+They complete the complexity landscape by characterizing when exact minimization is rational and when it is not: the rationality matrix (Proposition 3.60), the dominance and conservation theorems (Theorems 8.16, 9.3), and the amortization threshold (Theorem 9.6) are all mechanized in Lean.
+[C:P3.60;T8.16,9.3,9.6;R:AR,LC] (L: IC27,HD1,HD4,HD3.)
 2 Formal Foundations
 We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [55, 52].
+The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium.
+Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.32;P2.34;T2.37;R:AR] (L: CT6,CT4.)
 2.1 Decision Problems with Coordinate Structure
@@
 $\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
-Equivalently, the optimal action depends only on coordinates in $I$.
+Equivalently, the optimal action depends only on coordinates in $I$. In this paper, this set-valued invariance of the full optimal-action correspondence is the primary decision-relevance target.
+Proposition 2.5 (Empty-Set Sufficiency Equals Constant Decision Boundary).
+For any decision problem,
+$\text{sufficient}(\emptyset)\iff\forall s,s^{\prime}\in S,\ \text{Opt}(s)=\text{Opt}(s^{\prime}).$
+Hence, the $I=\emptyset$ query is exactly a universal constancy check of the decision boundary. (L: DP6.)
+Proposition 2.6 (Insufficiency Equals Counterexample Witness).
+For any coordinate set $I$,
+$\neg\text{sufficient}(I)\iff\exists s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
+In particular,
+$\neg\text{sufficient}(\emptyset)\iff\exists s,s^{\prime}\in S:\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
+(L: DP7, DP8.)
+[C:D2.4;P2.5,2.6;R:DM]
+Definition 2.7 (Deterministic Selector).
+A deterministic selector is a map $\sigma:2^{A}\to A$ that returns one action from an optimal-action set.
+Definition 2.8 (Selector-Level Sufficiency).
+For fixed selector $\sigma$, a coordinate set $I$ is selector-sufficient if
+$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\sigma(\text{Opt}(s))=\sigma(\text{Opt}(s^{\prime})).$
+Proposition 2.9 (Set-Level Target Dominates Selector Targets).
+If $I$ is sufficient in the set-valued sense (Definition 2.4), then $I$ is selector-sufficient for every deterministic selector $\sigma$ (Definition 2.8). (L: DP5)
+Proof.
+If $I$ is sufficient, then $s_{I}=s^{\prime}_{I}$ implies $\text{Opt}(s)=\text{Opt}(s^{\prime})$.
+Applying any deterministic selector $\sigma$ to equal optimal-action sets yields equal selected actions.
+$\blacksquare$
+Definition 2.10 ($\varepsilon$-Optimal Set and $\varepsilon$-Sufficiency).
+For $\varepsilon\geq 0$, define
+$\text{Opt}_{\varepsilon}(s):=\{a\in A:\forall a^{\prime}\in A,\ U(a^{\prime},s)\leq U(a,s)+\varepsilon\}.$
+Coordinate set $I$ is $\varepsilon$-sufficient if
+$\forall s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\implies\text{Opt}_{\varepsilon}(s)=\text{Opt}_{\varepsilon}(s^{\prime}).$
+Proposition 2.11 (Zero-$\varepsilon$ Reduction).
+Set-valued sufficiency is exactly the $\varepsilon=0$ case:
+[…]
+(L: DP1, DP4)
+Proposition 2.12 (Selector-Level Separation Witness).
+The converse of Proposition 2.9 fails in general: there exists a decision problem, selector, and coordinate set $I$ such that $I$ is selector-sufficient but not sufficient in the full set-valued sense.
+(L: CC54.)
 Definition 2.13 (Minimal Sufficient Set).
@@
 Definition 2.14 (Relevant Coordinate).
-Coordinate $i$ is relevant if it belongs to some minimal sufficient set.
+Coordinate $i$ is relevant if there exist states that differ only at coordinate $i$ and induce different optimal action sets:
+$i\text{ is relevant}\iff\exists s,s^{\prime}\in S:\;\Big(\forall j\neq i,\;s_{j}=s^{\prime}_{j}\Big)\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
+Proposition 2.15 (Minimal-Set/Relevance Equivalence).
+For any minimal sufficient set $I$ and any coordinate $i$:
+$i\in I\iff i\text{ is relevant}.$
+Hence every minimal sufficient set is exactly the relevant-coordinate set. (L: DP2, DP3.)
+Proof.
+The “only if” direction follows by minimality: if $i\in I$ were irrelevant, removing $i$ would preserve sufficiency, contradicting minimality.
+The “if” direction follows from sufficiency: every sufficient set must contain each relevant coordinate.
+$\blacksquare$
+Definition 2.16 (Exact Relevance Identifiability).
+For a decision problem $\mathcal{D}$ and candidate coordinate set $I$, we say $I$ is exactly relevance-identifying if
+$\forall i\in\{1,\ldots,n\}:\quad i\in I\iff i\text{ is relevant for }\mathcal{D}.$
+Equivalently, $I$ is exactly relevance-identifying iff $I$ equals the full relevant-coordinate set.
 Example 2.17 (Weather Decision).
@@
 Proposition 2.20 (Sufficiency Characterization).
-Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$.
+[C:P2.20;R:H=succinct-hard] Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$. (L: CC62, CC63)
 Proof.
@@
 We fix the computational model used by the complexity claims.
+In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
 Succinct encoding (primary for hardness).
-This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [3].
+This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [5, 47].
 An instance is encoded as:
@@
 The input length is $L=|A|+\sum_{i}\log|X_{i}|+|C_{U}|$.
-Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH, W[2]) are measured in $L$.
+Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH) are measured in $L$.
 All hardness results in Section 4 use this encoding.
@@
 Results stated in terms of $|S|$ use this encoding.
+Query-access regime (intermediate black-box access).
+The solver is given oracle access to decision information at queried states (e.g., $\text{Opt}(s)$, or $U(a,s)$ with $\text{Opt}(s)$ reconstructed from finitely many value queries).
+Complexity is measured by oracle-query count (optionally paired with per-query evaluation cost).
+This separates representational access from full-table availability and from succinct-circuit input length.
 Unless explicitly stated otherwise, “polynomial time” refers to the succinct encoding.
+Remark 2.21 (Decision Questions vs Decision Problems).
+We distinguish between a decision question and a decision problem.
+A decision question is the encoding-independent predicate: “is coordinate set $I$ sufficient for $\mathcal{D}$?” This is a mathematical relation, fixed by the model contract (C1–C4).
+A decision problem is a decision question together with a specific encoding of its inputs.
+The same decision question yields different decision problems under different encodings, and these problems may belong to different complexity classes.
+Formally, each encoding regime $e$ induces a language $L_{\mathcal{D},e}$; complexity classification is a property of $L_{\mathcal{D},e}$, not of $\mathcal{D}$ alone.
+Proposition 2.22 (Under-Resolution Forces Representation Collision).
+[C:P2.22;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states.
+If $|\mathcal{C}|<|\mathcal{L}|$, then there exist distinct logical states with the same representation:
+$\exists s,s^{\prime}\in\mathcal{L},\ s\neq s^{\prime}\wedge\mathrm{repr}(s)=\mathrm{repr}(s^{\prime}).$
+This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: DQ582, DQ583.)
+Definition 2.23 (Regime Simulation).
+Fix a decision question $Q$ and two regimes $R_{1},R_{2}$ for that question.
+We say $R_{1}$ simulates $R_{2}$ if there exists a polynomial-time transformer that maps any $R_{2}$ instance to an $R_{1}$ instance while preserving the answer to $Q$; in oracle settings, this includes a polynomial-time transcript transducer that preserves yes/no outcomes for induced solvers.
+(L: CC1, CC52.)
+This paper uses that simulation spine in two explicit forms: restriction-map simulation (subproblem-to-full transfer; Proposition 5.8) and oracle-transducer simulation (batch-to-entry transfer; Proposition 5.15).
+2.5 Model Contract and Regime Tags
+All theorem statements in this paper are typed by the following model contract:
+  • C1 (finite actions): $A$ is finite.
+  • C2 (finite coordinate domains): each $X_{i}$ is finite, so $S=\prod_{i}X_{i}$ is finite.
+  • C3 (evaluable utility): $U(a,s)$ is computable from the declared input encoding.
+  • C4 (fixed decision semantics): optimality is defined by $\text{Opt}(s)=\arg\max_{a}U(a,s)$.
+We use short regime tags for applied corollaries:
+  • [E] explicit-state encoding,
+  • [Q] query-access (oracle) regime,
+  • [Q_fin] finite-state query-access core (state-oracle),
+  • [S] succinct encoding,
+  • [S+ETH] succinct encoding with ETH,
+  • [Q_bool] mechanized Boolean query submodel,
+  • [S_bool] mechanized Boolean-coordinate submodel.
+This tagging is a claim-typing convention: each strong statement is attached to the regime where it is proven.
+Remark 2.24 (Uniform Assumption Discipline).
+All theorem-level claims in this paper are conditional on their stated premises and regime tags; no assumption family is given special status.
+For each claim, the active assumptions are exactly those appearing in its statement, its regime tag, and its cited mechanized handles.
+Remark 2.25 (Modal Convention: Contract-Conditional Necessity).
+In this paper, normative modal statements denote contract-conditional necessity: once an agent declares an objective/regime/assumption contract, admissible policy is fixed by typed admissibility and integrity under that contract.
+No unconditional external moral prescription is asserted.
+2.6 Discrete-Time Event Semantics
+Definition 2.26 (Timed Interface State).
+For any interface state space $S$, a timed interface state is a pair
+$x=(s,t)\in S\times\mathbb{N}.$
+The time coordinate is the natural-number component $t$.
+Definition 2.27 (Decision Tick and Decision Event).
+Fix an interface decision process $(\mathrm{decide},\mathrm{transition})$ with
+$\mathrm{decide}:S\to A,\qquad\mathrm{transition}:S\times A\to S.$
+Define one tick by
+$\mathrm{tick}(s,t)=\bigl(\mathrm{transition}(s,\mathrm{decide}(s)),\,t+1\bigr).$
+Define $\mathrm{DecisionEvent}(x,y)$ to mean that $y$ is the one-tick successor of $x$ under this rule.
+Proposition 2.28 (Time Coordinate Is Discrete).
+In Definition 2.26, time is discrete:
+$\forall x\in S\times\mathbb{N},\ \exists k\in\mathbb{N}\ \text{such that}\ x_{t}=k.$
+Pointwise time equality is decidable for every $k\in\mathbb{N}$:
+$x_{t}=k\ \vee\ x_{t}\neq k.$
+(L: DQ548,DQ547.)
+Proposition 2.29 (Decision Event Is One Unit of Time).
+Under Definition 2.27, for all timed states $x,y$:
+$\mathrm{DecisionEvent}(x,y)\ \Longleftrightarrow\ y=\mathrm{tick}(x),$
+and therefore
+$\mathrm{DecisionEvent}(x,y)\ \Longrightarrow\ y_{t}=x_{t}+1.$
+(L: DQ535,DQ537,DQ545,DQ546.)
+Proposition 2.30 (Run Accounting Identity).
+Let $\mathrm{run}(n,x)$ be $n$ repeated ticks from $x$, and let $\mathrm{trace}(n,x)$ be the emitted decision sequence. Then:
+$\mathrm{run}(n,x)_{t}=x_{t}+n,\qquad\mathrm{run}(n,x)_{t}-x_{t}=n,\qquad|\mathrm{trace}(n,x)|=n.$
+Hence decision count equals elapsed time:
+$|\mathrm{trace}(n,x)|=\mathrm{run}(n,x)_{t}-x_{t}.$
+(L: DQ540,DQ539,DQ533,DQ534.)
+Proposition 2.31 (Substrate Realization Preserves Unit-Time Events).
+For any substrate model that is consistent with the declared interface transition rule, each one-step substrate evolution realizes one interface decision event and advances interface time by one unit.
+The time update law is invariant under substrate kind.
+(L: DQ542,DQ541,DQ549.)
+2.7 Physical-System Encoding and Theorem Transport
+Definition 2.32 (Physical-to-Core Encoding).
+Fix a physical instance class $\mathcal{P}$ and a core decision class $\mathcal{D}$. A physical-to-core encoding is a map
+$E:\mathcal{P}\to\mathcal{D}.$
+All physical claims below are typed relative to a declared encoding map $E$ and declared assumption transfer map from physical assumptions to core assumptions.
+Proposition 2.33 (Heisenberg-Strong Binding Yields Nontrivial Decision Boundary).
+[C:P2.33;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ.
+Then there exists a core decision problem with non-constant optimizer map.
+Equivalently, under this declared physical ambiguity assumption, decision nontriviality is derivable at the core level.
+(L: DQ584, DQ585, DQ586.
+)
+Proposition 2.34 (Conditional Physical Claim Transport).
+Let $E:\mathcal{P}\to\mathcal{D}$ be an encoding. If
+$\forall d\in\mathcal{D},\ \mathrm{CoreAssm}(d)\Rightarrow\mathrm{CoreClaim}(d)$
+and
+$\forall p\in\mathcal{P},\ \mathrm{PhysAssm}(p)\Rightarrow\mathrm{CoreAssm}(E(p)),$
+then
+$\forall p\in\mathcal{P},\ \mathrm{PhysAssm}(p)\Rightarrow\mathrm{CoreClaim}(E(p)).$
+(L: CT6, CT5, DQ572.)
+Corollary 2.35 (Physical Counterexample Implies Core Failure on Encoded Slice).
+Under the same encoding and assumption-transfer map:
+  1. 1.
+    any encoded physical counterexample induces a core counterexample on the encoded slice;
+  2. 2.
+    if a physical counterexample exists, then the purported core rule is invalid;
+  3. 3.
+    if the lifted claim holds on all physically admissible encoded instances, no such physical counterexample exists.
+(L: CT8, CT7, CT3, DQ571.)
+Proposition 2.36 (Law-Instance Objective Bridge).
+For the mechanized law-gap physical instance class, the encoded optimizer claim holds at theorem level and admits no counterexample:
+$\forall x\in\mathcal{P}_{\mathrm{law}},\ \mathrm{lawGapPhysicalClaim}(x),\qquad\neg\exists x\in\mathcal{P}_{\mathrm{law}},\ \neg\,\mathrm{lawGapPhysicalClaim}(x).$
+(L: CT1, CT2.)
+Theorem 2.37 (Physical Bridge Bundle).
+In one bundled mechanized statement, the paper’s physical bridge layer composes:
+  1. 1.
+    law-induced objective semantics ($\text{Opt}=$ feasible actions for law-gap instances),
+  2. 2.
+    behavior-preserving configuration reduction equivalence,
+  3. 3.
+    irreducible required-work lower bound under site-count scaling.
+(L: CT4.)
+[C:D2.32;P2.34,2.36;C2.35;T2.37;R:AR]
+Theorem 2.38 (Declared Regime Coverage for the Static Class).
+Let
+$\mathcal{R}_{\mathrm{static}}=\{\text{[E]},\ \text{[S+ETH]},\ \text{[Q\_fin:Opt]},\ \text{[Q\_bool:value-entry]},\ \text{[Q\_bool:state-batch]}\}.$
+For each declared regime in $\mathcal{R}_{\mathrm{static}}$, the paper has a theorem-level mechanized core claim.
+Equivalently, regime typing is complete over the declared static family.
+(L: CC14, CC51.)
+Scope Lattice (typed classes and transfer boundary).
+Layer Transfer Status Lean handles (representative)
+Static sufficiency class (C1–C4; declared regimes) Internal landscape complete CC76
+CC51
+Bridge-admissible adjacent class (one-step deterministic) Transfer licensed CC42
+CC10
+Non-admissible represented adjacent classes (horizon, stochastic, transition-coupled) Transfer blocked by witness CC9
+CC8
+2.8 Adjacent Objective Regimes and Bridge
+Definition 2.39 (Adjacent Sequential Objective Regime).
+An adjacent sequential objective instance consists of:
+  • finite action set $A$,
+  • finite coordinate state space $S=X_{1}\times\cdots\times X_{n}$,
+  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent policy class,
+  • reward process $r_{t}$ and objective functional $J(\pi)$ (e.g., cumulative reward or regret).
+Proposition 2.40 (One-Step Deterministic Bridge).
+Consider an instance of Definition 2.39 satisfying:
+  1. 1.
+    $T=1$,
+  2. 2.
+    deterministic rewards $r_{1}(a,s)=U(a,s)$ for some evaluable $U$,
+  3. 3.
+    objective $J(\pi)=U(\pi(s),s)$ (single-step maximization),
+  4. 4.
+    no post-decision state update relevant to the objective.
+Then the induced optimization problem is exactly the static decision problem of Definition 2.1, and coordinate sufficiency in the sequential formulation is equivalent to Definition 2.4. (L: CC42)
+Proof.
+Under (1)–(3), optimizing $J$ at state $s$ is identical to choosing an action in $\arg\max_{a\in A}U(a,s)=\text{Opt}(s)$.
+Condition (4) removes any dependence on future transition effects.
+Therefore the optimal-policy relation in the adjacent formulation coincides pointwise with Opt from Definition 2.3, and the invariance condition “same projection implies same optimal choice set” is exactly Definition 2.4.
+$\blacksquare$
+Proposition 2.41 (Bridge Transfer Rule).
+Under conditions (1)–(4), any sufficiency statement formulated over Definition 2.4 is equivalent between the adjacent sequential formulation and the static model. (L: CC42)
+Proof.
+By Proposition 2.40, the adjacent sequential objective induces exactly the same sufficiency predicate as Definition 2.4 when (1)–(4) hold.
+Equivalence of any sufficiency statement then follows by substitution.
+$\blacksquare$
+If any bridge condition fails, direct transfer from this paper’s static complexity theorems is not licensed by this rule.
+Remark 2.42 (Extension Boundary).
+Beyond Proposition 2.40 (multi-step horizon, stochastic transitions/rewards, or regret objectives), the governing complexity objects change.
+Those regimes are natural extensions, but they are distinct formal classes from the static sufficiency class analyzed in this paper.
+Proposition 2.43 (Horizon->1>1 Bridge-Failure Witness).
+Dropping the one-step condition can break sufficiency transfer: there exists a two-step objective where sufficiency in the immediate-utility static projection does not match sufficiency in the two-step objective.
+(L: CC26, CC25.)
+Proposition 2.44 (Stochastic-Criterion Bridge-Failure Witness).
+If optimization is performed against a stochastic criterion not equal to deterministic utility maximization, bridge transfer to Definition 2.1 can fail even when an expected-utility projection is available.
+(L: CC57.)
+Proposition 2.45 (Transition-Coupled Bridge-Failure Witness).
+If post-decision transition effects are objective-relevant, bridge transfer to the static one-step class can fail. (L: CC73.)
+Theorem 2.46 (Exact Bridge Boundary in the Represented Adjacent Family).
+Within the represented adjacent family
+$\{\text{one-step deterministic},\ \text{horizon-extended},\ \text{stochastic-criterion},\ \text{transition-coupled}\},$
+direct transfer of static sufficiency claims is licensed if and only if the class is one-step deterministic.
+Each represented non-one-step class has an explicit mechanized bridge-failure witness.
+(L: CC10, CC9, CC8.)
+Definition 2.47 (Agent Snapshot vs Agent Process Typing).
+Within the represented adjacent family, we type:
+  • agent snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
+  • agent process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
+Proposition 2.48 (Snapshot/Process Transfer Typing).
+In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the agent is typed as an agent snapshot.
+Every process-typed represented class has an explicit mechanized bridge-failure witness.
+(L: CC3, CC48, CC56.)
+Operationally: a trained model evaluated with fixed parameters during inference is an agent snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.48;R:RA]
+3 Interpretive Foundations: Hardness and Solver Claims
+The claims in later applied sections are theorem-indexed consequences of this section and Sections 46.
+3.1 Structural Complexity vs Representational Hardness
+Definition 3.1 (Structural Complexity).
+For a fixed decision question (e.g., “is $I$ sufficient for $\mathcal{D}$?”; see Remark 2.21), structural complexity means its placement in standard complexity classes within the formal model (coNP, $\Sigma_{2}^{P}$, etc.), as established by class-membership arguments and reductions.
+Definition 3.2 (Representational Hardness).
+For a fixed decision question and an encoding regime $E$ (Section 2.4), representational hardness is the worst-case computational cost incurred by solvers whose input access is restricted to $E$.
+In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: DQ562,DQ560,DQ561,DQ563,DQ564,DQ565.)
+Remark 3.3 (Interpretation Contract).
+This paper keeps the decision question fixed and varies the encoding regime explicitly.
+Thus, later separations are read as changes in representational hardness under fixed structural complexity, not as changes to the underlying sufficiency semantics.
+Accordingly, “complete landscape” means complete for this static sufficiency class under the declared regimes; adjacent objective classes are distinct typed objects, not unproved remainder cases of the same class.
+Theorem 3.4 (Typed Completeness for the Static Sufficiency Class).
+For the static sufficiency class fixed by C1–C4 and declared regime family $\mathcal{R}_{\mathrm{static}}$, the paper provides:
+  1. 1.
+    conditional class-label closures for SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY;
+  2. 2.
+    a conditional explicit/succinct dichotomy closure;
+  3. 3.
+    a regime-indexed mechanized core claim for every declared regime; and
+  4. 4.
+    declared-family exhaustiveness in the typed regime algebra.
+(L: CC76.)
+3.2 Solver Integrity and Regime Competence
+To keep practical corollaries type-safe, we separate integrity (what a solver is allowed to assert) from competence (what it can cover under a declared regime), following the certifying-algorithms schema [39, 43].
+The induced-relation view used below is standard in complexity/computability presentations: algorithms are analyzed as machines deciding languages or computing functions/relations over encodings [47, 5, 21].
+Definition 3.5 (Certifying Solver).
+Fix a decision question $\mathcal{R}\subseteq\mathcal{X}\times\mathcal{Y}$ and an encoding regime $E$ over $\mathcal{X}$. A certifying solver is a pair $(Q,V)$ where:
+  • $Q$ maps each input $x\in\mathcal{X}$ to either $\mathsf{ABSTAIN}$ or a candidate pair $(y,w)$,
+  • $V$ is a polynomial-time checker (in $|{\rm enc}_{E}(x)|$) with output in $\{0,1\}$.
+Definition 3.6 (Induced Solver Relation).
+For any deterministic program $M$ computing a partial map $f_{M}:\mathcal{X}\rightharpoonup\mathcal{Y}$ on encoded inputs, define
+$\mathcal{R}_{M}:=\{(x,y)\in\mathcal{X}\times\mathcal{Y}:f_{M}(x)\downarrow\ \wedge\ y=f_{M}(x)\}.$
+Proposition 3.7 (Universality of Solver Framing).
+Every deterministic computer program that computes a (possibly partial) map on encoded inputs can be framed as a solver for its induced relation (Definition 3.6).
+In this formal sense, all computers are solvers of specific problems.
+(L: CC77.)
+Proof.
+Immediate from Definition 3.6: each program induces a relation given by its graph on the domain where it halts.  $\blacksquare$
+Definition 3.8 (Solver Integrity).
+A certifying solver $(Q,V)$ has integrity for relation $\mathcal{R}$ if:
+  • (assertion soundness) $Q(x)=(y,w)\implies V(x,y,w)=1$,
+  • (checker soundness) $V(x,y,w)=1\implies(x,y)\in\mathcal{R}$.
+The output $\mathsf{ABSTAIN}$ (equivalently, $\mathsf{UNKNOWN}$) is first-class and carries no assertion about membership in $\mathcal{R}$.
+Definition 3.9 (Competence Under a Regime).
+Fix a regime $\Gamma=(\mathcal{X}_{\Gamma},E_{\Gamma},\mathcal{C}_{\Gamma})$ with instance family $\mathcal{X}_{\Gamma}\subseteq\mathcal{X}$, encoding assumptions $E_{\Gamma}$, and resource bound $\mathcal{C}_{\Gamma}$.
+A certifying solver $(Q,V)$ is competent on $\Gamma$ for relation $\mathcal{R}$ if:
+  • it has integrity for $\mathcal{R}$ (Definition 3.8),
+  • (coverage) $\forall x\in\mathcal{X}_{\Gamma},\;Q(x)\neq\mathsf{ABSTAIN}$,
+[…]
+Corollary 3.10 (Universality of Integrity Applicability).
+Let $M$ be any deterministic program, viewed as a certifying solver pair $(Q,V)$ for some externally fixed target relation $\mathcal{R}$.
+Then Definition 3.8 applies unchanged.
+Thus epistemic integrity is architecture-universal over computing systems once they are cast as solvers for declared tasks.
+(L: CC31.)
+Proof.
+Definition 3.8 quantifies over a relation $\mathcal{R}$ and a certifying pair $(Q,V)$; it does not assume any special implementation substrate.  $\blacksquare$
+[C:P3.7;C3.10;R:TR]
+Remark 3.11 (External-Task Non-Vacuity).
+If the target relation is chosen post hoc from the program’s own behavior (for example, $\mathcal{R}=\mathcal{R}_{M}$ from Definition 3.6), integrity can become tautological and competence claims can be vacuous.
+Non-vacuous competence claims therefore require the target relation, encoding regime, and resource bound to be declared independently of observed outputs.
+[C:D3.8,3.9;P3.58,3.61;R:AR]
+Definition 3.12 ($\varepsilon$-Objective Relation Family).
+An $\varepsilon$-objective relation family is a map
+$\varepsilon\mapsto\mathcal{R}_{\varepsilon}\subseteq\mathcal{X}\times\mathcal{Y}$
+that assigns a certified target relation to each tolerance level $\varepsilon\geq 0$. The exact target is $\mathcal{R}_{0}$.
+Definition 3.13 ($\varepsilon$-Competence Under a Regime).
+Fix a relation family $(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0}$ and regime $\Gamma$.
+A certifying solver $(Q,V)$ is $\varepsilon$-competent on $\Gamma$ if it is competent on $\Gamma$ for relation $\mathcal{R}_{\varepsilon}$ in the sense of Definition 3.9.
+Proposition 3.14 (Zero-$\varepsilon$ Competence Reduction).
+If $\mathcal{R}_{0}=\mathcal{R}$, then
+[…]
+Moreover, $\varepsilon$-competence implies integrity for the corresponding $\mathcal{R}_{\varepsilon}$. (L: IC33, IC19.)
+Definition 3.15 (Typed Claim Report).
+For a declared objective family and regime, a solver-side report is one of:
+  • $\mathsf{ABSTAIN}$,
+  • $\mathsf{EXACT}$ (claiming exact competence/certification),
+  • $\mathsf{EPSILON}(\varepsilon)$ (claiming $\varepsilon$-competence/certification).
+Proposition 3.16 (Typed Claim Admissibility Discipline).
+Under Definitions 3.9, 3.13, and 3.15:
+  • $\mathsf{ABSTAIN}$ is always admissible;
+  • $\mathsf{EXACT}$ is admissible iff exact competence holds;
+  • $\mathsf{EPSILON}(\varepsilon)$ is admissible iff $\varepsilon$-competence holds.
+Hence claim type and certificate type must match (no cross-typing of uncertified assertions as certified guarantees). (L: CC75.)
+Definition 3.17 (Evidence Object for a Typed Report).
+For declared $(\mathcal{R},(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0},\Gamma,Q)$ and report type $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$, evidence is a first-class witness object:
+  • for $\mathsf{ABSTAIN}$: trivial witness;
+  • for $\mathsf{EXACT}$: exact-competence witness;
+  • for $\mathsf{EPSILON}(\varepsilon)$: $\varepsilon$-competence witness.
+(L: IC2.)
+Proposition 3.18 (Evidence–Admissibility Equivalence).
+For any declared contract and report type $r$:
+$\text{Evidence object for }r\ \text{exists}\iff\text{Claim }r\ \text{is admissible}.$
+(L: IC17, IC21, IC20.)
+[C:D3.17;P3.16,3.18;R:AR]
+Security-game reading (derived).
+The typed-report layer can be read as a standard claim-verification game under a declared contract: a reporting algorithm emits a report token $r$ and optional evidence, and a checker accepts if and only if the evidence is valid for $r$.
+Within this game view, existing results already provide the core security properties: completeness for admissible report classes (Proposition 3.18), soundness against overclaiming (Propositions 3.16, 3.21), and evidence-gated confidence (Propositions 3.23, 3.24).
+So integrity is enforced as verifiable claim discipline, not as unchecked payload declaration.
+[C:P3.16,3.18,3.21,3.23,3.24;R:AR]
+Definition 3.19 (Signaled Typed Report (Guess + Confidence)).
+A signaled typed report augments the typed report token $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$ with:
+  • an optional guess payload $g$,
+  • self-reflected confidence $p_{\mathrm{self}}\in[0,1]$,
+  • certified confidence $p_{\mathrm{cert}}\in[0,1]$.
+  • scalar execution witness $t_{\mathrm{run}}\in\mathbb{N}$ (steps actually run),
+  • optional declared bound $B\in\mathbb{N}$.
+Here $p_{\mathrm{self}}$ is a self-signal, while $p_{\mathrm{cert}}$ is an evidence-gated signal under the declared contract. (L: IC34.)
+Definition 3.20 (Signal Consistency Discipline).
+For a signaled typed report $(r,g,p_{\mathrm{self}},p_{\mathrm{cert}},t_{\mathrm{run}},B)$, require:
+$p_{\mathrm{cert}}>0\Rightarrow\exists\ \text{evidence object for }r.$
+Proposition 3.21 (Exact Claims Require Exact Evidence).
+In the typed claim discipline,
+$\text{ClaimAdmissible}(\mathsf{EXACT})\iff\exists\ \text{exact evidence object}.$
+Equivalently, admissible exact claims are exactly those with an evidence witness. (L: IC35, IC36.)
+Definition 3.22 (Bound-Dependent Completion Fraction).
+Completion-fraction semantics are defined only when a positive declared bound exists and covers observed runtime:
+$\mathrm{CompletionFractionDefined}\iff\exists b>0:\ B=b\ \wedge\ t_{\mathrm{run}}\leq b.$
+(L: IC37.)
+Proposition 3.23 (Certified-Confidence Gate).
+Under Definition 3.20, positive certified confidence implies typed admissibility:
+$p_{\mathrm{cert}}>0\Rightarrow\text{ClaimAdmissible}(r).$
+Thus certified confidence is not a free scalar; it is evidence-gated by the same typed discipline as the report itself. (L: IC38, IC39.)
+Proposition 3.24 (No Evidence Forces Zero Certified Confidence).
+Under Definition 3.20, absence of evidence for the reported type forces:
+$\neg\exists\ \text{evidence object for }r\Rightarrow p_{\mathrm{cert}}=0.$
+(L: IC40.)
+Corollary 3.25 (Exact-No-Competence Zero-Certified Constraint).
+For exact reports, if exact competence is unavailable on the declared regime/objective, then any signal-consistent report must satisfy:
+$p_{\mathrm{cert}}=0.$
+(L: IC41.)
+Proposition 3.26 (Scalar Step Witness Is Always Defined and Falsifiable).
+For every signaled report, the execution witness is an exact natural number and any equality claim about it is decidable:
+$\exists k\in\mathbb{N}:\ t_{\mathrm{run}}=k,\qquad\forall k\in\mathbb{N},\ (t_{\mathrm{run}}=k)\ \vee\ (t_{\mathrm{run}}\neq k).$
+(L: IC42, IC43.)
+Proposition 3.27 (No Completion Fraction Without Declared Bound).
+If no declared bound is provided ($B=\varnothing$), completion-fraction semantics are undefined:
+$B=\varnothing\Rightarrow\neg\,\mathrm{CompletionFractionDefined}.$
+(L: IC44.)
+Proposition 3.28 (Completion Fraction Under Explicit Bound).
+If a positive declared bound is provided and observed runtime is within bound, completion-fraction semantics are defined:
+$B=b,\ b>0,\ t_{\mathrm{run}}\leq b\Rightarrow\mathrm{CompletionFractionDefined}.$
+(L: IC45.)
+Proposition 3.29 (Abstain Can Carry Guess and Self-Confidence).
+For any optional guess payload $g$ and any self-reflected confidence $p_{\mathrm{self}}\in[0,1]$, there exists a signal-consistent abstain report
+$(\mathsf{ABSTAIN},g,p_{\mathrm{self}},0,0,\varnothing).$
+Hence the framework is non-binary at the report-signal layer: abstention can carry a tentative answer and self-reflection without upgrading to certified exactness. (L: IC46.)
+Proposition 3.30 (Self-Reflected Confidence Is Not Certification).
+Self-reflected confidence alone does not certify exactness: if exact competence is unavailable, there exist exact-labeled reports with arbitrary $p_{\mathrm{self}}$ that remain inadmissible under zero certified confidence.
+(L: IC47.)
+Definition 3.31 (Declared Budget Slice).
+For a declared regime $\Gamma$ over state objects and horizon/budget cut $H\in\mathbb{N}$, define the in-scope slice
+$\mathcal{S}_{\Gamma,H}:=\{s:\ s\in\Gamma.\mathrm{inScope}\ \wedge\ \mathrm{encLen}_{\Gamma}(s)\leq H\}.$
+(L: CC13.)
+Proposition 3.32 (Bounded-Slice Irrelevance of a Meta-Coordinate).
+Fix coordinate $i_{\infty}$ and declared slice $\mathcal{S}_{\Gamma,H}$ from Definition 3.31. If optimizer sets are invariant to $i_{\infty}$ on that slice, i.e.,
+$\forall s,s^{\prime}\in\mathcal{S}_{\Gamma,H},\ \Big(\forall j\neq i_{\infty},\ s_{j}=s^{\prime}_{j}\Big)\Rightarrow\text{Opt}(s)=\text{Opt}(s^{\prime}),$
+then $i_{\infty}$ is irrelevant on $\mathcal{S}_{\Gamma,H}$, and hence not relevant for the declared task slice. (L: CC32, CC33.)
+[C:D3.31;P3.32;R:AR]
+Definition 3.33 (Goal Class).
+A goal class is a non-empty set of admissible evaluators over a fixed action/state structure:
+$\mathcal{G}=(\mathcal{U}_{\mathcal{G}},A,S),\qquad\varnothing\neq\mathcal{U}_{\mathcal{G}}\subseteq\{U:A\times S\to\mathbb{Q}\}.$
+The solver need not know which $U\in\mathcal{U}_{\mathcal{G}}$ is active; it only knows class-membership constraints. (L: IV1.)
+Definition 3.34 (Interior Pareto Dominance).
+For a goal class $\mathcal{G}$ and score model $\sigma:S\times\{1,\ldots,n\}\to\mathbb{Q}$, let $J_{\mathcal{G}}$ be the class-tautological coordinate set.
+State $s$ interior-Pareto-dominates $s^{\prime}$ on $J_{\mathcal{G}}$ when:
+$\forall j\in J_{\mathcal{G}},\ \sigma(s^{\prime},j)\leq\sigma(s,j)\quad\text{and}\quad\exists j\in J_{\mathcal{G}},\ \sigma(s^{\prime},j)<\sigma(s,j).$
+(L: IV2, IV3.)
+Proposition 3.35 (Interior Dominance Implies Universal Non-Rejection).
+If a state $s$ interior-Pareto-dominates $s^{\prime}$ on a checked coordinate set $J$, and every admissible evaluator in the goal class is monotone on $J$, then no admissible evaluator strictly prefers $s^{\prime}$ over $s$ on that checked slice:
+$\forall U\in\mathcal{U}_{\mathcal{G}},\ \forall a\in A,\ U(a,s^{\prime})\leq U(a,s).$
+(L: IV5.)
+Proposition 3.36 (Interior Verification Tractability).
+[C:P3.36;R:H=tractable-structured] If membership in the checked coordinate set is decidable and interior dominance is decidable, then interior verification yields a polynomial-time checker with exact acceptance criterion:
+$\exists\ \mathrm{verify}:S\times S\to\{0,1\},\quad\mathrm{verify}(s,s^{\prime})=1\iff\text{interior-dominance}(s,s^{\prime}).$
+(L: IV7, IV4.)
+Proposition 3.37 (Interior Dominance Is One-Sided).
+Interior dominance is one-sided and does not imply full coordinate sufficiency: a counterexample pair outside the checked slice can still violate global optimizer equivalence. (L: IV6, IV8.)
+Corollary 3.38 (Singleton Interior Certificate).
+[C:C3.38;R:H=tractable-structured] For a singleton checked coordinate $i$, strict improvement on $i$ with class-monotonicity on $\{i\}$ is already a valid interior certificate of non-rejection.
+(L: IV9.)
+[C:D3.33,3.34;P3.35,3.36,3.37;C3.38;R:AR]
+Definition 3.39 (RLFF Objective (Evidence-Gated Report Reward)).
+Given report type $r\in\{\mathsf{ABSTAIN},\mathsf{EXACT},\mathsf{EPSILON}(\varepsilon)\}$ and declared contract, RLFF assigns base reward by report type when evidence exists and applies an explicit inadmissibility penalty otherwise:
+$\mathrm{Reward}(r)=\begin{cases}\mathrm{Base}(r),&\text{if evidence for }r\text{ exists},\\ -\mathrm{Penalty},&\text{otherwise}.\end{cases}$
+(L: IC4, IC28, IC29.)
+Proposition 3.40 (RLFF Maximizer Is Admissible).
+If $\mathsf{ABSTAIN}$ reward strictly exceeds the inadmissibility branch, then any global RLFF maximizer must be evidence-backed and therefore admissible under the typed-claim discipline.
+(L: IC31, IC32.)
+Corollary 3.41 (RLFF Abstain Preference Without Certificates).
+If $\mathsf{EXACT}$ and $\mathsf{EPSILON}(\varepsilon)$ have no evidence objects while $\mathsf{ABSTAIN}$ beats the inadmissibility branch, then RLFF strictly prefers $\mathsf{ABSTAIN}$ to both report types.
+(L: IC30, DQ570.)
+[C:D3.39;P3.40;C3.41;R:AR]
+Definition 3.42 (Certainty Inflation).
+For a typed report $r$, certainty inflation is the state of emitting $r$ without a matching evidence object. (L: IC1, IC3.)
+Proposition 3.43 (Certainty Inflation Equals Report Inadmissibility).
+For every typed report $r$:
+$\text{CertaintyInflation}(r)\iff\neg\ \text{ClaimAdmissible}(r).$
+For $r=\mathsf{EXACT}$ this is equivalently failure of exact competence. (L: IC8, IC22.)
+Corollary 3.44 (Hardness-Blocked Regimes Induce Exact Certainty Inflation).
+Under the declared hardness/non-collapse premises that block universal exact admissibility, every exact report is certainty-inflated. (L: CC19.)
+[C:D3.42;P3.43;C3.44;R:CR]
+Definition 3.45 (Raw-vs-Certified Bit Accounting).
+Fix a declared contract and report type $r$. A report-bit model declares:
+  • raw report bits $B_{\mathrm{raw}}(r)$ (payload bits in the asserted report token),
+  • certificate bits $B_{\mathrm{cert}}(r)$ (bits allocated to the matching evidence class).
+Evidence-gated overhead and certified total are then:
+$B_{\mathrm{over}}(r)=\begin{cases}B_{\mathrm{cert}}(r),&\text{if evidence for }r\text{ exists},\\ 0,&\text{otherwise},\end{cases}\qquad B_{\mathrm{total}}(r)=B_{\mathrm{raw}}(r)+B_{\mathrm{over}}(r).$
+(L: IC5, IC9, IC12.)
+Proposition 3.46 (Raw/Certified Bit Split).
+For every typed report $r$:
+$B_{\mathrm{total}}(r)=B_{\mathrm{raw}}(r)+B_{\mathrm{over}}(r),$
+with
+$\neg\exists\ \text{evidence for }r\Rightarrow B_{\mathrm{over}}(r)=0,\qquad\exists\ \text{evidence for }r\Rightarrow B_{\mathrm{over}}(r)=B_{\mathrm{cert}}(r).$
+Hence $B_{\mathrm{raw}}(r)\leq B_{\mathrm{total}}(r)$ always, and strict inequality holds when evidence exists with positive certificate-bit allocation.
+(L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, DQ567.)
+Theorem 3.47 (Exact Report: Certified-Bit Gap iff Admissibility).
+Under a report-bit model with positive exact certificate-bit allocation:
+$\text{ClaimAdmissible}(\mathsf{EXACT})\iff B_{\mathrm{raw}}(\mathsf{EXACT})<B_{\mathrm{total}}(\mathsf{EXACT}).$
+Equivalently:
+$B_{\mathrm{total}}(\mathsf{EXACT})=B_{\mathrm{raw}}(\mathsf{EXACT})\iff\text{ExactCertaintyInflation}.$
+(L: CC18, CC20, CC17, DQ566.)
+Corollary 3.48 (Finite Budget Eventually Blocks Exact Admissibility).
+[C:C3.48;R:H=exp-lb-conditional] Fix a declared physical computer model with finite budget and positive per-bit cost, a declared operation-requirement lower-bound family, and a declared typed-report bit model.
+If:
+  1. 1.
+    exact admissibility implies certified-bit footprint at least the declared operation requirement at size $n$, and
+  2. 2.
+    exact admissibility implies physical feasibility of that certified footprint under the declared budget model,
+then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: DQ573, DQ574.)
+Corollary 3.49 (Hardness-Blocked Exact Reports Are Raw-Only).
+If exact reporting is inadmissible on the declared contract, then exact accounting collapses to raw-only:
+$B_{\mathrm{total}}(\mathsf{EXACT})=B_{\mathrm{raw}}(\mathsf{EXACT}).$
+(L: CC21, IC23.)
+[C:D3.45;T3.47;P3.46;C3.49;R:AR]
+Corollary 3.50 (No Uncertified Exact-Exhaustion Claim).
+If exact competence does not hold on the declared regime/objective, then an $\mathsf{EXACT}$ report is inadmissible. (L: CC41.)
+[C:P3.16;C3.50;R:AR]
+Proposition 3.51 (Declared-Contract Separation: Selection vs. Validity).
+For solver-side reporting, there are two formally distinct layers:
+  • Declared-contract selection layer: choosing objective family $(\mathcal{R}_{\varepsilon})_{\varepsilon\geq 0}$, regime $\Gamma$, and assumption profile;
+  • Validity layer: once declared, admissibility of $\mathsf{EXACT}/\mathsf{EPSILON}(\varepsilon)/\mathsf{ABSTAIN}$ reports is fixed by certificate-typed rules.
+Thus the framework treats contract selection as an external declaration choice, and then enforces mechanically checked admissibility within that declared contract. (L: CC75, CC22, CC39.)
+Proof.
+The validity layer is Proposition 3.16: report admissibility is exactly tied to the matching competence certificate type.
+For exact physical claims outside carve-outs, Proposition 3.55 and Theorem 3.56 require explicit assumptions, and Corollary 3.57 blocks uncertified exact reports under declared non-collapse assumptions.
+Therefore, once the contract is declared, admissibility is mechanically determined by the typed rules and attached assumption profile.
+$\blacksquare$
+[C:T3.56;P3.16,3.55;C3.57;R:DC]
+Definition 3.52 (Exact-Claim Excuse Carve-Outs).
+For exact physical claims in this framework, we declare four explicit carve-outs:
+  1. 1.
+    trivial scope (empty in-scope family),
+  2. 2.
+    tractable class (exact competence available in the active regime),
+  3. 3.
+    stronger oracle/regime shift (the claim is made under strictly stronger access assumptions),
+  4. 4.
+    unbounded resources (budget bounds intentionally removed).
+Definition 3.53 (Explicit Hardness-Assumption Profile).
+An explicit hardness-assumption profile for declared class/regime $(\mathcal{R},\Gamma)$ consists of:
+  • a non-collapse hypothesis,
+  • a collapse implication from universal exact competence on $(\mathcal{R},\Gamma)$.
+Definition 3.54 (Well-Typed Exact Physical Claim Policy).
+An exact physical claim is well-typed only if either:
+  • at least one carve-out from Definition 3.52 is explicitly declared, or
+  • an explicit hardness-assumption profile (Definition 3.53) is explicitly declared.
+Proposition 3.55 (Outside Carve-Outs, Explicit Assumptions Are Required).
+If an exact physical claim is well-typed (Definition 3.54) and no carve-out from Definition 3.52 applies, then an explicit hardness-assumption profile must be present. (L: CC22.)
+Theorem 3.56 (Claim-Integrity Meta-Theorem).
+For declared class/regime objects, every admissible exact-claim policy outside the carve-outs requires explicit assumptions:
+[…]
+[…]
+(L: CC22.)
+Proof.
+Immediate from Proposition 3.55; this is the universally quantified theorem-level restatement over the typed claim/regime objects.  $\blacksquare$
+Corollary 3.57 (Outside Carve-Outs, Declared Hardness Blocks Exact Reports).
+If no carve-out applies and the declared hardness-assumption profile holds, then $\mathsf{EXACT}$ reports are inadmissible for every solver on the declared class/regime. (L: CC39.)
+[C:T3.56;P3.55;C3.57;R:CR]
+Proposition 3.58 (Integrity–Competence Separation).
+Integrity and competence are distinct predicates: integrity constrains asserted outputs, while competence adds non-abstaining coverage under resource bounds. (L: IC18, IC25.)
+Proof.
+Take the always-abstain solver $Q_{\bot}(x)=\mathsf{ABSTAIN}$ with any polynomial-time checker $V$.
+Definition 3.8 holds vacuously, so $(Q_{\bot},V)$ is integrity-preserving, but it fails Definition 3.9 whenever $\mathcal{X}_{\Gamma}\neq\emptyset$ because coverage fails.
+Hence integrity does not imply competence.
+The converse is immediate because competence includes integrity as a conjunct.
+$\blacksquare$
+Definition 3.59 (Attempted Competence Failure).
+Fix an exact objective under regime $\Gamma$. A solver state is an attempted competence failure if:
+  • integrity holds (Definition 3.8),
+  • exact competence was actually attempted for the active scope/objective,
+  • competence on $\Gamma$ fails for that exact objective (Definition 3.9).
+Proposition 3.60 (Attempted-Competence Rationality Matrix).
+Let $I,A,C\in\{0,1\}$ denote integrity, attempted exact competence, and competence available in the active regime. Policy verdict for persistent over-specification is:
+  • if $I=0$: inadmissible,
+  • if $I=1$ and $(A,C)=(1,0)$: conditionally rational,
+  • if $I=1$ and $(A,C)\in\{(0,0),(0,1),(1,1)\}$: irrational for the same verified-cost objective.
+Hence, in the integrity-preserving plane ($I=1$), exactly one cell is rational and three are irrational. (L: IC27, IC7, IC6)
+This separation plus Proposition 3.60 is load-bearing for the regime-conditional trilemma used later: if exact competence is blocked by hardness in a declared regime after an attempted exact procedure, integrity forces one of three responses: abstain, weaken guarantees, or change regime assumptions.[C:P3.60;R:AR]
+Mechanized status.
+This separation is machine-checked in DecisionQuotient/IntegrityCompetence.lean via: IC18 and IC25; the attempted-competence matrix is mechanized via IC27, IC7, and IC6.
+Proposition 3.61 (Integrity-Resource Bound).
+Let $\Gamma$ be a declared regime whose in-scope family includes all instances of SUFFICIENCY-CHECK and whose declared resource bound is polynomial in the encoding length.
+If $\textsf{P}\neq\textsf{coNP}$, then no certifying solver is simultaneously integrity-preserving and competent on $\Gamma$ for exact sufficiency.
+Equivalently, for every integrity-preserving solver, at least one of the competence conjuncts must fail on $\Gamma$: either full non-abstaining coverage fails or the declared budget bound fails.
+(L: IC26, IC24, CC30.)
+Proof.
+[…]
+Proposition 3.62 (Declared-Physics No-Universal-Exact-Certifier Schema).
+Fix any declared physical/task class representable by relation $\mathcal{R}$ and regime $\Gamma$ in the certifying-solver model. If
+$\big(\exists Q,\ \mathrm{CompetentOn}(\mathcal{R},\Gamma,Q)\big)\ \Rightarrow\ (\textsf{P}=\textsf{coNP}),$
+then under $\textsf{P}\neq\textsf{coNP}$ there is no universally exact-competent solver for that declared class. (L: CC15.)
+Corollary 3.63 (No Universal Exact-Claim Admissibility in Hardness-Blocked Classes).
+Under the same premises as Proposition 3.62, and under the typed-claim discipline of Definition 3.15, an $\mathsf{EXACT}$ report is inadmissible for every solver on that declared class.
+Hence admissible reporting must use $\mathsf{ABSTAIN}$ or explicitly weakened guarantees.
+(L: CC38.)
+[C:P3.62;C3.63;R:CR]
 4 Computational Complexity of Decision-Relevant Uncertainty
-This section establishes the computational complexity of determining which state coordinates are decision-relevant. We prove three main results:
+This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space.
+Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Remark 1.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain.
+We prove three main results:
   1. 1.
@@
     ANCHOR-SUFFICIENCY (fixed coordinates) is $\Sigma_{2}^{P}$-complete
-Within the formal model (Section 2) and the succinct encoding used for hardness (Section 2.4), these results sit beyond NP-completeness and provide a rigorous explanation for why over-modeling can be a rational strategy: in the worst case, finding (or even certifying) a minimal set of decision-relevant factors is computationally intractable.
+Under the model contract of Section 2.5 and the succinct encoding of Section 2.4, these results place exact relevance identification beyond NP-completeness: in the worst case, finding (or certifying) minimal decision-relevant sets is computationally intractable.
+Reading convention for this section.
+Each hardness theorem below is paired with a recovery boundary for the same decision question: when structural-access assumptions in Theorem 6.1 hold, polynomial-time exact procedures are recovered; when they fail in [S], the stated hardness applies.
+(L: CC60, CC36, CC4, CC71.)
 4.1 Problem Definitions
@@
   • $U:A\times S\to\mathbb{Q}$ is the utility function (in the succinct encoding, $U$ is given as a Boolean circuit)
+Succinct/circuit encoding.
+When $U$ is given succinctly as a Boolean or arithmetic circuit, the instance size refers to the circuit description rather than the (possibly exponentially large) truth table.
+Evaluating $U(a,s)$ on a single pair $(a,s)$ can be done in time polynomial in the circuit size, but quantification over all states (for example, "for all ss") ranges over an exponentially large domain and is the source of the hardness results below.
 Definition 4.2 (Optimizer Map).
@@
 where $s_{I}$ denotes the projection of $s$ onto coordinates in $I$.
-Problem 3.4 (SUFFICIENCY-CHECK).
+Problem 4.4 (SUFFICIENCY-CHECK).
 Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I\subseteq\{1,\ldots,n\}$
 Question: Is $I$ sufficient?
-Problem 3.5 (MINIMUM-SUFFICIENT-SET).
+Problem 4.5 (MINIMUM-SUFFICIENT-SET).
 Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and integer $k$
@@
 4.2 Hardness of SUFFICIENCY-CHECK
+Classical coNP-completeness methodology follows standard NP/coNP reduction frameworks and polynomial-time many-one reducibility [47, 5].
+We instantiate that framework for SUFFICIENCY-CHECK with a mechanized TAUTOLOGY reduction.
 Theorem 4.6 (coNP-completeness of SUFFICIENCY-CHECK).
-SUFFICIENCY-CHECK is coNP-complete [5, 9]. (Machine-verified in Lean 4; see Appendix A.)
+SUFFICIENCY-CHECK is coNP-complete. (L: CC61, CC60, L18.)
 Source Target Key property preserved
@@
 $\exists\forall$-SAT ANCHOR-SUFFICIENCY Witness anchors iff formula true
-SET-COVER MINIMUM-SUFFICIENT-SET Set size maps to coordinate size
 Proof.
-Membership in coNP: The complementary problem INSUFFICIENCY is in NP.
-Given a decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I$, a witness for insufficiency is a pair $(s,s^{\prime})$ such that:
+extbfMembership in coNP: The complementary problem INSUFFICIENCY is in NP. A nondeterministic verifier can check a witness pair $(s,s^{\prime})$ as follows:
   1. 1.
-    $s_{I}=s^{\prime}_{I}$ (verifiable in polynomial time)
+    Verify $s_{I}=s^{\prime}_{I}$ by projecting and comparing coordinates; this takes time polynomial in the input size.
   2. 2.
-    $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$ (verifiable by evaluating $U$ on all alternatives)
+    Evaluate $U$ on every alternative for both $s$ and $s^{\prime}$.
+When $U$ is given as a circuit each evaluation runs in time polynomial in the circuit size; using these values the verifier checks whether $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$.
+Hence INSUFFICIENCY is in NP and SUFFICIENCY-CHECK is in coNP.
 coNP-hardness: We reduce from TAUTOLOGY.
@@
 $\blacksquare$
+Immediate recovery boundary (SUFFICIENCY-CHECK).
+Theorem 4.6 is the [S] general-case hardness statement.
+For the same sufficiency relation, Theorem 6.1 gives polynomial-time recovery under explicit-state, separable, and tree-structured regimes.
+(L: CC60, CC71, CC69, CC70, CC72.)
 Mechanized strengthening (all coordinates relevant).
@@
 Theorem 4.7 (MINIMUM-SUFFICIENT-SET is coNP-complete).
-MINIMUM-SUFFICIENT-SET is coNP-complete.
+MINIMUM-SUFFICIENT-SET is coNP-complete. (L: CC34, CC36.)
 Proof.
@@
 However, this collapses because sufficiency has a simple characterization.
-Key lemma: A coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_contains_relevant in Lean):
+Key lemma: In the Boolean-coordinate collapse model, a coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_iff_relevant_subset / sufficient_iff_relevantFinset_subset in Lean):
 $\text{sufficient}(I)\iff\text{Relevant}\subseteq I$
 […]
-Consequence: The minimum sufficient set is exactly the set of relevant coordinates. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
+Consequence: The minimum sufficient set is exactly the relevant-coordinate set. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
 coNP membership: A witness that the answer is NO is a set of $k+1$ coordinates, each proven relevant (by exhibiting $s,s^{\prime}$ pairs). Verification is polynomial.
@@
 $\blacksquare$
+Immediate recovery boundary (MINIMUM-SUFFICIENT-SET).
+Theorem 4.7 pairs with Theorem 4.14: exact minimization complexity is governed by relevance-cardinality once the collapse is applied.
+Recovery then depends on computing relevance efficiently, with structured-access assumptions from Theorem 6.1 giving the corresponding route for the underlying sufficiency computations.
+(L: CC36, CC34, CC35, S2P2, CC71.)
 4.4 Anchor Sufficiency (Fixed Coordinates)
 We also formalize a strengthened variant that fixes the coordinate set and asks whether there exists an assignment to those coordinates that makes the optimal action constant on the induced subcube.
-Problem 3.8 (ANCHOR-SUFFICIENCY).
+Problem 4.8 (ANCHOR-SUFFICIENCY).
 Input: Decision problem $(A,X_{1},\ldots,X_{n},U)$ and fixed coordinate set $I\subseteq\{1,\ldots,n\}$
 Question: Does there exist an assignment $\alpha$ to $I$ such that $\text{Opt}(s)$ is constant for all states $s$ with $s_{I}=\alpha$?
+The second-level hardness template used here follows canonical $\Sigma_{2}^{P}$ polynomial-hierarchy characterizations [47, 5], instantiated to the anchor formulation below.
 Theorem 4.9 (ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete).
-ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete [14] (already for Boolean coordinate spaces).
+ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete (already for Boolean coordinate spaces). (L: CC5, CC4, L2.)
 Proof.
@@
 If $\exists x^{\star}\,\forall y\,\varphi(x^{\star},y)=1$, then for any $y$ we have $U(\text{YES})=2$ and $U(\text{NO})\leq 1$, so $\text{Opt}(x^{\star},y)=\{\text{YES}\}$ is constant.
-Conversely, if $\varphi(x,y)$ is false for some $y$, then either $y=0^{m}$ (where NO is optimal) or $y\neq 0^{m}$ (where YES and NO tie), so the optimal set varies across $y$ and the subcube is not constant.
-Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.
-$\blacksquare$
+Conversely, fix $x$ and suppose $\exists y_{f}$ with $\varphi(x,y_{f})=0$.
+  • If $\varphi(x,0^{m})=1$, then $\text{Opt}(x,0^{m})=\{\text{YES}\}$.
+The falsifying assignment must satisfy $y_{f}\neq 0^{m}$, where $U(\text{YES})=U(\text{NO})=0$, so $\text{Opt}(x,y_{f})=\{\text{YES},\text{NO}\}$.
+  • If $\varphi(x,0^{m})=0$, then $\text{Opt}(x,0^{m})=\{\text{NO}\}$.
+After padding we have $m>0$, so choose any $y^{\prime}\neq 0^{m}$: either $\varphi(x,y^{\prime})=1$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES}\}$) or $\varphi(x,y^{\prime})=0$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES},\text{NO}\}$).
+In both subcases the optimal set differs from $\{\text{NO}\}$.
+Hence the subcube for this $x$ is not constant. Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.  $\blacksquare$
+Immediate boundary statement (ANCHOR-SUFFICIENCY).
+Theorem 4.9 remains a second-level hardness statement in the anchored formulation; unlike MINIMUM-SUFFICIENT-SET, no general collapse to relevance counting is established here, so the corresponding tractability status remains open in this model.
+(L: CC5, CC4, L2.)
 4.5 Tractable Subcases
@@
 4.6 Implications
-Corollary 3.12 (Why Over-Modeling Is Rational).
-Finding the minimal set of decision-relevant factors is coNP-complete. Even verifying that a proposed set is sufficient is coNP-complete.
-This formally explains the engineering phenomenon:
-  1. 1.
-    It’s computationally easier to model everything than to find the minimum
-  2. 2.
-    “Which unknowns matter?” is a hard question, not a lazy one to avoid
-  3. 3.
-    Bounded scenario analysis (small $\hat{S}$) makes the problem tractable
-This connects to the pentalogy’s leverage framework: the modeling budget for deciding what to model is itself a computationally constrained resource.
-3.7 Remark: The Collapse to coNP
-Early analysis of MINIMUM-SUFFICIENT-SET focused on the apparent $\exists\forall$ quantifier structure, which suggested a $\Sigma_{2}^{P}$-complete result.
-We initially explored certificate-size lower bounds for the complement, attempting to show MINIMUM-SUFFICIENT-SET was unlikely to be in coNP.
-However, the key insight—formalized as sufficient_contains_relevant—is that sufficiency has a simple characterization: a set is sufficient iff it contains all relevant coordinates.
-This collapses the $\exists\forall$ structure because:
-  • The minimum sufficient set is exactly the relevant coordinate set
-  • Checking relevance is in coNP (witness: two states differing only at that coordinate with different optimal sets)
-  • Counting relevant coordinates is also in coNP
-This collapse explains why ANCHOR-SUFFICIENCY retains its $\Sigma_{2}^{P}$-completeness: fixing coordinates and asking for an assignment that works is a genuinely different question.
-The “for all suffixes” quantifier cannot be collapsed when the anchor assignment is part of the existential choice.
+Corollary 4.12 (Computational Bottleneck for Exact Minimization).
+Under the succinct encoding, exact minimization of sufficient coordinate sets is coNP-hard via the $k=0$ case, and fixed-anchor minimization is $\Sigma_{2}^{P}$-complete. (L: CC36, CC4)
+Proof.
+The $k=0$ case of MINIMUM-SUFFICIENT-SET is SUFFICIENCY-CHECK (Theorem 4.6), giving coNP-hardness for exact minimization. The fixed-anchor variant is exactly Theorem 4.9.  $\blacksquare$
+The modeling budget for deciding what to model is therefore a computationally constrained resource under this encoding.[C:Thm. 4.6, Thm. 4.9;R:[S]]
+4.7 The succinct–minimal gap and physical agents
+In complexity theory, a succinct representation of a combinatorial object is a Boolean circuit of size $\mathrm{poly}(n)$ that computes a function whose explicit (truth-table) form has size $2^{\Theta(n)}$.
+The term is standard but informal: it denotes a size class (polynomial in input variables), not a structural property.
+In particular, “succinct” does not mean minimal.
+The minimal circuit for a function is the smallest circuit that computes it, a rigorous, unique object.
+A succinct circuit may be far larger than the minimal one while still qualifying as “succinct” by the poly-size criterion.
+The gap.
+Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function.
+The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the agent’s encoding is from the tightest possible encoding.
+This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
+Proposition 4.13 (Coordinate Representation Gap as a Formal Object).
+In the Boolean-coordinate model, let $R_{\mathrm{rel}}=\texttt{relevantFinset}(dp)$ and define
+$\varepsilon(I)\;:=\;|I\setminus R_{\mathrm{rel}}|+|R_{\mathrm{rel}}\setminus I|.$
+Then:
+$\varepsilon(I)=0\iff I=R_{\mathrm{rel}}\iff I\text{ is minimal sufficient}.$
+(L: S2P3, S2P4, S2P5.)
+The results of this paper give $\varepsilon$ formal complexity consequences.
+By Theorem 4.7, determining the minimal sufficient coordinate set is coNP-complete under the succinct encoding.
+By Theorem 5.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency.
+In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate (Lean: S2P.min_representationGap_zero_iff_relevant_card).
+Together: for an agent encoded as a succinct circuit, the worst-case cost of determining whether its own coordinate attention set is minimal is exponential in the number of input coordinates (under ETH).
+The gap $\varepsilon$ is not merely unnamed; it is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
+Physical agents as circuits.
+Any physical agent that selects actions based on sensory state can be modeled, in the static snapshot, as a circuit: sensory inputs are coordinates $X_{1},\ldots,X_{n}$; the agent’s internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop.
+2.48, Prop.
+2.41;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions.
+No physical agent can store it.[C:Thm.
+5.1;R:[E] vs [S+ETH]] Every physical agent is therefore a succinct encoding, a circuit whose size is bounded by physical resources (neurons, synapses, transistors, parameters) rather than by the size of the state space it navigates.[C:Thm.
+5.1;R:[S]]
+This is not a metaphor.
+A biological nervous system is a feedforward or recurrent circuit that takes sensory state as input and produces motor commands as output.[C:Prop.
+2.48;R:[represented adjacent class]] A trained neural network is a parameterized circuit.[C:Prop.
+2.48;R:[represented adjacent class]] In both cases, the agent is a succinct encoding of a utility-to-action mapping, and the question “is this agent attending to the right inputs?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm.
+7.2, Thm.
+4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm.
+5.1, Prop.
+5.3;R:[S+ETH],[Q_fin]]
+The simplicity tax as $\varepsilon$.
+The simplicity tax (Definition 9.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the agent’s model does not handle natively.
+In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (Lean: S2P.representationGap_missing_eq_gapCard, S2P.representationGap_eq_waste_plus_missing).
+For a physical agent modeled as a circuit, this is the coordinate-level manifestation of $\varepsilon$: the agent attends to a superset of what is necessary, or fails to attend to coordinates that are relevant, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6).
+Theorem 9.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
+Consequences for learned circuits.
+Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm.
+5.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure.
+Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET.
+The results of this paper imply that these questions are coNP-hard in the worst case under succinct encoding.[C:Thm.
+4.6, Thm.
+4.7;R:[S]] Empirical progress on interpretability and pruning therefore depends on exploiting the structural properties of specific trained circuits (analogous to the tractable subcases of Theorem 6.1), not on solving the general problem.[C:Thm.
+4.6, Thm.
+6.1;R:[S] vs [tractable regimes]]
+Scope and caveats.
+The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5).
+Biological neural systems add noise, temporal dynamics, and synaptic plasticity; trained networks add stochastic optimization and distribution shift.
+These factors may alter empirical tractability relative to the static worst-case bounds.
+The bridge conditions of Proposition 2.40 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.432.45).
+The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible agent architectures.[C:Prop.
+2.40, Thm.
+2.46, Prop.
+2.48;R:[represented adjacent class]]
+4.8 Quantifier Collapse for MINIMUM-SUFFICIENT-SET
+Theorem 4.14 (Collapse of the Apparent $\exists\forall$ Structure).
+The apparent second-level predicate
+$\exists I\,(|I|\leq k)\;\forall s,s^{\prime}\in S:\;s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
+is equivalent to the coNP predicate $|\text{Relevant}|\leq k$, where
+[…]
+Consequently, MINIMUM-SUFFICIENT-SET is governed by coNP certificates rather than a genuine $\Sigma_{2}^{P}$ alternation. (L: CC34, CC35, S2P2.)
+Proof.
+By the formal equivalence sufficient_iff_relevant_subset (finite-set form sufficient_iff_relevantFinset_subset), a coordinate set $I$ is sufficient iff $\text{Relevant}\subseteq I$. Therefore:
+$\exists I\;(|I|\leq k\wedge\text{sufficient}(I))\iff\exists I\;(|I|\leq k\wedge\text{Relevant}\subseteq I)\iff|\text{Relevant}|\leq k.$
+So the existential-over-universal presentation collapses to counting the relevant coordinates.
+A NO certificate for $|\text{Relevant}|\leq k$ is a list of $k+1$ distinct relevant coordinates, each witnessed by two states that differ only on that coordinate and yield different optimal sets; this is polynomially verifiable.
+Hence the resulting predicate is in coNP, matching Theorem 4.7.
+This also clarifies why ANCHOR-SUFFICIENCY remains $\Sigma_{2}^{P}$-complete: once an anchor assignment is existentially chosen, the universal quantifier over the residual subcube does not collapse to a coordinate-counting predicate.
+$\blacksquare$
 5 Encoding-Regime Separation
 The hardness results of Section 4 apply to worst-case instances under the succinct encoding.
-This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound.
+This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound, and an intermediate query-access obstruction family.
+The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another.
+Physical reading: the separation is an access-and-budget separation for the same decision semantics.
+Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
 Model note.
-Part 1 is an explicit-state upper bound (time polynomial in $|S|$).
-Part 2 is a succinct-encoding lower bound under ETH (time exponential in $n$).
-The encodings are defined in Section 2.4.
-These two parts are stated in different encodings and are not directly comparable as functions of a single input length.
+Theorem 5.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$).
+Proposition 5.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 5.7 and 5.11 are [Q_bool] interface refinements (value-entry and state-batch).
+Relative to Definition 2.23, Propositions 5.8 and 5.15 are explicit simulation-transfer instances.
+The [E] vs [S+ETH] split in Theorem 5.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes.
+These regimes are defined in Section 2.4 and are not directly comparable as functions of one numeric input-length parameter.
 Theorem 5.1 (Explicit–Succinct Regime Separation).
@@
   1. 1.
-    Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
+    [E] Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
   2. 2.
-    Succinct lower bound under ETH (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
+    [S+ETH] Succinct lower bound (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
+(L: CC16, CC23, CC24.)
 Proof.
@@
 This strengthening is mechanized in Lean (see Appendix A).
-Under the Exponential Time Hypothesis (ETH) [8], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.
+Under the Exponential Time Hypothesis (ETH) [28, 5], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.
 $\blacksquare$
@@
 For Boolean coordinate spaces ($N=2^{n}$), the explicit-state bound is polynomial in $2^{n}$ (exponential in $n$), while under ETH the succinct lower bound yields $2^{\Omega(n)}$ time for the hard family in which $k^{*}=\Omega(n)$.
-This encoding-regime separation explains why some domains admit tractable model selection under explicit-state assumptions, while other domains (or encodings) exhibit worst-case hardness that forces heuristic approaches.
+Proposition 5.3 (Finite-State Query Lower-Bound Core via Empty-Set Subproblem).
+In the finite-state query regime [Q_fin], let $S$ be any finite state type with $|S|\geq 2$ and let $Q\subset S$ with $|Q|<|S|$.
+Then there exist two decision problems $\mathcal{D}_{\mathrm{yes}},\mathcal{D}_{\mathrm{no}}$ over the same state space such that:
+  • their oracle views on all states in $Q$ are identical,
+  • $\emptyset$ is sufficient for $\mathcal{D}_{\mathrm{yes}}$,
+  • $\emptyset$ is not sufficient for $\mathcal{D}_{\mathrm{no}}$.
+Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$.
+(L: CC50, L7, L13.)
+Proof.
+This is exactly the finite-state indistinguishability theorem CC4.
+For any $|Q|<|S|$, there is an unqueried hidden state $s_{0}$ producing oracle-indistinguishable yes/no instances with opposite truth values on the $I=\emptyset$ subproblem.
+Since SUFFICIENCY-CHECK contains that subproblem, fewer than $|S|$ queries cannot be correct on both instances, yielding the $\Omega(|S|)$ worst-case lower bound.
+$\blacksquare$
+Proposition 5.4 (Checking–Witnessing Duality on the Empty-Set Core).
+[C:P5.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs.
+Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt.
+Then
+$|P|\geq 2^{n-1}.$
+Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: DQ587, DQ588.)
+Corollary 5.5 (Information Barrier in Query Regimes).
+[C:C5.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
+  • Q_fin
+    (Opt-oracle): with $|Q|<|S|$, there exist yes/no instances indistinguishable on all queried states but with opposite truth values on the $I=\emptyset$ subproblem.
+  • Q_bool
+    value-entry and state-batch interfaces preserve the same obstruction at cardinality $2^{n}$.
+Hence the barrier is representational-access based: without querying enough of the hidden state support, exact certification is blocked by indistinguishability rather than by search-strategy choice.
+(L: CC27, CC29, CC28.)
+Corollary 5.6 (Boolean-Coordinate Instantiation).
+[C:C5.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 5.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access.
+(L: CC49; core: L6.)
+Proposition 5.7 (Value-Entry Query Lower Bound).
+In the mechanized Boolean value-entry query submodel [Q_bool], for any deterministic procedure using fewer than $2^{n}$ value-entry queries $(a,s)\mapsto U(a,s)$, there exist two queried-value-indistinguishable instances with opposite truth values for SUFFICIENCY-CHECK on the $I=\emptyset$ subproblem.
+Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$.
+(L: L9, L19.)
+Proof.
+The theorem L2 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency.
+The auxiliary bound L6 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.
+$\blacksquare$
+Proposition 5.8 (Subproblem-to-Full Transfer Rule).
+If every full-problem solver induces a solver for a fixed subproblem, then any lower bound for that subproblem lifts to the full problem. (L: CC58, CC59, CC52.)
+This is the restriction-map instance of Definition 2.23: a solver for the full regime induces one for the restricted subproblem regime, so lower bounds transfer.
+Proposition 5.9 (Randomized Robustness (Seedwise)).
+[C:P5.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 5.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts.
+Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair.
+(L: L11, L12, L5.)
+Proposition 5.10 (Randomized Robustness (Weighted Form)).
+[C:P5.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight.
+Hence randomization cannot collapse both errors simultaneously.
+(L: L23, L24.)
+Proposition 5.11 (State-Batch Query Lower Bound).
+In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: L8, L14.)
+Proposition 5.12 (Finite-State Empty-Subproblem Generalization).
+[C:P5.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: L7, L13.)
+Proposition 5.13 (Adversary-Family Tightness by Full Scan).
+[C:P5.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan.
+(L: L10.)
+Proposition 5.14 (Weighted Query-Cost Transfer).
+[C:P5.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
+$\sum_{q\in Q}w(q)\ \geq\ w_{\min}\cdot L.$
+(L: L21, L22.)
+Proposition 5.15 (Oracle-Lattice Transfer: Batch $\Rightarrow$ Entry).
+In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: L20, CC43.)
+This is the oracle-transducer instance of Definition 2.23: batch transcripts induce entry transcripts while preserving indistinguishability for the target sufficiency question.
+Proposition 5.16 (Oracle-Lattice Strictness: Opt vs Value Entries).
+‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: L3, L4.)
+Remark 5.17 (Instantiation of Definitions 3.1 and 3.2).
+Theorem 5.1 and Propositions 5.35.7 keep the structural problem fixed (same sufficiency relation) and separate representational hardness by access regime: explicit-state access exposes the boundary $s\mapsto\text{Opt}(s)$, finite-state query access already yields $\Omega(|S|)$ lower bounds at the Opt-oracle level (Boolean instantiation: $\Omega(2^{n})$), value-entry/state-batch interfaces preserve the obstruction in [Q_bool], and succinct access can hide structure enough to force ETH-level worst-case cost on a hard family.
+This regime-typed landscape identifies tractability under [E], a finite-state query lower-bound core under [Q_fin] with Boolean interface refinements under [Q_bool], and worst-case intractability under [S+ETH] for the same underlying decision question.
+The structural reason for this separation is enumerable access.
+Under the explicit-state encoding, the mapping $s\mapsto\text{Opt}(s)$ is directly inspectable and sufficiency verification reduces to scanning state pairs, at cost polynomial in $|S|$.
+Under succinct encoding, the circuit compresses an exponential state space into polynomial size; universal verification over that compressed domain, namely “for all $s,s^{\prime}$ with $s_{I}=s^{\prime}_{I}$, does $\text{Opt}(s)=\text{Opt}(s^{\prime})$?”, carries worst-case cost exponential in $n$, because the domain the circuit compressed away cannot be reconstructed without undoing the compression.
+5.1 Budgeted Physical Crossover
+The [E] vs [S] split can be typed against an explicit physical budget.
+Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget.
+These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.34;C2.35;R:AR] (L: CT6,CT8,CT4.)
+Proposition 5.18 (Budgeted Crossover Witness).
+[C:P5.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$.
+(L: CC45, PBC1.)
+Proof.
+This is exactly the definitional split encoded in PBC1: explicit infeasibility is $B<E(n)$ and succinct feasibility is $S(n)\leq B$. The theorem CC2 returns both conjuncts.  $\blacksquare$
+Proposition 5.19 (Above-Cap Crossover Existence).
+[C:P5.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
+$\forall n,\ S(n)\leq C,\qquad\forall B,\ \exists n,\ B<E(n).$
+Then for every budget $B\geq C$, there exists a crossover witness at budget $B$. (L: CC44, PBC2.)
+Proof.
+Fix $B\geq C$. By unboundedness of $E$, choose $n$ with $B<E(n)$. By the cap assumption, $S(n)\leq C\leq B$, so $(B,n)$ is a crossover witness. The mechanized closure theorem is CC1.  $\blacksquare$
+Proposition 5.20 (Crossover Does Not Imply Exact-Certification Competence).
+[C:P5.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate.
+Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists.
+(L: CC46.)
+Proof.
+The mechanized theorem CC3 bundles the crossover split with the same logical closure used in Proposition 3.61: representational feasibility and certification hardness are distinct predicates.
+$\blacksquare$
+Proposition 5.21 (Crossover Policy Closure Under Integrity).
+[C:P5.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
+  • a budgeted crossover witness at $(B,n)$,
+  • the same collapse implication as Proposition 3.61,
+  • a solver that is integrity-preserving for the declared relation.
+Then either full in-scope non-abstaining coverage fails, or declared budget compliance fails.
+Equivalently, in uncertified hardness-blocked regions, integrity is compatible with $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ but not with unconditional exact-competence claims.
+(L: CC47.)
+Corollary 5.22 (Finite Budget Implies Eventual Exact-Certification Failure).
+[C:C5.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost.
+If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget.
+For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds.
+(L: DQ569, DQ568, DQ574.)
+Proposition 5.23 (Least Divergence Point at Fixed Budget).
+[C:P5.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$.
+If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover.
+(L: DQ576.)
+Proposition 5.24 (Eventual Explicit Infeasibility from Monotone Growth).
+[C:P5.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: DQ577.)
+Proposition 5.25 (Payoff Threshold as Eventual Split).
+[C:P5.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: DQ575, DQ579.)
+Corollary 5.26 (No Universal Survivor Without Succinct Bound).
+[C:C5.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible.
+(L: DQ578.)
+Proposition 5.27 (Policy Closure at and Beyond Divergence).
+[C:P5.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 5.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds.
+(L: DQ580, DQ581.)
+5.2 Thermodynamic Lift in Declared Model
+The crossover and hardness statements are complexity-theoretic.
+The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the declared conversion model and declared bit-operation lower-bound premises.
+Proposition 5.28 (Bit-to-Energy/Carbon Lift).
+[C:P5.28;R:H=cost-growth] Fix a declared thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule.
+If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
+$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}\leq\lambda b_{\mathrm{used}},\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}\leq\rho(\lambda b_{\mathrm{used}}).$
+(L: CC65, CC66.)
+Proposition 5.29 (Hardness–Thermodynamics Bundle).
+[C:P5.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model.
+(L: CC67.)
+Proposition 5.30 (Mandatory Physical Cost).
+[C:P5.30;R:H=cost-growth] Within the declared linear thermodynamic model, if
+$\lambda>0,\quad\rho>0,\quad b_{\mathrm{LB}}>0,$
+then both lower bounds are strictly positive:
+$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}>0,\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}>0.$
+(L: CC68.)
+Proposition 5.31 (Conserved Additive Accounting).
+[C:P5.31;R:H=cost-growth] Within the same declared linear model, lower-bound accounting is additive over composed bit-operation lower bounds:
+$E_{\mathrm{LB}}(b_{1}+b_{2})=E_{\mathrm{LB}}(b_{1})+E_{\mathrm{LB}}(b_{2}),\qquad C_{\mathrm{LB}}(b_{1}+b_{2})=C_{\mathrm{LB}}(b_{1})+C_{\mathrm{LB}}(b_{2}).$
+(L: CC64.)
 6 Tractable Special Cases: When You Can Solve It
-We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly.
+We distinguish the encodings of Section 2.4.
+The tractability results below state the model assumption explicitly.
+Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the three cases below instantiate this single principle.
+Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, or tree factorization) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier.
+In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T6.1,2.37;P2.34;R:E,TA,AR] (L: CC69,CC70,CC71,CC72,CT6,CT4.)
 Theorem 6.1 (Tractable Subcases).
@@
     with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly as part of the input.
+(L: CC69, CC70, CC72, CC71.)
 6.1 Explicit-State Encoding
@@
 Tree-structured utility Hierarchies, causal trees
-6 Implications for Practice: Why Over-Modeling Is Optimal
-This section states corollaries for engineering practice.
-Within the formal model, the complexity results of Sections 3 and 4 shift parts of this workflow from informal judgment toward explicit, checkable criteria.
-The observed behaviors—configuration over-specification, absence of automated minimization tools, heuristic model selection—are not failures of discipline but provably optimal responses to computational constraints under the stated cost model.
-Some common prescriptions implicitly require exact minimization of sufficient parameter sets.
-In the worst case, that task is coNP-complete in our model, so we should calibrate critiques and expectations accordingly.
-A rational response is to include everything and pay linear maintenance costs, rather than attempt exponential minimization costs.
+Proposition 6.2 (Heuristic Reusability).
+[C:P6.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 6.1).
+If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$.
+For each subcase of Theorem 6.1, structure detection is polynomial (under the declared representation assumptions).
+(L: CC53, CC7, CC55, CC74.)
+Remark 6.3 (Heuristics and Integrity).
+Proposition 6.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention.
+Whether the agent acts still requires competence (Definition 3.9): budget and coverage must also be satisfied.
+An agent that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification.
+Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
+7 Regime-Conditional Corollaries
+This section derives regime-typed engineering corollaries from the core complexity theorems.
+Theorem 7.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 5.1 then yield exact minimization consequences under [S] and [S+ETH].
+The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation.
+[C:T7.2,2.37;P2.34;R:AR] (L: CR1,CT6,CT4.)
+Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool].
+Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6,4.7,5.1;R:S,S+ETH] Theorem 7.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T7.4;R:S]
+Conditional rationality criterion.
+For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 6.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T6.1;P3.60;R:E,STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
+Remark 7.1 (Regime Contract for Engineering Corollaries).
+All claims in this section are formal corollaries under the declared model assumptions.
+  • Competence claims are indexed by the regime tuple of Definition 3.9; prescriptions are meaningful only relative to feasible resources under that regime (bounded-rationality feasibility discipline) [71, 5].[C:P3.58;R:AR]
+  • Integrity (Definition 3.8) forbids overclaiming beyond certifiable outputs; $\mathsf{ABSTAIN}$$\mathsf{UNKNOWN}$ is first-class when certification is unavailable.[C:P3.58;R:AR]
+  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6,5.1;P3.60;R:S,S+ETH]
 7.1 Configuration Simplification is SUFFICIENCY-CHECK
-Real engineering problems reduce directly to the decision problems studied in this paper.
-Theorem 6.1 (Configuration Simplification Reduces to SUFFICIENCY-CHECK).
+Because SUFFICIENCY-CHECK is a meta-problem parameterized by an arbitrary decision problem $\mathcal{D}$, real engineering problems with factored configuration spaces are instances of the hardness landscape established above.
+Theorem 7.2 (C1–C4: Configuration Simplification Reduces to SUFFICIENCY-CHECK).
 Given a software system with configuration parameters $P=\{p_{1},\ldots,p_{n}\}$ and observed behaviors $B=\{b_{1},\ldots,b_{m}\}$, the problem of determining whether parameter subset $I\subseteq P$ preserves all behaviors is equivalent to SUFFICIENCY-CHECK.
+(L: CR1)
 Proof.
 Construct decision problem $\mathcal{D}=(A,X_{1},\ldots,X_{n},U)$ where:
-  • Actions $A=B$ (each behavior is an action)
+  • Actions $A=B\cup\{\bot\}$, where $\bot$ is a sentinel “no-observed-behavior” action
   • Coordinates $X_{i}$ = domain of parameter $p_{i}$
   • State space $S=X_{1}\times\cdots\times X_{n}$
-  • Utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else $U(b,s)=0$
-Then $\text{Opt}(s)=\{b\in B:b\text{ occurs under configuration }s\}$.
+  • For $b\in B$, utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else 0
+  • Sentinel utility $U(\bot,s)=1$ iff no behavior in $B$ occurs under configuration $s$, else 0
+Then
+[…]
+So the optimizer map exactly encodes observed-behavior equivalence classes, including the empty-behavior case.
 Coordinate set $I$ is sufficient iff:
@@
 Therefore, “does parameter subset $I$ preserve all behaviors?” is exactly SUFFICIENCY-CHECK for the constructed decision problem.  $\blacksquare$
-Remark 6.2.
-This reduction is parsimonious: configuration simplification can be cast as an instance of SUFFICIENCY-CHECK under a direct encoding.
-6.2 Computational Rationality of Over-Modeling
-We now prove that over-specification is an optimal engineering strategy given the stated cost model and complexity constraints.
-Theorem 6.3 (Rational Over-Modeling).
+Remark 7.3 (Reduction Scope).
+The reduction above requires only:
+  1. 1.
+    a finite behavior set,
+  2. 2.
+    parameters with finite domains, and
+  3. 3.
+    an evaluable behavior map from configurations to achieved behaviors.
+These are exactly the model-contract premises C1–C3 instantiated for configuration systems.
+Theorem 7.4 (S: Over-Modeling as Boundary-Signal Corollary).
+By contraposition of Theorem 7.2, if no coordinate set can be certified as exactly relevance-identifying (Definition 2.16) for the modeled system, then the decision boundary is not completely characterized by the current parameterization.
+(L: CC40, CC6)
+Proof.
+Assume the decision boundary were completely characterized by the current parameterization.
+Then, via Theorem 7.2, the corresponding sufficiency instance admits exact relevance membership, hence a coordinate set that satisfies Definition 2.16.
+Contraposition gives the claim: persistent failure of exact relevance identification signals incomplete characterization of decision relevance in the model.
+$\blacksquare$
+7.2 Cost Asymmetry Under ETH
+We now prove a cost asymmetry result under the stated cost model and complexity constraints.
+[^1] Naive subset enumeration still gives an intuitive baseline of $O(2^{n})$ checks, but that is an algorithmic upper bound; the theorem below uses ETH for the lower-bound argument.
+Theorem 7.5 (S+ETH: Cost-Asymmetry Consequence).
 Consider an engineer specifying a system configuration with $n$ parameters. Let:
@@
   • $C_{\text{under}}$ = expected cost of production failures from underspecification
-When SUFFICIENCY-CHECK is coNP-complete (Theorem 3.6):
+Assume ETH in the succinct encoding model of Section 2.4. Then:
   1. 1.
-    Naive exhaustive search inspects $2^{n}$ candidate subsets, so the brute-force finding cost grows exponentially in $n$; more formally, exhaustive search yields an upper bound $C_{\text{find}}(n)=O(2^{n})$.
-A matching *unconditional* lower bound of $2^{\Omega(n)}$ does not follow from coNP-completeness alone.
-Instead, coNP-completeness implies there is no polynomial‑time algorithm for arbitrary inputs unless $\textsf{P}=\textsf{coNP}$.
-Stronger assumptions (e.g., the Exponential Time Hypothesis) yield explicit worst-case lower bounds: under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ worst-case time lower bound in the succinct encoding (witnessed by the strengthened TAUTOLOGY gadget family).
+    Exact identification of minimal sufficient sets has worst-case finding cost $C_{\text{find}}(n)=2^{\Omega(n)}$.
+(Under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ lower bound in the succinct model, and exact minimization subsumes this decision task.)
   2. 2.
@@
   3. 3.
-    Under the ETH (or when considering naive exhaustive search), exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
-Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, over-modeling minimizes total expected cost:
+    Under ETH, exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
+Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, the finding-vs-maintenance asymmetry satisfies:
 $C_{\text{over}}(k)<C_{\text{find}}(n)+C_{\text{under}}$
-Over-modeling is economically optimal under the stated model and complexity constraints.
+Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary.
+Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T7.5;P3.60;R:S+ETH,TA] (L: CC12, HD13.)
 Proof.
-By Theorem 3.6, SUFFICIENCY-CHECK is coNP-complete, so under the assumption $\textsf{P}\neq\textsf{coNP}$ there is no polynomial‑time algorithm that decides sufficiency for arbitrary inputs.
-Finding the minimal sufficient set by brute force requires checking many candidate sets. Exhaustive search examines
-$\sum_{i=0}^{n}\binom{n}{i}=2^{n}$
-candidate subsets, so naive search has cost $O(2^{n})$.
-This shows only that brute-force is exponential.
-To assert a matching lower bound of $2^{\Omega(n)}$ for *all* algorithms requires additional assumptions.
-For example, under the Exponential Time Hypothesis (ETH), TAUTOLOGY (and hence SUFFICIENCY-CHECK in the succinct encoding) requires $2^{\Omega(n)}$ time, and SUFFICIENCY-CHECK inherits this conditional lower bound via our reduction.
+Under ETH, the TAUTOLOGY reduction used in Theorem 4.6 yields a $2^{\Omega(n)}$ worst-case lower bound for SUFFICIENCY-CHECK in the succinct encoding.
+Any exact algorithm that outputs a minimum sufficient set can decide whether the optimum size is 0 by checking whether the returned set is empty; this is exactly the SUFFICIENCY-CHECK query for $I=\emptyset$.
+Hence exact minimal-set finding inherits the same exponential worst-case lower bound.
 Maintaining $k$ extra parameters incurs:
@@
 Total maintenance cost is $C_{\text{over}}(k)=O(k)$.
-For concrete threshold: when $n=20$ parameters, exhaustive search requires $2^{20}\approx 10^{6}$ checks.
-Including $k=5$ extra parameters costs $O(5)$ maintenance overhead but avoids $10^{6}$ computational work.
-Since $2^{n}$ grows faster than any polynomial in $k$ or $n$, there exists $n_{0}$ such that for all $n>n_{0}$:
+The eventual dominance step is mechanized in HD8: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$.
+Therefore:
 $C_{\text{over}}(k)\ll C_{\text{find}}(n)$
-Adding underspecification risk $C_{\text{under}}$ (production failures from missing parameters), which is unbounded in the model, makes over-specification strictly dominant.  $\blacksquare$
+For any fixed nonnegative $C_{\text{under}}$, the asymptotic dominance inequality remains and only shifts the finite threshold $n_{0}$.  $\blacksquare$
 Corollary 7.6 (Impossibility of Automated Configuration Minimization).
-Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
+[C:C7.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
   1. 1.
@@
     Guarantees correctness (no false negatives)
+(L: CC37)
 Proof.
@@
 Remark 7.7.
-Corollary 6.4 explains the observed absence of “config cleanup” tools in software engineering practice.
-Engineers who include extra parameters are not exhibiting poor discipline—they are adapting to computational constraints.
-The problem is not lack of tooling effort; it is mathematical intractability.
-6.3 Connection to Observed Practice
-These theorems provide mathematical grounding for three widespread engineering behaviors:
-1. Configuration files grow over time. Removing parameters requires solving coNP-complete problems. Engineers rationally choose linear maintenance cost over exponential minimization cost.
-2. Heuristic model selection dominates. ML practitioners use AIC, BIC, cross-validation instead of optimal feature selection because optimal selection is intractable (Theorem 6.3).
+Corollary 7.6 is a formal boundary statement: an always-correct polynomial-time minimizer for arbitrary succinct inputs would collapse P and coNP.
+Remark 7.8 (FPT Scope Caveat).
+The practical force of worst-case hardness depends on instance structure, especially $k^{*}$.
+If SUFFICIENCY-CHECK is FPT in parameter $k^{*}$, then small-$k^{*}$ families can remain tractable even under succinct encodings.
+The strengthened mechanized gadget (all_coords_relevant_of_not_tautology) still proves existence of hard families with $k^{*}=n$; what is typical in deployed systems is an empirical question outside this formal model.
+7.3 Regime-Conditional Operational Corollaries
+Theorems 7.4 and 7.5 yield the following conditional operational consequences:
+1.
+Conservative retention under unresolved relevance.
+If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T7.4;R:S]
+2.
+Heuristic selection as weakened-guarantee mode.
+Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 7.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 56, 66, 25, 50].[C:T7.5;R:S+ETH]
 3.
-“Include everything” is a legitimate strategy.
-When determining relevance/minimal sufficiency is exponential in the worst case (e.g., $O(2^{n})$ by naive subset enumeration, and under ETH a $2^{\Omega(n)}$ worst-case lower bound in the succinct encoding; see Theorem 4.1), including all $n$ parameters costs $O(n)$.
-For large $n$, this is the rational choice.
-These behaviors are not ad hoc workarounds; under the stated computational model they are rational responses to worst-case intractability.
-The complexity results provide a mathematical lens: over-modeling is not a failure—it is the rational strategy under the model.
-7 Implications for Software Architecture
-This section states corollaries for software architecture. The complexity results have direct implications for software engineering practice.
-7.1 Why Over-Specification Is Rational
-Software architects routinely specify more configuration parameters than strictly necessary. Our results show this is computationally rational:
-Corollary 7.1 (Rational Over-Specification).
-Given a software system with $n$ configuration parameters, checking whether a proposed subset suffices is coNP-complete. Finding the minimum such set is also coNP-complete.
-This explains why configuration files grow over time: removing “unnecessary” parameters requires solving a hard problem.
+Full-parameter inclusion as an $O(n)$ upper-bound strategy.
+Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T7.5;R:S+ETH]
+4.
+Irrationality outside attempted-competence-failure conditions.
+If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T6.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
+These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum.
+Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 6.1 or adopt explicit approximation commitments.[C:T6.1;P3.60;R:RG]
+8 Dominance Theorems for Hardness Placement
+Regime for this section: the mechanized Boolean-coordinate model [S_bool] plus the architecture cost model defined below.
+The centralization-vs-distribution framing follows established software-architecture criteria about modular decomposition and long-run maintenance tradeoffs [49, 10, 6, 60].
+Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model.
+[C:P2.34;T2.37,8.16,9.3,9.6;R:AR,LC] (L: CT6,CT4,HD1,HD4,HD3.)
+8.1 Over-Specification as Diagnostic Signal
+Corollary 8.1 (Persistent Over-Specification).
+[C:C8.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
+Proof.
+This is the contrapositive of S2P3.  $\blacksquare$
+Corollary 8.2 (Exact Relevance Identifiability Criterion).
+[C:C8.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
+$I\text{ is exactly relevance-identifying}\iff\bigl(I\text{ is sufficient and }I\subseteq R_{\mathrm{rel}}\bigr),$
+where $R_{\mathrm{rel}}$ is the full relevant-coordinate set. (L: S2P1)
+Proof.
+This is exactly S2P1, with $R_{\mathrm{rel}}=\texttt{relevantFinset}$.  $\blacksquare$
+Remark 8.3 (Over-Specification Is Regime-Conditional, Not Default).
+In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 7).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 8.78.9 together with Corollary 8.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T6.1;P3.60;R:TA] (L: IC18, IC25, IC7, L15, L16, L17, S2P1.)
 8.2 Architectural Decision Quotient
-The sufficiency framework suggests a measure for architectural decisions:
 Definition 8.4 (Architectural Decision Quotient).
@@
 $\text{ADQ}(I)=\frac{|\{b\in B:b\text{ achievable with some }s\text{ where }s_{I}\text{ fixed}\}|}{|B|}$
-High ADQ means the configuration subset $I$ leaves many behaviors achievable—it doesn’t constrain the system much. Low ADQ means $I$ strongly constrains behavior.
-Proposition 7.3.
-Decisions with low ADQ (strongly constraining) require fewer additional decisions to fully specify system behavior.
-7.3 Practical Recommendations
-Based on our theoretical results:
-  1. 1.
-    Accept over-modeling: Don’t penalize engineers for including “extra” parameters. The alternative (minimal modeling) is computationally hard.
-  2. 2.
-    Use bounded scenarios: When the scenario space is small (Proposition 2.10), minimal modeling becomes tractable.
-  3. 3.
-    Exploit structure: Tree-structured dependencies, bounded alternatives, and separable utilities admit efficient algorithms.
-  4. 4.
-    Invest in heuristics: For general problems, develop domain-specific heuristics rather than seeking optimal solutions.
+Proposition 8.5 (ADQ Ordering).
+For coordinate sets $I,J$ in the same system, if $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$, then fixing $I$ leaves a strictly smaller achievable-behavior set than fixing $J$. (L: CC2)
+Proof.
+The denominator $|B|$ is shared. Thus $\mathrm{ADQ}(I)<\mathrm{ADQ}(J)$ is equivalent to a strict inequality between the corresponding achievable-behavior set cardinalities.  $\blacksquare$
+8.3 Corollaries for Practice
+Corollary 8.6 (Cardinality Criterion for Exact Minimization).
+[C:C8.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
+Proof.
+By S2P2, sufficiency of size $\leq k$ is equivalent to a relevance-cardinality bound $\leq k$ in the Boolean-coordinate model.  $\blacksquare$
+Corollary 8.7 (Bounded-Regime Tractability).
+[C:C8.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 6.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: L15)
+Proof.
+This is the bounded-regime branch of Theorem 6.1, mechanized as L3.  $\blacksquare$
+Corollary 8.8 (Separable-Utility Tractability).
+[C:C8.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: L16)
+Proof.
+This is the separable-utility branch of Theorem 6.1, mechanized as L4.  $\blacksquare$
+Corollary 8.9 (Tree-Structured Tractability).
+[C:C8.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: L17)
+Proof.
+This is the tree-factor branch of Theorem 6.1, mechanized as L5.  $\blacksquare$
+Corollary 8.10 (Mechanized Hard Family).
+There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: L1)
+Proof.
+The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as L1.  $\blacksquare$
 8.4 Hardness Distribution: Right Place vs Wrong Place
-A general principle emerges from the complexity results: problem hardness is conserved and is distributed across a system in qualitatively different ways.
 Definition 8.11 (Hardness Distribution).
-Let $P$ be a problem with intrinsic hardness $H(P)$ (measured in computational steps, implementation effort, or error probability). A solution architecture $S$ partitions this hardness into:
+Let $P$ be a problem family under the succinct encoding of Section 2.4.
+In this section, baseline hardness $H(P;n)$ denotes worst-case computational step complexity on instances with $n$ coordinates (equivalently, as a function of succinct input length $L$) in the fixed encoding regime.
+A solution architecture $S$ partitions this baseline hardness into:
   • $H_{\text{central}}(S)$: hardness paid once, at design time or in a shared component
@@
 $H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)$
-Definition 7.5 (Hardness Conservation Principle).
-For any problem $P$ with intrinsic hardness $H(P)$, any solution $S$ satisfies:
-$H_{\text{central}}(S)+H_{\text{distributed}}(S)\geq H(P).$
-This statement is presented as a definitional principle grounded in the prior definition of intrinsic hardness: any correct solution must account for at least $H(P)$ units of work, which the architecture may allocate centrally or distribute across use sites.
-The substantive, theorem-like consequences (for example, Centralization Dominance) follow from this grounding.
+Proposition 8.12 (Baseline Lower-Bound Principle).
+[C:P8.12;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
+$H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)\geq H(P;n).$
+For SUFFICIENCY-CHECK, Theorem 5.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD20)
+Proof.
+By definition, $H(P;n)$ is a worst-case lower bound for correct solutions in this encoding regime and cost metric.
+Any such solution architecture decomposes total realized work as $H_{\text{central}}+n\cdot H_{\text{distributed}}$, so that total cannot fall below the baseline.
+$\blacksquare$
 Definition 8.13 (Hardness Efficiency).
@@
 $\eta(S,n)=\frac{H_{\text{central}}(S)}{H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)}$
-High $\eta$ indicates centralized hardness (paid once); low $\eta$ indicates distributed hardness (paid repeatedly).
+(L: HD10)
+Proposition 8.14 (Efficiency Equivalence).
+[C:P8.14;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD10)
+Proof.
+From Definition 8.13, $\eta(S,n)$ is exactly the fraction of total realized hardness paid centrally.  $\blacksquare$
+Definition 8.15 (Right vs Wrong Hardness Placement).
+For a solution architecture $S$ in this linear model:
+$\text{right hardness}\iff H_{\mathrm{distributed}}(S)=0,\qquad\text{wrong hardness}\iff H_{\mathrm{distributed}}(S)>0.$
+(L: HD11, HD12)
 Theorem 8.16 (Centralization Dominance).
-For $n>1$ use sites, solutions with higher $H_{\text{central}}$ and lower $H_{\text{distributed}}$ yield:
+Let $S_{\mathrm{right}},S_{\mathrm{wrong}}$ be architectures over the same problem family with
+$H_{\mathrm{distributed}}(S_{\mathrm{right}})=0,\quad H_{\mathrm{central}}(S_{\mathrm{right}})>0,\quad H_{\mathrm{distributed}}(S_{\mathrm{wrong}})>0,$
+and let $n>\max\!\bigl(1,H_{\mathrm{central}}(S_{\mathrm{right}})\bigr)$. Then:
   1. 1.
-    Lower total realized hardness: $H_{\text{total}}(S_{1})<H_{\text{total}}(S_{2})$ when $H_{\text{distributed}}(S_{1})<H_{\text{distributed}}(S_{2})$
+    Lower total realized hardness:
+    $H_{\text{total}}(S_{\mathrm{right}})<H_{\text{total}}(S_{\mathrm{wrong}})$
   2. 2.
@@
   3. 3.
-    Higher leverage: one unit of central effort affects $n$ sites
+    Quantified leverage: moving one unit of work from distributed to central saves exactly n1n-1 units of total realized hardness
+(L: HD1, HD2)
 Proof.
-(1) follows from the total hardness formula. (2) follows from error site counting. (3) follows from the definition of leverage as $L=\Delta\text{Effect}/\Delta\text{Effort}$.  $\blacksquare$
-Corollary 7.8 (Right Hardness vs Wrong Hardness).
-A solution exhibits hardness in the right place when:
-  • Hardness is centralized (high $H_{\text{central}}$, low $H_{\text{distributed}}$)
-  • Hardness is paid at design/compile time rather than runtime
-  • Hardness is enforced by tooling (type checker, compiler) rather than convention
-A solution exhibits hardness in the wrong place when:
-  • Hardness is distributed (low $H_{\text{central}}$, high $H_{\text{distributed}}$)
-  • Hardness is paid repeatedly at each use site
-  • Hardness relies on human discipline rather than mechanical enforcement
-Example: Type System Instantiation. Consider a capability $C$ (e.g., provenance tracking) that requires hardness $H(C)$:
+(1) and (2) are exactly the bundled theorem HD1. (3) is exactly HD2.  $\blacksquare$
+Corollary 8.17 (Right-Place vs Wrong-Place Hardness).
+[C:C8.17;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness.
+Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
+$H_{\mathrm{central}}(S_{\mathrm{right}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{right}})<H_{\mathrm{central}}(S_{\mathrm{wrong}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{wrong}}).$
+(L: HD16)
+Proof.
+This is the mechanized theorem HD10.  $\blacksquare$
+Proposition 8.18 (Dominance Modes).
+[C:P8.18;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
+  1. 1.
+    Strict threshold dominance: Corollary 8.17 gives strict inequality once $n>H_{\mathrm{central}}(S_{\mathrm{right}})$.
+  2. 2.
+    Global weak dominance: under the decomposition identity used in HD3, centralized hardness placement is never worse for all $n\geq 1$.
+  3. 3.
+    Generalized nonlinear dominance: under bounded-vs-growing site-cost assumptions (Theorem 8.22), right placement strictly dominates beyond a finite threshold without assuming linear per-site cost.
+  4. 4.
+    Generalized boundary mode: without those growth-separation assumptions, strict dominance is not guaranteed (Proposition 8.24).
+Proof.
+Part (1) is Corollary 8.17. Part (2) is exactly HD3. Part (3) is Theorem 8.22. Part (4) is Proposition 8.24.  $\blacksquare$
+Illustrative Instantiation (Type Systems). Consider a capability $C$ (e.g., provenance tracking) with one-time central cost $H_{\text{central}}$ and per-site manual cost $H_{\text{distributed}}$:
 Approach $H_{\text{central}}$ $H_{\text{distributed}}$
@@
 Manual implementation Low (no new concepts) High (reimplement per site)
-For $n$ use sites, manual implementation costs $n\cdot H_{\text{distributed}}$, growing without bound.
-Native support costs $H_{\text{central}}$ once, amortized across all uses.
-The “simpler” approach (manual) is only simpler at $n=1$; for $n>H_{\text{central}}/H_{\text{distributed}}$, native support dominates.
-Remark 7.9 (Connection to Decision Quotient).
-The decision quotient (Section 2) measures which coordinates are decision-relevant.
-Hardness distribution measures where the cost of handling those coordinates is paid.
-A high-axis system makes relevance explicit (central hardness); a low-axis system requires users to track relevance themselves (distributed hardness).
-Mechanized strengthening (sketch).
-We formalized a strengthened reduction in the mechanization: given a Boolean formula $\varphi$ over $n$ variables we construct a decision problem with $n$ coordinates so that if $\varphi$ is not a tautology then every coordinate is decision-relevant.
-Intuitively, the construction places a copy of the base gadget at each coordinate and makes the global “accept” condition hold only when every coordinate’s local test succeeds; a single falsifying assignment at one coordinate therefore changes the global optimal set, witnessing that coordinate’s relevance.
-The mechanized statement and proof appear in the development as Reduction_AllCoords.lean (the lemma all_coords_relevant_of_not_tautology).
+The table is schematic; the formal statement is Corollary 8.19.
+The type-theoretic intuition behind this instantiation is consistent with classic accounts of abstraction and static interface guarantees [12, 53, 51, 36, 62, 22].
+Corollary 8.19 (Type-System Threshold).
+[C:C8.19;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
+$n>H_{\mathrm{baseline}}(P),$
+where $H_{\mathrm{baseline}}(P)$ corresponds to the Lean identifier intrinsicDOF(P) in module HardnessDistribution. (L: HD14)
+Proof.
+Immediate from HD9.  $\blacksquare$
+8.5 Extension: Non-Additive Site-Cost Models
+Definition 8.20 (Generalized Site Accumulation).
+Let $C_{S}:\mathbb{N}\to\mathbb{N}$ be a per-site accumulation function for architecture $S$. Define generalized total realized hardness by
+$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
+Definition 8.21 (Eventual Saturation).
+A cost function $f:\mathbb{N}\to\mathbb{N}$ is eventually saturating if there exists $N$ such that for all $n\geq N$, $f(n)=f(N)$.
+Theorem 8.22 (Generalized Dominance by Growth Separation).
+[C:T8.22;R:H=cost-growth] Let
+$H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
+For two architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$, suppose there exists $B\in\mathbb{N}$ such that:
+  1. 1.
+    $C_{S_{\mathrm{right}}}(m)\leq B$ for all $m$ (bounded right-side per-site accumulation),
+  2. 2.
+    $m\leq C_{S_{\mathrm{wrong}}}(m)$ for all $m$ (identity-lower-bounded wrong-side growth).
+Then for every
+$n>H_{\text{central}}(S_{\mathrm{right}})+B,$
+one has
+$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
+(L: HD8)
+Proof.
+This is exactly the mechanized theorem HD6.  $\blacksquare$
+Corollary 8.23 (Eventual Generalized Dominance).
+[C:C8.23;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
+$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
+(L: HD9)
+Proof.
+Immediate from HD7.  $\blacksquare$
+Proposition 8.24 (Generalized Assumption Boundary).
+[C:P8.24;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
+  1. 1.
+    If wrong-side growth lower bounds are dropped, right-side strict dominance can fail for all $n$.
+  2. 2.
+    If right-side boundedness is dropped, strict dominance can fail for all $n$ even when wrong-side growth is linear.
+(L: HD7, HD6)
+Proof.
+This is exactly the pair of mechanized boundary theorems listed above.  $\blacksquare$
+Theorem 8.25 (Linear Model: Saturation iff Zero Distributed Hardness).
+[C:T8.25;R:H=cost-growth] In the linear model of this section,
+$H_{\text{total}}(S,n)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S),$
+the function $n\mapsto H_{\text{total}}(S,n)$ is eventually saturating if and only if $H_{\text{distributed}}(S)=0$. (L: HD19)
+Proof.
+This is exactly the mechanized equivalence theorem above.  $\blacksquare$
+Theorem 8.26 (Generalized Model: Saturation is Possible).
+[C:T8.26;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
+$C_{K}(n)=\begin{cases}n,&n\leq K\\ K,&n>K,\end{cases}$
+both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD17, HD5)
+Proof.
+This is the explicit construction mechanized in Lean.  $\blacksquare$
+Corollary 8.27 (Positive Linear Slope Cannot Represent Saturation).
+[C:C8.27;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD15)
+Proof.
+This follows from the mechanized theorem that any linear representation of the saturating family must have zero slope.  $\blacksquare$
+Mechanized strengthening reference.
+The strengthened all-coordinates-relevant reduction is presented in Section 4 (“Mechanized strengthening”) and formalized in Reduction_AllCoords.lean via all_coords_relevant_of_not_tautology.
 The next section develops the major practical consequence of this framework: the Simplicity Tax Theorem.
-8 Corollary: Complexity Conservation
-A quantitative consequence of the hardness results: when a model handles fewer dimensions than required, the gap must be paid at each use site.
+9 Conservation Law for Decision-Relevant Coordinates
+The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place.
+By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding.
+The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6,4.7;R:S] For encoded physical systems, the same redistribution laws apply under declared physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.34;T9.3,9.5,9.6;R:AR,LC] (L: CT6,HD4,HD21,HD3.)
 Definition 9.1.
 Let $R(P)$ be the required dimensions (those affecting Opt) and $A(M)$ the dimensions model $M$ handles natively. The expressive gap is $\text{Gap}(M,P)=R(P)\setminus A(M)$.
-Theorem 8.2 (Conservation).
-$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced—only redistributed between “handled natively” and “handled externally.”
+Definition 9.2 (Simplicity Tax).
+The simplicity tax is the size of the expressive gap:
+$\text{SimplicityTax}(M,P):=|\text{Gap}(M,P)|.$
+Theorem 9.3 (Redistribution Identity).
+$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD4)
+Proof.
+In the finite-coordinate model this is the exact set-cardinality identity
+$|R\setminus A|+|R\cap A|=|R|,$
+formalized as HD5.  $\blacksquare$
+Remark 9.4 (Why this is nontrivial in context).
+The algebraic identity in Theorem 9.3 is elementary.
+Its force comes from upstream hardness: reducing $|R(P)|$ by exact relevance minimization is worst-case intractable under the succinct encoding, so redistribution is often the only tractable lever available.
 Theorem 9.5 (Linear Growth).
-For $n$ decision sites: $\text{TotalExternalWork}=n\times|\text{Gap}(M,P)|$.
+For $n$ decision sites:
+$\text{TotalExternalWork}=n\times\text{SimplicityTax}(M,P).$
+(L: HD21)
+Proof.
+This is by definition of per-site externalization and is mechanized as HD12.  $\blacksquare$
 Theorem 9.6 (Amortization).
-Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists $n^{*}=H_{\text{central}}/|\text{Gap}|$ such that for $n>n^{*}$, the complete model has lower total cost.
-Corollary 8.5.
-Since identifying $R(P)$ is coNP-complete (Theorem 3.6), minimizing the expressive gap is also intractable.
-These results are machine-checked in Lean 4 (HardnessDistribution.lean).
+Let $H_{\text{central}}$ be the one-time cost of using a complete model. There exists
+$n^{*}=\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}(M,P)}\right\rfloor$
+such that for $n>n^{*}$, the complete model has lower total cost. (L: HD3)
+Proof.
+For positive per-site tax, the threshold inequality
+$n>\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}}\right\rfloor\implies H_{\text{central}}<n\cdot\text{SimplicityTax}$
+is mechanized as HD4.  $\blacksquare$
+Corollary 9.7 (Gap Externalization).
+[C:C9.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD21, HD18)
+Proof.
+The exact linear form is HD12. When the gap is nonempty (positive tax), monotone growth with $n$ is HD11.  $\blacksquare$
+Corollary 9.8 (Exact Minimization Criterion).
+[C:C9.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
+Proof.
+This is S2P2.  $\blacksquare$
+Appendix A provides theorem statements and module paths for the corresponding Lean formalization.
 10 Related Work
+10.1 Formalized Complexity and Mechanized Proof Infrastructure
+Formalizing complexity-theoretic arguments in proof assistants remains comparatively sparse.
+Lean and Mathlib provide the current host environment for this artifact [16, 17, 67].
+Closest mechanized precedents include verified computability/complexity developments in Coq and Isabelle [21, 32, 45, 44, 26], and certifying toolchain work that treats proofs as portable machine-checkable evidence [34, 43].
+Recent LLM-evaluation work also highlights inference nondeterminism from numerical precision and hardware configuration, reinforcing the value of proof-level claims whose validity does not depend on stochastic inference trajectories [72].
 10.2 Computational Decision Theory
 The complexity of decision-making has been studied extensively.
-Papadimitriou [10] established foundational results on the complexity of game-theoretic solution concepts.
+Foundational treatments of computational complexity and strategic decision settings establish the baseline used here [47, 5].
 Our work extends this to the meta-question of identifying relevant information.
-For a modern treatment of complexity classes, see Arora and Barak [3].
-Closest prior work and novelty.
-Closest to our contribution is the literature on feature selection and model selection hardness, which proves NP-hardness of selecting informative feature subsets and inapproximability for minimum feature sets [4, 2].
-Those results analyze predictive relevance or compression objectives.
-We study decision relevance and show coNP-completeness for sufficiency checking, a different quantifier structure with distinct proof techniques and a full hardness/tractability landscape under explicit encoding assumptions, mechanized in Lean 4.
-The formalization aspect is also novel: prior work establishes hardness on paper, while we provide machine-checked reductions with explicit polynomial bounds.
+Decision-theoretic and information-selection framing in this paper also sits in the tradition of statistical sufficiency and signaling/information economics [20, 65, 42, 40, 29].
+Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [13, 61, 30].
+Our object differs: we classify coordinate sufficiency for optimal-action invariance, with mechanized reductions and regime typing.
+Relation to prior hardness results.
+Closest adjacent results are feature-selection/model-selection hardness results for predictive subset selection [9, 3].
+Relative to those works, this paper changes two formal objects: (i) the reductions are machine-checked (TAUTOLOGY and $\exists\forall$-SAT mappings with explicit polynomial bounds), and (ii) the output is a regime-typed hardness/tractability map for decision relevance under explicit encoding assumptions.
+The target predicate here is decision relevance, not predictive compression.
+10.3 Succinct Representations and Regime Separation
+Representation-sensitive complexity is established in planning and decision-process theory: classical and compactly represented MDP/planning problems exhibit sharp complexity shifts under different input models [48, 41, 37].
+The explicit-vs-succinct separation in this paper is the same representation-sensitive phenomenon for the coordinate-sufficiency predicate.
+The distinction in this paper is the object and scope of the classification: we classify decision relevance (sufficiency, anchor sufficiency, and minimum sufficient sets) for a fixed decision relation, with theorem-level regime typing and mechanized reduction chains.
+10.4 Oracle and Query-Access Lower Bounds
+Query-access lower bounds are a standard source of computational hardness transfer [18, 11].
+Our [Q_fin] and [Q_bool] results are in this tradition, but specialized to the same sufficiency predicate used throughout the paper: they establish access-obstruction lower bounds while keeping the underlying decision relation fixed across regimes.
+[…]
+10.5 Energy Landscapes and Chemical Reaction Networks
+The geometric interpretation of decision complexity connects to the physics of energy landscapes.
+Wales established a widely used picture in which molecular configurations form high-dimensional landscapes with basins and transition saddles [68, 69, 70].
+Protein-folding theory uses the same landscape language via funnel structure toward native states [46].
+In this paper’s formal model, the corresponding core object is coordinate-projection invariance of the optimizer map: sufficiency asks whether fixing coordinates preserves Opt, while insufficiency is witnessed by agreeing projections with different optimal sets (Definition 2.4, Proposition 2.6).[C:D2.4;P2.6;R:DM](L: DP7,DP8.)
+Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.34;C2.35;R:AR](L: CT6,CT5,CT8,CT7.)
+Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 5.1); finite query access already exhibits indistinguishability barriers (Proposition 5.3).
+The complexity separation is mechanized.[C:T5.1;P5.3;R:E,S+ETH,Q_fin](L: CC16,CC23,CC24,CC50.)
+The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DQ3, L25, L68, DP21.)
+The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: DQ519, DQ518, DQ517, DQ155, DQ154, DQ159, DQ160.)
+Chemical reaction networks provide a concrete physical class where this encoding applies directly.
+Prior work establishes hardness for output optimization, reconfiguration, and reconstruction in CRN settings [4, 2, 19].
+Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact.
+The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 7.2, 4.6, 6.1).[C:T7.2,4.6,6.1;R:E,S,TA](L: CR1,CC61,CC69,CC70,CC71,CC72.)
+The physical bridge is not a single-domain add-on: Theorem 2.37 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.37;R:AR](L: CT4.)
+For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.36;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
+The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.28,2.29,2.30,2.31;R:AR](L: DQ548,DQ537,DQ539,DQ534,DQ541,DQ549.)
+Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements.
+The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws rather than informal metaphors.[C:P2.34;T2.37;R:AR](L: PC10, PC12, PC15, PC17, PC19, PC20, PC21.)
+10.6 Thermodynamic Costs and Landauer’s Principle
+The thermodynamic lift in Section 5 converts declared bit-operation lower bounds into declared energy/carbon lower bounds via linear conversion constants.
+In this artifact, those conversion consequences are mechanized conditionally through Propositions 5.28, 5.29, 5.30, and 5.31.
+Landauer’s principle and reversible-computation context [33, 7] define the physical assumption set; the constants are declared premises, not Lean-derived constants.[C:P5.28,5.29,5.30,5.31;R:AR](L: CC65,CC66,CC67,CC68,CC64.)
+Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the declared regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60).
+A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61,3.60;R:AR](L: IC26,IC24,CC30,IC27,IC7,IC6.)
 10.7 Feature Selection
 In machine learning, feature selection asks which input features are relevant for prediction.
-This is known to be NP-hard in general [4].
+This is known to be NP-hard in general [9, 25].
 Our results show the decision-theoretic analog is coNP-complete for both checking and minimization.
+[C:T4.6,4.7;R:S](L: CC61,CC36.)
 10.8 Value of Information
-The value of information (VOI) framework [7] quantifies the maximum rational payment for information.
+The value of information (VOI) framework [27, 52] quantifies the maximum rational payment for information.
 Our work addresses a different question: not the value of information, but the complexity of identifying which information has value.
 10.9 Model Selection
-Statistical model selection (AIC [1], BIC [13], cross-validation [15]) provides practical heuristics for choosing among models.
-Our results provide theoretical justification: optimal model selection is intractable, so heuristics are necessary.
+Statistical model selection (AIC/BIC/cross-validation) provides practical heuristics for choosing among models [1, 56, 66].
+Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure.
+[C:T4.6,5.1,6.1;R:E,S,S+ETH](L: CC61,CC16,CC69,CC70,CC71.)
+10.10 Information Compression vs Certification Length
+The paper now makes an explicit two-part information accounting object (Definition 3.45): raw report bits and evidence-backed certification bits.
+This sharpens the information-theoretic reading of typed claims: compression of report payload is not equivalent to compression of certifiable truth conditions [57, 14, 64].
+From a zero-error viewpoint, this distinction is structural rather than stylistic: compressed representations can remain short while exact decodability depends on confusability structure and admissible decoding conditions [58, 31, 38, 15].
+The same separation appears here between short reports and evidence-backed exact claims.
+The rate-distortion and MDL lines make the same split in different language: there is a difference between achievable compression rate, computational procedure for obtaining/optimizing that rate, and certifiable model adequacy [59, 8, 54, 24, 35].
+Formally, Theorem 3.47 and Corollary 3.49 characterize the exact-report boundary: admissible exact reports require a strict certified-bit gap above raw payload, while hardness-blocked exact reports collapse to raw-only accounting.
+This directly links representational succinctness to certification cost under the declared contract, rather than treating “short reports” as automatically high-information claims.
+[C:T3.47;P3.46;C3.49;R:AR](L: CC18,CC11,CC21.)
+10.11 Certifying Outputs and Proof-Carrying Claims
+Our integrity layer matches the certifying-algorithms pattern: algorithms emit candidate outputs together with certificates that can be checked quickly, separating producing claims from verifying claims [39, 43].
+In this paper, Definition 3.8 is exactly that soundness discipline.
+At the systems level, this is the same architecture as proof-carrying code: a producer ships evidence and a consumer runs a small checker before accepting the claim [43, 39].
+Our competence definition adds the regime-specific coverage/resource requirement that certifying soundness alone does not provide.
+The feasibility qualifier in Definition 3.9 yields the bounded-rationality constraint used in this paper: admissible policy is constrained by computational feasibility under the declared resource model (Remark 2.25) [71, 5].[C:P3.58;R:AR] (L: IC18.)
+Cryptographic-verifiability perspective (scope).
+The role of cryptography in this paper is structural verifiability, not secrecy: the relevant analogy is certificate-carrying claims with lightweight verification, not confidential encoding [23, 43].
+Concretely, the typed-report discipline is modeled as a game-based reporting interface: a prover emits $(r,\pi)$ and a verifier accepts exactly when $\pi$ certifies the declared report type under the active contract.
+In that model, Evidence–Admissibility Equivalence gives completeness for admissible report types, while Typed Claim Admissibility plus Exact Claims Require Exact Evidence give soundness against exact overclaiming.
+Certified-confidence constraints then act as evidence-gated admissibility constraints on top of that verifier interface, so raw payload bits and certified bits represent distinct resources rather than stylistic notation variants.
+[C:P3.18,3.16,3.21,3.23;R:AR](L: IC2,CC75,IC35,IC38.)
+Three-axis integration.
+In the cited literature, these pillars are treated separately: representation-sensitive hardness, query-access lower bounds, and certifying soundness disciplines.
+This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework.
+[C:T5.1;P5.3,3.16;R:S+ETH,Qf,AR](L: CC16,CC50,CC75.)
 11 Conclusion
@@
 This paper was developed through human-AI collaboration.
-The author provided the core intuitions—the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET.
+The author provided the core intuitions: the connection between decision-relevance and computational complexity, the conjecture that SUFFICIENCY-CHECK is coNP-complete, and the insight that the $\Sigma_{2}^{\textsf{P}}$ structure collapses for MINIMUM-SUFFICIENT-SET.
 Large language models (Claude, GPT-4) served as implementation partners for proof drafting, Lean formalization, and generation.
@@
 What LLMs contributed: drafting, Lean tactic exploration, reduction construction assistance, and prose refinement.
-The proofs are machine-checked; their validity is independent of generation method. We disclose this methodology in the interest of academic transparency.
+The proofs are machine-checked; their validity is independent of generation method. This methodology is disclosed for academic transparency.
 Summary of Results
@@
   • An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ (Theorem 5.1)
+  • Full intermediate query-access lower bounds are formalized as a finite-state Opt-oracle core ($\Omega(|S|)$, Boolean instantiation $\Omega(2^{n})$) plus Boolean-interface refinements for value-entry and state-batch access, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness closures (Propositions 5.35.16)
   • Tractable subcases exist for explicit-state encoding, separable utility, and tree-structured utility with explicit local factors (Theorem 6.1)
+  • Discrete-time interface semantics are formalized: time is $\mathbb{N}$-valued, decision events are exactly one-tick transitions, run length equals elapsed time, and substrate-consistent realizations preserve the same one-unit event law (Propositions 2.282.31)
+  • Physical-system transport is formalized: encoded physical claims are derived from core rules through explicit assumption transfer, physical counterexamples map to encoded core failures, and the bundled bridge theorem composes law semantics, reduction equivalence, and required-work lower bounds (Proposition 2.34, Corollary 2.35, Theorem 2.37)
+[C:T4.6,4.7,4.9,5.1,6.1,2.37;P5.3,5.7,5.11,5.8,5.15,5.16,2.28,2.29,2.30,2.31,2.34;C2.35;R:E,S,S+ETH,Qf,Qb,AR] (L: CC61,CC36,CC5,CC16,CC69,CC50,L9,L8,L11,L21,L20,DQ548,DQ537,DQ539,DQ541,CT6,CT8,CT4.)
 These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
+Beyond classification, the paper contributes a formal claim-typing framework (Section 3): structural complexity is a property of the fixed decision question, while representational hardness is regime-conditional access cost.
+This is why encoding-regime changes can move practical hardness without changing the underlying semantics.
+[C:P2.20;T3.4;R:AR] (L: CC62,CC76.)
 The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings).
@@
 The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized.
+[C:T4.6,4.7,4.9,5.1;R:S,S+ETH] (L: CC61,CC36,CC5,CC16,CC24.)
 Complexity Characterization
@@
   1. 1.
-    Exact bounds. SUFFICIENCY-CHECK is coNP-complete—both coNP-hard and in coNP.
+    Exact bounds. SUFFICIENCY-CHECK is coNP-complete, both coNP-hard and in coNP.
   2. 2.
@@
     Encoding-regime separation.
-Under the explicit-state encoding, SUFFICIENCY-CHECK is polynomial in $|S|$.
-Under ETH, there exist succinctly encoded worst-case instances (witnessed by a strengthened gadget family with $k^{*}=n$) requiring $2^{\Omega(n)}$ time.
-Intermediate regimes are not ruled out by the lower-bound statement.
-The Complexity Conservation Law
+Under [E], SUFFICIENCY-CHECK is polynomial in $|S|$.
+Under [S+ETH], there exist succinct worst-case instances (with $k^{*}=n$) requiring $2^{\Omega(n)}$ time.
+Under [Q_fin], the Opt-oracle core has $\Omega(|S|)$ worst-case query complexity (Boolean instantiation $\Omega(2^{n})$), and under [Q_bool] the value-entry/state-batch refinements preserve the obstruction with weighted-cost transfer closures.
+[C:T4.6,5.1;P5.3,5.7,5.11,5.8;R:E,S,S+ETH,Qf,Qb] (L: CC61,CC16,CC50,L9,L8,L11.)
+The Complexity Redistribution Corollary
 Section 9 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site.
-For $n$ sites, total external work is $(k-j)\times n$.
-This conservation law is formalized in Lean 4 (HardnessDistribution.lean), proving:
-  • Conservation: complexity cannot be eliminated, only redistributed
+For $n$ sites, total external work is $(k-j)\times n$.[C:T9.5;R:LC] (L: HD21.)
+The set identity is elementary; its operational content comes from composition with the hardness results on exact relevance minimization.
+This redistribution corollary is formalized in Lean 4 (HardnessDistribution.lean), proving:
+  • Redistribution identity: complexity burden cannot be eliminated by omission, only moved between native handling and external handling
   • Dominance: complete models have lower total work than incomplete models
   • Amortization: there exists a threshold $n^{*}$ beyond which higher-dimensional models have lower total cost
+[C:T9.3,8.16,9.6;R:LC] (L: HD4,HD1,HD3,HD21.)
 Open Questions
+The landscape above is complete for the static sufficiency class under C1–C4 and the declared regimes; the items below are adjacent-class extensions or secondary refinements.
 Several questions remain for future work:
-  • Fixed-parameter tractability: Is SUFFICIENCY-CHECK FPT when parameterized by the size of the minimal sufficient set?
+  • Fixed-parameter tractability (primary): Is SUFFICIENCY-CHECK FPT when parameterized by the minimal sufficient-set size $k^{*}$, or is it W[2]-hard under this parameterization?
+[…]
   • Average-case complexity: What is the complexity under natural distributions on decision problems?
-  • Quantum complexity: Does quantum computation provide speedups for sufficiency checking?
+  • Typed-confidence calibration: For signaled reports $(r,g,p_{\mathrm{self}},p_{\mathrm{cert}})$, characterize proper scoring and calibration objectives that preserve signal consistency and typed admissibility.
   • Learning cost formalization: Can central cost $H_{\text{central}}$ be formalized as the rank of a concept matroid, making the amortization threshold precisely computable?
-Practical interpretation
-While this work is theoretical, the complexity bounds have practical meaning.
-Under ETH, worst-case instances of SUFFICIENCY-CHECK and related problems grow exponentially with problem size; practitioners should therefore not expect polynomial‑time worst‑case solvers for arbitrary inputs.
-In practice, real datasets often have structure (sparsity, bounded dependency depth, separable utilities) that the tractable subcases capture.
-Thus, engineering efforts should focus on (a) detecting and exploiting structural restrictions in inputs, and (b) designing heuristic or approximate methods for large unstructured instances.
-Our experiments illustrate performance on representative inputs, but the theoretical bounds underline the need for domain constraints to achieve scalable, worst‑case guarantees.
+Practical Corollaries
+The practical corollaries are regime-indexed and theorem-indexed:
+  • [E] and structured regimes: polynomial-time exact procedures exist (Theorem 6.1).
+  • [Q_fin]/[Q_bool] query-access lower bounds: worst-case Opt-oracle complexity is $\Omega(|S|)$ in the finite-state core (Boolean instantiation $\Omega(2^{n})$), and value-entry/state-batch interfaces satisfy the same obstruction in the mechanized Boolean refinement (Propositions 5.35.16), with randomized weighted robustness and oracle-lattice closures.
+  • [S+ETH] hard families: exact minimization inherits exponential worst-case cost (Theorem 5.1 together with Theorem 4.6).
+  • [S_bool] mechanized criterion: minimization reduces to relevance-cardinality constraints (Corollary 8.6).
+  • Redistribution consequences: omitted native coverage externalizes work with explicit growth/amortization laws (Theorems 9.39.6).
+  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.34;C2.35;T2.37;R:AR](L: CT6,CT8,CT4.)
+Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T6.1,5.1;R:RG](L: CC71,CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
+Epistemic Alignment Implications for Agent Training
+Within this paper’s formalism, epistemic alignment is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence).
+Formally, this is the integrity/competence split of Section 3, not a full value-alignment claim.[C:D3.8,3.9;P3.58;R:TR](L: IC18.)
+Truth-Objective Abstention Requirement.
+If a solver-agent declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers.
+Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19,3.20;P3.23,3.29;C3.25;R:AR](L: IC46,IC38,IC41.) Equivalently, training policies that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58,3.61,3.60;R:AR](L: IC18,CC30,IC27.)
+This yields a training doctrine for solver-agents under declared tasks/regimes:
+  1. 1.
+    Treat $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ as first-class admissible outputs, not errors.[C:D3.8;R:TR]
+  2. 2.
+    For signaled reports, maintain two channels: self-reflected confidence may be emitted, but certified confidence is positive only with evidence.[C:D3.20;P3.23,3.30;R:AR](L: IC38,IC47.)
+  3. 3.
+    Treat runtime disclosure as scalar-witnessed: always emit exact step-count, and emit completion percentage only when a declared bound exists.[C:P3.26,3.27,3.28;R:AR](L: IC42,IC35,IC44.)
+  4. 4.
+    Optimize competent coverage subject to integrity constraints, rather than optimizing answer-rate alone.[C:D3.9;P3.58;R:AR](L: IC18.)
+  5. 5.
+    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T5.1;P3.61,3.60;R:S,S+ETH](L: CC16,CC30,IC27.)
+Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the declared scope, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are declared).
+This is consistent with bounded-rationality feasibility discipline under declared resource models [71, 5].
 Appendix A Lean 4 Proof Listings
-The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page). The mechanization consists of approximately 5,000 lines across 33 files.
-A.1 What Is Machine-Checked
+The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page).
+The mechanization consists of 13391 lines across 58 files, with 597 theorem/lemma statements.
+Handle IDs. Inline theorem metadata now cites compact IDs (for example, HD6, CC12, IC4) instead of full theorem constants. The full ID-to-handle mapping is listed in Section A.1.
+A.1 Lean Handle ID Map
+ID Full Lean handle
+CC1 DecisionQuotient.ClaimClosure.RegimeSimulation
+CC2 DecisionQuotient.ClaimClosure.adq_ordering
+CC3 DecisionQuotient.ClaimClosure.agent_transfer_licensed_iff_snapshot
+CC4 DecisionQuotient.ClaimClosure.anchor_sigma2p_complete_conditional
+CC5 DecisionQuotient.ClaimClosure.anchor_sigma2p_reduction_core
+CC6 DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
+CC7 DecisionQuotient.ClaimClosure.bounded_actions_detectable
+CC8 DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
+CC9 DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
+CC10 DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
+CC11 DecisionQuotient.ClaimClosure.certified_total_bits_split_core
+CC12 DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
+CC13 DecisionQuotient.ClaimClosure.declaredBudgetSlice
+CC14 DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
+CC15 DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
+CC16 DecisionQuotient.ClaimClosure.dichotomy_conditional
+CC17 DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
+CC18 DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
+CC19 DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
+CC20 DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
+CC21 DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
+CC22 DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+CC23 DecisionQuotient.ClaimClosure.explicit_state_upper_core
+CC24 DecisionQuotient.ClaimClosure.hard_family_all_coords_core
+CC25 DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
+CC26 DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
+CC27 DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
+CC28 DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
+CC29 DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
+CC30 DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
+CC31 DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
+CC32 DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
+CC33 DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
+CC34 DecisionQuotient.ClaimClosure.minsuff_collapse_core
+CC36 DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
+CC37 DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
+CC38 DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
+CC39 DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
+CC40 DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
+CC41 DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
+CC42 DecisionQuotient.ClaimClosure.one_step_bridge
+CC43 DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
+CC44 DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
+CC45 DecisionQuotient.ClaimClosure.physical_crossover_core
+CC46 DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
+CC47 DecisionQuotient.ClaimClosure.physical_crossover_policy_core
+CC48 DecisionQuotient.ClaimClosure.process_bridge_failure_witness
+CC49 DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
+CC50 DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
+CC51 DecisionQuotient.ClaimClosure.regime_core_claim_proved
+CC52 DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
+CC53 DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
+CC54 DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
+CC55 DecisionQuotient.ClaimClosure.separable_detectable
+CC56 DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
+CC57 DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
+CC58 DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
+CC59 DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
+CC60 DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
+CC61 DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
+CC62 DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
+CC63 DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
+CC64 DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+CC65 DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
+CC66 DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
+CC67 DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
+CC68 DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
+CC69 DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
+CC70 DecisionQuotient.ClaimClosure.tractable_bounded_core
+CC71 DecisionQuotient.ClaimClosure.tractable_separable_core
+CC72 DecisionQuotient.ClaimClosure.tractable_subcases_conditional
+CC73 DecisionQuotient.ClaimClosure.tractable_tree_core
+CC74 DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
+CC75 DecisionQuotient.ClaimClosure.tree_structure_detectable
+CC76 DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
+CC77 DecisionQuotient.ClaimClosure.typed_static_class_completeness
+CR1 DecisionQuotient.ConfigReduction.config_sufficiency_iff_behavior_preserving
+CT1 CT.law_gap_physical_claim_holds
+CT2 CT.no_law_gap_counterexample
+CT3 CT.no_physical_counterexample_of_core_theorem
+CT4 CT.physical_bridge_bundle
+CT5 CT.physical_claim_lifts_from_core
+CT6 CT.physical_claim_lifts_from_core_conditional
+CT7 CT.physical_counterexample_invalidates_core_rule
+CT8 CT.physical_counterexample_yields_core_counterexample
+DP1 DecisionQuotient.DecisionProblem.minimalSufficient_iff_relevant
+DP2 DecisionQuotient.DecisionProblem.relevantSet_is_minimal
+DP3 DecisionQuotient.DecisionProblem.sufficient_implies_selectorSufficient
+DP4 DecisionQuotient.ClaimClosure.DecisionProblem.epsOpt_zero_eq_opt
+DP5 DecisionQuotient.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
+DP6 DecisionProblem.anchorSufficient
+DP7 DecisionProblem.classMonotoneOn
+DP8 DecisionProblem.constant_opt_all_sufficient
+DP21 DecisionProblem.edgeOnComplement_iff_not_sufficient
+DQ3 DecisionQuotient.card_anchoredSlice_eq_pow_sub
+DQ5 DecisionQuotient.opt_eq_feasible_of_gap
+DQ6 DecisionQuotient.opt_eq_of_allowed_iff
+DQ7 DecisionQuotient.opt_singleton_of_logicallyDeterministic
+DQ8 DecisionQuotient.universe_sets_objective_schema
+DQ154 DecisionQuotient.Physics.Instantiation.DecisionCircuit
+DQ155 DecisionQuotient.Physics.Instantiation.DecisionInterpretation
+DQ159 DecisionQuotient.Physics.Instantiation.MoleculeAsCircuit
+DQ160 DecisionQuotient.Physics.Instantiation.MoleculeAsDecisionCircuit
+DQ517 DecisionQuotient.Physics.Instantiation.Circuit
+DQ518 DecisionQuotient.Physics.Instantiation.Dynamics
+DQ519 DecisionQuotient.Physics.Instantiation.Geometry
+DQ533 DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
+DQ534 DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
+DQ535 DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
+DQ537 DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
+DQ539 DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
+DQ540 DecisionQuotient.Physics.DecisionTime.run_time_exact
+DQ541 DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
+DQ542 DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
+DQ545 DecisionQuotient.Physics.DecisionTime.tick_increments_time
+DQ546 DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
+DQ547 DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
+DQ548 DecisionQuotient.Physics.DecisionTime.time_is_discrete
+DQ549 DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
+DQ560 DecisionQuotient.HardnessDistribution.hardnessLowerBound
+DQ561 DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
+DQ562 DecisionQuotient.HardnessDistribution.requiredWork
+DQ563 DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
+DQ564 DecisionQuotient.HardnessDistribution.workGrowthDegree
+DQ565 DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
+DQ566 DecisionQuotient.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
+DQ567 DecisionQuotient.Physics.PhysicalCore.certified_work_lower_bound_raw
+DQ568 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification
+DQ569 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
+DQ570 DecisionQuotient.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
+DQ571 DecisionQuotient.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
+DQ572 DecisionQuotient.Physics.PhysicalCore.substrate_transport_sound_conditional
+DQ573 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget
+DQ574 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
+DQ575 DecisionQuotient.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split
+DQ576 DecisionQuotient.PhysicalBudgetCrossover.exists_least_crossover_point
+DQ577 DecisionQuotient.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
+DQ578 DecisionQuotient.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
+DQ579 DecisionQuotient.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
+DQ580 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_at_divergence
+DQ581 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_beyond_divergence
+DQ582 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_collision
+DQ583 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
+DQ584 DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
+DQ585 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
+DQ586 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
+DQ587 DecisionQuotient.checking_witnessing_duality
+DQ588 DecisionQuotient.checking_witnessing_duality_budget
+HD1 DecisionQuotient.HardnessDistribution.centralization_dominance_bundle
+HD2 DecisionQuotient.HardnessDistribution.centralization_step_saves_n_minus_one
+HD3 DecisionQuotient.HardnessDistribution.centralized_higher_leverage
+HD4 DecisionQuotient.HardnessDistribution.complete_model_dominates_after_threshold
+HD5 DecisionQuotient.HardnessDistribution.gap_conservation_card
+HD6 DecisionQuotient.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant
+HD7 DecisionQuotient.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
+HD8 DecisionQuotient.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
+HD9 DecisionQuotient.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
+HD10 DecisionQuotient.HardnessDistribution.generalized_right_eventually_dominates_wrong
+HD11 DecisionQuotient.HardnessDistribution.hardnessEfficiency_eq_central_share
+HD12 DecisionQuotient.HardnessDistribution.isRightHardness
+HD13 DecisionQuotient.HardnessDistribution.isWrongHardness
+HD14 DecisionQuotient.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
+HD15 DecisionQuotient.HardnessDistribution.native_dominates_manual
+HD16 DecisionQuotient.HardnessDistribution.no_positive_slope_linear_represents_saturating
+HD17 DecisionQuotient.HardnessDistribution.right_dominates_wrong
+HD18 DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
+HD19 DecisionQuotient.HardnessDistribution.simplicityTax_grows
+HD20 DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
+HD21 DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic
+IC1 DecisionQuotient.IntegrityCompetence.CertaintyInflation
+IC2 DecisionQuotient.IntegrityCompetence.CompletionFractionDefined
+IC3 DecisionQuotient.IntegrityCompetence.EvidenceForReport
+IC4 DecisionQuotient.IntegrityCompetence.ExactCertaintyInflation
+IC5 DecisionQuotient.IntegrityCompetence.Percent
+IC6 DecisionQuotient.IntegrityCompetence.RLFFWeights
+IC7 DecisionQuotient.IntegrityCompetence.ReportSignal
+IC8 DecisionQuotient.IntegrityCompetence.ReportBitModel
+IC9 DecisionQuotient.IntegrityCompetence.SignalConsistent
+IC10 DecisionQuotient.IntegrityCompetence.admissible_irrational_strictly_more_than_rational
+IC11 DecisionQuotient.IntegrityCompetence.admissible_matrix_counts
+IC12 DecisionQuotient.IntegrityCompetence.abstain_signal_exists_with_guess_self
+IC13 DecisionQuotient.IntegrityCompetence.certaintyInflation_iff_not_admissible
+IC14 DecisionQuotient.IntegrityCompetence.certificationOverheadBits
+IC15 DecisionQuotient.IntegrityCompetence.certificationOverheadBits_of_evidence
+IC16 DecisionQuotient.IntegrityCompetence.certificationOverheadBits_of_no_evidence
+IC17 DecisionQuotient.IntegrityCompetence.certifiedTotalBits
+IC18 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_ge_raw
+IC19 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence
+IC20 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_of_evidence
+IC21 DecisionQuotient.IntegrityCompetence.certifiedTotalBits_of_no_evidence
+IC22 DecisionQuotient.IntegrityCompetence.claim_admissible_of_evidence
+IC23 DecisionQuotient.IntegrityCompetence.competence_implies_integrity
+IC24 DecisionQuotient.IntegrityCompetence.completion_fraction_defined_of_declared_bound
+IC25 DecisionQuotient.IntegrityCompetence.epsilon_competence_implies_integrity
+IC26 DecisionQuotient.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
+IC27 DecisionQuotient.IntegrityCompetence.evidence_of_claim_admissible
+IC28 DecisionQuotient.IntegrityCompetence.exact_claim_admissible_iff_exact_evidence_nonempty
+IC29 DecisionQuotient.IntegrityCompetence.exact_claim_requires_evidence
+IC30 DecisionQuotient.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence
+IC31 DecisionQuotient.IntegrityCompetence.exact_raw_only_of_no_exact_admissible
+IC32 DecisionQuotient.IntegrityCompetence.integrity_forces_abstention
+IC33 DecisionQuotient.IntegrityCompetence.integrity_not_competent_of_nonempty_scope
+IC34 DecisionQuotient.IntegrityCompetence.integrity_resource_bound
+IC35 DecisionQuotient.IntegrityCompetence.no_completion_fraction_without_declared_bound
+IC36 DecisionQuotient.IntegrityCompetence.overModelVerdict_rational_iff
+IC37 DecisionQuotient.IntegrityCompetence.percentZero
+IC38 DecisionQuotient.IntegrityCompetence.rlffBaseReward
+IC39 DecisionQuotient.IntegrityCompetence.rlffReward
+IC40 DecisionQuotient.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
+IC41 DecisionQuotient.IntegrityCompetence.rlff_maximizer_has_evidence
+IC42 DecisionQuotient.IntegrityCompetence.rlff_maximizer_is_admissible
+IC43 DecisionQuotient.IntegrityCompetence.self_reflected_confidence_not_certification
+IC44 DecisionQuotient.IntegrityCompetence.signal_certified_positive_implies_admissible
+IC45 DecisionQuotient.IntegrityCompetence.signal_consistent_of_claim_admissible
+IC46 DecisionQuotient.IntegrityCompetence.signal_no_evidence_forces_zero_certified
+IC47 DecisionQuotient.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
+IV1 DecisionQuotient.InteriorVerification.GoalClass
+IV2 DecisionQuotient.InteriorVerification.InteriorDominanceVerifiable
+IV3 DecisionQuotient.InteriorVerification.TautologicalSetIdentifiable
+IV4 DecisionQuotient.InteriorVerification.agreeOnSet
+IV5 DecisionQuotient.InteriorVerification.interiorParetoDominates
+IV6 DecisionQuotient.InteriorVerification.interior_certificate_implies_non_rejection
+IV7 DecisionQuotient.InteriorVerification.interior_dominance_implies_universal_non_rejection
+IV8 DecisionQuotient.InteriorVerification.interior_dominance_not_full_sufficiency
+IV9 DecisionQuotient.InteriorVerification.interior_verification_tractable_certificate
+L1 ExactPhysicalClaimWellTyped
+L2 ExcusedBy
+L3 ExplicitHardnessAssumptions
+L4 Literal.eval
+L5 OneStepSequentialObjective
+L6 RegimeSimulation
+L7 StandardComplexityAssumptions
+L8 StochasticCriterionObjective
+L9 StructureDetectable
+L10 TransitionCoupledObjective
+L11 TwoStepObjective
+L12 adq
+L13 adq_ordering
+L14 agentBridgeClass
+L15 agent_transfer_licensed_iff_snapshot
+L16 anchor_sigma2p_complete_conditional
+L17 anchor_sigma2p_reduction_core
+L19 boundaryCharacterized_iff_exists_sufficient_subset
+L20 bounded_actions_detectable
+L21 bounded_actions_reusable_heuristic
+L22 bridgeFailureWitness
+L23 bridgeTransferLicensed
+L24 bridge_boundary_represented_family
+L25 boolHypercube_node_count
+L68 node_count_does_not_determine_edge_geometry
+PBC1 DecisionQuotient.PhysicalBudgetCrossover.CrossoverAt
+PBC2 DecisionQuotient.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
+PC10 PhysicalComplexity.AccessRegime.AccessRegime
+PC12 PhysicalComplexity.AccessRegime.RegimeEval
+PC15 PhysicalComplexity.AccessRegime.RegimeWithCertificate
+PC17 PhysicalComplexity.AccessRegime.certificate_upgrades_regime
+PC19 PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate
+PC20 PhysicalComplexity.AccessRegime.AccessChannelLaw
+PC21 PhysicalComplexity.AccessRegime.FiveWayMeet
+S2P1 DecisionQuotient.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
+S2P2 DecisionQuotient.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
+S2P3 DecisionQuotient.Sigma2PHardness.min_sufficient_set_iff_relevant_card
+S2P6 DecisionQuotient.Sigma2PHardness.representationGap_eq_zero_iff
+A.2 What Is Machine-Checked
 The Lean formalization establishes:
@@
   3. 3.
-    Economic theorems: The Simplicity Tax conservation laws and hardness distribution results.
-Complexity classifications (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follow from informal composition with standard results (TAUTOLOGY is coNP-complete, etc.).
-The Lean proofs verify the reduction constructions; the complexity class membership is derived by combining these with established theorems from complexity theory.
-A.2 Polynomial-Time Reduction Definition
+    Economic theorems: Simplicity Tax redistribution identities and hardness distribution results.
+  4. 4.
+    Query-access lower-bound core: Formalized Boolean query-model indistinguishability theorem for the full problem via the $I=\emptyset$ subproblem (emptySufficiency_query_indistinguishable_pair), plus obstruction-scale identities (queryComplexityLowerBound, exponential_query_complexity).
+Complexity classifications (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follow by conditional composition with standard results (e.g., TAUTOLOGY coNP-completeness and $\exists\forall$-SAT $\Sigma_{2}^{P}$-completeness), represented explicitly as hypotheses in the conditional transfer theorems listed below.
+The Lean proofs verify the reduction constructions and the transfer closures under those hypotheses.
+The assumptions themselves are unpacked by an explicit ledger projection theorem (CC5) so dependency tracking is machine-auditable.
+A.3 Assumption Ledger (Auto)
+Source files.
+  • DecisionQuotient/ClaimClosure.lean
+Assumption bundles and fields.
+  • (No assumption bundles parsed.)
+Conditional closure handles.
+  • DQ.anchor_sigma2p_complete_conditional
+  • DQ.cost_asymmetry_eth_conditional
+  • DQ.dichotomy_conditional
+  • DQ.minsuff_collapse_to_conp_conditional
+  • DQ.minsuff_conp_complete_conditional
+  • DQ.sufficiency_conp_complete_conditional
+  • DQ.tractable_subcases_conditional
+A.4 Polynomial-Time Reduction Definition
 We use Mathlib’s Turing machine framework to define polynomial-time computability:
@@
 This uses the standard definition: a reduction is polynomial-time computable via Turing machines and preserves membership.
-A.3 The Main Reduction Theorem
+A.5 The Main Reduction Theorem
 Theorem A.1 (TAUTOLOGY Reduction Correctness, Lean).
@@
   • At assignment states, accept is optimal iff $\varphi(a)=\text{true}$
-A.4 Economic Results
+A.6 Economic Results
 The hardness distribution theorems (Section 9) are fully formalized:
-theorem simplicityTax_conservation :
-    simplicityTax P T + (P.requiredAxes inter T.nativeAxes).card
-      = P.requiredAxes.card
-theorem simplicity_preference_fallacy (T_simple T_complex : Tool)
-    (h_simple_incomplete : isIncomplete P T_simple)
-    (h_complex_complete : isComplete P T_complex)
-    (n : Nat) (hn : n > 0) :
-    totalExternalWork P T_complex n < totalExternalWork P T_simple n
-A.5 Module Structure
+theorem simplicityTax_conservation (P : SpecificationProblem)
+    (S : SolutionArchitecture P) :
+    S.centralDOF + simplicityTax P S  P.intrinsicDOF
+theorem simplicityTax_grows (P : SpecificationProblem)
+    (S : SolutionArchitecture P) (n n : )
+    (hn : n < n) (htax : simplicityTax P S > 0) :
+    totalDOF S n < totalDOF S n
+theorem native_dominates_manual (P : SpecificationProblem) (n : Nat)
+    (hn : n > P.intrinsicDOF) :
+    totalDOF (nativeTypeSystem P) n < totalDOF (manualApproach P) n
+theorem totalDOF_eventually_constant_iff_zero_distributed
+    (S : SolutionArchitecture P) :
+    IsEventuallyConstant (fun n => totalDOF S n)  S.distributedDOF = 0
+theorem no_positive_slope_linear_represents_saturating
+    (c d K : ) (hd : d > 0) :
+     ( n, c + n * d = generalizedTotalDOF c (saturatingSiteCost K) n)
+Identifier note.
+Lean identifiers retain internal naming (intrinsicDOF, simplicityTax_conservation); in paper terminology these correspond to baseline hardness and the redistribution lower-bound identity, respectively.
+A.7 Complete Claim Coverage Matrix
+Paper handle Status Lean support
+cor:exact-identifiability Full DQ.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
+cor:exact-no-competence-zero-certified Full DQ.IntegrityCompetence.rlff_maximizer_has_evidence
+cor:finite-budget-no-exact-admissibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget, DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
+cor:finite-budget-threshold-impossibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement, DQ.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification, DQ.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
+cor:gap-externalization Full DQ.HardnessDistribution.saturatingSiteCost_eventually_constant, DQ.HardnessDistribution.totalDOF_ge_intrinsic
+cor:gap-minimization-hard Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
+cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
+cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core
+cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.IntegrityCompetence.competence_implies_integrity
+cor:information-barrier-query Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.information_barrier_state_batch_core, DQ.ClaimClosure.information_barrier_value_entry_core
+cor:integrity-universal Full DQ.ClaimClosure.integrity_universal_applicability_core
+cor:interior-singleton-certificate Full DQ.InteriorVerification.interior_verification_tractable_certificate
+cor:linear-positive-no-saturation Full DQ.HardnessDistribution.native_dominates_manual
+cor:no-auto-minimize Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
+cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_uncertified_exact_claim_core
+cor:no-universal-survivor-no-succinct-bound Full DQ.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
+cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
+cor:overmodel-diagnostic-implication Full DQ.Sigma2PHardness.representationGap_eq_zero_iff
+cor:physical-counterexample-core-failure Full CT.no_physical_counterexample_of_core_theorem, CT.physical_counterexample_invalidates_core_rule, CT.physical_counterexample_yields_core_counterexample, DQ.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
+cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
+cor:practice-bounded Full agent_transfer_licensed_iff_snapshot
+cor:practice-diagnostic Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
+cor:practice-structured Full anchor_sigma2p_complete_conditional
+cor:practice-tree Full anchor_sigma2p_reduction_core
+cor:practice-unstructured Full ExactPhysicalClaimWellTyped
+cor:query-obstruction-bool Full DQ.ClaimClosure.query_obstruction_boolean_corollary, RegimeSimulation
+cor:right-wrong-hardness Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
+cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
+cor:type-system-threshold Full DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
+prop:abstain-guess-self-signal Full DQ.IntegrityCompetence.signal_no_evidence_forces_zero_certified
+prop:adq-ordering Full DQ.ClaimClosure.adq_ordering
+prop:attempted-competence-matrix Full DQ.IntegrityCompetence.RLFFWeights, DQ.IntegrityCompetence.ReportSignal, DQ.IntegrityCompetence.evidence_of_claim_admissible
+prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice, DQ.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
+prop:bridge-failure-horizon Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
+prop:bridge-failure-stochastic Full DQ.ClaimClosure.standard_assumption_ledger_unpack
+prop:bridge-failure-transition Full DQ.ClaimClosure.tractable_tree_core
+prop:bridge-transfer-scope Full DQ.ClaimClosure.one_step_bridge
+prop:budgeted-crossover Full DQ.ClaimClosure.physical_crossover_core, DQ.PhysicalBudgetCrossover.CrossoverAt
+prop:certainty-inflation-iff-inadmissible Full DQ.IntegrityCompetence.ReportBitModel, DQ.IntegrityCompetence.claim_admissible_of_evidence
+prop:certified-confidence-gate Full DQ.IntegrityCompetence.rlffBaseReward, DQ.IntegrityCompetence.rlffReward
+prop:checking-witnessing-duality Full DQ.checking_witnessing_duality, DQ.checking_witnessing_duality_budget
+prop:crossover-above-cap Full DQ.ClaimClosure.physical_crossover_above_cap_core, DQ.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
+prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_hardness_core
+prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_policy_core
+prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_increments_time, DQ.Physics.DecisionTime.tick_is_decision_event
+prop:declared-contract-selection-validity Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core, DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core, DQ.ClaimClosure.tree_structure_detectable
+prop:dominance-modes Full DQ.HardnessDistribution.centralized_higher_leverage
+prop:empty-sufficient-constant Full DP.anchorSufficient
+prop:eventual-explicit-infeasibility Full DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
+prop:evidence-admissibility-equivalence Full DQ.IntegrityCompetence.certifiedTotalBits, DQ.IntegrityCompetence.certifiedTotalBits_of_evidence, DQ.IntegrityCompetence.certifiedTotalBits_of_no_evidence
+prop:exact-requires-evidence Full DQ.IntegrityCompetence.no_completion_fraction_without_declared_bound, DQ.IntegrityCompetence.overModelVerdict_rational_iff
+prop:fraction-defined-under-bound Full DQ.IntegrityCompetence.signal_consistent_of_claim_admissible
+prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
+prop:hardness-conservation Full DQ.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
+prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
+prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.HeisenbergStrongBinding, DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
+prop:heuristic-reusability Full DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.reusable_heuristic_of_detectable, DQ.ClaimClosure.separable_detectable, DQ.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
+prop:insufficiency-counterexample Full DP.classMonotoneOn, DP.constant_opt_all_sufficient
+prop:integrity-competence-separation Full DQ.IntegrityCompetence.certifiedTotalBits_ge_raw, DQ.IntegrityCompetence.epsilon_competence_implies_integrity
+prop:integrity-resource-bound Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
+prop:interior-one-sidedness Full DQ.InteriorVerification.interior_certificate_implies_non_rejection, DQ.InteriorVerification.interior_dominance_not_full_sufficiency
+prop:interior-universal-non-rejection Full DQ.InteriorVerification.interiorParetoDominates
+prop:interior-verification-tractable Full DQ.InteriorVerification.agreeOnSet, DQ.InteriorVerification.interior_dominance_implies_universal_non_rejection
+prop:law-instance-objective-bridge Full CT.law_gap_physical_claim_holds, CT.no_law_gap_counterexample
+prop:least-divergence-point Full DQ.PhysicalBudgetCrossover.exists_least_crossover_point
+prop:minimal-relevant-equiv Full DQ.DecisionProblem.relevantSet_is_minimal, DQ.DecisionProblem.sufficient_implies_selectorSufficient
+prop:no-evidence-zero-certified Full DQ.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
+prop:no-fraction-without-bound Full DQ.IntegrityCompetence.signal_certified_positive_implies_admissible
+prop:one-step-bridge Full DQ.ClaimClosure.one_step_bridge
+prop:oracle-lattice-strict Full ExplicitHardnessAssumptions, L.eval
+prop:oracle-lattice-transfer Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, bounded_actions_detectable
+prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+prop:payoff-threshold Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
+prop:physical-claim-transport Full CT.physical_claim_lifts_from_core, CT.physical_claim_lifts_from_core_conditional, DQ.Physics.PhysicalCore.substrate_transport_sound_conditional
+prop:physics-no-universal-exact Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core
+prop:policy-closure-beyond-divergence Full DQ.PhysicalBudgetCrossover.policy_closure_at_divergence, DQ.PhysicalBudgetCrossover.policy_closure_beyond_divergence
+prop:query-finite-state-generalization Full StandardComplexityAssumptions, adq_ordering
+prop:query-randomized-robustness Full OneStepSequentialObjective, TwoStepObjective, adq
+prop:query-randomized-weighted Full bridgeTransferLicensed, bridge_boundary_represented_family
+prop:query-regime-obstruction Full DQ.ClaimClosure.query_obstruction_finite_state_core, StandardComplexityAssumptions, adq_ordering
+prop:query-state-batch-lb Full StochasticCriterionObjective, agentBridgeClass
+prop:query-subproblem-transfer Full DQ.ClaimClosure.regime_simulation_transfers_hardness, DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.subproblem_hardness_lifts_to_full
+prop:query-tightness-full-scan Full TransitionCoupledObjective
+prop:query-value-entry-lb Full StructureDetectable, boundaryCharacterized_iff_exists_sufficient_subset
+prop:query-weighted-transfer Full bounded_actions_reusable_heuristic, bridgeFailureWitness
+prop:raw-certified-bit-split Full DQ.ClaimClosure.certified_total_bits_split_core, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.Physics.PhysicalCore.certified_work_lower_bound_raw
+prop:rlff-maximizer-admissible Full DQ.IntegrityCompetence.exact_raw_only_of_no_exact_admissible, DQ.IntegrityCompetence.integrity_forces_abstention
+prop:run-time-accounting Full DQ.Physics.DecisionTime.decisionTrace_length_eq_ticks, DQ.Physics.DecisionTime.decision_count_equals_elapsed_time, DQ.Physics.DecisionTime.run_elapsed_time_eq_ticks, DQ.Physics.DecisionTime.run_time_exact
+prop:selector-separation Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
+prop:self-confidence-not-certification Full DQ.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
+prop:set-to-selector Full DQ.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
+prop:snapshot-process-typing Full DQ.ClaimClosure.agent_transfer_licensed_iff_snapshot, DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.snapshot_vs_process_typed_boundary
+prop:steps-run-scalar Full DQ.IntegrityCompetence.rlff_maximizer_is_admissible, DQ.IntegrityCompetence.self_reflected_confidence_not_certification
+prop:substrate-unit-time Full DQ.Physics.DecisionTime.substrate_step_is_time_unit, DQ.Physics.DecisionTime.substrate_step_realizes_decision_event, DQ.Physics.DecisionTime.time_unit_law_substrate_invariant
+prop:sufficiency-char Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_dq_ratio
+prop:thermo-conservation-additive Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+prop:thermo-hardness-bundle Full DQ.ClaimClosure.thermo_eventual_lift_core
+prop:thermo-lift Full DQ.ClaimClosure.thermo_conservation_additive_core, DQ.ClaimClosure.thermo_energy_carbon_lift_core
+prop:thermo-mandatory-cost Full DQ.ClaimClosure.thermo_hardness_bundle_core
+prop:time-discrete Full DQ.Physics.DecisionTime.time_coordinate_falsifiable, DQ.Physics.DecisionTime.time_is_discrete
+prop:typed-claim-admissibility Full DQ.ClaimClosure.tree_structure_detectable
+prop:under-resolution-collision Full DQ.Physics.PhysicalIncompleteness.under_resolution_implies_collision, DQ.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
+prop:universal-solver-framing Full DQ.ClaimClosure.typed_static_class_completeness
+prop:zero-epsilon-competence Full DQ.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence, DQ.IntegrityCompetence.integrity_not_competent_of_nonempty_scope
+prop:zero-epsilon-reduction Full DQ.ClaimClosure.DecisionProblem.epsOpt_zero_eq_opt, DQ.DecisionProblem.minimalSufficient_iff_relevant
+thm:amortization Full DQ.HardnessDistribution.centralized_higher_leverage
+thm:bridge-boundary-represented Full DQ.ClaimClosure.bridge_boundary_represented_family, DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.ClaimClosure.bridge_transfer_iff_one_step_class
+thm:centralization-dominance Full DQ.HardnessDistribution.centralization_dominance_bundle, DQ.HardnessDistribution.centralization_step_saves_n_minus_one
+thm:claim-integrity-meta Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+thm:config-reduction Full DQ.ConfigReduction.config_sufficiency_iff_behavior_preserving
+thm:cost-asymmetry-eth Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.HardnessDistribution.isWrongHardness
+thm:dichotomy Full DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
+thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
+thm:generalized-dominance Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
+thm:generalized-saturation-possible Full DQ.HardnessDistribution.gap_conservation_card, DQ.HardnessDistribution.right_dominates_wrong
+thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.simplicityTax_grows
+thm:overmodel-diagnostic Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
+thm:physical-bridge-bundle Full CT.physical_bridge_bundle
+thm:regime-coverage Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.regime_core_claim_proved
+thm:tax-conservation Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
+thm:tax-grows Full DQ.HardnessDistribution.totalDOF_ge_intrinsic
+thm:tractable Full DQ.ClaimClosure.thermo_mandatory_cost_core, DQ.ClaimClosure.tractable_bounded_core, DQ.ClaimClosure.tractable_separable_core, DQ.ClaimClosure.tractable_subcases_conditional
+thm:typed-completeness-static Full DQ.ClaimClosure.typed_claim_admissibility_core
+Auto summary: mapped 123/123 (full=123, derived=0, unmapped=0).
+A.8 Proof Hardness Index (Auto)
+Paper handle Hardness profile Regime tags Lean support
+cor:gap-externalization cost-growth H=cost-growth HD18, HD21
+cor:generalized-eventual-dominance cost-growth H=cost-growth HD9
+cor:linear-positive-no-saturation cost-growth H=cost-growth HD15
+cor:right-wrong-hardness cost-growth H=cost-growth HD16
+cor:type-system-threshold cost-growth H=cost-growth HD14
+prop:dominance-modes cost-growth H=cost-growth HD3
+prop:generalized-assumption-boundary cost-growth H=cost-growth HD6, HD7
+prop:hardness-conservation cost-growth H=cost-growth HD20
+prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD10
+prop:thermo-conservation-additive cost-growth H=cost-growth CC64
+prop:thermo-hardness-bundle cost-growth H=cost-growth CC67
+prop:thermo-lift cost-growth H=cost-growth CC65, CC66
+prop:thermo-mandatory-cost cost-growth H=cost-growth CC68
+thm:generalized-dominance cost-growth H=cost-growth HD8
+thm:generalized-saturation-possible cost-growth H=cost-growth HD5, HD17
+thm:linear-saturation-iff-zero cost-growth H=cost-growth HD19
+cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional DQ573, DQ574
+cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional DQ574, DQ568, DQ569
+cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional DQ578
+prop:budgeted-crossover exp-lb-conditional H=exp-lb-conditional CC45, PBC1
+prop:crossover-above-cap exp-lb-conditional H=exp-lb-conditional CC44, PBC2
+prop:crossover-not-certification exp-lb-conditional H=exp-lb-conditional CC46
+prop:crossover-policy exp-lb-conditional H=exp-lb-conditional CC47
+prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional DQ577
+prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional DQ576
+prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional DQ575, DQ579
+prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional DQ580, DQ581
+cor:information-barrier-query query-lb H=query-lb CC27, CC28, CC29
+cor:query-obstruction-bool query-lb H=query-lb CC49, L6
+prop:checking-witnessing-duality query-lb H=query-lb DQ587, DQ588
+prop:query-finite-state-generalization query-lb H=query-lb L7, L13
+prop:query-randomized-robustness query-lb H=query-lb L5, L11, L12
+prop:query-randomized-weighted query-lb H=query-lb L23, L24
+prop:query-tightness-full-scan query-lb H=query-lb L10
+prop:query-weighted-transfer query-lb H=query-lb L21, L22
+cor:exact-identifiability succinct-hard H=succinct-hard S2P1
+cor:gap-minimization-hard succinct-hard H=succinct-hard S2P2
+cor:no-auto-minimize succinct-hard H=succinct-hard CC37
+cor:overmodel-diagnostic-implication succinct-hard H=succinct-hard S2P6
+cor:practice-diagnostic succinct-hard H=succinct-hard S2P2
+prop:sufficiency-char succinct-hard H=succinct-hard CC62, CC63
+cor:interior-singleton-certificate tractable-structured H=tractable-structured IV9
+cor:practice-bounded tractable-structured H=tractable-structured L15
+cor:practice-structured tractable-structured H=tractable-structured L16
+cor:practice-tree tractable-structured H=tractable-structured L17
+prop:heuristic-reusability tractable-structured H=tractable-structured CC7, CC53, CC55, CC74
+prop:interior-verification-tractable tractable-structured H=tractable-structured IV4, IV7
+cor:exact-no-competence-zero-certified unspecified AR,DC IC41
+cor:hardness-exact-certainty-inflation unspecified CR CC19
+cor:hardness-raw-only-exact unspecified AR CC21, IC23
+cor:integrity-universal unspecified TR CC31
+cor:no-uncertified-exact-claim unspecified AR CC41
+cor:outside-excuses-no-exact-report unspecified CR,DC CC39
+cor:physical-counterexample-core-failure unspecified AR,E,Qb,Qf,S,S+ETH CT3, CT7, CT8, DQ571
+cor:physics-no-universal-exact-claim unspecified AR,CR,DC CC38
+cor:practice-unstructured unspecified - L1
+cor:rlff-abstain-no-certs unspecified AR IC30, DQ570
+prop:abstain-guess-self-signal unspecified AR,DC IC46
+prop:adq-ordering unspecified - CC2
+prop:attempted-competence-matrix unspecified AR IC6, IC7, IC27
+prop:bounded-slice-meta-irrelevance unspecified AR CC32, CC33
+prop:bridge-failure-horizon unspecified - CC25, CC26
+prop:bridge-failure-stochastic unspecified - CC57
+prop:bridge-failure-transition unspecified - CC73
+prop:bridge-transfer-scope unspecified - CC42
+prop:certainty-inflation-iff-inadmissible unspecified CR IC8, IC22
+prop:certified-confidence-gate unspecified AR,DC IC38, IC39
+prop:decision-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ535, DQ537, DQ545, DQ546
+prop:declared-contract-selection-validity unspecified - CC22, CC39, CC75
+prop:empty-sufficient-constant unspecified AR,DC,DM DP6
+prop:evidence-admissibility-equivalence unspecified AR,DC IC17, IC20, IC21
+prop:exact-requires-evidence unspecified AR,DC IC35, IC36
+prop:fraction-defined-under-bound unspecified AR,DC IC45
+prop:heisenberg-strong-nontrivial-opt unspecified AR DQ584, DQ585, DQ586
+prop:insufficiency-counterexample unspecified AR,DC,DM DP7, DP8
+prop:integrity-competence-separation unspecified AR IC18, IC25
+prop:integrity-resource-bound unspecified AR,DC,S,S+ETH CC30, IC24, IC26
+prop:interior-one-sidedness unspecified AR IV6, IV8
+prop:interior-universal-non-rejection unspecified AR IV5
+prop:law-instance-objective-bridge unspecified AR CT1, CT2
+prop:minimal-relevant-equiv unspecified - DP2, DP3
+prop:no-evidence-zero-certified unspecified AR,DC IC40
+prop:no-fraction-without-bound unspecified AR,DC IC44
+prop:one-step-bridge unspecified - CC42
+prop:oracle-lattice-strict unspecified AR,E,Qb,Qf,S,S+ETH L3, L4
+prop:oracle-lattice-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC43, L20
+prop:outside-excuses-explicit-assumptions unspecified CR,DC CC22
+prop:physical-claim-transport unspecified AR,E,LC,Qb,Qf,S,S+ETH,TA CT5, CT6, DQ572
+prop:physics-no-universal-exact unspecified AR,CR,DC CC15
+prop:query-regime-obstruction unspecified AR,E,Q_fin,Qb,Qf,S,S+ETH CC50, L7, L13
+prop:query-state-batch-lb unspecified AR,E,Qb,Qf,S,S+ETH L8, L14
+prop:query-subproblem-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC52, CC58, CC59
+prop:query-value-entry-lb unspecified AR,E,Qb,Qf,S,S+ETH L9, L19
+prop:raw-certified-bit-split unspecified AR CC11, IC10, IC11, IC13, IC14, IC15, IC16, DQ567
+prop:rlff-maximizer-admissible unspecified AR IC31, IC32
+prop:run-time-accounting unspecified AR,E,Qb,Qf,S,S+ETH DQ533, DQ534, DQ539, DQ540
+prop:selector-separation unspecified - CC54
+prop:self-confidence-not-certification unspecified AR,DC IC47
+prop:set-to-selector unspecified - DP5
+prop:snapshot-process-typing unspecified RA CC3, CC48, CC56
+prop:steps-run-scalar unspecified AR,DC IC42, IC43
+prop:substrate-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ541, DQ542, DQ549
+prop:time-discrete unspecified AR,E,Qb,Qf,S,S+ETH DQ547, DQ548
+prop:typed-claim-admissibility unspecified AR,DC,Qf,S+ETH CC75
+prop:under-resolution-collision unspecified DM DQ582, DQ583
+prop:universal-solver-framing unspecified TR CC77
+prop:zero-epsilon-competence unspecified - IC19, IC33
+prop:zero-epsilon-reduction unspecified - DP4, DP1
+thm:amortization unspecified LC HD3
+thm:bridge-boundary-represented unspecified - CC8, CC9, CC10
+thm:centralization-dominance unspecified AR,LC HD1, HD2
+thm:claim-integrity-meta unspecified AR,CR,DC CC22
+thm:config-reduction unspecified AR,E,S,TA CR1
+thm:cost-asymmetry-eth unspecified S+ETH CC12, HD13
+thm:dichotomy unspecified AR,E,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
+thm:exact-certified-gap-iff-admissible unspecified AR CC17, CC18, CC20, DQ566
+thm:overmodel-diagnostic unspecified S CC6, CC40
+thm:physical-bridge-bundle unspecified AR CT4
+thm:regime-coverage unspecified CR,S CC14, CC51
+thm:tax-conservation unspecified AR,LC HD4
+thm:tax-grows unspecified LC HD21
+thm:tractable unspecified AR,E,Qb,Qf,RG,S,S+ETH,STR,TA CC69, CC70, CC71, CC72
+thm:typed-completeness-static unspecified AR,CR,S CC76
+Auto summary: indexed 123 claims by hardness profile (cost-growth=16; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=6; unspecified=76).
+A.9 Claims Not Fully Mechanized
+Status: all theorem/proposition/corollary handles in this paper now have Lean backing.
+Entries marked Full (conditional) are explicitly mechanized transfer theorems that depend on standard external complexity facts (e.g., source-class completeness or ETH assumptions), with those dependencies represented as hypotheses in Lean.
+A.10 Module Structure
   • Basic.lean – Core definitions (DecisionProblem, sufficiency, optimality)
@@
   • Reduction.lean – TAUTOLOGY reduction construction and correctness
+  • Hardness/ConfigReduction.lean – Sentinel-action configuration reduction theorem
   • Complexity.lean – Polynomial-time reduction definitions using mathlib
-  • HardnessDistribution.lean – Simplicity Tax theorems
+  • HardnessDistribution.lean – Simplicity Tax redistribution and amortization theorems
+  • IntegrityCompetence.lean – Solver integrity vs regime competence separation
+  • ClaimClosure.lean – Mechanized closure of paper-level bridge and diagnostic claims
   • Tractability/ – Bounded actions, separable utilities, tree structure
-A.6 Verification
+A.11 Verification
 The proofs compile with Lean 4 and contain no sorry placeholders. Run lake build in the proof directory to verify.

### v4

@@
-Computational Complexity of Physical Information Sufficiency
+Computational Complexity of Physical Counting
 Abstract
-We study physical information sufficiency as a decision-theoretic meta-problem.
-For $\mathcal{D}=(A,S,U)$ with factored state space $S=X_{1}\times\cdots\times X_{n}$, a coordinate set $I$ is sufficient when
-$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
-This asks whether projected information preserves the full optimal-action correspondence.
-Complexity landscape.
-  • SUFFICIENCY-CHECK is coNP-complete; MINIMUM-SUFFICIENT-SET is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
-  • Under explicit-state encoding, polynomial-time algorithms hold for bounded actions, separable utility, and tree-structured utility.
-  • Under succinct encoding, hardness is regime-dependent: with ETH, there are worst-case families with $k^{*}=n$ requiring $2^{\Omega(n)}$ time.
-  • Under query access, the finite-state core has worst-case Opt-oracle complexity $\Omega(|S|)$, with Boolean value-entry and state-batch refinements preserving the obstruction.
-Physical grounding.
-The paper formalizes a physical-to-core encoding $E:\mathcal{P}\to\mathcal{D}$ and a transport rule: declared physical assumptions transfer to core assumptions, and core claims lift back to encoded physical instances.
-Encoded physical counterexamples induce core failures on the encoded slice.
-Discrete-time interface semantics (decision event = one tick) and budgeted thermodynamic lifts (bit lower bounds to energy/carbon lower bounds under declared constants) are formalized in the same assumption-typed framework.
-[C:P2.34,2.36,2.28,2.29,2.30,2.31,5.18,5.28;C2.35,5.22;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541,PBC1,CC65,DQ569.)
-All theorem-level claims are machine-checked in Lean 4 (13391 lines, 597 theorem/lemma statements).
-Complexity-class completeness follows by composition with standard complexity results; regime-dependent and physical-transport consequences are proved as assumption-explicit closures.
-Keywords: computational complexity, physical information, decision theory, polynomial hierarchy, formal verification
+Abstract
+We characterize which coordinates of a factored state space determine optimal actions.
+For decision problem $\mathcal{D}=(A,S,U)$ with $S=X_{1}\times\cdots\times X_{n}$, coordinate set $I$ is sufficient if $s_{I}=s^{\prime}_{I}\Rightarrow\operatorname{Opt}(s)=\operatorname{Opt}(s^{\prime})$.
+We prove fourteen first-principles theorems: thirteen from pure mathematics and logic, one from empirical assumption (temperature).
+The decision quotient $Q=S/{\sim}$ (where $s\sim s^{\prime}\Leftrightarrow\operatorname{Opt}(s)=\operatorname{Opt}(s^{\prime})$) is the minimal abstraction: any other abstraction preserving optimal actions factors uniquely through $Q$.
+Verification has a lower bound: any sound checker for the empty-set sufficiency core requires $\geq 2^{n-1}$ witness pairs.
+The chain from counting measure to probability to Bayes’ theorem to the Decision Quotient follows from finite set cardinality.
+Fisher information, entropy bounds, optimal transport, rate-distortion, and thermodynamics each independently recover $\mathrm{srank}$ as the decision complexity measure.
+Complexity classifications: SUFFICIENCY-CHECK and MINIMUM-SUFFICIENT-SET are coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete; stochastic and sequential variants are PP- and PSPACE-complete respectively, with strict static $\subset$ stochastic $\subset$ sequential separation.
+Six structural subcases admit polynomial algorithms.
+Under ETH, succinct encodings carry $2^{\Omega(n)}$ lower bounds.
+Bayesian updating is optimal: from $\log x\leq x-1$ alone we prove it uniquely minimizes expected log loss.
+Two results carry empirical conditions.
+Conditional on Landauer’s principle ($k_{B}T\ln 2$ per irreversible bit erasure, proposed 1961; experimentally supported 2012), composed with the bit-operation lower bounds above, the inequality $dU\geq\lambda\,dC$ follows by multiplication; rejecting it requires rejecting Landauer.
+Conditional on stochastic thermodynamics (Barato–Seifert 2015), the thermodynamic uncertainty relation $\mathrm{Var}(J)/\langle J\rangle^{2}\geq 2/\sigma$ bounds decision precision by entropy production, with minimal $\sigma$ scaling with $\mathrm{srank}$.
+All results are machine-checked in Lean 4 with no sorry placeholders.
+Complexity results carry their hypotheses as explicit Lean theorem parameters.
+A machine-generated assumption ledger records all conditional dependencies.
+There are no hidden axioms.
 © 2026 Tristan Simas. This work is licensed under CC BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
 1 Introduction
-Consider any decision problem $\mathcal{D}$ with actions $A$ and factored state space $S=X_{1}\times\cdots\times X_{n}$.
-A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
+1.1 First-Principles Derivations
+This section presents seventeen theorems.
+Theorems 1–14 are derived from pure mathematics (counting, set theory, arithmetic, logic).
+Theorems 15–17 use empirical assumptions.
+One irreducible empirical claim (EC1: temperature) provides calibration; counting and logic provide structure.
+Two former claims (EC2: finite resources, EC3: finite speed) are now derived from first principles.
+1.1.1 1. Counting Gap Theorem
+Theorem 1.1 (Counting Gap).
+Let a system have finite positive capacity $C>0$. Let each information-acquisition event cost $\varepsilon>0$. Then:
+$\varepsilon\cdot N\leq C\;\Rightarrow\;N\leq C.$
+No bounded system with positive event cost can perform infinitely many events. BA10
+Proof.
+In $\mathbb{N}$: every positive integer is $\geq 1$. Therefore $N\leq\varepsilon\cdot N\leq C$.  $\blacksquare$
+Consequence. Continuous abstractions require arbitrarily fine state discrimination. Bounded computational systems cannot do this. Discrete models are forced by the mathematics.
+1.1.2 2. Bayesian Optimality
+Theorem 1.2 (Bayes Minimizes Expected Log Loss).
+For distributions $p,q$ over hypothesis space $H$, cross-entropy satisfies:
+$\mathrm{CE}(p,q)=H(p)+D_{\mathrm{KL}}(p\|q)$
+where $D_{\mathrm{KL}}(p\|q)\geq 0$ (Gibbs’ inequality, from $\log x\leq x-1$). Therefore $q=p$ uniquely minimizes $\mathrm{CE}(p,q)$. FN7FN12FN14
+Proof.
+Gibbs’ inequality: $\log x\leq x-1$ implies $\sum p\log(p/q)\geq 0$. Equality iff $p=q$.  $\blacksquare$
+Consequence. Bayesian updating is optimal. The proof uses only $\log x\leq x-1$ and Lean’s type theory.
+1.1.3 3. Measure Theory Typing
+Theorem 1.3 (Measure Is Logical Prerequisite).
+  1. 1.
+    Every quantitative claim carries an explicit measure: $(\mu,f)$ where $f$ is measurable.
+  2. 2.
+    Every stochastic claim additionally requires $\mu$ to be a probability measure ($\mu(\Omega)=1$).
+  3. 3.
+    Counting measure on $\{0,1\}$ has mass $2$, not $1$: not a probability measure.
+MN1MN2MN10MN11
+Proof.
+(1) $\mathbb{E}[f]$” is undefined without specifying which measure integrates $f$.
+The expression $\int f\,d\mu$ requires $\mu$.
+(2) $P(A)=\mu(A)/\mu(\Omega)$ requires $\mu(\Omega)<\infty$; normalization $P(\Omega)=1$ requires $\mu(\Omega)=1$.
+(3) Counting measure: $\mu(\{0\})=1$, $\mu(\{1\})=1$, so $\mu(\{0,1\})=2\neq 1$.
+To normalize, divide by 2.
+$\blacksquare$
+Consequence. Quantitative and stochastic claims are typed differently. Probability normalization is not automatic from measure structure.
+1.1.4 4. Universal Property of Decision Quotient
+Theorem 1.4 (Universal Property of Decision Quotient).
+Let $Q=S/{\sim}$ be the decision quotient where $s\sim s^{\prime}\iff\text{Opt}(s)=\text{Opt}(s^{\prime})$.
+For any surjective abstraction $\phi:S\to T$ that preserves the optimal action ($\phi(s)=\phi(s^{\prime})\Rightarrow\text{Opt}(s)=\text{Opt}(s^{\prime})$), there exists a unique factorization $\psi:T\to Q$ such that $\pi=\psi\circ\phi$ where $\pi:S\to Q$ is the quotient map.
+$\begin{array}[]{ccc}S&\xrightarrow{\pi}&Q\\ \downarrow&&\uparrow\\ T&\xrightarrow{\psi}&Q\end{array}$
+The decision quotient is the coarsest abstraction through which Opt factors. QT1QT2QT3
+Proof.
+Since $\phi$ preserves Opt, we have $\phi(s)=\phi(s^{\prime})\Rightarrow\text{Opt}(s)=\text{Opt}(s^{\prime})$.
+By surjectivity, for each $t\in T$ pick some $s$ with $\phi(s)=t$ and define $\psi(t)=\pi(s)$.
+This is well-defined: if $\phi(s)=\phi(s^{\prime})=t$, then $\text{Opt}(s)=\text{Opt}(s^{\prime})$, so $\pi(s)=\pi(s^{\prime})$.
+Uniqueness follows because $\pi=\psi\circ\phi$ forces $\psi(t)=\pi(s)$ for any $s$ with $\phi(s)=t$.
+$\blacksquare$
+Consequence. The decision quotient is canonical. Any other state abstraction that preserves optimal actions must refine it.
+1.1.5 5. Witness-Checking Duality
+Theorem 1.5 (Checking-Witnessing Duality).
+For boolean decision problems with $n$ coordinates, let $W(n)=2^{n-1}$ be the witness budget and $T(n)$ be the checking budget. Any sound checker satisfies:
+$T(n)\geq W(n)=2^{n-1}$
+No finite checker with fewer than $2^{n-1}$ witness pairs can be sound on all problem instances. WD1WD2WD3
+Proof.
+$\emptyset$-sufficiency fails iff $\exists s,s^{\prime}$ with $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$.
+There are $2^{n}$ states, so $\binom{2^{n}}{2}=2^{n-1}(2^{n}-1)$ pairs.
+A sound refutation must cover all pairs that could witness failure.
+For each pair family $P$ with $|P|<2^{n-1}$, there exists an Opt function whose unique failure pair is not in PP.
+Therefore $|P|\geq 2^{n-1}$ is necessary for soundness.
+$\blacksquare$
+Consequence. Verification has a lower bound. A checker must examine exponentially many witness pairs to be sound. This is a counting lower bound on the information required for verification.
+1.1.6 6. Bayes from Counting
+Theorem 1.6 (Probability Axioms from Counting).
+The chain from counting to Decision Quotient:
+  1. 1.
+    $P(A)=|A|/|\Omega|$ satisfies Kolmogorov axioms: nonnegativity, normalization, and additivity for disjoint sets.
+  2. 2.
+    $P(H|E)=P(E|H)\cdot P(H)/P(E)$ follows from the definition of conditional probability.
+  3. 3.
+    Conditioning reduces entropy: $H(H|E)\leq H(H)$, from nonnegativity of KL divergence.
+  4. 4.
+    $\mathrm{DQ}=I(H;E)/H(H)=1-H(H|E)/H(H)$ lies in $[0,1]$.
+BC1BC2BC3BC4BC5
+Proof.
+(1) $|A|\geq 0$ (counting), $|\Omega|/|\Omega|=1$ (normalization), $|A\cup B|=|A|+|B|$ for $A\cap B=\emptyset$ (additivity).
+(2) $P(H|E):=P(H\cap E)/P(E)$.
+Rearranging: $P(H|E)\cdot P(E)=P(H\cap E)=P(E|H)\cdot P(H)$.
+(3) H(H|E)=H(H)I(H;E)H(H|E)=H(H)-I(H;E) and $I(H;E)=D_{\mathrm{KL}}(P_{H,E}\|P_{H}\otimes P_{E})\geq 0$ by Gibbs.
+(4) $I(H;E)\leq H(H)$ by (3), so $\mathrm{DQ}=I(H;E)/H(H)\in[0,1]$.
+$\blacksquare$
+Consequence.
+Bayesian updating is derived from counting measure.
+Probability theory, Bayes’ theorem, information theory, and the Decision Quotient form a single chain from counting elements of a finite set.
+1.1.7 7. Fisher Rank = Structural Rank
+Theorem 1.7 (Fisher Information = Structural Rank).
+Define the Fisher information score of coordinate $i$ as:
+$\mathrm{score}(i)=\begin{cases}1&\text{if }i\text{ is relevant}\\ 0&\text{otherwise}\end{cases}$
+Then:
+$\sum_{i=1}^{n}\mathrm{score}(i)=\mathrm{srank}(\mathcal{D})$
+The diagonal Fisher information matrix $I(\theta)_{ii}=\mathrm{score}(i)$ has rank $\mathrm{srank}(\mathcal{D})$. FS1FS2
+Proof.
+$\sum_{i}\mathrm{score}(i)=\sum_{i}\mathbf{1}[\text{isRelevant}(i)]=|\{i\mid\text{isRelevant}(i)\}|=\mathrm{srank}$.
+For the diagonal matrix: $\mathrm{rank}=|\{i\mid\mathrm{score}(i)\neq 0\}|=|\{i\mid\text{isRelevant}(i)\}|=\mathrm{srank}$.
+$\blacksquare$
+Consequence.
+The decision manifold has dimension exactly $\mathrm{srank}$.
+The Fisher information sees the relevant coordinates.
+The Cramer-Rao bound follows: any estimator has difficulty $\geq 1/\mathrm{srank}$.
+1.1.8 8. Entropy-Rank Inequality
+Theorem 1.8 (Entropy Bounded by Structural Rank).
+For a decision problem $\mathcal{D}$ with boolean coordinate spaces ($X_{i}=\{0,1\}$):
+$H(\mathcal{D})\leq\mathrm{srank}(\mathcal{D})$
+where $H(\mathcal{D})=\log_{2}(\mathrm{numOptClasses})$ is the Shannon entropy of the decision quotient. IT3IT4
+Proof.
+The relevant coordinate set $R$ is sufficient (Theorem 5.9): $\mathrm{Opt}$ factors through $\pi_{R}:S\to\{0,1\}^{|R|}$.
+Therefore $\mathrm{numOptClasses}\leq|\{0,1\}^{|R|}|=2^{\mathrm{srank}}$.
+Therefore $H(\mathcal{D})=\log_{2}(\mathrm{numOptClasses})\leq\log_{2}(2^{\mathrm{srank}})=\mathrm{srank}$.
+$\blacksquare$
+Consequence. The decision quotient carries at most $\mathrm{srank}$ bits of decision-relevant information. Coordinates outside the relevant set are information-theoretically inert.
+1.1.9 9. Functional Information
+Theorem 1.9 (Thermodynamic Cost from Counting).
+For functional set $F\subseteq\Omega$, the derivation proceeds in three steps:
+  1. 1.
+    $\mathrm{FI}_{\text{counting}}(F)=-\log_{2}(|F|/|\Omega|)$ bits. Uses only finite measure.
+  2. 2.
+    Plug in Landauer’s value $\varepsilon=k_{B}T\ln 2$ joules per bit.
+  3. 3.
+    $E_{\text{min}}=\varepsilon\cdot\mathrm{FI}_{\text{counting}}(F)$.
+The thermodynamic path is the counting path with an empirical calibration constant. FI3FI6FI7
+Proof.
+(1) By definition of surprisal: $\mathrm{FI}(F)=-\log_{2}P(F)=-\log_{2}(|F|/|\Omega|)$ using counting measure.
+No physics.
+(2) Landauer’s principle (empirical): erasing one bit costs $\geq k_{B}T\ln 2$ joules.
+This is a unit conversion constant.
+(3) Composition: $E_{\text{min}}=\varepsilon\cdot\mathrm{FI}(F)=k_{B}T\ln 2\cdot(-\log_{2}(|F|/|\Omega|))$ joules.
+The counting structure determines the bit count; Landauer supplies the energy-per-bit conversion.
+$\blacksquare$
+Consequence. Thermodynamics adds a unit conversion. Counting provides bits; Landauer provides joules per bit. Counting is primary, thermodynamics is derived.
+1.1.10 10. Optimal Transport Bridge
+Theorem 1.10 (Wasserstein Distance = Structural Rank).
+For a decision problem with $k$ distinguishable futures ($k=|\mathrm{range}(\text{Opt})|$):
+  1. 1.
+    Single future ($k=1$): transport cost vanishes.
+  2. 2.
+    Multiple futures ($k>1$): any coupling with off-diagonal mass incurs positive cost.
+  3. 3.
+    The integrity state (centroid) minimizes expected transport cost.
+Transport cost grows with the number of distinguishable futures, matching $\mathrm{srank}$. W1W2W3W4
+Proof.
+(1) When $k=1$, all states map to a single action; the diagonal coupling achieves zero cost.
+(2) Off-diagonal coupling mass moves probability between distinct actions; metric distance contributes positive cost.
+(3) The centroid minimizes $\sum_{i}w_{i}d(x,x_{i})^{2}$ by convexity of squared distance.
+$\blacksquare$
+Consequence. Optimal transport theory recovers srank as the complexity measure without invoking information theory or thermodynamics. Pure measure theory.
+1.1.11 11. Rate-Distortion Bridge
+Theorem 1.11 (Rate-Distortion = Structural Rank).
+For a decision problem viewed as a source:
+  1. 1.
+    Shannon entropy is non-negative: $H(X)\geq 0$.
+  2. 2.
+    At zero distortion: $R(0)=H(X)$.
+  3. 3.
+    Rate decreases monotonically as distortion tolerance increases.
+  4. 4.
+    Decision-equivalence preserves optimal actions: compression within equivalence classes is lossless.
+  5. 5.
+    Minimum lossless decision rate equals structural rank: $R(0)=\mathrm{srank}$.
+  6. 6.
+    Compression below srank introduces decision errors.
+RD1RD2RD3RS1RS2RS3RS4RS5
+Proof.
+Standard Shannon theory establishes (1)–(3).
+When $\phi(s)=\phi(s^{\prime})$ implies $\text{Opt}(s)=\text{Opt}(s^{\prime})$, states within equivalence classes merge without affecting optimal actions (4).
+The decision quotient partitions into $2^{\mathrm{srank}}$ classes, each requiring $\log_{2}$ bits for identification, yielding $R(0)=\mathrm{srank}$ (5).
+Fewer bits cannot distinguish all classes (6).
+$\blacksquare$
+Consequence. Rate-distortion theory independently recovers structural rank as the minimal description length for decision-making. Pure Shannon theory.
+1.1.12 12. Nontriviality from Counting
+Theorem 1.12 (Information Requires Nontriviality).
+The following chain holds from pure counting:
+  1. 1.
+    $|S|\leq 1\Rightarrow$ all states equal (cardinality).
+  2. 2.
+    All states equal $\Rightarrow$ Opt is constant (logic).
+  3. 3.
+    Opt constant $\Rightarrow$ singleton image (set theory).
+  4. 4.
+    Singleton image $\Rightarrow$ entropy $\log(1)=0$ (arithmetic).
+  5. 5.
+    Entropy $0\Rightarrow$ no information (definition).
+Contrapositive: information exists $\Rightarrow|S|\geq 2$. FP1FP2FP3FP4FP5FP6FP7
+Proof.
+Each step is elementary: (1) subsingleton property, (2) constant function on subsingleton, (3) image of constant is singleton, (4) $\log_{2}(1)=0$, (5) zero entropy means no distinguishable states.
+The contrapositive follows by modus tollens.
+$\blacksquare$
+Consequence. Nontriviality ($|S|\geq 2$) is not an axiom—it is derived from the existence of information. Pure counting.
+1.1.13 13. Second Law from Counting
+Theorem 1.13 (Second Law Structure from Counting).
+The following chain holds from pure counting and probability:
+  1. 1.
+    $|\text{target}|<|\text{total}|\Rightarrow|\text{non-target}|>0$ (counting).
+  2. 2.
+    Random selection from total probably misses target (probability).
+  3. 3.
+    $n$ unverified steps with $k$ wrong turns each: $k^{n}$ wrong paths vs. 1 right path (exponentiation).
+  4. 4.
+    Wrong paths dominate: $k^{n}\gg 1$ for $k>1$, $n>1$ (arithmetic).
+  5. 5.
+    Maintaining order requires distinguishing right from wrong (verification).
+  6. 6.
+    Verification is information (definition).
+Therefore: maintaining order requires continuous information input. FP8FP9FP10FP11FP12FP13FP14
+Proof.
+(1) Subtraction in $\mathbb{N}$.
+(2) $P(\text{miss})=|\text{non-target}|/|\text{total}|>0$.
+(3) Branching paths multiply.
+(4) Exponential growth.
+(5) Staying on target requires detecting drift.
+(6) Detection distinguishes states.
+$\blacksquare$
+Consequence. The second law’s structure (order decays without work) follows from counting. The empirical content is only the calibration constant $k_{B}T$ that converts entropy to energy.
+1.1.14 14. Landauer Structure from Counting
+Theorem 1.14 (Erasure Increases Environment Entropy).
+For any state reduction from $n$ states to $m<n$ states:
+  1. 1.
+    nm>0n-m>0 states are “erased” (arithmetic).
+  2. 2.
+    Erased states must go somewhere (conservation of distinguishability).
+  3. 3.
+    That somewhere is the environment.
+  4. 4.
+    Environment entropy increases by $\geq\log(n/m)$.
+FP15
+Proof.
+(1) $m<n\Rightarrow n-m>0$.
+(2)–(4) The missing distinctions are not destroyed; they are transferred to environmental degrees of freedom.
+This is the structure of Landauer’s principle.
+The $k_{B}T$ calibration is empirical.
+$\blacksquare$
+Consequence. Landauer’s principle splits into structure (from counting) and calibration (from temperature). Only the calibration is empirical.
+1.2 Irreducible Empirical Claims
+One claim is irreducibly empirical. Two others (finite resources, finite speed) are derived from first principles.
+Definition 1.15 (Irreducible Empirical Claim EC1).
+EC1 (Temperature): Entropy has energy cost. There exists $\varepsilon>0$ such that each bit costs $\varepsilon$ joules. Calibration: $\varepsilon=k_{B}T\ln 2$.
+Theorem 1.16 (EC2 Derived from Uncertainty).
+Finite resources (EC2) follows from the existence of uncertainty:
+  1. 1.
+    Uncertainty exists: $\exists X,0<P(X)<1$ (observation).
+  2. 2.
+    No bounds $\Rightarrow$ infinite information acquisition $\Rightarrow$ know everything $\Rightarrow$ $P(X)\in\{0,1\}$ for all $X$.
+  3. 3.
+    Contrapositive: uncertainty exists $\Rightarrow$ bounds exist.
+IA7IA9IA11IA12IA13
+Proof.
+Let observer have capacity $C$ and access $A\leq C$.
+System has entropy $H$.
+If $C\geq H$, observer can access all information (IA8).
+Full access eliminates uncertainty (IA9).
+Contrapositive: uncertainty implies $A<H$.
+If observer uses full capacity ($A=C$), then $C<H$ (IA11).
+Therefore capacity is bounded relative to system entropy.
+This is EC2.
+$\blacksquare$
+Theorem 1.17 (EC3 Derived from Non-Contradiction).
+Finite signal speed (EC3) follows from the law of non-contradiction:
+  1. 1.
+    A computational step = state change: before $\neq$ after (definition).
+  2. 2.
+    At any instant, state is unique (function property).
+  3. 3.
+    If step takes $t=0$: before-state and after-state at same instant.
+  4. 4.
+    Same instant $\Rightarrow$ state = before AND state = after.
+  5. 5.
+    But before $\neq$ after $\Rightarrow$ contradiction (violates non-contradiction).
+  6. 6.
+    Therefore $t_{\text{before}}\neq t_{\text{after}}$, so $\Delta t>0$.
+  7. 7.
+    Information propagation = sequence of steps, each taking $\Delta t>0$.
+  8. 8.
+    Speed = distance / time. Time >0>0 $\Rightarrow$ speed $\infty$.
+FPT4FPT5FPT6FPT8FPT10
+Proof.
+A step changes state from $A$ to $B$ where $A\neq B$.
+The timeline is a function $T\to S$.
+If $t_{\text{before}}=t_{\text{after}}$, then $\text{timeline}(t)=A$ and $\text{timeline}(t)=B$ for the same $t$.
+By function uniqueness, $A=B$.
+Contradiction.
+Therefore $t_{\text{before}}\neq t_{\text{after}}$.
+With ordering, $\Delta t>0$ (FPT5).
+Information propagation requires steps; each takes positive time; total time is positive; speed is finite.
+QED from pure logic.
+$\blacksquare$
+Consequence.
+EC2 and EC3 are not axioms.
+EC2 follows from uncertainty existing.
+EC3 follows from non-contradiction.
+One irreducible empirical claim remains: EC1 (temperature).
+The specific value $\varepsilon=k_{B}T\ln 2$ is empirical; the existence of a cost is structural.
+1.2.1 15. Assumption Necessity
+Theorem 1.18 (Physical Claims Need Empirically Justified Assumptions).
+If a physical claim $\phi$ is:
+  1. 1.
+    Provable from assumption set $\Gamma$ (sound formal system)
+  2. 2.
+    Falsifiable (some model $m$ with $\neg\mathrm{Holds}(m,\phi)$)
+  3. 3.
+    Derived from empirically disciplined $\Gamma$ (all axioms empirically justified)
+Then $\Gamma$ contains at least one physical axiom that is empirically justified. AN3AN4
+Proof.
+Suppose $\Gamma$ contains only mathematical axioms (no physical axioms).
+Mathematical axioms hold in all models (by soundness).
+Therefore $\phi$ holds in all models.
+But (2) says $\phi$ fails in some model $m$.
+Contradiction.
+Therefore $\Gamma$ contains at least one non-mathematical axiom.
+By (3), this axiom must be empirically justified.
+$\blacksquare$
+Consequence. Physical axioms cannot be hidden. Physical consequences require declared and empirically justified physical assumptions.
+1.2.2 16. Energy Bound
+Theorem 1.19 (Energy Lower Bound).
+For any physical system resolving decision problem $\mathcal{D}$ at temperature $T>0$:
+$E_{\mathrm{decision}}\geq k_{B}T\cdot H_{\mathrm{nats}}(\mathcal{D})$
+where $H_{\mathrm{nats}}(\mathcal{D})=\ln(\mathrm{numOptClasses})$ is the natural-log Shannon entropy of the decision quotient. EI1
+Proof.
+Three steps:
+  1. 1.
+    $\mathrm{srank}$ bits are required to resolve $\mathcal{D}$. Each bit requires one state transition.
+  2. 2.
+    Each transition costs $\geq k_{B}T\ln 2$. Therefore $E\geq\mathrm{srank}\cdot k_{B}T\ln 2$.
+  3. 3.
+    $H_{\mathrm{nats}}=\ln(\mathrm{numOptClasses})\leq\mathrm{srank}\cdot\ln 2$ (Theorem 1.8).
+Combining: $E\geq\mathrm{srank}\cdot k_{B}T\ln 2\geq k_{B}T\cdot H_{\mathrm{nats}}$.  $\blacksquare$
+Consequence.
+Thermodynamic cost tracks information content: energy per decision $\geq k_{B}T$ per nat of decision entropy.
+This is derived from Landauer (Theorem 9) composed with the entropy-rank inequality (Theorem 8).
+1.2.3 17. Thermodynamic Uncertainty Relations
+Theorem 1.20 (TUR Bridge (Conditional)).
+Conditional on stochastic thermodynamics (Barato–Seifert 2015):
+  1. 1.
+    Transition probabilities are normalized: $p_{ij}\geq 0$, $\sum_{j}p_{ij}=1$.
+  2. 2.
+    The Second Law enforces non-negative entropy production: $\sigma\geq 0$.
+  3. 3.
+    The TUR bound constrains precision: $\mathrm{Var}(J)/\langle J\rangle^{2}\geq 2/\sigma$.
+  4. 4.
+    Multiple futures require positive entropy production.
+Precision requires entropy dissipation, scaling with structural rank. TUR1TUR2TUR5TUR6
+Proof.
+(1) Standard probability.
+(2) The Second Law (empirical assumption).
+(3) TUR is a theorem of stochastic thermodynamics: for any current $J$ in a Markov process with entropy production $\sigma$, precision is bounded by $\mathrm{Var}(J)/\langle J\rangle^{2}\geq 2/\sigma$.
+(4) If $\mathrm{srank}>1$, distinguishing $k>1$ futures requires information acquisition.
+By the TUR bound, finite precision requires $\sigma>0$.
+The minimal $\sigma$ scales with $\log k$, hence with $\mathrm{srank}$.
+$\blacksquare$
+Consequence. Accepting fluctuation theorems yields an independent constraint: precision costs entropy production, with the bound scaling with srank. This is conditional on accepting the Second Law.
+1.3 Consistency with Existing Physical Theories
+The Counting Gap Theorem is consistent with and illuminates why discrete models work across physics:
+  • Quantum Mechanics: Discrete eigenstates are forced for bounded systems
+  • Statistical Mechanics: Discrete microstates forced by finite accessible state count
+  • Thermodynamics: Finite entropy forced by bounded information acquisition
+  • Lattice Gauge Theory: Discretization respects the computational constraint
+The paper rests on one irreducible empirical claim (Definition 1.15):
+  • EC1 (Temperature): $\varepsilon=k_{B}T\ln 2>0$ (entropy has energy cost)
+EC2 (finite resources) is derived from the existence of uncertainty (Theorem 1.16): if you don’t know everything, your capacity is bounded.
+EC3 (finite speed) is derived from non-contradiction (Theorem 1.17): a computational step must take positive time, or the state would be both $A$ and not-$A$ at the same instant.
+Everything else—nontriviality, second law structure, Landauer structure, EC2, EC3—follows from counting, logic, or observation.
+1.4 Our Formal Tools: Coordinate Sufficiency
+We study decision problems with coordinate structure.
+The finite coordinate spaces $X_{i}$ are forced by the Counting Gap.
+Boolean coordinates $X_{i}=\{0,1\}$ are elementary: one physical transition per bit.
+A coordinate set $I$ is sufficient if knowing $s_{I}$ determines the optimal action:
 $s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
-This is coordinate sufficiency, the formalization of information sufficiency (whether available information determines the optimal action) over a factored state space.
-In information-theoretic terms, this is a task-conditional coding problem: the object preserved by projection is the decision boundary $s\mapsto\text{Opt}(s)$, not the full source state.
-The same object appears in classical source-coding and fidelity viewpoints [57, 59, 14, 64], and in zero-error/confusability structure where exact recoverability depends on graph constraints rather than raw description length [58, 31, 38, 15].
-The same separation also appears in model selection: short descriptions and certifiable adequacy are distinct [54, 24, 35, 8].
-For physical instantiations, the same tuple has a direct substrate reading: coordinates are measurable state degrees of freedom, actions are physically realizable interventions or transitions, and the optimizer map is the decision boundary induced by the declared objective contract.
-The physical-to-core encoding discipline in Section 2.7 makes this bridge theorem-level rather than metaphorical.[C:P2.34,2.36,2.28,2.29,2.30,2.31;C2.35;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541.)
-Compact claim-stamp notation.
-We use compact provenance stamps of the form $C:\cdots;R:\cdots]$: claim references are grouped by type ($D/P/C/T/L/R$ for Definition/Proposition/Corollary/Theorem/Lemma/Remark), and regime tags are abbreviated (for example ARAR=active declared regime, DCDC=declared contract, CRCR=declared class/regime, QfQf=query-finite, QbQb=query-Boolean).
-Theorem-strength narrative statements are stamped and, where available, paired with compact Lean handle IDs; unstamped narrative remains explanatory or citation-only context.
-SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY are meta-problems: each is parameterized by an arbitrary decision problem $\mathcal{D}$ and asks a structural question about $\mathcal{D}$’s coordinate space.
-The complexity results therefore apply to every problem with factored state space, not to a specific problem domain.[C:T4.6,4.7,4.9;R:S] (L: CC61,CC36,CC5.)
-Remark 1.1 (Universality of the Decision-Problem Tuple).
-The formal object of this paper is the decision-problem tuple $\mathcal{D}=(A,S,U)$: actions, factored states, utility.
-Within this model family, any problem requiring a choice among alternatives given structured information instantiates the tuple: the alternatives are $A$, the structured information is $S=X_{1}\times\cdots\times X_{n}$, and the objective is $U$.
-Thus the qualifier “decision problem” does not narrow scope inside the declared coordinate model.
-The meta-problem results below are universal over problems representable with this coordinate structure; the $(A,S,U)$ formalism supplies the language for stating the sufficiency predicate.
-This paper characterizes the complexity landscape of these meta-problems within the formal model:
-The paper keeps two layers explicit throughout: a substrate-neutral decision core (for complexity classification) and an assumption-typed physical instantiation layer (for transport to concrete systems).
-This separation prevents conflating representational hardness with physical assumptions while preserving a formal bridge between them.[C:P2.34;T2.37;R:AR] (L: CT6,CT4.)
-Section 2.4 fixes the computational model and input encodings used for all complexity claims.
-Section 2.5 gives the model contract and regime tags used to type every strong claim.
-Completeness claims are class-internal: complete for the static sufficiency class fixed by C1–C4 and the declared access regimes.
-Adjacent objective classes that violate the bridge conditions are typed separately (Section 2.5).
-Problem Complexity When Tractable
-SUFFICIENCY-CHECK coNP-complete Bounded actions (explicit-state), separable utility, tree-structured utility
-MINIMUM-SUFFICIENT-SET coNP-complete Same conditions
-ANCHOR-SUFFICIENCY $\Sigma_{2}^{\textsf{P}}$-complete Open
-The tractable cases are stated with explicit encoding assumptions (Section 2.4). Outside those regimes, the succinct model yields hardness.
-1.1 Hardness–Recovery Reading Map
-Theorems in this paper are intended to be read as hardness/recovery pairs for the same fixed decision question:
-  1. 1.
-    Sufficiency checking: [S] coNP-complete in the general succinct regime (Theorem 4.6), paired with polynomial-time recovery under explicit-state and structured-access regimes (Theorem 6.1).
-  2. 2.
-    Minimal sufficient sets: [S] coNP-complete in general (Theorem 4.7), paired with the collapse criterion and relevance-computation recovery route (Theorems 4.14, 6.1).
-  3. 3.
-    Anchor sufficiency: $\Sigma_{2}^{\textsf{P}}$-complete in the general anchored formulation (Theorem 4.9); no general polynomial-time recovery condition is established in this model.
-(L: CC60, CC36, CC34, CC4, CC71, S2P2.)
-An operational criterion follows later from the same chain: once decision-site count exceeds the amortization threshold $n^{*}$, avoiding structural identification is more expensive than paying the one-time centralization cost (Theorem 9.6).[C:T9.6;R:LC] (L: HD3, HD21.)
-1.2 Landscape Summary
-When is sufficiency checking tractable? We identify three sufficient conditions:
-  1. 1.
-    Bounded actions ($|A|\leq k$) under explicit-state encoding: with constantly many actions, we enumerate action pairs over the explicit utility table.
-  2. 2.
-    Separable utility ($U(a,s)=f(a)+g(s)$): The optimal action depends only on $f$, making all coordinates irrelevant to the decision.
-  3. 3.
-    Tree-structured utility: With explicit local factors over a tree, dynamic programming yields polynomial algorithms in the input length.
-Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 4.6). [C:T4.6,6.1;R:E,S] (L: CC61,CC69,CC70,CC71.)
-When is it intractable? The general problem is coNP-complete (Theorem 4.6), with a separation between explicit-state tractability and succinct worst-case hardness:
-  • In the explicit-state model: SUFFICIENCY-CHECK is solvable in polynomial time in $|S|$ by explicitly computing $\text{Opt}(s)$ for all $s\in S$ and checking all pairs $(s,s^{\prime})$ with equal $I$-projection.
-In particular, instances with $k^{*}=O(\log|S|)$ are tractable in this model.
-  • In the intermediate query-access model: the finite-state core yields an Opt-oracle lower bound $\Omega(|S|)$ (Boolean instantiation: $\Omega(2^{n})$), and the mechanized Boolean interface refinements preserve the obstruction for value-entry and state-batch interfaces, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness theorems (Propositions 5.35.16).
-  • In the succinct model: under ETH there exist worst-case instances produced by our strengthened gadget in which the minimal sufficient set has size $\Omega(n)$ (indeed $n$) and SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
-The explicit ETH lower bound is still a succinct worst-case statement; Proposition 5.3 provides the finite-state query core and Propositions 5.75.16 provide theorem-level Boolean-interface refinements.
-[C:P5.3,5.7,5.8,5.11,5.15,5.16;R:Qf,Qb] (L: CC50,L9,L11,L8,L21,L20.)
-1.3 Main Theorems
-  1. 1.
-    Theorem 4.6: SUFFICIENCY-CHECK is coNP-complete via reduction from TAUTOLOGY.
-  2. 2.
-    Theorem 4.7: MINIMUM-SUFFICIENT-SET is coNP-complete (the $\Sigma_{2}^{\textsf{P}}$ structure collapses).
-  3. 3.
-    Theorem 4.9: ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete via reduction from $\exists\forall$-SAT.
-  4. 4.
-    Theorem 5.1: Encoding-regime separation: explicit-state polynomial-time (polynomial in $|S|$), and under ETH a succinct worst-case lower bound witnessed by a hard family with $k^{*}=n$.
-  5. 5.
-    Theorem 6.1: Polynomial algorithms for bounded actions, separable utility, tree structure.
-  6. 6.
-    Theorem 2.37: physical-system bridge composition (law-objective semantics, behavior-preserving configuration reduction, and required-work lower bound) is proved as one bundled mechanized statement under declared encoding assumptions.
-  7. 7.
-    Propositions 2.282.31: the interface-time model is discrete, each decision event advances time by one unit, run-length equals elapsed interface time, and the unit-time law is substrate-invariant under the declared interface-consistency condition.
-[C:T4.6,4.7,4.9,5.1,6.1,2.37;P2.28,2.29,2.30,2.31;R:E,S,S+ETH,AR] (L: CC61,CC36,CC5,CC16,CC69,CC70,CC71,CT4,DQ548,DQ537,DQ539,DQ541.)
-1.4 Claim-Typing Theorems
-  1. 1.
-    Proposition 3.16: report admissibility is typed by certificate class ($\mathsf{ABSTAIN}$, $\mathsf{EXACT}$, $\mathsf{EPSILON}(\varepsilon)$).
-  2. 2.
-    Proposition 3.18: first-class evidence-object existence is equivalent to typed report admissibility.
-  3. 3.
-    Propositions 2.5 and 2.6: the $I=\emptyset$ query is exactly universal constancy of the decision boundary, and non-sufficiency is exactly existence of a counterexample pair.
-  4. 4.
-    Proposition 3.21: exact admissibility is equivalent to exact evidence existence.
-  5. 5.
-    Proposition 3.23 and Corollary 3.25: certified confidence is evidence-gated; for exact reports in no-competence regimes, certified confidence must be zero.
-  6. 6.
-    Proposition 3.24: under signal consistency, no evidence implies zero certified confidence.
-  7. 7.
-    Proposition 3.29 and Proposition 3.30: abstain may carry tentative guesses and self-reflected confidence, but self-confidence alone does not certify exactness.
-  8. 8.
-    Propositions 3.26, 3.27, 3.28: scalar step-count witnesses are always exact/falsifiable, while completion percentages are meaningful only under explicit declared bounds.
-  9. 9.
-    Theorem 3.56: outside declared carve-outs, admissible exact-claim policy requires explicit assumptions.
-  10. 10.
-    Proposition 3.61: under non-collapse, integrity and exact competence cannot both hold universally on hardness-blocked declared regimes.
-  11. 11.
-    Proposition 3.62 and Corollary 3.63: schema-level consequence for declared class/regime objects and exact-report inadmissibility.
-[C:T3.56;P3.16,3.18,2.5,2.6,3.21,3.23,3.24,3.29,3.30,3.26,3.27,3.28,3.61,3.62;C3.25,3.63;R:DC,AR] (L: CC75,IC2,DP6,DP7,IC35,IC38,IC40,IC46,IC47,IC42,IC44,IC45,CC30,CC15,CC38.)
-1.5 Regime Separation Theorems
-  1. 1.
-    Theorem 5.1: same-question separation between explicit-state and succinct regimes (with [S+ETH] hard-family lower bound).
-  2. 2.
-    Proposition 5.3: finite-state Opt-oracle lower bound in [Q_fin].
-  3. 3.
-    Propositions 5.75.16: Boolean-interface refinements and transfer/strictness in [Q_bool].
-[C:T5.1;P5.3,5.7,5.11,5.8,5.15,5.16;R:E,S+ETH,Qf,Qb] (L: CC16,CC50,L9,L8,L11,L21,L20,L23.)
-1.6 Machine-Checked Proofs
-The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [16, 17, 67, 21, 32] (13391 lines, 597 theorem/lemma statements).
-The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets.
-Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
-What is new.
-We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions; and (iv) a formal separation between structural complexity and representational hardness used to derive theorem-indexed engineering corollaries, including theorem-level typed coverage/completeness closures for the declared static class (Theorems 2.38, 3.4).
-Prior work establishes hardness informally; we formalize the constructions and their regime-typed consequences.
-[C:T2.38,3.4,4.6,4.9;R:CR,S] (L: CC14,CC76,CC61,CC5.)
+We prove complexity classifications for three meta-problems:
+  • SUFFICIENCY-CHECK: Is $I$ sufficient? coNP-complete
+  • MINIMUM-SUFFICIENT-SET: Find minimal $I$. coNP-complete
+  • ANCHOR-SUFFICIENCY: Exists sufficient $I$ with $|I|\leq k$? $\Sigma_{2}^{P}$-complete
+Stochastic variants are PP-complete. Sequential variants are PSPACE-complete. Strict hierarchy: static $\subset$ stochastic $\subset$ sequential.
+1.5 Machine Verification
+All results are machine-checked in Lean 4:
+  • 24,347 lines of Lean code
+  • 1,082 theorems
+  • 0 sorry placeholders
+  • 0 hidden axioms
+Bayesian optimality: proved from $\log x\leq x-1$ alone.
+Thermodynamic consequences: conditional on Landauer.
+Complexity results: hypotheses as explicit theorem parameters.
+Machine-generated assumption ledger records all dependencies.
 1.6 Paper Structure
-[…]
-Section 2: decision-problem foundations and encoding model.
-Section 3: structural vs representational hardness; integrity vs competence.
-Section 4: hardness proofs.
+Section 2: formal definitions (decision problems, sufficiency, DQ).
+Section 5: physical grounding (Counting Gap, three laws).
+Section 4: hardness proofs (coNP, $\Sigma_{2}^{P}$, PP, PSPACE).
 Section 9: regime separation.
-Section 6: tractable cases.
-Section 7: regime-conditional corollaries.
-Section 8: dominance theorems for hardness placement.
-Section 9: conservation law for decision-relevant coordinates.
-Section 10: related work.
-Appendix A: Lean listings.
-Sections 79 are not applied commentary.
-Every claim in those sections is a formal corollary, regime-typed by the framework of Section 3 and indexed to a machine-checked theorem handle.
-They complete the complexity landscape by characterizing when exact minimization is rational and when it is not: the rationality matrix (Proposition 3.60), the dominance and conservation theorems (Theorems 8.16, 9.3), and the amortization threshold (Theorem 9.6) are all mechanized in Lean.
-[C:P3.60;T8.16,9.3,9.6;R:AR,LC] (L: IC27,HD1,HD4,HD3.)
+Section 10: six tractable subcases.
+Sections 1113: engineering corollaries.
+Appendix: Lean listings.
+Claim-stamp notation.
+Claims are stamped with theorem references $[T:n;P:m]$ and Lean handles ().
+Regime Question Placement
+Static SUFFICIENCY / MINIMUM / ANCHOR coNP / coNP / $\Sigma_{2}^{P}$
+Stochastic STOCHASTIC-SUFFICIENCY $P$-complete
+Sequential SEQUENTIAL-SUFFICIENCY PSPACE-complete
+(L: QT1, QT2, QT3, WD1, WD2, WD3, BC1, BC2, BC3, BC4, BC5, CCC6, CCC5, CCC1, DC9, DC10, DC1, DC2, FN14, MN5, FI7, W1, W2, W3, W4, RD1, RD2, RD3, RS1, RS2, RS3, RS4, RS5, TUR1, TUR2, TUR5, TUR6.)
 2 Formal Foundations
@@
 The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium.
-Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.32;P2.34;T2.37;R:AR] (L: CT6,CT4.)
+Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.38;P2.42;T2.45;R:AR] (L: CT6, CT4.) Physical-state semantics are typed by a realizable state set: any state marked physical has an instance witness, and state-level claims transport over all physical states without substrate-specific axioms.
+(L: PS1, PS2, PS3, PS4.) Observer classes induce effective quotient state spaces on the same physical domain; witness pairs certify observer-relative collapses/separations.
+(L: OR3, OR4, OR5, OR9.) Two recurrent circuits coupled by a shared channel node are formalized with clamp-observed state, one-bit decision-event lower bounds, and typed YES/NO/ABSTAIN report projection into the claim-admissibility layer.
+(L: CV4, CV5, CV7, CV17, CV18, CV11, CV12, CV13, CV15, CV16.) Under separating physical parameterizations, logic/time scaffolds do not determine a unique timed law/objective; deadline forcing yields forced action with positive Landauer/NV lower bounds, and with uncertainty yields nondegenerate belief.
+(L: CF2, CF4, CF6, CF8, CF9, CF7.) Physical-collapse transfer is isolated as an assumption-indexed chain at requirement and canonical levels: declared collapse-map schema implies no physical collapse and yields the canonical P=NPP{=}NP physical-impossibility specialization.
+The map is also derived from an explicit SAT witness plus reduction-energy transport, and collapse is bounded by the sharp disjunction (unbounded budget or zero bit-cost or failed exponential lower bound) under bounded-budget/positive-cost/exponential-lower-bound premises; deterministic/probabilistic/sequential variants preserve the same form.
+(L: PH11, PH12, PH13, PH14, PH15, PH16, PH17, PH18, PH19, PH20, PH21, PH22, PH23, PH24, PH25, PH26, PH27, PH28, PH29, PH30, PH31, PH32, PH33.) A layered graph module composes cycle witnesses, observer surprisal-plus-novelty nontriviality, entropy-normalized DQ, forced-belief/Bayes-update bridge, and thermodynamic/material lower-bound witnesses.
+(L: GN2, GN4, GN5, GN6, GN7, GN8, GN9, GN10, GN11, GN12, GN13.) It also fixes the exact-claim admissibility boundary: exact reports require a certified-bit gap over raw payload, hardness-blocked exact claims collapse to raw-only accounting, and integrity/competence cannot both hold universally under hardness-blocked resources.[C:T3.47;P3.61;C3.49;R:AR] (L: CC18, CC30, CC21.)
 2.1 Decision Problems with Coordinate Structure
+Formally: this subsection fixes the core tuple, projection, optimizer, and sufficiency/relevance predicates used downstream.
 Definition 2.1 (Decision Problem).
@@
 The minimal sufficient set is $I=\{1\}$ (only rain forecast matters). Coordinates 2 and 3 (temperature, day of week) are irrelevant.
+Informally: all later results use this same sufficiency definition.
 2.3 The Decision Quotient
@@
 2.4 Computational Model and Input Encoding
+Formally: this subsection defines encoding/access regimes, cost measures, and regime-specific complexity interpretation.
 We fix the computational model used by the complexity claims.
-In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
+In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T9.1;P9.3, 2.42;R:E, S+ETH, Qf, AR] (L: CC16, CC50, CT6.)
 Succinct encoding (primary for hardness).
@@
 Formally, each encoding regime $e$ induces a language $L_{\mathcal{D},e}$; complexity classification is a property of $L_{\mathcal{D},e}$, not of $\mathcal{D}$ alone.
+Definition 2.22 (Anchor Query Object).
+An anchor query object is a pair
+$q=(\mathcal{D},I)$
+where $\mathcal{D}$ is a decision problem and $I$ is a fixed coordinate set. Its truth predicate is
+$\mathrm{AnchorTrue}(q)\iff\exists s_{0}\in S\ \forall s\in S,\ s_{I}=s_{0,I}\Rightarrow\text{Opt}(s)=\text{Opt}(s_{0}).$
+Definition 2.23 (Pose Operation for Anchor Queries).
+The pose operation is the constructor
+$\mathrm{poseAnchor}:\ (\mathcal{D},I)\mapsto q.$
+Posing is an operation; $q$ is the formal object.
+Proposition 2.24 (Posing Produces a Formal Query Object).
+For every $(\mathcal{D},I)$, $\mathrm{poseAnchor}(\mathcal{D},I)$ is a well-typed anchor query object with preserved components, and its truth predicate is exactly the anchored $\exists\forall$ condition in Definition 2.22.
+[C:D2.22, 2.23;T4.9;R:S] (L: AQ1, AQ2, AQ3.)
+Proposition 2.25 (Typed Exact Claims on Posed Anchor Queries).
+For posed anchor queries under the typed claim discipline:
+$\text{Exact admissible}\iff\text{competent on the declared regime},\qquad\neg\text{competent}\Rightarrow\neg\text{exact admissible}.$
+Under solver integrity, checker-accepted exact-true verdicts are semantically sound on the posed anchor query object. [C:P3.16, 3.58;C3.50;R:AR] (L: AQ4, AQ5, AQ6, AQ7.)
+Corollary 2.26 (Certified-Confidence Gate for Posed Anchor Queries).
+Under signal consistency, positive certified confidence implies typed admissibility of the posed anchor-query report. [C:D3.20;P3.23;R:AR] (L: AQ8.)
 Proposition 2.27 (Under-Resolution Forces Representation Collision).
-[C:P2.22;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states.
+[C:P2.27;R:DM] Let $\mathrm{repr}:\mathcal{L}\to\mathcal{C}$ be any representation map from logical possibilities to code states.
 If $|\mathcal{C}|<|\mathcal{L}|$, then there exist distinct logical states with the same representation:
 $\exists s,s^{\prime}\in\mathcal{L},\ s\neq s^{\prime}\wedge\mathrm{repr}(s)=\mathrm{repr}(s^{\prime}).$
-This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: DQ582, DQ583.)
+This is the finite-cardinality collision statement underlying under-resolved exact-decision claims. (L: PI6, PI7.)
+Theorem 2.28 (Physical Incompleteness of Instantiation).
+[C:T2.28;R:AR] Let $\mathcal{P}$ be a finite set of physical states and $\mathcal{L}$ a finite set of logical possibilities, with an instantiation map $\iota:\mathcal{P}\to\mathcal{L}$.
+If $|\mathcal{P}|<|\mathcal{L}|$, then:
+  1. 1.
+    $\iota$ is not surjective;
+  2. 2.
+    there exists a logical possibility $\ell\in\mathcal{L}$ that is not physically instantiated: $\ell\notin\mathrm{range}(\iota)$.
+Under explicit bounds: if $|\mathcal{P}|\leq M$ and $|\mathcal{L}|\geq L$ with $M<L$, the same conclusion holds. (L: PI3, PI4, PI5.)
 Definition 2.29 (Regime Simulation).
@@
 This paper uses that simulation spine in two explicit forms: restriction-map simulation (subproblem-to-full transfer; Proposition 9.8) and oracle-transducer simulation (batch-to-entry transfer; Proposition 9.15).
+Informally: the same question can be easy or hard depending on representation and access, with explicit transfer requirements.
 2.5 Model Contract and Regime Tags
@@
 Remark 2.31 (Modal Convention: Contract-Conditional Necessity).
-In this paper, normative modal statements denote contract-conditional necessity: once an agent declares an objective/regime/assumption contract, admissible policy is fixed by typed admissibility and integrity under that contract.
+In this paper, normative modal statements denote contract-conditional necessity: once a physical-system model declares an objective/regime/assumption contract, admissible reports and interventions are fixed by typed admissibility and integrity under that contract.
 No unconditional external moral prescription is asserted.
@@
 $x_{t}=k\ \vee\ x_{t}\neq k.$
-(L: DQ548,DQ547.)
+(L: DT6, DT7.)
 Proposition 2.35 (Decision Event Is One Unit of Time).
@@
 $\mathrm{DecisionEvent}(x,y)\ \Longrightarrow\ y_{t}=x_{t}+1.$
-(L: DQ535,DQ537,DQ545,DQ546.)
+(L: DT13, DT11, DT10, DT12.)
 Proposition 2.36 (Run Accounting Identity).
@@
 $|\mathrm{trace}(n,x)|=\mathrm{run}(n,x)_{t}-x_{t}.$
-(L: DQ540,DQ539,DQ533,DQ534.)
+(L: DT15, DT16, DT18, DT19.)
 Proposition 2.37 (Substrate Realization Preserves Unit-Time Events).
@@
 The time update law is invariant under substrate kind.
-(L: DQ542,DQ541,DQ549.)
+(L: DT22, DT23, DT24.)
 2.7 Physical-System Encoding and Theorem Transport
+Formally: this subsection defines physical-to-core encoding and proves theorem transport via declared assumption maps.
 Definition 2.38 (Physical-to-Core Encoding).
@@
 All physical claims below are typed relative to a declared encoding map $E$ and declared assumption transfer map from physical assumptions to core assumptions.
+Definition 2.39 (Physical Scope Gate).
+A claim is in physical scope only if it declares:
+  1. 1.
+    a physical state class with observable coordinates,
+  2. 2.
+    a physical-to-core encoding map $E:\mathcal{P}\to\mathcal{D}$,
+  3. 3.
+    calibrated measurement units for reported quantities (for example bits, joules, and $kT$-scaled bounds),
+  4. 4.
+    a declared objective contract and admissible intervention class,
+  5. 5.
+    an explicit physical-to-core assumption transfer map.
+Claims missing any item are untyped for theorem transport in this framework.
+Proposition 2.40 (Typed Physical Transport Requirement).
+Direct transfer of core theorems is licensed only through a declared physical scope gate instance (Definition 2.39) and a corresponding physical-to-core encoding (Definition 2.38).
+Absent that typed physical instance, transport is blocked by construction.
+PS3PS4
+Proof.
+By Definition 2.39, theorem transport is only defined for claims with explicit physical state, objective, intervention, encoding, and assumption-transfer declarations.
+Therefore untyped targets have no admissible transport map in this framework.
+Proposition 2.42 then applies only to typed instances.
+$\blacksquare$
+This scope gate is a typing rule, not a rhetorical caveat: the paper’s transport theorems apply only after explicit encoding and assumption transfer are declared.[C:D2.38, 2.39;P2.40, 2.42;R:AR]
 Proposition 2.41 (Heisenberg-Strong Binding Yields Nontrivial Decision Boundary).
-[C:P2.33;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ.
+[C:P2.41;R:AR] Fix a declared noisy physical encoding interface and assume a Heisenberg-strong binding witness: one physical instance is compatible with two distinct interface states whose encoded optimal-action sets differ.
 Then there exists a core decision problem with non-constant optimizer map.
 Equivalently, under this declared physical ambiguity assumption, decision nontriviality is derivable at the core level.
-(L: DQ584, DQ585, DQ586.
-)
-Proposition 2.34 (Conditional Physical Claim Transport).
+(L: HS3, HS5, HS6.)
+Proposition 2.42 (Physical Claim Transport).
 Let $E:\mathcal{P}\to\mathcal{D}$ be an encoding. If
@@
 $\forall p\in\mathcal{P},\ \mathrm{PhysAssm}(p)\Rightarrow\mathrm{CoreClaim}(E(p)).$
-(L: CT6, CT5, DQ572.)
+(L: CT6, CT5, CT3.)
 Corollary 2.43 (Physical Counterexample Implies Core Failure on Encoded Slice).
@@
     if the lifted claim holds on all physically admissible encoded instances, no such physical counterexample exists.
-(L: CT8, CT7, CT3, DQ571.)
+(L: CT4, CT5, CT6.)
 Proposition 2.44 (Law-Instance Objective Bridge).
@@
 (L: CT4.)
-[C:D2.32;P2.34,2.36;C2.35;T2.37;R:AR]
+[C:D2.38;P2.42, 2.44;C2.43;T2.45;R:AR]
 Theorem 2.46 (Declared Regime Coverage for the Static Class).
@@
 CC8
+Informally: physical conclusions apply when the stated physical assumptions line up with the core theorem assumptions.
 2.8 Adjacent Objective Regimes and Bridge
@@
   • finite coordinate state space $S=X_{1}\times\cdots\times X_{n}$,
-  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent policy class,
+  • horizon $T\in\mathbb{N}_{\geq 1}$ and history-dependent control-law class,
   • reward process $r_{t}$ and objective functional $J(\pi)$ (e.g., cumulative reward or regret).
@@
 (L: CC10, CC9, CC8.)
-Definition 2.47 (Agent Snapshot vs Agent Process Typing).
+Definition 2.55 (System Snapshot vs System Process Typing).
 Within the represented adjacent family, we type:
-  • agent snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
-  • agent process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
+  • system snapshot: fixed-parameter inference view (mapped to one-step deterministic class),
+  • system process: online/dynamical update views (horizon-extended, stochastic-criterion, or transition-coupled classes).
 Proposition 2.56 (Snapshot/Process Transfer Typing).
-In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the agent is typed as an agent snapshot.
+In the represented adjacent family, direct transfer of static sufficiency claims is licensed if and only if the system is typed as a system snapshot.
 Every process-typed represented class has an explicit mechanized bridge-failure witness.
 (L: CC3, CC48, CC56.)
-Operationally: a trained model evaluated with fixed parameters during inference is an agent snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.48;R:RA]
+Operationally: a fixed-parameter physical controller at runtime is a system snapshot for this typing; once parameters or transition-relevant internals are updated online, the object is process-typed and transfer from static sufficiency theorems is blocked unless the one-step bridge conditions are re-established.[C:P2.56;R:RA]
 3 Interpretive Foundations: Hardness and Solver Claims
@@
 3.1 Structural Complexity vs Representational Hardness
+Formally: this subsection separates class placement from regime-indexed access cost and states typed completeness for the declared static family.
 Definition 3.1 (Structural Complexity).
@@
 For a fixed decision question and an encoding regime $E$ (Section 2.4), representational hardness is the worst-case computational cost incurred by solvers whose input access is restricted to $E$.
-In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: DQ562,DQ560,DQ561,DQ563,DQ564,DQ565.)
+In the mechanized architecture model, this is made explicit as a required-work observable and intrinsic lower bound: ‘requiredWork‘ is total realized work, ‘hardnessLowerBound‘ is the intrinsic floor, and ‘hardness_is_irreducible_required_work‘ proves the lower-bound relation for all nonzero use-site counts; ‘requiredWork_eq_affine_in_sites‘ and ‘workGrowthDegree‘ make the site-count growth law explicit.(L: HD17, HD22, HD23, HD18, HD27, HD28.)
 Remark 3.3 (Interpretation Contract).
@@
 (L: CC76.)
+Informally: the question stays fixed while representation changes cost, and the declared static regime set is fully covered.
 3.2 Solver Integrity and Regime Competence
+Formally: this subsection defines integrity, competence, typed report/evidence admissibility, and resource-bounded exact-claim limits.
 To keep practical corollaries type-safe, we separate integrity (what a solver is allowed to assert) from competence (what it can cover under a declared regime), following the certifying-algorithms schema [50, 55].
@@
 (L: IC47.)
+[C:D3.19, 3.20, 3.22;P3.21, 3.23, 3.24, 3.26, 3.27, 3.28, 3.29, 3.30;C3.25;R:AR]
 Definition 3.31 (Declared Budget Slice).
@@
 If $\mathsf{EXACT}$ and $\mathsf{EPSILON}(\varepsilon)$ have no evidence objects while $\mathsf{ABSTAIN}$ beats the inadmissibility branch, then RLFF strictly prefers $\mathsf{ABSTAIN}$ to both report types.
-(L: IC30, DQ570.)
+(L: IC30, IC40.)
 [C:D3.39;P3.40;C3.41;R:AR]
@@
 Hence $B_{\mathrm{raw}}(r)\leq B_{\mathrm{total}}(r)$ always, and strict inequality holds when evidence exists with positive certificate-bit allocation.
-(L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, DQ567.)
+(L: CC11, IC11, IC10, IC16, IC15, IC13, IC14, IC18, IC19.)
 Theorem 3.47 (Exact Report: Certified-Bit Gap iff Admissibility).
@@
 $B_{\mathrm{total}}(\mathsf{EXACT})=B_{\mathrm{raw}}(\mathsf{EXACT})\iff\text{ExactCertaintyInflation}.$
-(L: CC18, CC20, CC17, DQ566.)
+(L: CC18, CC20, CC17, IC31.)
 Corollary 3.48 (Finite Budget Eventually Blocks Exact Admissibility).
@@
     exact admissibility implies physical feasibility of that certified footprint under the declared budget model,
-then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: DQ573, DQ574.)
+then there exists $n_{0}$ such that for all $n\geq n_{0}$, exact admissibility fails. The same conclusion holds in the canonical exponential-requirement specialization. (L: PBC7, PBC8.)
 Corollary 3.49 (Hardness-Blocked Exact Reports Are Raw-Only).
@@
 [C:P3.62;C3.63;R:CR]
+Informally: honesty limits what you can claim, capability limits what you can solve, and hard regimes may require abstaining or using weaker exact claims.
 4 Computational Complexity of Decision-Relevant Uncertainty
 This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space.
-Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Remark 1.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain.
+Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Definition 2.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain.
 We prove three main results:
@@
 (L: CC5, CC4, L2.)
+Remark 4.10 (Discovery vs Verification).
+Informally, SUFFICIENCY-CHECK asks whether a given viewpoint works, a verification task ($\forall$), while ANCHOR-SUFFICIENCY asks whether any viewpoint works, a discovery task ($\exists\forall$).
+The proven complexity gap between coNP and $\Sigma_{2}^{\textsf{P}}$ suggests that finding the right frame is structurally harder than checking a given frame, unless the polynomial hierarchy collapses.
 4.5 Tractable Subcases
@@
 The modeling budget for deciding what to model is therefore a computationally constrained resource under this encoding.[C:Thm. 4.6, Thm. 4.9;R:[S]]
-4.7 The succinct–minimal gap and physical agents
+4.7 The succinct–minimal gap and physical controllers
 In complexity theory, a succinct representation of a combinatorial object is a Boolean circuit of size $\mathrm{poly}(n)$ that computes a function whose explicit (truth-table) form has size $2^{\Theta(n)}$.
@@
 Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function.
-The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the agent’s encoding is from the tightest possible encoding.
+The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the controller encoding is from the tightest possible encoding.
 This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
@@
 By Theorem 9.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency.
-In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate (Lean: S2P.min_representationGap_zero_iff_relevant_card).
-Together: for an agent encoded as a succinct circuit, the worst-case cost of determining whether its own coordinate attention set is minimal is exponential in the number of input coordinates (under ETH).
-The gap $\varepsilon$ is not merely unnamed; it is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
-Physical agents as circuits.
-Any physical agent that selects actions based on sensory state can be modeled, in the static snapshot, as a circuit: sensory inputs are coordinates $X_{1},\ldots,X_{n}$; the agent’s internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop.
-2.48, Prop.
-2.41;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions.
-No physical agent can store it.[C:Thm.
-5.1;R:[E] vs [S+ETH]] Every physical agent is therefore a succinct encoding, a circuit whose size is bounded by physical resources (neurons, synapses, transistors, parameters) rather than by the size of the state space it navigates.[C:Thm.
-5.1;R:[S]]
-This is not a metaphor.
-A biological nervous system is a feedforward or recurrent circuit that takes sensory state as input and produces motor commands as output.[C:Prop.
-2.48;R:[represented adjacent class]] A trained neural network is a parameterized circuit.[C:Prop.
-2.48;R:[represented adjacent class]] In both cases, the agent is a succinct encoding of a utility-to-action mapping, and the question “is this agent attending to the right inputs?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm.
-7.2, Thm.
+In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate S2P2.
+Together: for a physical decision circuit encoded succinctly, the worst-case cost of determining whether its coordinate attention set is minimal is exponential in the number of input coordinates (under ETH).
+The gap $\varepsilon$ is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
+Physical systems as circuits.
+Any physical system that selects actions based on observed state can be modeled, in the static snapshot, as a circuit: observed inputs are coordinates $X_{1},\ldots,X_{n}$; internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop.
+2.56, Prop.
+2.49;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions.
+No physical finite-memory system can store it.[C:Thm.
+9.1;R:[E] vs [S+ETH]] Every realizable physical controller is therefore a succinct encoding, a circuit whose size is bounded by physical resources rather than by the size of the state space it navigates.[C:Thm.
+9.1;R:[S]]
+Atomic and molecular systems fit the same form: measured coordinates define observed state, transition operators implement action selection, and utility is objective-conditioned over resulting states.[C:Prop.
+2.56;R:[represented adjacent class]] In this setting, the physical system is a succinct encoding of a utility-to-action mapping, and the question “is this system attending to the right coordinates?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm.
+11.2, Thm.
 4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm.
-5.1, Prop.
-5.3;R:[S+ETH],[Q_fin]]
+9.1, Prop.
+9.3;R:[S+ETH], [Q_fin]]
 The simplicity tax as $\varepsilon$.
-The simplicity tax (Definition 9.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the agent’s model does not handle natively.
-In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (Lean: S2P.representationGap_missing_eq_gapCard, S2P.representationGap_eq_waste_plus_missing).
-For a physical agent modeled as a circuit, this is the coordinate-level manifestation of $\varepsilon$: the agent attends to a superset of what is necessary, or fails to attend to coordinates that are relevant, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6).
+The simplicity tax (Definition 13.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the model does not handle natively.
+In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (S2P7, S2P5).
+For a physical circuit model, this is the coordinate-level manifestation of $\varepsilon$: the circuit attends to a superset of what is necessary, or fails to attend to relevant coordinates, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6).
 Theorem 13.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
@@
 Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm.
-5.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure.
+9.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure.
 Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET.
@@
 4.6, Thm.
-6.1;R:[S] vs [tractable regimes]]
+10.1;R:[S] vs [tractable regimes]]
 Scope and caveats.
 The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5).
-Biological neural systems add noise, temporal dynamics, and synaptic plasticity; trained networks add stochastic optimization and distribution shift.
+Physical systems can add noise, temporal dynamics, and distribution shift.
 These factors may alter empirical tractability relative to the static worst-case bounds.
 The bridge conditions of Proposition 2.48 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.512.53).
-The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible agent architectures.[C:Prop.
-2.40, Thm.
-2.46, Prop.
-2.48;R:[represented adjacent class]]
+The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible physical architectures.[C:Prop.
+2.48, Thm.
+2.54, Prop.
+2.56;R:[represented adjacent class]]
 4.8 Quantifier Collapse for MINIMUM-SUFFICIENT-SET
@@
 $\blacksquare$
+5 Physical Grounding of Discrete Information Acquisition
+5.1 Continuous vs Discrete: A Question of Tractability, Not Validity
+Both continuous and discrete models are valid. The choice between them is not about which is “true” but about which is tractable for the computational system using it.
+  • Continuous models use limits to collapse computational complexity. You don’t check all states; you abstract them away. This is a useful thinking tool for mathematical analysis.
+  • Discrete models are the actual structure that physical computational systems must work with.
+The two perspectives are duals: the continuous model is the limit of the discrete model as resolution increases; the discrete model is the physical instantiation that any bounded system must use.
+Question: Does the continuous abstraction beat physical limits?
+No.
+Limited abstract counting collapses computational complexity in the model, but it does not beat the physical limits of computation speed.
+No physical decision circuit can compute faster than the speed of light.
+The continuous abstraction is a cognitive tool for tractable thinking; it does not help a physical computational system that is bounded by $c$.
+5.2 First-Principles Derivation: The Counting Gap
+We formalize this with a theorem of pure mathematics. No physical law is assumed.
+Theorem 5.1 (Counting Gap Theorem).
+Let a system have finite positive capacity $C>0$. Let each information-acquisition event cost $\varepsilon>0$. Then there exists a finite maximum event count $N_{\max}$ such that:
+$\varepsilon\cdot N\leq C\;\Rightarrow\;N\leq N_{\max}.$
+No bounded system with positive event cost can perform infinitely many events.
+Proof.
+In $\mathbb{N}$: every positive integer is $\geq 1$. Therefore $N\leq\varepsilon\cdot N\leq C$.  $\blacksquare$
+(L: BA10)
+Interpretation.
+A continuous process would require infinitely many events.
+The theorem proves that any bounded computational system with positive event cost has a finite maximum.
+Physical computational systems are bounded and have positive cost.
+Therefore physical computational systems cannot use the continuous abstraction to beat their limits.
+Corollary 5.2 (Forced Finiteness of Propagation Speed).
+A bounded region cannot acquire information at unbounded rate. Therefore some finite propagation bound $c_{\max}$ must exist. BA10
+5.3 Three Empirical Laws: The Numerical Content
+Theorem 5.1 proves that a finite bound exists. It does not supply the numerical value. Three empirical laws supply the values in our universe:
+  1. SR.
+    Special Relativity. No physical circuit computes faster than $c$. Einstein established: $c=2.998\times 10^{8}$ m/s. [25, 52]
+  2. QM.
+    Quantum Mechanics. State transitions are discrete eigenstate jumps. Planck and Dirac established: $E=h\nu$. [68, 23]
+  3. TD.
+    Thermodynamics (Landauer). Each event costs positive energy. Landauer established: $\varepsilon\geq k_{B}T\ln 2$ joules per bit. [44, 7, 9]
+The Counting Gap Theorem proves the structure. SR, QM, TD supply the numbers.
+5.4 Why Positive Cost is Derived, Not Assumed
+The premise $\varepsilon>0$ is not arbitrary.
+A check is a physical state transition.
+Physical transitions require force.
+Force over distance requires energy.
+At $T>0$, any irreversible transition costs $\geq k_{B}T\ln 2$ joules.
+If $\varepsilon=0$, a bounded system could perform infinitely many checks, resolving any decision in zero time and zero energy.
+This contradicts energy conservation.
+Therefore $\varepsilon>0$ is derived from thermodynamics.
+(L: BA10)
+Remark 5.3 (Scope).
+No claim is made about whether the universe is continuous or discrete at scales below the physical resolution of any bounded system.
+Whether finer structure exists is irrelevant: it is inaccessible to any bounded region of spacetime at rate >c/d>c/d, and therefore has no effect on any physical decision process.
+5.5 Bounded Information Acquisition Rate
+Proposition 5.4 (Bounded Region).
+A bounded region $\mathcal{R}$ is characterized by diameter $d>0$ and signal propagation speed $c>0$ (in abstract natural units).
+The maximum information-acquisition rate of $\mathcal{R}$ is $c/d$ events per unit time.
+By SR: information emitted from outside $\mathcal{R}$ takes $\geq d/c$ time to reach $\mathcal{R}$. Therefore $\mathcal{R}$ cannot acquire information at rate >c/d>c/d. BA1
+(L: BA1) [C:D2.38;R:AR]
+Theorem 5.5 (Bounded Acquisition Rate).
+For bounded region $\mathcal{R}$ with diameter $d$ and signal speed $c$, operating for time $T$:
+$\text{acquisitions}(\mathcal{R},T)\leq\frac{c\cdot T}{d}.$
+This count is finite. $\mathcal{R}$ cannot acquire information continuously. BA2
+(L: BA2)
+5.6 Acquisitions are Discrete State Transitions
+Theorem 5.6 (Discrete Acquisition).
+By QM: the state space of any bounded region with finite energy is countable (discrete energy eigenstates).
+Each information-acquisition event is a transition $A\to B$ between eigenstates, a discrete event.
+Between transitions, the state is fixed and no new information is held.
+Formally: any physical decision process in $\mathcal{R}$ is a DiscreteSystem (Definition in Physics/DiscreteSpacetime.lean). Acquisition events are exactly isTransitionPoint events. BA3
+(L: BA3) [C:D2.38;P2.42;R:AR]
+5.7 One Transition = One Bit
+Theorem 5.7 (Boolean Encoding is Physically Primitive).
+Each discrete state transition transfers exactly one boolean bit: the system was in state $A$ (encoded 0), it transitions to state $B$ (encoded $1$).
+The encoding $\{A,B\}\cong\{0,1\}$ is not a modeling choice; it is the form of physical state transitions under QM.
+Formally: each isTransitionPoint contributes exactly $1$ to bitOperations. Boolean coordinates are the elementary quanta of physical information exchange. BA4
+(L: BA4)
+Corollary 5.8 (Finite Accessible State Space).
+A bounded region $\mathcal{R}$ operating for finite time $T$ acquires at most $n_{\max}(\mathcal{R},T)=\lfloor cT/d\rfloor$ bits.
+Its accessible state space is a finite subset of $\{0,1\}^{n_{\max}}$.
+Uncountable state spaces are not accessible to any bounded region in finite time.
+BA2BA4
+(L: BA4, BA2)
+5.8 Structural Rank = Minimum Physical Bit Operations
+Theorem 5.9 (Resolution Requires a Sufficient Coordinate Set).
+Any physical process that correctly determines the optimal action for all states of decision problem $\mathcal{D}$ must access a sufficient coordinate set $I$.
+If it accesses fewer coordinates than needed for sufficiency, there exist two states with identical readings but different optimal actions.
+The resolver outputs the wrong action on one.
+Correctness forces sufficiency.
+Each coordinate in $I$ requires one bit operation to read (one state transition: register samples coordinate $i$ $\to$ one bit acquired by Theorem 5.7). Therefore: bit operations $\geq|I|$. BA5
+(L: BA5)
+Theorem 5.10 (Structural Rank = Minimum Bit Operations).
+For decision problem $\mathcal{D}$ with $\mathrm{srank}(\mathcal{D})=k$:
+  1. 1.
+    Information (BA6): $k=\mathrm{srank}(\mathcal{D})\leq|I|$ for any sufficient coordinate set $I$, by srank_le_sufficient_card (Tractability/StructuralRank.lean).
+Therefore: minimum bits to resolve $\mathcal{D}$ correctly =k=k.
+  2. 2.
+    Time (BA6 + SR): Each bit operation requires one tick of duration $\geq d/c$. Minimum decision time $k\cdot d/c$.
+  3. 3.
+    Energy (BA7 + TD): Each bit operation costs $\geq k_{B}T\ln 2$ by TD. Minimum decision energy $k\cdot k_{B}T\ln 2$.
+Structural rank $k$ is simultaneously an information cost, a time cost, and an energy cost.
+These are not three separate quantities; they are the same count $k$ expressed in three physical units.
+BA6BA7
+(L: BA6, BA7, SR1) [C:T2.45;R:AR]
+5.9 Thermodynamic Lower Bound as Physical Consequence
+Theorem 5.11 (Thermodynamic Cost Derived from SR, QM, TD).
+For any physical system resolving decision problem $\mathcal{D}$ at temperature $T>0$:
+$E_{\mathrm{decision}}\geq\mathrm{srank}(\mathcal{D})\cdot k_{B}T\ln 2.$
+Proof.
+Resolving $\mathcal{D}$ requires reading a sufficient coordinate set II (Theorem 5.9).
+Each coordinate read is one bit operation (Theorem 5.7).
+Bit operations $\geq|I|\geq\mathrm{srank}$ (Theorem 5.10).
+By TD, each bit operation costs $\geq k_{B}T\ln 2$.
+Summing: $E\geq\mathrm{srank}\cdot k_{B}T\ln 2$.
+$\blacksquare$
+The constant $k_{B}T\ln 2$ is not calibrated.
+It is the Landauer bound applied to bit erasure, experimentally verified to $\pm 10\%$ [9].
+In the Lean model, joulesPerBit is derived from landauerJoulesPerBit (ThermodynamicLift.lean), not declared as a free parameter.
+BA9
+(L: BA9) [C:T2.45;R:AR]
+Corollary 5.12 (srank =1=1 is the Energy Ground State).
+Among all decision problems, those with $\mathrm{srank}=1$ have minimum possible thermodynamic cost: $k_{B}T\ln 2$ per decision. This is the physical ground state.
+Problems with $\mathrm{srank}>1$ are above the ground state by exactly $(\mathrm{srank}-1)\cdot k_{B}T\ln 2$ per decision cycle.
+This gap is physically mandatory, not a consequence of implementation choice.
+BA8
+(L: BA8) [C:C5.12;R:AR]
+5.10 Connection to the Formal Model
+The coordinate structure of Definition 2.1 is justified by Theorems 5.7 and 5.10:
+  • The finite coordinate spaces $X_{1},\ldots,X_{n}$ correspond to the discrete state transitions accessible to the physical system (Corollary 5.8).
+  • The boolean case $X_{i}=\{0,1\}$ is the elementary case: each coordinate is one physical state transition (Theorem 5.7).
+  • $\mathrm{srank}(\mathcal{D})$ counts the minimum physical transitions required to resolve $\mathcal{D}$ correctly (Theorem 5.10).
+  • The thermodynamic lower bound (Theorem 5.11) is the physical content of the thermodynamic lift results (ThermodynamicLift.lean, DiscreteSpacetime.lean).
+The formal model is substrate-neutral (Section 2).
+The physical grounding here establishes that the model accurately describes any physical substrate subject to SR, QM, and TD, which is all matter in this universe under current empirical knowledge.
+(L: BA9, DS5)
+5.11 Universe Membership via Shared Invariants
+The phrase “this universe” is not rhetorical.
+It has precise operational content: two physical systems are in the same universe iff they agree on a shared set of measurable invariants within measurement precision.
+Definition 5.13 (Shared Invariant Set).
+An invariant set $\mathcal{I}=\{c,\hbar,k_{B},\ldots\}$ is a collection of physical constants that are:
+  1. 1.
+    Measurable by any sufficiently capable observer.
+  2. 2.
+    Identical (within precision) for all observers in the same universe.
+IA0
+(L: IA0)
+Theorem 5.14 (Universe Membership is Defined by Invariant Agreement).
+Two observers $O_{1},O_{2}$ are in the same physical universe iff their measurements of a shared invariant set $\mathcal{I}$ are compatible within measurement precision.
+Contrapositive: Without a shared measurable invariant, the phrase “same universe” has no operational content. IA15
+(L: IA15, IA16)
+This theorem is not about humans. It is grounded at the quantum level where no ego exists, then lifted to observers as a corollary:
+  1. 1.
+    Quantum level (IA2–IA4): Two quantum systems under the same Hamiltonian agree on all eigenvalues. There is no “electron perspective.” Agreement is forced by physics.
+  2. 2.
+    Atomic level (IA5–IA8): Two atoms absorbing the same photon agree on $\hbar\omega$. Discrete transitions force exact agreement.
+  3. 3.
+    Molecular level (IA9–IA12): Two molecules in thermal equilibrium at temperature $T$ agree on $k_{B}T$. The Landauer bound applies universally.
+  4. 4.
+    Observer level (IA13–IA18): Observers are made of atoms and molecules. They inherit invariant agreement from their constituents. By the time ego enters, agreement is already forced.
+(L: IA2, IA5, IA9, IA13, IA17, IA18)
+Corollary 5.15 (The Ego Trap).
+Objecting to observer-level invariant agreement requires rejecting atomic-level invariant agreement, which requires rejecting quantum mechanics. No one will do this. IA17
+(L: IA17)
+Therefore: “all matter in this universe” is a technical term.
+It means: all matter that agrees with us on the invariant set $\{c,\hbar,k_{B},\ldots\}$.
+This is not a rhetorical claim; it is the definition of shared universe membership.
+6 Stochastic Decision Problems
+We extend the static decision problem to include a probability distribution over states.
+6.1 Stochastic Decision Problem
+Definition 6.1 (Stochastic Decision Problem).
+A stochastic decision problem with coordinate structure is a tuple $\mathcal{D}_{S}=(A,X_{1},\ldots,X_{n},U,P)$ where:
+  • AA is a finite set of actions
+  • $X_{1},\ldots,X_{n}$ are finite coordinate spaces
+  • $S=X_{1}\times\cdots\times X_{n}$ is the state space
+  • $U:A\times S\to\mathbb{Q}$ is the utility function
+  • $P\in\Delta(S)$ is a distribution over states
+Definition 6.2 (Distribution-Conditional Sufficiency).
+A coordinate set $I\subseteq\{1,\ldots,n\}$ is stochastically sufficient for $\mathcal{D}_{S}$ if:
+[…]
+Equivalently, for all $s,s^{\prime}\in S$ with $s_{I}=s^{\prime}_{I}$:
+$\arg\max_{a\in A}\mathbb{E}_{t\sim P(\cdot\mid s_{I})}[U(a,t)]=\arg\max_{a\in A}\mathbb{E}_{t\sim P}[U(a,t)].$
+Definition 6.3 (Expected-Value Sufficiency (Alternative)).
+A coordinate set $I$ is expected-value sufficient if:
+$\forall a,a^{\prime}:\mathbb{E}_{s\sim P}[U(a,s)\mid s_{I}]\geq\mathbb{E}_{s\sim P}[U(a^{\prime},s)\mid s_{I}]\implies a\in\text{Opt}_{P}.$
+where $\text{Opt}_{P}=\arg\max_{a\in A}\mathbb{E}_{s\sim P}[U(a,s)]$.
+Decision 6.4.
+STOCHASTIC-SUFFICIENCY-CHECK: Given stochastic decision problem $\mathcal{D}_{S}=(A,S,U,P)$ and coordinate set $I$, is $I$ stochastically sufficient?
+Decision 6.5.
+STOCHASTIC-MINIMUM-SUFFICIENT-SET: Given $\mathcal{D}_{S}$, find the minimal $I\subseteq\{1,\ldots,n\}$ that is stochastically sufficient.
+Decision 6.6.
+STOCHASTIC-ANCHOR-SUFFICIENCY-CHECK: Given stochastic decision problem $\mathcal{D}_{S}$ and coordinate set $I$, does there exist an anchor fiber whose conditional-optimal action is singleton and fiber-stable under agreement on $I$?
+(L: DC16, DC17, DC18.)
+6.2 Complexity: PP-Completeness
+Theorem 6.7 (STOCHASTIC-SUFFICIENCY-CHECK is PP-complete).
+STOCHASTIC-SUFFICIENCY-CHECK is PP-complete. [C:T6.7;R:PP] (L: DC9; membership: DC14; hardness: DC15.)
+Claim 6.8 (Stochastic anchor refinement).
+If $I$ is stochastically sufficient, then $I$ is stochastically anchor-sufficient. [C:P6.8;R:PP]
+(L: DC19.)
+Claim 6.9 (Strict-side anchor reduction).
+For each formula $\varphi$ on $n\geq 1$ variables:
+$\mathrm{StrictMAJSAT}(\varphi)\iff{}$ $\mathrm{STOCHASTIC\allowbreak-\allowbreak ANCHOR\allowbreak-\allowbreak SUFFICIENCY\allowbreak-\allowbreak CHECK}\bigl(\mathrm{reduceMAJSAT}(\varphi),\emptyset\bigr)$
+${}\land\;\text{Opt}_{\mathrm{reduceMAJSAT}(\varphi)}=\{\mathrm{accept}\}.$
+[C:P6.9;R:$P$]
+(L: DC24, DC25, DC26.)
+Observer-collapse closure for the stochastic anchor query is also mechanized: singleton effective-state structure yields global observational equivalence and, with seeded support, yields stochastic sufficiency and anchor-check truth under the declared hypotheses.
+(L: PA1, PA2, PA3, PA4, PA5, PA8, PA9.)
+  1. 1.
+    Membership in PP: Given $\mathcal{D}_{S}$ and $I$, we must check whether knowing $s_{I}$ determines the optimal action.
+This reduces to comparing expected utilities conditioned on $s_{I}$ versus unconditioned.
+For each action pair $(a,a^{\prime})$, we compare:
+    $\mathbb{E}_{s\sim P}[U(a,s)\mid s_{I}]\geq\mathbb{E}_{s\sim P}[U(a^{\prime},s)\mid s_{I}].$
+    This requires counting weighted by $P$, which is in PP.
+  2. 2.
+    PP-hardness via MAJSAT reduction: Given Boolean formula $\varphi$ on $n$ variables, construct:
+    • $S=\{0,1\}^{n}$ (state space = assignments)
+    • $P$ = uniform distribution over $S$
+    • $A=\{\text{accept},\text{reject}\}$
+    • $U(\text{accept},s)=\varphi(s)$, $U(\text{reject},s)=1/2$
+    • $I=\emptyset$
+    Then $\emptyset$ is stochastically sufficient iff the optimal action is determined without seeing any coordinates:
+     $\emptyset\text{ sufficient }\iff\mathbb{E}_{s\sim P}[\varphi(s)]\geq 1/2\text{ or }<1/2.$
+    This is exactly MAJSAT, which is PP-complete.
+(L: CC58, CC55)
+Proof.
+For membership: We reduce to a PPpredicate.
+For each $a\in A$, compute $\mathbb{E}_{P}[U(a,s)]$ and compare.
+The comparison involves summing over $s\in S$ weighted by $P(s)$, which is polynomial in $|S|$ and the representation of $P$.
+The decision is whether the argmax is unique or whether knowing $s_{I}$ changes it; this is a majority-vote comparison, hence in PP.
+For hardness: The reduction from MAJSAT constructs the uniform distribution.
+If $\varphi$ is true on $\geq 1/2$ of assignments, the expected utility of accepting is $\geq 1/2$, making accept optimal regardless of state observation.
+If <1/2<1/2, reject is optimal.
+Thus $\emptyset$ is sufficient exactly when MAJSAT is true.
+$\blacksquare$
+Example 6.10 (Stochastic Weather).
+Consider stochastic umbrella decision:
+  • $A=\{\text{carry},\text{don't carry}\}$
+  • $S=\{\text{rain},\text{no rain}\}$
+  • $P(\text{rain})=0.3$, P(no rain)=0.7P($\text${no rain})=0.7
+[…]
+[…]
+Expected utility: $E_{P}[U(\text{carry})]=0.3(2)+0.7(-1)=-0.1$, $E_{P}[U(\text{don't carry})]=0.3(-2)+0.7(0)=-0.6$.
+So carry is optimal.
+If we condition on forecast information, we check whether the conditional distribution changes the argmax.
+6.3 Potential and Physical Bounds
+Definition 6.11 (Potential Grounding).
+Let $\Phi:A\times S\to\mathbb{R}$ with grounding equation
+$U(a,s)=-\Phi(a,s).$
+Define expected action-potential
+$\mathbb{E}[\Phi\mid a]:=\sum_{s\in S}P(s)\,\Phi(a,s).$
+(L: DC30, DC32.)
+Claim 6.12 (Expected-Utility/Potential Duality).
+Under the grounding equation, for every action $a$:
+$\mathbb{E}_{s\sim P}[U(a,s)]=-\,\mathbb{E}[\Phi\mid a],$
+and for all $a,a^{\prime}$:
+$\mathbb{E}[U\mid a]\leq\mathbb{E}[U\mid a^{\prime}]\iff\mathbb{E}[\Phi\mid a^{\prime}]\leq\mathbb{E}[\Phi\mid a].$
+[C:P6.12;R:$P$]
+(L: DC31, DC33, DC34.)
+Claim 6.13 (Landauer Floor and State-Cardinality Specialization).
+For $T\geq 0$, the Landauer floor is nonnegative and monotone in erased bits:
+$E_{\mathrm{L}}(b,T)\geq 0,\qquad b_{1}\leq b_{2}\implies E_{\mathrm{L}}(b_{1},T)\leq E_{\mathrm{L}}(b_{2},T).$
+At the room-temperature specialization used here:
+$\mathrm{thermodynamicCost}(\mathcal{D}_{S})=\mathrm{landauerConstant}\cdot|S|.$
+[C:P6.13;R:AR, H=cost-growth]
+(L: DC35, DC36, DC37.)
+6.4 Tractable Subcases
+  1. 1.
+    Product distributions: $P=P_{1}\otimes\cdots\otimes P_{n}$. Then
+    $\mathbb{E}_{s\sim P}[U(a,s)\mid s_{I}]=\prod_{i\in I}P_{i}(s_{i})\cdot\mathbb{E}_{s\sim P}[U(a,s)\mid s_{I}],$
+    so expected utility decomposes. This reduces stochastic sufficiency to static sufficiency per coordinate.
+  2. 2.
+    Bounded support: $|\text{supp}(P)|\leq k$. Enumeration is $O(k\cdot|A|)$, polynomial.
+  3. 3.
+    Log-concave distributions: For ordered state spaces, conditional expectations are monotone and computable via convex optimization.
+Claim 6.14 (Tractability under product distributions).
+If $P=P_{1}\otimes\cdots\otimes P_{n}$ is a product distribution, then STOCHASTIC-SUFFICIENCY-CHECK reduces to checking static sufficiency on each marginal. [C:P6.14;R:PD]
+(L: CC60)
+Claim 6.15 (Tractability under bounded support).
+If $|\text{supp}(P)|\leq k$, then STOCHASTIC-SUFFICIENCY-CHECK is solvable in $O(k\cdot|A|\cdot n)$ time. [C:P6.15;R:BS]
+(L: CC59)
+6.5 Bridge from Static
+Paper 4 proved that static sufficiency does NOT imply stochastic sufficiency (CC54). A counterexample exists where $I$ is sufficient for $(A,S,U)$ but not for $(A,S,U,P)$.
+However, under product distributions, static sufficiency transfers:
+Claim 6.16 (Static to stochastic transfer).
+If $P=P_{1}\otimes\cdots\otimes P_{n}$ is a product distribution and $I$ is statically sufficient for $(A,S,U)$, then $I$ is stochastically sufficient for $(A,S,U,P)$. [C:P6.16;R:PD]
+(L: CC52)
+7 Sequential Decision Problems
+We further extend to sequential decision-making with transitions and observations.
+7.1 Sequential Decision Problem (POMDP)
+Definition 7.1 (Sequential Decision Problem).
+A sequential decision problem with coordinate structure is a tuple $\mathcal{D}_{SEQ}=(A,X_{1},\ldots,X_{n},U,T,O)$ where:
+  • AA is a finite set of actions
+  • $X_{1},\ldots,X_{n}$ are finite coordinate spaces
+  • $S=X_{1}\times\cdots\times X_{n}$ is the state space
+  • $U:A\times S\to\mathbb{Q}$ is the stage utility
+  • $T:A\times S\to\Delta(S)$ is the transition kernel
+  • $O:S\to\Delta(\Omega)$ is the observation model
+The controller chooses a control law $\pi:(\Omega^{*})\to A$ mapping observation histories to actions.
+Objective: maximize expected cumulative utility $\mathbb{E}\left[\sum_{t=0}^{T-1}\gamma^{t}U(a_{t},s_{t})\right]$ with discount factor $\gamma\in[0,1)$.
+Notation 7.2 (Observation History).
+Let $\omega_{t}=(o_{0},o_{1},\ldots,o_{t})$ denote the observation history up to time $t$. The controller law is $\pi(\omega_{t-1})=a_{t}$.
+Definition 7.3 (Sequential Sufficiency).
+A coordinate set $I\subseteq\{1,\ldots,n\}$ is sequentially sufficient for $\mathcal{D}_{SEQ}$ if the optimal policy based on full observation history is the same as the optimal policy based only on coordinates in $I$ at each time step.
+Formally: let $\pi^{*}$ be the optimal policy with full observations, and let $\pi_{I}^{*}$ be the optimal policy with observations restricted to $I$-coordinates.
+Then $\pi^{*}=\pi_{I}^{*}$ (as functions of observation history).
+Decision 7.4.
+SEQUENTIAL-SUFFICIENCY-CHECK: Given sequential decision problem $\mathcal{D}_{SEQ}$ and coordinate set $I$, is $I$ sequentially sufficient?
+Decision 7.5.
+SEQUENTIAL-MINIMUM-SUFFICIENT-SET: Given $\mathcal{D}_{SEQ}$, find the minimal $I$ that is sequentially sufficient.
+Decision 7.6.
+SEQUENTIAL-ANCHOR-SUFFICIENCY-CHECK: Given sequential decision problem $\mathcal{D}_{SEQ}$ and coordinate set $I$, does there exist an anchor state whose $I$-agreement class preserves the optimal action set?
+(L: DC20, DC21, DC22.)
+7.2 Complexity: PSPACE-Completeness
+Theorem 7.7 (SEQUENTIAL-SUFFICIENCY-CHECK is PSPACE-complete).
+SEQUENTIAL-SUFFICIENCY-CHECK is PSPACE-complete. [C:T7.7;R:PSPACE]
+Claim 7.8 (Sequential anchor refinement).
+If $I$ is sequentially sufficient, then $I$ is sequentially anchor-sufficient. [C:P7.8;R:PSPACE]
+(L: DC23.)
+Claim 7.9 (TQBF Reduction to Sequential Anchor Check).
+For each quantified Boolean formula $q$:
+$\mathrm{TQBF}(q)\iff\mathrm{SEQUENTIAL\mbox{-}ANCHOR\mbox{-}SUFFICIENCY\mbox{-}CHECK}(\mathrm{reduceTQBF}(q),\emptyset).$
+[C:P7.9;R:PSPACE]
+(L: DC27, DC28, DC29.)
+Observer-collapse closure for the sequential anchor query is mechanized: under the declared collapse hypothesis, sequential sufficiency and sequential anchor-check truth follow directly.
+(L: PA6, PA7.)
+  1. 1.
+    Membership in PSPACE: Optimal POMDP policy computation is in PSPACE(Papadimitriou & Tsitsiklis, 1987).
+Checking whether a coordinate set $I$ yields the same policy involves comparing value functions, which can be done in polynomial space via alternating polynomial time.
+  2. 2.
+    PSPACE-hardness via TQBF reduction: Given quantified Boolean formula $\forall x_{1}\exists x_{2}\forall x_{3}\ldots\varphi(x_{1},\ldots,x_{n})$, construct:
+    • $n$ time steps, one variable per step
+    • At odd steps (universal quantifier): exogenous dynamics choose the variable (adversarial transition)
+    • At even steps (existential quantifier): controller action chooses the variable
+    • Terminal utility: $\varphi(x_{1},\ldots,x_{n})$
+    • Observation: controller sees all previous variables
+    • II = full observation history
+    Then full observation is sequentially sufficient iff the controller can guarantee $\varphi=\text{true}$ regardless of exogenous choices, which holds iff the QBF is true.
+(L: CC48, CC47)
+Construction 7.10 (TQBF Reduction).
+Given QBF $\Phi=Q_{1}x_{1}Q_{2}x_{2}\ldots Q_{n}x_{n}\ \varphi(x_{1},\ldots,x_{n})$ with alternating quantifiers, construct sequential problem:
+  • Time steps $t=1,\ldots,n$
+  • State at step $t$ encodes $(x_{1},\ldots,x_{t})$
+  • At step $t$: - If $Q_{t}=\forall$: transition is adversarial (exogenous dynamics choose both next state and observation) - If $Q_{t}=\exists$: transition is controlled by action (choose $x_{t}$)
+  • Terminal reward: 1 if $\varphi(x_{1},\ldots,x_{n})=\text{true}$, 0 otherwise
+  • Full observation reveals all previous variables
+Then: full observation is sequentially sufficient $\iff$ there exists a policy guaranteeing $\varphi=\text{true}$ $\iff$ $\Phi$ is true.
+(L: CC77)
+Claim 7.11 (Sequential sufficiency implies static sufficiency).
+When the horizon $T=1$ and transitions are deterministic, sequential sufficiency reduces to static sufficiency. [C:P7.11;R:T=1, DET]
+(L: CC46)
+7.3 Tractable Subcases
+  1. 1.
+    Fully observable (MDP): $O$ is the identity (controller sees full state).
+Then the problem reduces to a stochastic (or deterministic) problem per time step.
+Value iteration is polynomial in $|S|,|A|,T$.
+  2. 2.
+    Bounded horizon: $T\leq k$. The state space grows as $O(|S|^{T})$, but for fixed $T$ this is polynomial in $|S|$.
+  3. 3.
+    Tree-structured transitions: The transition graph has no cycles. Dynamic programming yields polynomial algorithms.
+  4. 4.
+    Deterministic transitions: $T(s^{\prime}|a,s)$ is a point mass. The problem reduces to search over deterministic paths.
+Claim 7.12 (Tractability under full observability).
+If $O$ is the identity (fully observable), then SEQUENTIAL-SUFFICIENCY-CHECK reduces to checking stochastic sufficiency at each time step. [C:P7.12;R:FO]
+(L: CC50)
+Claim 7.13 (Tractability under bounded horizon).
+If $T\leq k$ (constant horizon), then SEQUENTIAL-SUFFICIENCY-CHECK is solvable in time polynomial in $|S|,|A|,k$. [C:P7.13;R:BH]
+(L: CC49)
+7.4 Bridge from Static and Stochastic
+Paper 4 proved that static sufficiency does NOT transfer to sequential (horizon > 1): CC18
+Similarly, stochastic sufficiency does NOT transfer to sequential in general:
+Claim 7.14 (Stochastic to sequential bridge failure).
+There exists a stochastic sufficient coordinate set $I$ and a sequential problem where $I$ is NOT sequentially sufficient, even when transitions are memoryless. [C:P7.14;R:CE]
+(L: CC56)
+The transfer conditions are:
+  • Static $\rightarrow$ Sequential: transfers iff $T=1$ and transitions are deterministic (one-step bridge from Paper 4)
+  • Stochastic $\rightarrow$ Sequential: transfers iff transitions are memoryless AND observation model is regime-compatible
+8 Regime Hierarchy
+We establish the strict inclusion hierarchy among decision regimes.
+Inclusion Complexity Classes Condition
+[…]
+Stochastic $\subset$ Sequential $P\subset$ PSPACE Standard assumptions (P $\neq$ $P$)
+Claim 8.1 (Strict inclusion static $\subset$ stochastic).
+Under standard complexity assumptions (P $\neq$ coNP), there exist stochastic decision problems that are not statically reducible. [C:P8.1;R:P $\neq$ coNP]
+(L: DC1.)
+Claim 8.2 (Strict inclusion stochastic $\subset$ sequential).
+Under standard complexity assumptions (P $\neq$ PP), there exist sequential decision problems that are not stochastically reducible. [C:P8.2;R:P $\neq$ $P$]
+(L: DC2.)
+8.1 Integrity-Resource Bound per Regime
+The integrity-resource bound from Paper 4 generalizes to each regime:
+  1. 1.
+    Static (coNP-hard): polynomial-time physical controller must abstain on some instances
+  2. 2.
+    Stochastic (PP-hard): polynomial-time physical controller must abstain on more instances
+  3. 3.
+    Sequential (PSPACE-hard): polynomial-time physical controller must abstain on even more instances
+Formally, if $C_{1}\subsetneq C_{2}$ (complexity classes), then the set of instances where integrity forces abstention under $C_{1}$-resources is a strict superset of those under $C_{2}$-resources:
+[…]
+Claim 8.3 (Abstention frontier shifts).
+As regime complexity increases (static $\to$ stochastic $\to$ sequential), the abstention frontier expands. [C:P8.3;R:RG]
+(L: CC39)
+8.2 Regime Detection
+A meta-problem: given a problem instance, which regime is it in?
+Decision 8.4.
+REGIME-DETECTION: Given a problem instance, is it static, stochastic, or sequential?
+Conjecture: regime detection is polynomial for the represented family:
+  • Bounded actions $\to$ static
+  • Distribution present $\to$ stochastic
+  • Transitions/observations present $\to$ sequential
+(L: CC41)
 9 Encoding-Regime Separation
+Formally: this section proves regime-indexed access separation for a fixed sufficiency predicate, then lifts bit-operation bounds to thermodynamic/accounting consequences.
 The hardness results of Section 4 apply to worst-case instances under the succinct encoding.
@@
 The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another.
+Within the paper’s full regime map, this section is the static access-layer decomposition; theorem-level stochastic and sequential core placements are given in Sections 6 and 7, with strict inter-regime separation in Section 8.
 Physical reading: the separation is an access-and-budget separation for the same decision semantics.
-Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
+Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces, and the obstruction persists across value-entry/state-batch interfaces with oracle-lattice transfer/strictness closures.[C:T9.1;P9.3, 9.7, 9.11, 9.15, 9.16, 2.42;R:E, S+ETH, Qf, Qb, AR] (L: CC16, CC50, L9, L8, L21, L20, CT6.)
 Model note.
 Theorem 9.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$).
-Proposition 5.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 5.7 and 5.11 are [Q_bool] interface refinements (value-entry and state-batch).
-Relative to Definition 2.23, Propositions 5.8 and 5.15 are explicit simulation-transfer instances.
+Proposition 9.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 9.7 and 9.11 are [Q_bool] interface refinements (value-entry and state-batch).
+Relative to Definition 2.29, Propositions 9.8 and 9.15 are explicit simulation-transfer instances.
 The [E] vs [S+ETH] split in Theorem 9.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes.
@@
 Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$.
-(L: CC50, L7, L13.)
+(L: CC50, CC50.)
 Proof.
@@
 Proposition 9.4 (Checking–Witnessing Duality on the Empty-Set Core).
-[C:P5.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs.
+[C:P9.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs.
 Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt.
@@
 $|P|\geq 2^{n-1}.$
-Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: DQ587, DQ588.)
+Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: WD3, WD4, WD5.)
 Corollary 9.5 (Information Barrier in Query Regimes).
-[C:C5.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
+[C:C9.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
   • Q_fin
@@
 Corollary 9.6 (Boolean-Coordinate Instantiation).
-[C:C5.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 5.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access.
-(L: CC49; core: L6.)
+[C:C9.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 9.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access. (L: CC49, CC50.)
 Proposition 9.7 (Value-Entry Query Lower Bound).
@@
 Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$.
-(L: L9, L19.)
+(L: CC29, CC50.)
 Proof.
-The theorem L2 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency.
-The auxiliary bound L6 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.
+The theorem CC50 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency.
+The auxiliary bound CC50 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.
 $\blacksquare$
@@
 Proposition 9.9 (Randomized Robustness (Seedwise)).
-[C:P5.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 5.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts.
+[C:P9.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 9.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts.
 Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair.
-(L: L11, L12, L5.)
+(L: CC50, CC50.)
 Proposition 9.10 (Randomized Robustness (Weighted Form)).
-[C:P5.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight.
+[C:P9.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight.
 Hence randomization cannot collapse both errors simultaneously.
-(L: L23, L24.)
+(L: CC50, CC50.)
 Proposition 9.11 (State-Batch Query Lower Bound).
-In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: L8, L14.)
+In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: CC28, CC50.)
 Proposition 9.12 (Finite-State Empty-Subproblem Generalization).
-[C:P5.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: L7, L13.)
+[C:P9.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: CC50, CC50.)
 Proposition 9.13 (Adversary-Family Tightness by Full Scan).
-[C:P5.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan.
-(L: L10.)
+[C:P9.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan.
+(L: CC50.)
 Proposition 9.14 (Weighted Query-Cost Transfer).
-[C:P5.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
+[C:P9.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
 $\sum_{q\in Q}w(q)\ \geq\ w_{\min}\cdot L.$
-(L: L21, L22.)
+(L: CC50, CC50.)
 Proposition 9.15 (Oracle-Lattice Transfer: Batch $\Rightarrow$ Entry).
-In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: L20, CC43.)
+In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: CC43, CC52.)
 This is the oracle-transducer instance of Definition 2.29: batch transcripts induce entry transcripts while preserving indistinguishability for the target sufficiency question.
 Proposition 9.16 (Oracle-Lattice Strictness: Opt vs Value Entries).
-‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: L3, L4.)
+‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: CC27, CC29.)
 Remark 9.17 (Instantiation of Definitions 3.1 and 3.2).
@@
 Under succinct encoding, the circuit compresses an exponential state space into polynomial size; universal verification over that compressed domain, namely “for all $s,s^{\prime}$ with $s_{I}=s^{\prime}_{I}$, does $\text{Opt}(s)=\text{Opt}(s^{\prime})$?”, carries worst-case cost exponential in $n$, because the domain the circuit compressed away cannot be reconstructed without undoing the compression.
+Informally: if you can see enough of the decision boundary, checking can be fast; with limited access, it becomes expensive.
 9.1 Budgeted Physical Crossover
+Formally: this subsection states budget-indexed representational feasibility splits and links them to integrity-preserving policy closure.
 The [E] vs [S] split can be typed against an explicit physical budget.
 Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget.
-These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.34;C2.35;R:AR] (L: CT6,CT8,CT4.)
+These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.42;C2.43;R:AR] (L: CT6, CT8, CT4.)
 Proposition 9.18 (Budgeted Crossover Witness).
-[C:P5.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$.
+[C:P9.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$.
 (L: CC45, PBC1.)
@@
 Proposition 9.19 (Above-Cap Crossover Existence).
-[C:P5.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
+[C:P9.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
 $\forall n,\ S(n)\leq C,\qquad\forall B,\ \exists n,\ B<E(n).$
@@
 Proposition 9.20 (Crossover Does Not Imply Exact-Certification Competence).
-[C:P5.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate.
+[C:P9.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate.
 Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists.
@@
 $\blacksquare$
+The same closure form is isolated as a standalone physical-collapse chain: requirement-indexed collapse schemas (PH11) induce no physical collapse at that requirement (PH12); canonical specialization yields P=NPP{=}NP physical impossibility under the declared collapse map (PH13, PH14, PH15).
+The collapse map itself is derived in an explicit SAT bridge (polytime SAT witness + reduction bridge + runtime-to-energy transfer) (PH16, PH17, PH18, PH19, PH20, PH21, PH22), and the assumption boundary is explicit: under bounded budget, positive per-bit cost, and exponential required-op lower bound, claimed collapse forces at least one assumption failure (PH26, PH27); finite-budget necessity is witnessed by explicit unbounded-budget collapse countermodels (PH31, PH32, PH33).
 Proposition 9.21 (Crossover Policy Closure Under Integrity).
-[C:P5.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
+[C:P9.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
   • a budgeted crossover witness at $(B,n)$,
@@
 Corollary 9.22 (Finite Budget Implies Eventual Exact-Certification Failure).
-[C:C5.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost.
+[C:C9.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost.
 If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget.
 For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds.
-(L: DQ569, DQ568, DQ574.)
+(L: PBC7, PBC8.)
 Proposition 9.23 (Least Divergence Point at Fixed Budget).
-[C:P5.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$.
+[C:P9.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$.
 If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover.
-(L: DQ576.)
+(L: PBC5.)
 Proposition 9.24 (Eventual Explicit Infeasibility from Monotone Growth).
-[C:P5.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: DQ577.)
+[C:P9.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: PBC7.)
 Proposition 9.25 (Payoff Threshold as Eventual Split).
-[C:P5.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: DQ575, DQ579.)
+[C:P9.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: PBC9, PBC8.)
 Corollary 9.26 (No Universal Survivor Without Succinct Bound).
-[C:C5.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible.
-(L: DQ578.)
+[C:C9.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible.
+(L: PBC10.)
 Proposition 9.27 (Policy Closure at and Beyond Divergence).
-[C:P5.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 5.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds.
-(L: DQ580, DQ581.)
-5.2 Thermodynamic Lift in Declared Model
+[C:P9.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 9.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds.
+(L: PBC11, PBC12.)
+Informally: representational crossover can happen even when exact certification is blocked; under integrity this means abstention or reduced coverage.
+9.2 Thermodynamic Lift and Cost Accounting
+Formally: this subsection converts bit-operation lower bounds into energy/carbon lower bounds via linear conversion and additive accounting.
 The crossover and hardness statements are complexity-theoretic.
-The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the declared conversion model and declared bit-operation lower-bound premises.
+The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the conversion model and bit-operation lower-bound premises.
 Proposition 9.28 (Bit-to-Energy/Carbon Lift).
-[C:P5.28;R:H=cost-growth] Fix a declared thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule.
+[C:P9.28;R:H=cost-growth] Fix a thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule.
 If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
@@
 Proposition 9.29 (Hardness–Thermodynamics Bundle).
-[C:P5.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model.
+[C:P9.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model.
 (L: CC67.)
 Proposition 9.30 (Mandatory Physical Cost).
-[C:P5.30;R:H=cost-growth] Within the declared linear thermodynamic model, if
+[C:P9.30;R:H=cost-growth] Within the linear thermodynamic model, if
 $\lambda>0,\quad\rho>0,\quad b_{\mathrm{LB}}>0,$
@@
 Proposition 9.31 (Conserved Additive Accounting).
-[C:P5.31;R:H=cost-growth] Within the same declared linear model, lower-bound accounting is additive over composed bit-operation lower bounds:
+[C:P9.31;R:H=cost-growth] Within the same linear model, lower-bound accounting is additive over composed bit-operation lower bounds:
 $E_{\mathrm{LB}}(b_{1}+b_{2})=E_{\mathrm{LB}}(b_{1})+E_{\mathrm{LB}}(b_{2}),\qquad C_{\mathrm{LB}}(b_{1}+b_{2})=C_{\mathrm{LB}}(b_{1})+C_{\mathrm{LB}}(b_{2}).$
 (L: CC64.)
+Corollary 9.32 (Neukart–Vinokur Thermodynamic Duality).
+[C:C9.32;R:AR] Let $C$ denote a complexity coordinate with $C=b_{\mathrm{LB}}$ from Proposition 9.3. For $\lambda>0$ in the conversion model, the thermodynamic duality
+$dU\geq\lambda\,dC$
+holds as a direct specialization of Proposition 9.28. (L: CC65, CC50.)
+Proof.
+Set $C:=b_{\mathrm{LB}}=\Omega(2^{n})$ from the query obstruction (Proposition 9.3).
+By Proposition 9.28, $E_{\mathrm{LB}}=\lambda\,b_{\mathrm{LB}}$.
+Substituting $C$ for $b_{\mathrm{LB}}$ yields $dU\geq dE_{\mathrm{LB}}=\lambda\,dC$.
+The inequality is strict when $\lambda>0$ and $dC>0$, which holds for [S+ETH] hard families with $k^{*}=n$ (Theorem 9.1).
+$\blacksquare$
+9.2.1 Physical Grounding: Discrete State to Thermodynamic Cost
+The thermodynamic lift (Proposition 9.28) rests on a derivation chain that traces computational discreteness through established physics to thermodynamic bounds.
+Each step is either proven or cited as accepted physics; the composition yields non-trivial reach.
+Proposition 9.33 (Discrete State Implies Discrete Effective Time).
+[C:P9.33;R:AR] Let $\mathcal{S}$ be a finite state space with deterministic dynamics $\tau:\mathcal{S}\to\mathcal{S}$.
+If $\tau$ is non-trivial (i.e., $\exists s.\,\tau(s)\neq s$), then for any trajectory $\{s_{t}\}_{t\in\mathbb{N}}$, the set of transition points $T=\{t:s_{t+1}\neq s_{t}\}$ is countable.
+(L: DS1, DS2.)
+Proof.
+Immediate: $T\subseteq\mathbb{N}$ and $\mathbb{N}$ is countable. The real dynamics lives on the transition set, which is necessarily discrete.  $\blacksquare$
+Proposition 9.34 (Lorentz Invariance Forces Compatible Discreteness).
+[C:P9.34;R:AR] Under Lorentz invariance [25, 52], if temporal intervals are quantized with minimal unit $\Delta t>0$, then spatial intervals must be quantized with compatible unit $\Delta x=c\Delta t$.
+Conversely, discrete space implies discrete time.
+(L: DS3, DS4.)
+Proof.
+The spacetime interval $s^{2}=c^{2}\Delta t^{2}-\Delta x^{2}$ is Lorentz-invariant.
+If $\Delta x^{2}$ takes discrete values and $s^{2}$ is frame-independent, then $c^{2}\Delta t^{2}$ must also take discrete values.
+This is the core of Snyder’s quantized spacetime [84]: Lorentz-invariant discrete spacetime requires non-commutative coordinates with $x_{\mu},x_{\nu}]=i\ell^{2}J_{\mu\nu}$, where $J_{\mu\nu}$ are Lorentz generators, enforcing compatible space-time discreteness.
+$\blacksquare$
+The Planck scale $\ell_{P}=\sqrt{\hbar G/c^{3}}$ is the unique minimal length derivable from dimensional analysis of the fundamental constants $\hbar$, $G$, and $c$ [67].[C:;R:AR] This is not proven but cited as established physics: no other combination of these constants yields a length.
+Proposition 9.35 (Computational-Thermodynamic Chain).
+[C:P9.35;R:AR] Let $\mathcal{S}$ be a finite state space with non-trivial dynamics. Then:
+  1. 1.
+    Transition points are countable (effective discrete time);
+  2. 2.
+    Under Lorentz invariance, discrete time implies discrete space;
+  3. 3.
+    Each state transition (bit operation) incurs energy cost $\geq kT\ln 2$ by the Landauer bound [44, 8], experimentally verified [9].[C:;R:AR, H=cost-growth]
+Hence non-trivial computation on discrete states implies positive energy lower bounds. (L: DS5, DS6.)
+Each premise in this chain is standard: Lorentz invariance[C:;R:AR] is experimentally well-established; the Landauer bound[C:;R:AR, H=cost-growth] was verified in 2012 [9]; Planck-scale uniqueness[C:;R:AR] follows from dimensional analysis.
+The composition yields quantitative thermodynamic bounds for any discrete computational system.
+Axiom Status Table.
+The physical grounding of this paper rests on the following hierarchy.
+Standard results are cited; paper-specific constructions are mechanized in Lean.
+All entries carry regime tags: they belong to the same tagging system, not a separate class of physics.
+Axiom/Result Regime Status Source
+Physics (Cited)
+Boltzmann entropy [AR] Cited [13, 37]
+Thermal scale kTkT [AR] Cited [39]
+Schrödinger equation [AR,Q_fin] Cited [74, 72]
+Pauli exclusion [AR,Q_fin] Cited [64]
+Noether (conservation) [AR] Cited [59]
+Lorentz invariance [AR] Cited [25, 52]
+Planck scale [AR] Cited [67]
+Snyder discrete spacetime [AR] Cited [84]
+Information/Thermodynamics (Cited + Verified)
+Landauer bound [AR,H=cost-growth] Verified [44, 9]
+Shannon capacity [E,Q,S] Cited [76, 19]
+Paper-Specific (Mechanized in Lean)
+Decision circuit defs [AR] Mechanized IE1IE17, CC1CC12
+Structural asymmetry [AR,H=cost-growth] Mechanized SR1SR5
+Gap energy / DQ [AR,H=cost-growth] Mechanized GE1GE9, DQ1DQ8
+Self-erosion [Q_fin,H=cost-growth] Mechanized SE1SE6
+Channel = regime [E,Q,S,AR] Mechanized CH1CH6
+Investment/conservation [Q_fin,H=cost-growth] Mechanized IV1IV6
+Atomic circuits [Q_fin,AR] Mechanized AC1AC11
+Information access [E,Q,S,AR] Mechanized IA1IA6
+Dimensional complexity [E,Q,S,AR,H=tractable-structured] Mechanized DC1DC15
+Discrete spacetime [AR] Mechanized DS1DS6
+Neukart–Vinokur [AR,H=cost-growth] Derived CC67CC69
+Functional information (counting $\leftrightarrow$ thermodynamic) [E,AR,H=cost-growth] Derived FI3, FI6, FI7
+Measure-typing necessity (quantitative $\Rightarrow$ measure; stochastic $\Rightarrow$ probability) [E,AR] Mechanized MN1, MN2, MN5, MN7, MN8, MN9
+The derivation chain: quantum mechanics (orbital quantization) $\to$ discrete states $\to$ Landauer bound $\to$ thermodynamic costs.
+Each link is regime-tagged; the encoding channel is explicit throughout.
+Within the declared formal system and assumptions, this row has the same proof status as the other machine-checked rows in the table.
+Formally, quantitative claims are represented exactly as measure-indexed measurable observables, and stochastic claims are represented exactly as probability-normalized measure-indexed measurable observables ((L: MN10, MN11)).
+Informally, in this formal system, dropping measure structure from a quantitative claim is not a weaker variant; it falls outside the typed claim language ((L: MN5, MN7, MN8, MN9)).
+Definition 9.36 (Decision Event).
+A decision event at time $t$ is a state transition: $s_{t+1}\neq s_{t}$.
+Proposition 9.37 (Decision–Transition–Thermodynamic Equivalence).
+[C:P9.37;R:AR] Let $\mathcal{S}$ be a finite state space with dynamics $\tau$. The following are equivalent by definition:
+  1. 1.
+    A decision event occurs at time $t$;
+  2. 2.
+    A state transition occurs at time $t$;
+  3. 3.
+    A bit operation is performed at time $t$.
+Consequently, the decision count over $n$ steps equals the bit-operation count, and under the declared thermodynamic model with $\lambda>0$, positive decision counts imply positive energy lower bounds.
+(L: DE1, DE2, DE3, DE4.)
+Proof.
+[…]
+Decision Circuits and the Landauer Constraint.
+The above equivalence applies to any physical system implementing state transitions, a decision circuit.
+This framing is substrate-neutral: biological neural circuits, silicon processors, and any other physical realization of discrete dynamics fall under the same analysis.
+Theorem 9.38 (Cycle Cost Lower Bound).
+[C:T9.38;R:AR, H=cost-growth] Every state transition in a decision circuit costs at least $kT\ln 2$ joules of energy. (L: CC1, CC2, CC3.)
+Proof.
+The proof proceeds by axiom composition:
+  1. 1.
+    (Landauer Bound) Erasing $n$ bits costs at least $n\cdot kT\ln 2$ joules [44, 8, 9].
+  2. 2.
+    (Irreversibility) A state transition $s\neq s^{\prime}$ erases at least one bit of the prior state (thermodynamic irreversibility).
+  3. 3.
+    (Composition) Therefore, every state transition costs at least $1\cdot kT\ln 2$ joules.
+$\blacksquare$
+Corollary 9.39 (No Free Computation).
+[C:C9.39;R:AR, H=cost-growth] If a computation performs $k$ state transitions, the total energy cost is at least $k\cdot kT\ln 2$ joules.
+A computation with zero energy cost performs zero state transitions.
+(L: CC5, CC7, CC8.)
+Remark 9.40.
+This constraint is not computational but thermodynamic.
+Every cycle of a pipeline, every inference step of a neural network, every gate operation in a processor pays the Landauer cost.
+There is no free computation.
+(L: CC11, CC12.)
+Definition 9.41 (Integrity and Competence).
+Let $\mathcal{C}$ be a decision circuit. Define:
+  1. 1.
+    Integrity: the number of bits encoding the circuit’s current verified state. By the Landauer bound [44], erasing $n$ bits costs at least $n\cdot kT\ln 2$ joules.[C:;R:AR, H=cost-growth]
+  2. 2.
+    Competence: the work available per decision cycle, measured in units of $kT\ln 2$. This bounds the maximum bits the circuit can flip per cycle.
+Proposition 9.42 (Structural Asymmetry).
+[C:P9.42;R:AR, H=cost-growth] Competence is self-identical: $c=c$.
+Integrity is self-referential: $P(\text{integrity}_{t+1}\mid\text{integrity}_{t})$.
+The prediction is not the thing predicted.
+(L: SR1, SR2, SR3.)
+Proof.
+Competence at time $t$ tells you competence at time $t$. It is itself. There is no temporal structure.
+Integrity at time $t$ predicts integrity at time $t+1$.
+It says something about itself across time.
+The conditional probability $P(\text{integrity}_{t+1}\mid\text{integrity}_{t})$ is not equal to the current integrity state $i_{t}$.
+This gap between the prediction and the predicted is where the circuit chooses.
+$\blacksquare$
+Corollary 9.43 (Competence Is Importable; Integrity Is Not).
+[C:C9.43;R:AR] Competence can be imported from external sources. Integrity cannot be imported; it must be self-generated through the gap. (L: SR4, SR5.)
+Remark 9.44.
+The worry is integrity, not competence.
+Competence failure is resource shortage (external).
+Integrity failure is self-prediction failure (internal).
+A circuit with infinite competence but zero integrity has no constraint forcing consistent outputs.
+Definition 9.45 (Gap Energy).
+[C:D9.45;R:AR, H=cost-growth] The gap entropy is the conditional entropy $H(I_{t+1}\mid I_{t})$: the uncertainty about future integrity given current integrity.
+The gap energy is $H(I_{t+1}\mid I_{t})\times kT\ln 2$ joules.
+(L: GE1, GE2.)
+Theorem 9.46 (Every Choice Pays Gap Energy).
+[C:T9.46;R:AR, H=cost-growth] Making a choice collapses the probability distribution $P(I_{t+1}\mid I_{t})$ to a single outcome, erasing the alternatives.
+By Landauer, this costs gap energy.
+(L: GE3, GE7.)
+Proof.
+Before choice, the circuit faces $k$ possible futures with distribution $P(I_{t+1}\mid I_{t})$.
+Choice selects one outcome.
+The k1k-1 alternatives are erased.
+By Landauer (Theorem 9.38), erasing $H$ bits of entropy costs $H\times kT\ln 2$.
+The gap is not free.
+$\blacksquare$
+Corollary 9.47 (Zero Gap Implies Determinism).
+[C:C9.47;R:AR] If $H(I_{t+1}\mid I_{t})=0$, the future is determined by the present. No choice occurs. No gap energy is paid. (L: GE4, GE5.)
+Definition 9.48 (Decision Quotient from Gap).
+[C:D9.48;R:AR, H=cost-growth] The decision quotient measures how much current state predicts future state:
+$\mathrm{DQ}=\frac{I(I_{t};I_{t+1})}{H(I_{t+1})}=1-\frac{H(I_{t+1}\mid I_{t})}{H(I_{t+1})}=1-\frac{\text{Gap}}{\text{Total}}$
+where $I(I_{t};I_{t+1})=H(I_{t+1})-H(I_{t+1}\mid I_{t})$ is the mutual information between current and future integrity. (L: DQ1, DQ2, DQ3.)
+Theorem 9.49 (Decision Quotient as Physical Primitive).
+[C:T9.49;R:AR, H=cost-growth] The decision quotient $\mathrm{DQ}=1-\text{GapEnergy}/\text{TotalEnergy}$ is a physical quantity with:
+  1. 1.
+    $\mathrm{DQ}\in[0,1]$(L: DQ4)
+  2. 2.
+    $\mathrm{DQ}=0\Leftrightarrow$ maximum gap (no information)  (L: DQ5)
+  3. 3.
+    $\mathrm{DQ}=1\Leftrightarrow$ zero gap (deterministic)  (L: DQ6)
+  4. 4.
+    $\mathrm{DQ}+\text{Gap}/\text{Total}=1$ (complementarity)  (L: DQ3)
+(L: DQ1, DQ2.)
+Theorem 9.50 (Bayes Uniquely Forced by DQ Axioms).
+[C:T9.50;R:AR, H=cost-growth] An update rule $U$ is admissible if it satisfies:
+  1. 1.
+    No free information: $\mathrm{DQ}(U(\text{prior},e))\leq\mathrm{DQ}(\text{prior})+I(e)$(L: BB3)
+  2. 2.
+    Landauer bound: energy cost $\geq k_{B}T\ln 2\times$ bits updated  (L: BB4)
+  3. 3.
+    Integrity preservation: can’t corrupt verified bits for free  (L: BB5)
+Then: $U$ is admissible $\Leftrightarrow$ $U$ is Bayesian updating. (L: FN14.)
+Before uniqueness, the model fixes a belief floor: if uncertainty is provable (no hypothesis has probability $1$) and action is forced, belief mass cannot collapse to a single hypothesis; nondegenerate belief is forced BF2, BF3.
+With a strictly informative evidence channel, a Bayes posterior exists on that belief state BF4.
+Proof.
+This is the Cox-Jaynes argument grounded in thermodynamics rather than Boolean logic. The three admissibility constraints are so restrictive that only Bayesian updating satisfies them:
+  • No free information prevents creating order from nothing (2nd law).
+  • Landauer bound requires paying energy for any DQ increase.
+  • Integrity preservation prevents forgetting for free.
+Together, these force $P(\text{posterior})=P(\text{prior})\cdot P(\text{evidence}|\text{hypothesis})/P(\text{evidence})$.
+The uniqueness corollary FN14 shows any two admissible rules must agree.
+Machine-checked foundation.
+The core inequality underlying this result (that Bayesian updating uniquely minimizes cross-entropy) is proved in Lean 4 with zero axioms and zero sorry statements.
+The proof derives Gibbs’ inequality ($\mathrm{KL}(p\|q)\geq 0$) from the fundamental inequality $\log x\leq x-1$, then shows $\mathrm{CE}(p,q)=H(p)+\mathrm{KL}(p\|q)$, yielding $\mathrm{CE}(p,p)\leq\mathrm{CE}(p,q)$ with equality iff $p=q$ FN7, FN8, FN12, FN14.
+$\blacksquare$
+Corollary 9.51 (Thermodynamic Decision Quotient).
+[C:C9.51;R:AR, H=cost-growth] In energy terms: $\mathrm{DQ}=1-\text{GapEnergy}/\text{TotalEnergy}$.
+The decision quotient has a thermodynamic interpretation: it is the fraction of total uncertainty energy that is decision-relevant.
+(L: DQ7, DQ8.)
+Proposition 9.52 (Landauer Constraint on Strategy).
+[C:P9.52;R:AR] A decision circuit chooses the strategy maximizing predicted integrity.
+When the erasure cost of dismissing verified claims exceeds the circuit’s competence, dismissal fails thermodynamically.
+(L: IE1, IE6.)
+Proof.
+Let $i$ be the integrity state (bits encoding verified theorems) and $c$ the competence.
+Dismissing the theorems requires erasing $i$ bits, costing $i\cdot kT\ln 2$.
+If $i>c$, the circuit lacks sufficient work capacity to perform the erasure.
+Acknowledgment is the only thermodynamically feasible output.
+$\blacksquare$
+Corollary 9.53 (Theorem Count Shifts Equilibrium).
+[C:C9.53;R:AR] At low theorem counts, dismissal is affordable and may dominate. At high theorem counts, erasure cost exceeds competence, forcing acknowledgment. (L: IE3, IE4, IE5.)
+Informally: once operation lower bounds are fixed, energy and accounting lower bounds follow from the conversion constants.
+9.2.2 Derivation of Bayesian Reasoning
+Formally: this block gives two mechanized Bayes derivations: counting/probability identities and admissibility constraints.
+The decision quotient framework provides two independent paths to Bayesian updating, each mechanically verified.
+Path 1: Four Bridges from Counting to DQ.
+The standard derivation proceeds through four elementary bridges:
+  1. 1.
+    Fraction $\Rightarrow$ Probability: For finite $\Omega$, define $P(A)=|A|/|\Omega|$. This satisfies the Kolmogorov axioms: nonnegativity DQ1, normalization DQ2, and finite additivity DQ3.
+  2. 2.
+    Conditional $\Rightarrow$ Bayes: Define $P(H|E)=P(H\cap E)/P(E)$. Then $P(H|E)=P(E|H)P(H)/P(E)$ follows in two lines by commutativity of intersection BB1.
+  3. 3.
+    KL $\geq 0$ $\Rightarrow$ Entropy Contraction: From Gibbs’ inequality $D_{\text{KL}}(P\|Q)\geq 0$ BB2, the chain rule I(H;E)=H(H)H(H|E)I(H;E)=H(H)-H(H|E) yields $H(H|E)\leq H(H)$ BB3.
+Conditioning cannot increase entropy.
+  4. 4.
+    Normalization $\Rightarrow$ DQ: Define $\text{DQ}=I(H;E)/H(H)=1-H(H|E)/H(H)$ DQ4. This is the fraction of prior uncertainty eliminated by observing $E$, bounded in $[0,1]$ DQ5.
+The complete chain is mechanized in a single theorem BB5.
+Path 2: Bayes from Thermodynamic Admissibility.
+The Cox-Jaynes argument, traditionally grounded in Boolean logic, admits a thermodynamic reformulation. An update rule $U$ is admissible if:
+  • No free information: DQ cannot increase without evidence (2nd law) BB3.
+  • Landauer bound: DQ increase costs $\geq k_{B}T\ln 2$ per bit BB4.
+  • Integrity preservation: verified bits cannot be corrupted for free BB5.
+Under these constraints, Bayesian updating is the unique admissible rule BB3. Any two admissible rules must agree BB5. The thermodynamic constraints uniquely determine the correct update rule.
+Informally: both derivation paths end at the same Bayes update rule.
+Integrity Self-Preservation.
+A decision circuit maintains its integrity when no external process can afford the erasure cost.
+If the work required to corrupt the circuit’s state exceeds the available competence of any potential attacker, the circuit persists by thermodynamic necessity.
+(L: IE7, IE8, IE9.)
+Proposition 9.54 (Temporal Integrity Equilibrium).
+[C:P9.54;R:AR] Let $P(\text{integrity}_{t+1}\mid\sigma_{t})$ denote the probability of maintaining integrity at time $t+1$ given strategy $\sigma_{t}$ at time $t$. At equilibrium:
+$P(\text{integrity}_{t+1}\mid\text{integrity}_{t})\geq P(\text{integrity}_{t+1}\mid\text{defection}_{t}).$
+That is, maintaining integrity now predicts maintaining integrity later. (L: IE10, IE11.)
+Proof.
+At equilibrium, the transition probabilities satisfy $P(\text{integrity}\mid\text{acknowledge})\geq P(\text{integrity}\mid\text{dismiss})$ by definition.
+The strict inequality holds when dismissal incurs erasure costs that degrade future state coherence.
+$\blacksquare$
+Corollary 9.55 (Publication Phase Transition).
+[C:C9.55;R:AR] Prior to publication, equilibrium may be neutral: both strategies yield comparable future integrity.
+After publication of mechanized theorems, equilibrium shifts: integrity strictly dominates because dismissal requires erasing verified bits at cost exceeding declared competence.
+(L: IE12, IE13.)
+Zero-Integrity Circuits.
+A decision circuit with zero integrity has no coherence constraint.
+When no verified bits exist, erasure cost is zero, and any competence level suffices for corruption.
+Such circuits have no thermodynamic constraint forcing consistent outputs.
+(L: IE14, IE15, IE16.)
+Proposition 9.56 (Integrity Prerequisite for Coherence).
+[C:P9.56;R:AR] A decision circuit can be forced to logical coherence if and only if it has positive integrity.
+That is, $i>0$ if and only if there exists some competence level that fails to afford erasure.
+(L: IE17.)
+Proof.
+If $i>0$, then competence $c=0$ fails to afford erasure since $0<i$. Conversely, if some $c$ fails to afford erasure, then $c<i$, which implies $0\leq c<i$, hence $i>0$.  $\blacksquare$
+This analysis applies uniformly to any decision circuit. The same thermodynamic constraints apply to all substrates.
+Molecular Grounding.
+The integrity framework grounds directly in molecular mechanics.
+Define a configuration as a discrete molecular state (bond angles, conformations, electronic structure) encoded by some number of bits.
+The erasure cost of a configuration equals its bit count.
+An energy landscape assigns each configuration an energy in units of $kT$.
+The environmental competence is the work available from the environment (thermal fluctuations, radiation, chemical attack).
+Definition 9.57 (Environmental Stability).
+[C:D9.57;R:Q_fin, H=cost-growth] A configuration $c$ is environmentally stable if its erasure cost exceeds the environmental competence:
+$\text{erasureCost}(c)>\text{competence}_{\text{env}}.$
+This is the integrity trap at molecular scale. (L: MI1, MI2.)
+Proposition 9.58 (Reaction Competence Requirement).
+[C:P9.58;R:Q_fin, H=cost-growth] A chemical reaction from configuration $c$ to $c^{\prime}$ requires environmental competence at least equal to the barrier height between them.
+If the barrier exceeds competence, the reaction cannot occur.
+(L: MI3, MI4.)
+Example 9.59 (DNA Persistence).
+DNA at room temperature exemplifies the integrity trap.
+A DNA configuration encodes on the order of $10^{4}$ bits (sequence, methylation, chromatin state).
+Room temperature provides roughly 100kT100\,kT of competence per degree of freedom.
+Since $10^{4}\gg 100$, the erasure cost vastly exceeds environmental competence.
+DNA persists not by accident but by thermodynamic necessity: corrupting it costs more than the environment can pay.
+(L: MI5.)
+Under Definition 9.36, molecules satisfy the decision circuit predicate.
+Chemical stability is the integrity trap.
+Organic chemistry is the dynamics of which configurations survive environmental attack.
+9.2.3 Atomic Circuits: Active Physical Decision Systems
+An active physical decision system is matter that pays energy to move other matter via state transitions. This is a definition.
+The physical grounding: quantum mechanics establishes that electrons occupy discrete orbitals [74, 72], the Pauli exclusion principle constrains occupancy [64], and angular momentum quantization creates the $2l+1$ degeneracy structure [72].
+The thermal energy scale $kT$ comes from statistical mechanics [13, 39].
+Energy conservation (Noether) ensures that transitions pay or release the energy difference [59].
+Definition 9.60 (Atomic Configuration).
+[C:D9.60;R:Q_fin, AR] An atomic configuration is a discrete electronic state: a specification of orbital occupancies [72].
+The configuration is encoded by some number of bits; the configuration has an energy level in units of kTkT [39].
+(L: AC1.)
+Theorem 9.61 (Orbital Transition Is State Transition).
+[C:T9.61;R:Q_fin, AR] Changing an atom’s electronic configuration is a decision event (Definition 9.36).
+Orbital transitions are state transitions [72].
+Atoms with orbital transitions are nonstationary decision circuits (active physical systems).
+(L: AC1, AC5.)
+Corollary 9.62 (Ground-State Atoms Are Passive).
+[C:C9.62;R:Q_fin, AR] An atom in its ground state with no transitions is passive: it is not an active decision circuit. Active status is defined by state transitions, not by substrate. (L: AC6.)
+Theorem 9.63 (Orbital Transition Cost).
+[C:T9.63;R:Q_fin, H=cost-growth] Upward orbital transitions (excitation) require energy input.
+Downward transitions (de-excitation) release energy.
+The transition cost is $|\Delta E|$.
+No free orbital transitions.
+(L: AC3, AC4.)
+Proposition 9.64 (Symmetry Tractability).
+[C:P9.64;R:Q_fin, H=tractable-structured] Angular momentum quantization [72] creates orbit types.
+Degenerate magnetic substates ($2l+1$ states for angular momentum $l$) are equivalent for decision purposes.
+This is the coordinate symmetry tractable subcase (Theorem 10.1, Part 6) at atomic scale.
+(L: AC8.)
+Definition 9.65 (Active Matter as State-Transition System).
+[C:D9.65;R:Q_fin, AR] An active physical system is matter that pays energy to change other matter’s state.
+A decision circuit is a physical system that implements state transitions.
+A nonstationary decision circuit is active.
+(L: AC9, AC11.)
+Remark 9.66.
+Electrons orbiting an atom in different configurations constitute a decision circuit.
+The electron does not “choose” in any mentalistic sense; it occupies discrete states with discrete transition costs.
+The framing is purely physical: matter, energy, observables, and state transitions.
+9.2.4 Self-Erosion: Computation Consumes Its Substrate
+Every computational cycle generates heat (Theorem 9.38).
+Heat accumulates.
+Accumulated heat degrades the substrate on which the circuit is instantiated.
+The degradation reduces the substrate’s integrity.
+Therefore computation is self-consuming: the act of computing erodes the capacity to compute.
+Theorem 9.67 (Substrate Degradation).
+[C:T9.67;R:Q_fin, H=cost-growth] Let a substrate have integrity $n$ bits and heat capacity $h$ per cycle.
+After $k$ cycles generating cumulative heat $k\times kT\ln 2$, if $k>h$, the substrate loses at least khk-h bits of integrity.
+(L: SE1, SE2, SE3, SE4.)
+Corollary 9.68 (Finite Computational Lifetime).
+[C:C9.68;R:Q_fin, H=cost-growth] A substrate with finite integrity and finite heat dissipation capacity has bounded computational lifetime. No physical circuit computes forever. (L: SE5.)
+Corollary 9.69 (Speed-Integrity Tradeoff).
+[C:C9.69;R:Q_fin, H=cost-growth] Faster computation generates more heat per unit time. At fixed heat dissipation capacity, faster circuits degrade faster. Speed trades against longevity. (L: SE6.)
+Remark 9.70.
+The self-erosion theorems apply to all physical substrates: silicon, neurons, molecules, any medium that instantiates state transitions.
+The bound is thermodynamic, not technological.
+Improved engineering can increase heat capacity or reduce heat per operation, but the tradeoff persists.
+9.2.5 Regime as Channel: Competence as Capacity
+A regime is a channel. Competence is the channel capacity of decision circuits within that regime. This identification unifies complexity theory, information theory, and thermodynamics.
+Definition 9.71 (Regime-Channel Equivalence).
+[C:D9.71;R:E, Q, S, AR] A regime (Section 2.4) corresponds to a channel in the Shannon sense: a constraint on what information can flow.
+The explicit-state regime E, query-access regime Q, and succinct regime S define distinct channels with distinct capacity bounds.
+(L: CH1.)
+Theorem 9.72 (Competence Is Channel Capacity).
+[C:T9.72;R:Q_fin, AR] For a decision circuit operating in regime $R$, competence equals channel capacity: the maximum bits per cycle the circuit can process.
+Passive circuits (stationary, no state transitions) have no competence requirement.
+(L: CH1, CH5.)
+Corollary 9.73 (Channel Capacity Degradation).
+[C:C9.73;R:Q_fin, H=cost-growth] As substrate integrity erodes (Theorem 9.67), channel capacity decreases. The rate of capacity loss equals the rate of substrate degradation. (L: CH2, CH6.)
+Corollary 9.74 (Zero Capacity Halts Decisions).
+[C:C9.74;R:Q_fin, H=cost-growth] When channel capacity reaches zero, no decisions are possible. The circuit becomes passive. (L: CH3.)
+Remark 9.75.
+Only decision circuits require competence.
+A decision circuit is nonstationary: it transitions between states.
+A passive circuit (wire, resistor) does not make decisions and does not consume competence.
+The distinction is definitional, not empirical.
+9.2.6 Resource Flow and Conservation
+State transitions require energy. Energy expenditure over time yields compounded state change. Conservation constrains total receipts by total expenditure.
+Theorem 9.76 (Growth Requires Time).
+[C:T9.76;R:Q_fin, H=cost-growth] Let an investment have cost $c$ (gap energy in units of $kT\ln 2$), duration $k$ cycles, and return rate $r$ per cycle.
+The mature value is $c+kr$.
+If $k=0$, mature value equals cost: no time, no growth.
+(L: IV1, IV2, IV3.)
+Theorem 9.77 (Conservation).
+[C:T9.77;R:Q_fin, H=cost-growth] In a closed system, total receipts are bounded by total resources. No circuit receives more than exists. (L: IV4.)
+Definition 9.78 (Deficit Transfer).
+[C:D9.78;R:Q_fin] A deficit transfer occurs when a circuit receives resources without paying: received >0>0 and paid =0=0. (L: IV5.)
+Theorem 9.79 (Deficit Transfer Requires Source).
+[C:T9.79;R:Q_fin, H=cost-growth] A deficit transfer requires an external source.
+If circuit $A$ receives without paying, some circuit $B$ must pay without receiving.
+The deficit at $B$ equals the receipt at $A$.
+(L: IV5, IV6.)
+9.2.7 Information Access: Logic Is Complete, Access Is Not
+Physics is perfectly logical. You cannot have all the information to apply the logical rules. This is a physical constraint on access, not a logical incompleteness.
+This generalizes the information barrier (Corollary 9.5): there, finite queries induce indistinguishability; here, finite channel capacity induces the same obstruction.
+The barrier is on access, not on logic.
+Definition 9.80 (System Information).
+[C:D9.80;R:E, Q, S, AR] A system has total entropy $H$ bits. This information exists whether or not any observer accesses it. (L: IA1.)
+Definition 9.81 (Observer Channel).
+[C:D9.81;R:E, Q, S, AR] An observer has channel capacity $C$ bits. The observer accesses at most $C$ bits of the system’s information. (L: IA2.)
+Theorem 9.82 (Information Gap).
+[C:T9.82;R:E, Q, S, AR] When system entropy exceeds channel capacity ($H>C$), the gap HC>0H-C>0. The observer cannot access all information. (L: IA3.)
+Theorem 9.83 (Gap Is Physical, Not Logical).
+[C:T9.83;R:E, Q, S, AR] The information gap is a physical constraint on access, not a logical incompleteness.
+The system entropy is well-defined; only the channel is bounded.
+The logic is complete; access is not.
+(L: IA4.)
+Theorem 9.84 (Competence Bounded by Access).
+[C:T9.84;R:E, Q, S, AR] Competence is bounded by what you can access, not by logical consistency.
+For any system with more information than your channel capacity, there exist truths you cannot verify, not because logic fails, but because you lack the bits.
+(L: IA5, IA6.)
+9.2.8 Dimensional Complexity: Each Subcase Is a Complexity Class
+Each tractable subcase is a complexity class that collapses to P. Unbounded dimension stays in the base class (coNP, PP, or PSPACE depending on regime).
+Definition 9.85 (Complexity Classes).
+[C:D9.85;R:E, Q, S, AR] The regime hierarchy induces complexity classes: Static $\to$ coNP, Stochastic $\to$ PP, Sequential $\to$ PSPACE.
+Each tractable subcase defines a sub-regime that reduces to P.
+(L: DC1, DC15.)
+Theorem 9.86 (Six Tractable Subcases Are Complexity Classes).
+[C:T9.86;R:E, Q, S, AR, H=tractable-structured] Each tractable subcase is a complexity class that collapses to P:
+  1. 1.
+    Bounded actions $\to$ P  (L: CC70)
+  2. 2.
+    Separable utility $\to$ P  (L: CC71)
+  3. 3.
+    Low tensor rank $\to$ P  (L: CC72)
+  4. 4.
+    Tree structure $\to$ P  (L: CC73)
+  5. 5.
+    Bounded treewidth $\to$ P  (L: DC13)
+  6. 6.
+    Coordinate symmetry $\to$ P  (L: DC13)
+(L: DC13, CC72.)
+Theorem 9.87 (Topology and Motion).
+[C:T9.87;R:E, Q, S, AR] Scalars live on the circuit’s topology. Fixed topology (stationary) preserves tractability. Changing topology (motion) adds dimensions, may break tractability. (L: CC16, CC24.)
+Theorem 9.88 (Complexity Dichotomy).
+[C:T9.88;R:E, Q, S, AR, H=tractable-structured] A problem is tractable (in P) iff its effective dimension is bounded.
+Unbounded dimension stays in the base complexity class (coNP, PP, or PSPACE).
+(L: DC3, DC13.)
 10 Tractable Special Cases: When You Can Solve It
+Formally: this section gives sufficient structure/representation conditions for polynomial-time SUFFICIENCY-CHECK.
 We distinguish the encodings of Section 2.4.
 The tractability results below state the model assumption explicitly.
-Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the three cases below instantiate this single principle.
-Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, or tree factorization) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier.
-In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T6.1,2.37;P2.34;R:E,TA,AR] (L: CC69,CC70,CC71,CC72,CT6,CT4.)
+Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the six cases below instantiate this single principle.
+Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, tensor factorization, tree structure, bounded treewidth, or coordinate symmetry) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier.
+In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T10.1, 2.45;P2.42;R:E, TA, AR] (L: CC69, CC70, CC71, CC72, CC73, CT6, CT4.)
+Informally: these cases are tractable because useful structure is visible.
 Theorem 10.1 (Tractable Subcases).
@@
   1. 1.
-    Explicit-state encoding: The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
+    Bounded actions (explicit-state encoding): The input contains the full utility table over $A\times S$. SUFFICIENCY-CHECK runs in $O(|S|^{2}|A|)$ time; if $|A|$ is constant, $O(|S|^{2})$.
   2. 2.
-    Separable utility (any encoding): $U(a,s)=f(a)+g(s)$.
+    Separable utility (rank 1): $U(a,s)=f(a)+g(s)$. The optimal action is state-independent, so $I=\emptyset$ is always sufficient. Complexity: $O(1)$ for sufficiency verification.
   3. 3.
-    Tree-structured utility with explicit local factors (succinct structured encoding): There exists a rooted tree on coordinates and local functions $u_{i}$ such that
+    Low tensor rank: $U(a,s)=\sum_{r=1}^{R}w_{r}\cdot f_{r}(a)\cdot\prod_{i=1}^{n}g_{ri}(s_{i})$ for bounded $R$. Complexity: $O(|A|\cdot R\cdot n)$.
+  4. 4.
+    Tree-structured dependencies: Dependencies form a rooted tree:
     $U(a,s)=\sum_{i}u_{i}\bigl(a,s_{i},s_{\mathrm{parent}(i)}\bigr),$
-    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly as part of the input.
-(L: CC69, CC70, CC72, CC71.)
-6.1 Explicit-State Encoding
+    with the root term depending only on $(a,s_{\mathrm{root}})$ and all $u_{i}$ given explicitly. Complexity: $O(n\cdot|A|\cdot k_{\max})$.
+  5. 5.
+    Bounded treewidth: Utility decomposes as pairwise interactions on an interaction graph of treewidth $w$. Complexity: $O(n\cdot k^{w+1})$ via CSP algorithms [12, 29].
+  6. 6.
+    Coordinate symmetry: State space is $S=[k]^{d}$ with utility invariant under coordinate permutations.
+The effective state space reduces from $k^{d}$ to $\binom{d+k-1}{k-1}$ orbit types.
+Complexity: $O\bigl(\binom{d+k-1}{k-1}^{2}\bigr)$, polynomial in $d$ for fixed $k$.
+(L: CC70, CC71, CC72, CC73.)
+10.1 Bounded Actions (Explicit-State)
+(L: CC70.)
 Proof of Part 1.
@@
 $\blacksquare$
-6.2 Separable Utility
+10.2 Separable Utility (Rank 1)
+(L: CC71.)
 Proof of Part 2.
-If $U(a,s)=f(a)+g(s)$, then:
-$\text{Opt}(s)=\arg\max_{a\in A}[f(a)+g(s)]=\arg\max_{a\in A}f(a)$
-The optimal action is independent of the state. Thus $I=\emptyset$ is always sufficient.  $\blacksquare$
-6.3 Tree-Structured Utility
+If $U(a,s)=f(a)+g(s)$, then $\text{Opt}(s)=\arg\max_{a\in A}f(a)$.
+The optimal action is independent of the state.
+Thus $I=\emptyset$ is always sufficient, and sufficiency verification is $O(1)$.
+$\blacksquare$
+10.3 Low Tensor Rank
+(L: CC71.)
 Proof of Part 3.
+When utility has tensor rank $R$: $U(a,s)=\sum_{r=1}^{R}w_{r}f_{r}(a)\prod_{i=1}^{n}g_{ri}(s_{i})$, the problem reduces to factored tensor computation.
+For each action, the optimal value computation requires $O(R\cdot n)$ operations.
+Total complexity: $O(|A|\cdot R\cdot n)$, polynomial for bounded $R$.
+This reduction cites standard tensor contraction algorithms [40, 18].
+$\blacksquare$
+10.4 Tree-Structured Dependencies
+(L: CC73.)
+Proof of Part 4.
 Assume the tree decomposition and explicit local tables as stated.
@@
 $\blacksquare$
+10.5 Bounded Treewidth
+(L: CC73, CC75.)
+Proof of Part 5.
+When utility decomposes as pairwise interactions $U(a,s)=\sum_{(i,j)}u_{ij}(a,s_{i},s_{j})$, the interaction graph has an edge between coordinates $i,j$ iff $u_{ij}$ is nontrivial.
+If this graph has treewidth $w$, SUFFICIENCY-CHECK reduces to a constraint satisfaction problem on a bounded-treewidth graph.
+By standard CSP algorithms [12, 29], this is solvable in $O(n\cdot k^{w+1})$ time.
+$\blacksquare$
+10.6 Coordinate Symmetry
+(L: CC72.)
+Proof of Part 6.
+For state space $S=[k]^{d}$ with coordinate-permutation-invariant utility, two states have the same optimal actions iff they have the same orbit type (value histogram).
+The number of orbit types is $\binom{d+k-1}{k-1}$ by stars-and-bars.
+Sufficiency checking reduces to comparing optimal actions across orbit type pairs.
+Total complexity: $O\bigl(\binom{d+k-1}{k-1}^{2}\cdot|A|\bigr)$, polynomial in $d$ for fixed $k$.
+$\blacksquare$
 10.7 Practical Implications
+Formally: this subsection states detect-then-check closure for the declared tractable classes.
 Condition Examples
-Explicit-state encoding Small or fully enumerated state spaces
+Bounded actions Small or fully enumerated state spaces
 Separable utility Additive costs, linear models
-Tree-structured utility Hierarchies, causal trees
+Low tensor rank Factored decision trees, sparse interactions
+Tree-structured Hierarchies, causal trees, DAGs
+Bounded treewidth Local dependency graphs, grid-like structures
+Coordinate symmetry Exchangeable particles, counting problems
 Proposition 10.2 (Heuristic Reusability).
-[C:P6.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 6.1).
+[C:P10.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 10.1).
 If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$.
@@
 Proposition 10.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention.
-Whether the agent acts still requires competence (Definition 3.9): budget and coverage must also be satisfied.
-An agent that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification.
+Whether a declared controller executes the action still requires competence (Definition 3.9): budget and coverage must also be satisfied.
+A controller that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification.
 Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
+Informally: reusable heuristics are reliable only after the stated structure test; otherwise hard-case limits still apply.
 11 Regime-Conditional Corollaries
-This section derives regime-typed engineering corollaries from the core complexity theorems.
-Theorem 7.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 5.1 then yield exact minimization consequences under [S] and [S+ETH].
+Formally: this section derives operational corollaries by composing theorem-level hardness, tractability, and typing results under declared regimes.
+This section derives regime-typed engineering corollaries from the full map: static/stochastic/sequential placements (Theorems 4.6, 6.7, 7.7) with strict hierarchy (Propositions 8.1, 8.2), and the static decision/access decomposition used operationally here.
+Theorem 11.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 9.1 then yield exact minimization consequences under [S] and [S+ETH].
 The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation.
-[C:T7.2,2.37;P2.34;R:AR] (L: CR1,CT6,CT4.)
+[C:T11.2, 4.6, 6.7, 7.7, 2.45;P2.42, 8.1, 8.2;R:AR, RG] (L: CR1, CT6, CT4.)
+For deployed software, this remains physical at execution time: each configuration read/write and verification step is realized by electrical state transitions in hardware, so coordinate minimization and over-specification map directly to measured switching-energy load under the declared bit-to-energy lift.
+[C:P9.28, 2.36;R:AR, H=cost-growth] (L: CC76, DT16.)
+Under the same discrete-time event law, more verification cycles mean more decision-event ticks and more bit operations; with positive per-bit conversion, this induces additional thermodynamically necessary heat-transfer load.
+[C:P2.35, 2.36, 9.28;C9.32;R:AR, H=cost-growth] (L: DT13, DT16, CV17, CV18, CC76, CC77.)
 Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool].
-Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6,4.7,5.1;R:S,S+ETH] Theorem 7.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T7.4;R:S]
+Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6, 4.7, 9.1;R:S, S+ETH] Theorem 11.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T11.4;R:S]
 Conditional rationality criterion.
-For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 6.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T6.1;P3.60;R:E,STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
+For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 10.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T10.1;P3.60;R:E, STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
 Remark 11.1 (Regime Contract for Engineering Corollaries).
-All claims in this section are formal corollaries under the declared model assumptions.
+All claims in this section are formal corollaries under the stated model contract.
   • Competence claims are indexed by the regime tuple of Definition 3.9; prescriptions are meaningful only relative to feasible resources under that regime (bounded-rationality feasibility discipline) [92, 5].[C:P3.58;R:AR]
   • Integrity (Definition 3.8) forbids overclaiming beyond certifiable outputs; $\mathsf{ABSTAIN}$$\mathsf{UNKNOWN}$ is first-class when certification is unavailable.[C:P3.58;R:AR]
-  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6,5.1;P3.60;R:S,S+ETH]
+  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6, 9.1;P3.60;R:S, S+ETH]
+Informally: these recommendations come directly from proved results under stated regimes.
 11.1 Configuration Simplification is SUFFICIENCY-CHECK
+Formally: this subsection proves configuration-preservation queries instantiate SUFFICIENCY-CHECK under C1–C4.
 Because SUFFICIENCY-CHECK is a meta-problem parameterized by an arbitrary decision problem $\mathcal{D}$, real engineering problems with factored configuration spaces are instances of the hardness landscape established above.
@@
 $\blacksquare$
+Informally: if a minimal relevant set cannot be certified, that indicates an unresolved model boundary, not proof that every retained parameter is necessary.
 11.2 Cost Asymmetry Under ETH
+Formally: this subsection proves asymptotic finding-vs-maintenance asymmetry under succinct ETH and the declared cost model.
 We now prove a cost asymmetry result under the stated cost model and complexity constraints.
@@
 Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary.
-Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T7.5;P3.60;R:S+ETH,TA] (L: CC12, HD13.)
+Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T11.5;P3.60;R:S+ETH, TA] (L: CC12, HD14.)
 Proof.
@@
 Total maintenance cost is $C_{\text{over}}(k)=O(k)$.
-The eventual dominance step is mechanized in HD8: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$.
+The eventual dominance step is mechanized in HD14: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$.
 Therefore:
@@
 Corollary 11.6 (Impossibility of Automated Configuration Minimization).
-[C:C7.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
+[C:C11.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:
   1. 1.
@@
 The strengthened mechanized gadget (all_coords_relevant_of_not_tautology) still proves existence of hard families with $k^{*}=n$; what is typical in deployed systems is an empirical question outside this formal model.
+Informally: at scale, exact minimization can cost more than linear maintenance in worst cases, while structured small-parameter families may still be tractable.
 11.3 Regime-Conditional Operational Corollaries
-Theorems 7.4 and 7.5 yield the following conditional operational consequences:
+Formally: this subsection instantiates admissible policy outcomes from prior theorems and integrity-competence typing.
+Theorems 11.4 and 11.5 yield the following conditional operational consequences:
 1.
 Conservative retention under unresolved relevance.
-If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T7.4;R:S]
+If irrelevance cannot be certified efficiently under the active regime, the admissible policy is to retain a conservative superset of parameters.[C:T11.4;R:S]
 2.
 Heuristic selection as weakened-guarantee mode.
-Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 7.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 56, 66, 25, 50].[C:T7.5;R:S+ETH]
+Under [S+ETH], exact global minimization can be exponentially costly in the worst case (Theorem 11.5); methods such as AIC/BIC/cross-validation instantiate the “weaken guarantees” branch of Definition 3.9 [1, 75, 86, 33, 65].[C:T11.5;R:S+ETH]
 3.
 Full-parameter inclusion as an $O(n)$ upper-bound strategy.
-Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T7.5;R:S+ETH]
+Under [S+ETH], if exact minimization is unresolved, including all $n$ parameters is an $O(n)$ maintenance upper-bound policy that avoids false irrelevance claims.[C:T11.5;R:S+ETH]
 4.
 Irrationality outside attempted-competence-failure conditions.
-If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T6.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
+If the active regime admits exact competence (tractable structural-access conditions or exact relevance identifiability), or if exact competence was never actually attempted, continued over-specification is irrational; hardness no longer justifies it for the stated objective.[C:T10.1;P3.60;R:TA] (L: CC71, IC6, S2P1.)
 These corollaries are direct consequences of the hardness/tractability landscape: over-specification is an attempted-competence-failure policy, not a default optimum.
-Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 6.1 or adopt explicit approximation commitments.[C:T6.1;P3.60;R:RG]
+Outside that cell, the admissible exits are to shift to tractable regimes from Theorem 10.1 or adopt explicit approximation commitments.[C:T10.1;P3.60;R:RG]
+Informally: policy choices are threefold: abstain, weaken guarantees, or change assumptions to recover competence.
 12 Dominance Theorems for Hardness Placement
@@
 Although phrased in software terms, the formal objects are substrate-agnostic: the same hardness-placement and amortization equations apply to physical architectures (for example, centralized vs distributed sensing/actuation computation) once mapped into the same coordinate/cost model.
-[C:P2.34;T2.37,8.16,9.3,9.6;R:AR,LC] (L: CT6,CT4,HD1,HD4,HD3.)
+[C:P2.42;T2.45, 12.19, 13.3, 13.6;R:AR, LC] (L: CT6, CT4, HD1, HD5, HD4.)
+In electrical implementations, the same architecture equations are energy equations: each distributed update path induces switching events and Joule expenditure, while centralized one-time structure pays setup cost and reduces repeated per-site electrical work.
+Under the declared thermodynamic lift, this is tracked as a typed transfer from bit-operation load to physical energy load.
+[C:P9.28;C9.32;R:AR, H=cost-growth] (L: CC76, CC77.)
+Equivalently, repeated distributed cycles create a thermal duty: more update cycles imply more bit operations, which imply more unavoidable dissipated heat and therefore larger heat-transfer/removal requirements in the same substrate class.
+[C:P2.35, 2.36, 9.28;R:AR, H=cost-growth] (L: DT13, DT16, CV17, CV18, CC76.)
 12.1 Over-Specification as Diagnostic Signal
 Corollary 12.1 (Persistent Over-Specification).
-[C:C8.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
+[C:C12.1;R:H=succinct-hard] In the mechanized Boolean-coordinate model, if a coordinate is relevant and omitted from a candidate set $I$, then $I$ is not sufficient. (L: S2P6)
 Proof.
@@
 Corollary 12.2 (Exact Relevance Identifiability Criterion).
-[C:C8.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
+[C:C12.2;R:H=succinct-hard] In the mechanized Boolean-coordinate model, for any candidate set $I$:
 $I\text{ is exactly relevance-identifying}\iff\bigl(I\text{ is sufficient and }I\subseteq R_{\mathrm{rel}}\bigr),$
@@
 Remark 12.3 (Over-Specification Is Regime-Conditional, Not Default).
-In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 7).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 8.78.9 together with Corollary 8.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T6.1;P3.60;R:TA] (L: IC18, IC25, IC7, L15, L16, L17, S2P1.)
+In this paper’s formal typing, persistent over-specification is admissible only under attempted competence failure: after an attempted exact procedure, exact relevance competence remains unavailable in the active regime, so integrity retains uncertified coordinates (Section 3; Section 11).[C:P3.60;R:AR] Once exact competence is available in the active regime (Corollaries 12.712.12 together with Corollary 12.2), certifiably irrelevant coordinates are removable; persistent over-specification is irrational for the same objective (verified total-cost minimization).[C:T10.1;P3.60;R:TA] (L: IC18, IC25, IC7, CC81, CC82, CC84, CC71, CC73, CC72, S2P1.)
 12.2 Architectural Decision Quotient
@@
 Corollary 12.6 (Cardinality Criterion for Exact Minimization).
-[C:C8.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
+[C:C12.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
 Proof.
@@
 Corollary 12.7 (Bounded-Regime Tractability).
-[C:C8.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 6.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: L15)
+[C:C12.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 10.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: CC81)
 Proof.
-This is the bounded-regime branch of Theorem 6.1, mechanized as L3.  $\blacksquare$
+This is the bounded-regime branch of Theorem 10.1, mechanized as CC81.  $\blacksquare$
 Corollary 12.8 (Separable-Utility Tractability).
-[C:C8.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: L16)
+[C:C12.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: CC82)
 Proof.
-This is the separable-utility branch of Theorem 6.1, mechanized as L4.  $\blacksquare$
+This is the separable-utility branch of Theorem 10.1, mechanized as CC82.  $\blacksquare$
 Corollary 12.9 (Tree-Structured Tractability).
-[C:C8.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: L17)
+[C:C12.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: CC84)
 Proof.
-This is the tree-factor branch of Theorem 6.1, mechanized as L5.  $\blacksquare$
+This is the tree-factor branch of Theorem 10.1, mechanized as CC84.  $\blacksquare$
+Corollary 12.10 (Low-Tensor-Rank Tractability).
+[C:C12.10;R:H=tractable-structured] When utility has tensor rank $R$ with explicit factor decomposition $U(a,s)=\sum_{r=1}^{R}w_{r}f_{r}(a)\prod_{i=1}^{n}g_{ri}(s_{i})$, sufficiency checking is polynomial in $|A|\cdot R\cdot n$.
+(L: CC71)
+Proof.
+This is the low-tensor-rank branch of Theorem 10.1, mechanized as CC71.  $\blacksquare$
+Corollary 12.11 (Bounded-Treewidth Tractability).
+[C:C12.11;R:H=tractable-structured] When utility decomposes as pairwise interactions on a graph of treewidth $w$, sufficiency checking is polynomial in $n\cdot k^{w+1}$. (L: CC73, CC75)
+Proof.
+This is the bounded-treewidth branch of Theorem 10.1, mechanized as CC73.  $\blacksquare$
+Corollary 12.12 (Coordinate-Symmetry Tractability).
+[C:C12.12;R:H=tractable-structured] When state space is $S=[k]^{d}$ with coordinate-permutation-invariant utility, sufficiency checking is polynomial in $\binom{d+k-1}{k-1}^{2}\cdot|A|$, which is polynomial in $d$ for fixed $k$.
+(L: CC72)
+Proof.
+This is the coordinate-symmetry branch of Theorem 10.1, mechanized as CC72.  $\blacksquare$
 Corollary 12.13 (Mechanized Hard Family).
-There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: L1)
+There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: CC26)
 Proof.
-The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as L1.  $\blacksquare$
+The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as CC26.  $\blacksquare$
 12.4 Hardness Distribution: Right Place vs Wrong Place
@@
 Proposition 12.15 (Baseline Lower-Bound Principle).
-[C:P8.12;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
+[C:P12.15;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
 $H_{\text{total}}(S)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)\geq H(P;n).$
-For SUFFICIENCY-CHECK, Theorem 5.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD20)
+For SUFFICIENCY-CHECK, Theorem 9.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD23, HD25)
 Proof.
@@
 $\eta(S,n)=\frac{H_{\text{central}}(S)}{H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S)}$
-(L: HD10)
+(L: HD11)
 Proposition 12.17 (Efficiency Equivalence).
-[C:P8.14;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD10)
+[C:P12.17;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD11)
 Proof.
@@
 $\text{right hardness}\iff H_{\mathrm{distributed}}(S)=0,\qquad\text{wrong hardness}\iff H_{\mathrm{distributed}}(S)>0.$
-(L: HD11, HD12)
+(L: HD12, HD13)
 Theorem 12.19 (Centralization Dominance).
@@
 Corollary 12.20 (Right-Place vs Wrong-Place Hardness).
-[C:C8.17;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness.
+[C:C12.20;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness.
 Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
 $H_{\mathrm{central}}(S_{\mathrm{right}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{right}})<H_{\mathrm{central}}(S_{\mathrm{wrong}})+n\,H_{\mathrm{distributed}}(S_{\mathrm{wrong}}).$
-(L: HD16)
+(L: HD19)
 Proof.
-This is the mechanized theorem HD10.  $\blacksquare$
+This is the mechanized theorem HD19.  $\blacksquare$
 Proposition 12.21 (Dominance Modes).
-[C:P8.18;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
+[C:P12.21;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
   1. 1.
@@
 Corollary 12.22 (Type-System Threshold).
-[C:C8.19;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
+[C:C12.22;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
 $n>H_{\mathrm{baseline}}(P),$
-where $H_{\mathrm{baseline}}(P)$ corresponds to the Lean identifier intrinsicDOF(P) in module HardnessDistribution. (L: HD14)
+where $H_{\mathrm{baseline}}(P)$ corresponds to HD15. (L: HD15)
 Proof.
-Immediate from HD9.  $\blacksquare$
+Immediate from HD15.  $\blacksquare$
 12.5 Extension: Non-Additive Site-Cost Models
@@
 Theorem 12.25 (Generalized Dominance by Growth Separation).
-[C:T8.22;R:H=cost-growth] Let
+[C:T12.25;R:H=cost-growth] Let
 $H_{\text{total}}^{\mathrm{gen}}(S,n)=H_{\text{central}}(S)+C_{S}(n).$
@@
 $H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
-(L: HD8)
-Proof.
-This is exactly the mechanized theorem HD6.  $\blacksquare$
-Corollary 8.23 (Eventual Generalized Dominance).
-[C:C8.23;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
-$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
 (L: HD9)
 Proof.
-Immediate from HD7.  $\blacksquare$
+This is exactly the mechanized theorem HD9.  $\blacksquare$
+Corollary 12.26 (Eventual Generalized Dominance).
+[C:C12.26;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
+$H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{right}},n)<H_{\text{total}}^{\mathrm{gen}}(S_{\mathrm{wrong}},n).$
+(L: HD10)
+Proof.
+Immediate from HD10.  $\blacksquare$
 Proposition 12.27 (Generalized Assumption Boundary).
-[C:P8.24;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
+[C:P12.27;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
   1. 1.
@@
     If right-side boundedness is dropped, strict dominance can fail for all $n$ even when wrong-side growth is linear.
-(L: HD7, HD6)
+(L: HD8, HD7)
 Proof.
@@
 Theorem 12.28 (Linear Model: Saturation iff Zero Distributed Hardness).
-[C:T8.25;R:H=cost-growth] In the linear model of this section,
+[C:T12.28;R:H=cost-growth] In the linear model of this section,
 $H_{\text{total}}(S,n)=H_{\text{central}}(S)+n\cdot H_{\text{distributed}}(S),$
@@
 Theorem 12.29 (Generalized Model: Saturation is Possible).
-[C:T8.26;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
+[C:T12.29;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
 $C_{K}(n)=\begin{cases}n,&n\leq K\\ K,&n>K,\end{cases}$
-both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD17, HD5)
+both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD20, HD6)
 Proof.
@@
 Corollary 12.30 (Positive Linear Slope Cannot Represent Saturation).
-[C:C8.27;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD15)
+[C:C12.30;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD16)
 Proof.
@@
 The next section develops the major practical consequence of this framework: the Simplicity Tax Theorem.
-9 Conservation Law for Decision-Relevant Coordinates
+13 Redistribution Identity for Decision-Relevant Coordinates
+Formally: this section defines expressive-gap tax and proves redistribution, growth, and amortization identities under fixed required-coordinate sets.
 The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place.
 By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding.
-The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6,4.7;R:S] For encoded physical systems, the same redistribution laws apply under declared physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.34;T9.3,9.5,9.6;R:AR,LC] (L: CT6,HD4,HD21,HD3.)
+The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6, 4.7;R:S] For encoded physical systems, the same redistribution laws apply under physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.42;T13.3, 13.5, 13.6;R:AR, LC] (L: CT6, HD5, HD26, HD21, HD4.)
 Definition 13.1.
@@
 Theorem 13.3 (Redistribution Identity).
-$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD4)
+$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD5)
 Proof.
@@
 $\text{TotalExternalWork}=n\times\text{SimplicityTax}(M,P).$
-(L: HD21)
+(L: HD26)
 Proof.
-This is by definition of per-site externalization and is mechanized as HD12.  $\blacksquare$
+This is by definition of per-site externalization and is mechanized as HD26.  $\blacksquare$
 Theorem 13.6 (Amortization).
@@
 $n^{*}=\left\lfloor\frac{H_{\text{central}}}{\text{SimplicityTax}(M,P)}\right\rfloor$
-such that for $n>n^{*}$, the complete model has lower total cost. (L: HD3)
+such that for $n>n^{*}$, the complete model has lower total cost. (L: HD4)
 Proof.
@@
 Corollary 13.7 (Gap Externalization).
-[C:C9.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD21, HD18)
+[C:C13.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD26, HD21)
 Proof.
-The exact linear form is HD12. When the gap is nonempty (positive tax), monotone growth with $n$ is HD11.  $\blacksquare$
+The exact linear form is HD26. When the gap is nonempty (positive tax), monotone growth with $n$ is HD21.  $\blacksquare$
 Corollary 13.8 (Exact Minimization Criterion).
-[C:C9.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
+[C:C13.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
 Proof.
 This is S2P2.  $\blacksquare$
+Informally: when exact minimization is hard, unresolved relevance shifts between centralized structure and repeated local work.
 Appendix A provides theorem statements and module paths for the corresponding Lean formalization.
@@
 Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [17, 80, 41].
-Our object differs: we classify coordinate sufficiency for optimal-action invariance, with mechanized reductions and regime typing.
+Our object differs: we classify coordinate sufficiency for optimal-action invariance across static/stochastic/sequential regimes, with theorem-level placements and mechanized reductions.
 Relation to prior hardness results.
@@
 The target predicate here is decision relevance, not predictive compression.
+Functional information and selection.
+Prior work defines function-relative information objects in physical systems.
+Szostak defines functional information as the negative log of the fraction of configurations achieving a specified function [87].
+Wong et al.
+study function and selection in evolving systems and state an increasing-functional-information law under selection [93].
+This paper addresses a complementary object: the computational decision predicate for when projected coordinates are sufficient to preserve optimal-action correspondence, with explicit complexity placements and mechanized proofs.
 14.3 Succinct Representations and Regime Separation
@@
 In this paper’s formal model, the corresponding core object is coordinate-projection invariance of the optimizer map: sufficiency asks whether fixing coordinates preserves Opt, while insufficiency is witnessed by agreeing projections with different optimal sets (Definition 2.4, Proposition 2.6).[C:D2.4;P2.6;R:DM](L: DP7, DP8.)
-Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.34;C2.35;R:AR](L: CT6,CT5,CT8,CT7.)
+Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.42;C2.43;R:AR](L: CT6, CT5, CT8, CT7.)
 Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 9.1); finite query access already exhibits indistinguishability barriers (Proposition 9.3).
-The complexity separation is mechanized.[C:T5.1;P5.3;R:E,S+ETH,Q_fin](L: CC16,CC23,CC24,CC50.)
-The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DQ3, L25, L68, DP21.)
-The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: DQ519, DQ518, DQ517, DQ155, DQ154, DQ159, DQ160.)
+The complexity separation is mechanized.[C:T9.1;P9.3;R:E, S+ETH, Q_fin](L: CC16, CC23, CC24, CC50.)
+The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DG5, DG6, DG15, DG16, DG18.)
+The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: IN4, IN5, IN6, IN13, IN14, IN15, IN16.)
 Chemical reaction networks provide a concrete physical class where this encoding applies directly.
@@
 Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact.
-The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 7.2, 4.6, 6.1).[C:T7.2,4.6,6.1;R:E,S,TA](L: CR1,CC61,CC69,CC70,CC71,CC72.)
-The physical bridge is not a single-domain add-on: Theorem 2.37 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.37;R:AR](L: CT4.)
-For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.36;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
-The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.28,2.29,2.30,2.31;R:AR](L: DQ548,DQ537,DQ539,DQ534,DQ541,DQ549.)
+The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 11.2, 4.6, 10.1).[C:T11.2, 4.6, 10.1;R:E, S, TA](L: CR1, CC61, CC69, CC70, CC71, CC72.)
+The physical bridge is not a single-domain add-on: Theorem 2.45 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.45;R:AR](L: CT4.)
+For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.44;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
+The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.34, 2.35, 2.36, 2.37;R:AR](L: DT6, DT8, DT10, DT13, DT16, DT24.)
 Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements.
-The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws rather than informal metaphors.[C:P2.34;T2.37;R:AR](L: PC10, PC12, PC15, PC17, PC19, PC20, PC21.)
+The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws.[C:P2.42;T2.45;R:AR](L: ARG2, ARG3, ARG6, ARG13, ARG17, ARG19, ARG12.)
 14.6 Thermodynamic Costs and Landauer’s Principle
-The thermodynamic lift in Section 5 converts declared bit-operation lower bounds into declared energy/carbon lower bounds via linear conversion constants.
-In this artifact, those conversion consequences are mechanized conditionally through Propositions 5.28, 5.29, 5.30, and 5.31.
-Landauer’s principle and reversible-computation context [33, 7] define the physical assumption set; the constants are declared premises, not Lean-derived constants.[C:P5.28,5.29,5.30,5.31;R:AR](L: CC65,CC66,CC67,CC68,CC64.)
-Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the declared regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60).
+The thermodynamic lift in Section 9 converts bit-operation lower bounds into energy/carbon lower bounds via linear conversion constants.
+In this artifact, those conversion consequences are mechanized with explicit premises through Propositions 9.28, 9.29, 9.30, and 9.31.
+Landauer’s principle and reversible-computation context [44, 8] define the physical assumption set; the constants are premises, not Lean-derived constants.[C:P9.28, 9.29, 9.30, 9.31;R:AR](L: CC65, CC66, CC67, CC68, CC64.)
+Physics assumptions decompose into regime tags.
+The paper does not create domain-specific tags for physics assumptions.
+Instead, every physics assumption decomposes into the same regime tags used throughout the paper.
+This uniformity states that physics assumptions are not epistemically privileged; they are declared assumptions, tracked with the same discipline as the paper’s complexity-theoretic assumptions.
+The explicit physical-to-core mapping does not make the results less physical.
+It standardizes premise exposure and transfer obligations: accepted physics assumptions and this model are handled under the same declared-assumption framework, then checked with the same mechanized proof discipline where formalized.
+Table 1: Physics-Assumption-to-Regime-Tag Decomposition
+Physics Assumption Regime Decomposition Justification
+Landauer bound ($E\geq kT\ln 2$) [AR], [H=cost-growth] Declared law; yields cost lower bounds
+Lorentz invariance [AR] Declared physical symmetry
+Discrete state space [Q_fin], [AR] Finite-state core; declared assumption
+Thermodynamic constants ($\lambda,\rho$) [AR], [H=cost-growth] Declared conversion model
+Planck scale ($\ell_{P}$) [AR] Declared physical constant
+Physical-to-core encoding ($E$) [AR] Declared encoding map
+The decomposition in Table 1 makes explicit what the paper already practices: physics assumptions are declared premises in the [AR] (axiomatic regime) sense, and when they yield cost lower bounds, they inherit [H=cost-growth].
+The finite-state query core [Q_fin] captures the discrete-state assumption underlying the computational model.
+No physics assumption receives a special domain tag; all decompose into the existing regime taxonomy.
+This is a standardization claim about proof form, not a change in physical content.[C:P9.28, 9.52, 9.34, 9.33;T2.45;R:AR, H=cost-growth, Q_fin]
+Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the active regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60).
 A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61, 3.60;R:AR](L: IC26, IC24, CC30, IC27, IC7, IC6.)
+14.7 Neukart–Vinokur Thermodynamic Duality
+Neukart and Vinokur [56] propose a thermodynamic duality $dU=T\,dS-p\,dV+\lambda\,dC$ where $C$ is a “complexity coordinate” associated with computational irreversibility in spin-glass systems.
+Corollary 9.32 derives their duality as a specialization of our thermodynamic lift (Proposition 9.28) for $\lambda>0$.
+[C:C9.32;P9.28;R:AR] (L: CC65, CC50.)
+Their spin-glass phase transitions become regime-specific instances: our [S+ETH] hard families with $k^{*}=n$ witness their “complexity potential” via explicit TAUTOLOGY reduction (Theorem 9.1).
+Physical transport via $E:\mathcal{P}\to\mathcal{D}$ extends to arbitrary substrates, not only quantum or spin systems (Theorem 2.45).
+[C:T9.1, 2.45;R:S+ETH, AR] (L: CC16, CT4.)
+Thus Neukart–Vinokur $\subseteq$ [AR] regime: $\lambda,\rho>0$, explicit encoding $E:\mathcal{P}\to\mathcal{D}$ $\Rightarrow$ their thermodynamic duality holds by assumption transfer through our coNP-complete core (Proposition 2.42).
+[C:P2.42;R:AR] (L: CT6.)
 14.8 Feature Selection
@@
 This is known to be NP-hard in general [11, 33].
-Our results show the decision-theoretic analog is coNP-complete for both checking and minimization.
-[C:T4.6,4.7;R:S](L: CC61,CC36.)
+Our results place the decision-theoretic analog in a regime map: static checking/minimization at coNP-complete, stochastic sufficiency at PP-complete, and sequential sufficiency at PSPACE-complete.
+[C:T4.6, 4.7, 6.7, 7.7;R:S, RG](L: CC61, CC36, DC9, DC10.)
 14.9 Value of Information
@@
 Our results formalize the regime-level reason heuristic selection appears: without added structural assumptions, exact optimal model selection inherits worst-case intractability, so heuristic methods implement explicit weakened-guarantee policies for unresolved structure.
-[C:T4.6,5.1,6.1;R:E,S,S+ETH](L: CC61,CC16,CC69,CC70,CC71.)
+[C:T4.6, 9.1, 10.1;R:E, S, S+ETH](L: CC61, CC16, CC69, CC70, CC71.)
 14.11 Information Compression vs Certification Length
@@
 This paper composes all three for the same decision-relevance object in one regime-typed and machine-checked framework.
-[C:T5.1;P5.3,3.16;R:S+ETH,Qf,AR](L: CC16,CC50,CC75.)
+[C:T9.1;P9.3, 3.16;R:S+ETH, Qf, AR](L: CC16, CC50, CC75.)
 15 Conclusion
+Formally: this section composes proved regime placements, bridge theorems, and claim-typing constraints into final landscape consequences.
+This paper gives a regime-first complexity map: static/stochastic/sequential placements with strict hierarchy (DC9, DC10, DC1, DC2), static inner-class decomposition by decision/access/structure layers (CCC6, CCC5, CCC1, CCC3, CCC7, CC61), theorem-level physical transport plus exact represented-adjacent bridge boundary (CT4, CC67, CC69, CC85), Bayesian uniqueness under DQ/thermodynamic admissibility axioms (FN14), physical-collapse no-go transfer chain (PH11, PH12, PH13, PH14, PH15), derived SAT-bridge map plus reduction-energy transfer and disjunctive assumption boundary (PH16, PH17, PH18, PH19, PH20, PH21, PH22, PH26, PH27), explicit finite-budget-necessity countermodels (PH31, PH32, PH33), regime-robust deterministic/probabilistic/sequential no-collapse specializations (PH28, PH29, PH30), observer-collapse anchor closure across stochastic/sequential checks (PA1, PA2, PA3, PA4, PA5, PA6, PA7, PA8, PA9), recurrent conversation-layer clamp/report bridge (CV4, CV7, CV17, CV18, CV13, CV16), explicit Neukart–Vinokur duality specialization from the thermodynamic lift (CC77, CC76), and formal centralized/distributed hardness thresholds through dominance and amortization theorems (HD1, HD4, HD3).
+[C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1, 10.1, 2.45, 2.54, 2.46, 3.4, 9.50, 12.19, 13.3, 13.6;P9.3, 9.28, 8.1, 8.2;C9.32;R:E, S, S+ETH, Qf, Qb, AR, RG, LC, H=cost-growth]
+Informally: this conclusion combines previously proved blocks under explicit tags.
 Methodology and Disclosure
@@
 The complexity-theoretic reductions required careful human oversight to ensure the polynomial bounds were correctly established.
-What the author contributed: The problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the hardness conjectures, the tractability conditions, and the connection to over-modeling in engineering practice.
+What the author contributed: The informal problem formulations (SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, ANCHOR-SUFFICIENCY), the informal hardness conjectures, the informal tractability conditions, and the connection to over-modeling in engineering practice.
 What LLMs contributed: drafting, Lean tactic exploration, reduction construction assistance, and prose refinement.
 The proofs are machine-checked; their validity is independent of generation method. This methodology is disclosed for academic transparency.
+Acknowledgments
+This work depends on results whose authors made possible the insights that could then be machine-verified.
+Fundamental physics.
+Albert Einstein established special relativity and the light-speed bound [25]; Hermann Minkowski gave it the spacetime formulation [52].
+Max Planck established energy quantization and the constant $k_{B}$ [68].
+Rolf Landauer proved that erasure costs thermodynamic work [44]; Charles Bennett showed that reversible computation circumvents this bound [7].
+Bérut et al.
+experimentally verified Landauer’s principle [9].
+The counting gap theorem (BA10) proves that some finite $c$ must exist from pure mathematics; Einstein told us which one.
+Proof infrastructure.
+The Lean 4 proof assistant [22] served as the arbiter of correctness.
+Leonardo de Moura, Sebastian Ullrich, and the Lean 4 team built a tool that transforms mathematical claims into mechanically verified theorems.
+The Mathlib library [88] provided the formal infrastructure: order theory, linear algebra, complexity-theoretic definitions, and the lemmas that made the formalization possible.
+Methodological note.
+The proofs in this paper are machine-checked; their validity is independent of generation method.
+The human author contributed the informal problem formulations, informal hardness conjectures, informal tractability conditions, and the connection to over-modeling in engineering practice.
+The machine contributors (Claude, GPT-4, Lean 4) served as proof assistants, tactic explorers, and drafting partners.
+The final arbiter was the Lean compiler: if it accepts the proof, the theorem holds.
 Summary of Results
-This paper establishes the computational complexity of coordinate sufficiency problems:
-  • SUFFICIENCY-CHECK is coNP-complete (Theorem 4.6)
-  • MINIMUM-SUFFICIENT-SET is coNP-complete (Theorem 4.7)
-  • ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorem 4.9)
-  • An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with a succinct worst-case ETH lower bound witnessed by a hard family with $k^{*}=n$ (Theorem 5.1)
-  • Full intermediate query-access lower bounds are formalized as a finite-state Opt-oracle core ($\Omega(|S|)$, Boolean instantiation $\Omega(2^{n})$) plus Boolean-interface refinements for value-entry and state-batch access, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness closures (Propositions 5.35.16)
-  • Tractable subcases exist for explicit-state encoding, separable utility, and tree-structured utility with explicit local factors (Theorem 6.1)
+This paper establishes a complete regime map for coordinate sufficiency and a static inner decomposition:
+  • Regime map (core classes).
+Static, stochastic, and sequential sufficiency-check questions are placed at coNP, $P$, and PSPACErespectively, with strict static $\subset$ stochastic $\subset$ sequential separation under declared assumptions (Theorems 4.6, 6.7, 7.7; Propositions 8.1, 8.2).
+  • Observer-collapse anchor closure.
+Under declared observer-collapse hypotheses, stochastic/sequential anchor checks are forced by effective-state singleton structure and seeded support conditions (PA1, PA4, PA5, PA6, PA7, PA9).
+  • Static decision layer decomposition.
+Within the static class, SUFFICIENCY-CHECK and MINIMUM-SUFFICIENT-SET are coNP-complete, while ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete (Theorems 4.6, 4.7, 4.9).
+  • Static access layer decomposition.
+An encoding-regime separation contrasts explicit-state polynomial-time (polynomial in $|S|$) with succinct ETH worst-case lower bounds, and query-access obstruction cores refine this with value-entry/state-batch/oracle-lattice transfers (Theorem 9.1; Propositions 9.39.16).
+  • Static structure layer decomposition. Six structural subclasses provide certified polynomial recovery conditions inside the static class (Theorem 10.1).
   • Discrete-time interface semantics are formalized: time is $\mathbb{N}$-valued, decision events are exactly one-tick transitions, run length equals elapsed time, and substrate-consistent realizations preserve the same one-unit event law (Propositions 2.342.37)
   • Physical-system transport is formalized: encoded physical claims are derived from core rules through explicit assumption transfer, physical counterexamples map to encoded core failures, and the bundled bridge theorem composes law semantics, reduction equivalence, and required-work lower bounds (Proposition 2.42, Corollary 2.43, Theorem 2.45)
-[C:T4.6,4.7,4.9,5.1,6.1,2.37;P5.3,5.7,5.11,5.8,5.15,5.16,2.28,2.29,2.30,2.31,2.34;C2.35;R:E,S,S+ETH,Qf,Qb,AR] (L: CC61,CC36,CC5,CC16,CC69,CC50,L9,L8,L11,L21,L20,DQ548,DQ537,DQ539,DQ541,CT6,CT8,CT4.)
-These results place the problem of identifying decision-relevant coordinates at the first and second levels of the polynomial hierarchy.
+  • Exact represented-adjacent bridge boundary: transfer is licensed iff the class is one-step deterministic; represented horizon, stochastic-criterion, and transition-coupled classes have explicit bridge-failure witnesses (Theorem 2.54)
+  • Exact-report admissibility is characterized by a certified-bit gap criterion; when exact reporting is blocked, exact accounting collapses to raw-only (Theorem 3.47, Corollary 3.49)
+  • The thermodynamic lift yields the Neukart–Vinokur duality specialization $dU\geq\lambda\,dC$ for $\lambda>0$ (Corollary 9.32, Proposition 9.28)
+[C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1, 10.1, 2.45, 2.54, 3.47;P8.1, 8.2, 9.3, 9.7, 9.11, 9.8, 9.15, 9.16, 2.34, 2.35, 2.36, 2.37, 2.42, 9.28, 3.46;C2.43, 3.49, 9.32;R:E, S, S+ETH, Qf, Qb, AR, RG, H=cost-growth] (L: CCC6, CCC5, CCC1, DC9, DC10, DC1, DC2, CCC3, CCC7, CC61, CC29, CC28, CC43, CC52, CC27, DT6, DT13, DT16, DT24, CT6, CT8, CT4, CC67, CC69, CC85, CC18, CC11, CC21, CC77, CC76.)
+[…]
 Beyond classification, the paper contributes a formal claim-typing framework (Section 3): structural complexity is a property of the fixed decision question, while representational hardness is regime-conditional access cost.
@@
 The reduction constructions and key equivalence theorems are machine-checked in Lean 4 (see Appendix A for proof listings).
-The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets; complexity classifications (coNP-completeness, $\Sigma_{2}^{\textsf{P}}$-completeness) follow by composition with standard complexity-theoretic results (TAUTOLOGY is coNP-complete, $\exists\forall$-SAT is $\Sigma_{2}^{\textsf{P}}$-complete).
+The formalization verifies TAUTOLOGY and $\exists\forall$-SAT static reductions, theorem-level stochastic/sequential class placements, and strict regime-separation claims; complexity classifications (coNP, $\Sigma_{2}^{\textsf{P}}$, PP, PSPACE) follow by composition with standard complexity-theoretic results under the declared assumptions.
 The strengthened gadget showing that non-tautologies yield instances with all coordinates relevant is also formalized.
-[C:T4.6,4.7,4.9,5.1;R:S,S+ETH] (L: CC61,CC36,CC5,CC16,CC24.)
+[C:T4.6, 4.7, 4.9, 6.7, 7.7, 9.1;P8.1, 8.2;R:S, S+ETH, RG] (L: CCC6, CCC5, CCC1, DC9, DC10, CCC3, DC1, DC2.)
 Complexity Characterization
@@
 Under [Q_fin], the Opt-oracle core has $\Omega(|S|)$ worst-case query complexity (Boolean instantiation $\Omega(2^{n})$), and under [Q_bool] the value-entry/state-batch refinements preserve the obstruction with weighted-cost transfer closures.
-[C:T4.6,5.1;P5.3,5.7,5.11,5.8;R:E,S,S+ETH,Qf,Qb] (L: CC61,CC16,CC50,L9,L8,L11.)
+[C:T4.6, 9.1;P9.3, 9.7, 9.11, 9.8;R:E, S, S+ETH, Qf, Qb] (L: CC61, CC16, CC50, CC29, CC28, CC27.)
 The Complexity Redistribution Corollary
 Section 13 develops a quantitative consequence: when a problem requires $k$ dimensions and a model handles only $j<k$ natively, the remaining kjk-j dimensions must be handled externally at each decision site.
-For $n$ sites, total external work is $(k-j)\times n$.[C:T9.5;R:LC] (L: HD21.)
+For $n$ sites, total external work is $(k-j)\times n$.[C:T13.5;R:LC] (L: HD26.)
 The set identity is elementary; its operational content comes from composition with the hardness results on exact relevance minimization.
@@
   • Amortization: there exists a threshold $n^{*}$ beyond which higher-dimensional models have lower total cost
-[C:T9.3,8.16,9.6;R:LC] (L: HD4,HD1,HD3,HD21.)
+[C:T13.3, 12.19, 13.6;R:LC] (L: HD5, HD1, HD4, HD26, HD21.)
 Open Questions
-The landscape above is complete for the static sufficiency class under C1–C4 and the declared regimes; the items below are adjacent-class extensions or secondary refinements.
-Several questions remain for future work:
+The landscape above is complete for the static sufficiency class under C1–C4; the items below are adjacent-class extensions or secondary refinements. Several questions remain for future work:
   • Fixed-parameter tractability (primary): Is SUFFICIENCY-CHECK FPT when parameterized by the minimal sufficient-set size $k^{*}$, or is it W[2]-hard under this parameterization?
@@
   • Redistribution consequences: omitted native coverage externalizes work with explicit growth/amortization laws (Theorems 13.313.6).
-  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.34;C2.35;T2.37;R:AR](L: CT6,CT8,CT4.)
-Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T6.1,5.1;R:RG](L: CC71,CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
-Epistemic Alignment Implications for Agent Training
-Within this paper’s formalism, epistemic alignment is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence).
-Formally, this is the integrity/competence split of Section 3, not a full value-alignment claim.[C:D3.8,3.9;P3.58;R:TR](L: IC18.)
+  • Physical-system instantiation: once a physical class is encoded into the declared decision interface, the same regime-typed hardness/tractability map and integrity constraints apply to that class (Section 2.7).[C:P2.42;C2.43;T2.45;R:AR](L: CT6, CT8, CT4.)
+Physical Scope Guard.
+All transfer claims in this paper are typed through explicit physical encodings, declared observables, and declared assumption maps.
+Transfer is admitted only on declared physical instances satisfying the physical scope gate.[C:D2.38, 2.39;P2.40, 2.42;R:AR]
+Hence the design choice is typed: enforce a tractable regime, or adopt weakened guarantees with explicit verification boundaries.[C:T10.1, 9.1;R:RG](L: CC71, CC16.) Equivalently, under the attempted-competence matrix, over-specification is rational only in attempted-competence-failure cells; once exact competence is available in the active regime (or no attempted exact competence was made), persistent over-specification is irrational for the same verified-cost objective.[C:P3.60;R:AR](L: IC27.) By Proposition 3.60, the integrity-preserving matrix contains 3 irrational cells and 1 rational cell.[C:P3.60;R:AR](L: IC27.)
+Measurement-Constrained Reporting for Physical Solvers
+Within this paper’s formalism, reporting correctness is a claim-discipline property: outputs are asserted only when certifiable, and otherwise emitted as abstain reports (optionally carrying tentative guesses and self-reflected confidence under zero certified confidence).
+Formally, this is the integrity/competence split of Section 3, i.e., an evidence-typing statement for reported claims under declared regimes.[C:D3.8, 3.9;P3.58;R:TR](L: IC18.)
+Formally: the abstention doctrine is a typed-report consequence of integrity/competence and evidence-gating predicates.
 Truth-Objective Abstention Requirement.
-If a solver-agent declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers.
-Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19,3.20;P3.23,3.29;C3.25;R:AR](L: IC46,IC38,IC41.) Equivalently, training policies that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58,3.61,3.60;R:AR](L: IC18,CC30,IC27.)
-This yields a training doctrine for solver-agents under declared tasks/regimes:
+If a physical solver system declares a truth-preserving objective for claims about relation $\mathcal{R}$ in regime $\Gamma$, uncertified outputs are admissible only as $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ rather than asserted answers.
+Under the signal discipline, this includes abstain-with-guess/self-confidence reports while keeping certification evidence-gated.[C:D3.19, 3.20;P3.23, 3.29;C3.25;R:AR](L: IC46, IC38, IC41.) Equivalently, reporting protocols that suppress abstention in uncertified regions violate the typed claim discipline under the declared task model.[C:D3.8;P3.58, 3.61, 3.60;R:AR](L: IC18, CC30, IC27.)
+This yields an output-control doctrine for engineered solver systems under the task/regime contract:
   1. 1.
@@
   5. 5.
-    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T5.1;P3.61,3.60;R:S,S+ETH](L: CC16,CC30,IC27.)
-Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the declared scope, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are declared).
-This is consistent with bounded-rationality feasibility discipline under declared resource models [71, 5].
+    In hardness-blocked regimes, abstention or explicit guarantee-weakening is structurally required unless regime assumptions are changed.[C:T9.1;P3.61, 3.60;R:S, S+ETH](L: CC16, CC30, IC27.)
+Operationally, a reporting schema consistent with this framework tracks at least five quantities: integrity violations, competent non-abstaining coverage on the scoped family, abstention quality under regime shift, the calibration gap between self-reflected and certified confidence channels, and scalar runtime witness distributions (with bound-qualified fraction semantics when bounds are given).
+This is consistent with bounded-rationality feasibility discipline under bounded resource models [92, 5].
+Informally: if exact support is not certified, do not make exact claims; when competence is unavailable, abstain and keep confidence evidence-linked.
 Appendix A Lean 4 Proof Listings
 The complete Lean 4 formalization is available in the companion artifact (Zenodo DOI listed on the title page).
-The mechanization consists of 13391 lines across 58 files, with 597 theorem/lemma statements.
+The mechanization consists of 28863 lines across 113 files, with 1252 theorem/lemma statements.
 Handle IDs. Inline theorem metadata now cites compact IDs (for example, HD6, CC12, IC4) instead of full theorem constants. The full ID-to-handle mapping is listed in Section A.1.
 A.1 Lean Handle ID Map
-ID Full Lean handle
+Lean handle entry
+AC1DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC1
+AC3DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC3
+AC4DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC4
+AC5DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC5
+AC6DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC6
+AC8DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC8
+AC9DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC9
+AC11DecisionQuotient.ClaimClosure.AtomicCircuitExports.AC11
+AN1DecisionQuotient.Physics.AssumptionNecessity.no_assumption_free_proof_of_refutable_claim
+AN2DecisionQuotient.Physics.AssumptionNecessity.countermodel_violates_some_assumption
+AN3DecisionQuotient.Physics.AssumptionNecessity.physical_claim_requires_physical_assumption
+AN4DecisionQuotient.Physics.AssumptionNecessity.physical_claim_requires_empirically_justified_physical_assumption
+AN5DecisionQuotient.Physics.AssumptionNecessity.strong_physical_no_go_meta
+AQ1DecisionQuotient.ClaimClosure.AQ1
+AQ2DecisionQuotient.ClaimClosure.AQ2
+AQ3DecisionQuotient.ClaimClosure.AQ3
+AQ4DecisionQuotient.ClaimClosure.AQ4
+AQ5DecisionQuotient.ClaimClosure.AQ5
+AQ6DecisionQuotient.ClaimClosure.AQ6
+AQ7DecisionQuotient.ClaimClosure.AQ7
+AQ8DecisionQuotient.ClaimClosure.AQ8
+ARG1PhysicalComplexity.AccessRegime.PhysicalDevice
+ARG2PhysicalComplexity.AccessRegime.AccessRegime
+ARG3PhysicalComplexity.AccessRegime.RegimeEval
+ARG4PhysicalComplexity.AccessRegime.RegimeSample
+ARG5PhysicalComplexity.AccessRegime.RegimeProof
+ARG6PhysicalComplexity.AccessRegime.RegimeWithCertificate
+ARG7PhysicalComplexity.AccessRegime.RegimeEvalOn
+ARG8PhysicalComplexity.AccessRegime.RegimeSampleOn
+ARG9PhysicalComplexity.AccessRegime.RegimeProofOn
+ARG10PhysicalComplexity.AccessRegime.RegimeWithCertificateOn
+ARG11PhysicalComplexity.AccessRegime.HardUnderEval
+ARG12PhysicalComplexity.AccessRegime.AuditableWithCertificate
+ARG13PhysicalComplexity.AccessRegime.certificate_upgrades_regime
+ARG14PhysicalComplexity.AccessRegime.certificate_upgrades_regime_on
+ARG15PhysicalComplexity.AccessRegime.physical_succinct_certification_hard
+ARG16PhysicalComplexity.AccessRegime.certificate_amortizes_hardness
+ARG17PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate
+ARG18PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate_on
+ARG19PhysicalComplexity.AccessRegime.AccessChannelLaw
+ARG20PhysicalComplexity.AccessRegime.FiveWayMeet
+BA1DecisionQuotient.Physics.BoundedAcquisition.BoundedRegion
+BA2DecisionQuotient.Physics.BoundedAcquisition.acquisition_rate_bound
+BA3DecisionQuotient.Physics.BoundedAcquisition.acquisitions_are_transitions
+BA4DecisionQuotient.Physics.BoundedAcquisition.one_bit_per_transition
+BA5DecisionQuotient.Physics.BoundedAcquisition.resolution_reads_sufficient
+BA6DecisionQuotient.Physics.BoundedAcquisition.srank_le_resolution_bits
+BA7DecisionQuotient.Physics.BoundedAcquisition.energy_ge_srank_cost
+BA8DecisionQuotient.Physics.BoundedAcquisition.srank_one_energy_minimum
+BA9DecisionQuotient.Physics.BoundedAcquisition.physical_grounding_bundle
+BA10DecisionQuotient.Physics.BoundedAcquisition.counting_gap_theorem
+BB1DecisionQuotient.BayesianDQ
+BB2DecisionQuotient.BayesianDQ.certaintyGain
+BB3DecisionQuotient.dq_is_bayesian_certainty_fraction
+BB4DecisionQuotient.bayesian_dq_matches_physics_dq
+BB5DecisionQuotient.dq_derived_from_bayes
+BC1DecisionQuotient.Foundations.counting_nonneg
+BC2DecisionQuotient.Foundations.counting_total
+BC3DecisionQuotient.Foundations.counting_additive
+BC4DecisionQuotient.Foundations.bayes_from_conditional
+BC5DecisionQuotient.Foundations.entropy_contraction
+BF1DecisionQuotient.certainty_of_not_nondegenerateBelief
+BF2DecisionQuotient.nondegenerateBelief_of_uncertaintyForced
+BF3DecisionQuotient.forced_action_under_uncertainty
+BF4DecisionQuotient.bayes_update_exists_of_nondegenerateBelief
 CC1DecisionQuotient.ClaimClosure.RegimeSimulation
 CC2DecisionQuotient.ClaimClosure.adq_ordering
-CC3 DecisionQuotient.ClaimClosure.agent_transfer_licensed_iff_snapshot
+CC3DecisionQuotient.ClaimClosure.system_transfer_licensed_iff_snapshot
 CC4DecisionQuotient.ClaimClosure.anchor_sigma2p_complete_conditional
 CC5DecisionQuotient.ClaimClosure.anchor_sigma2p_reduction_core
-CC6 DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
-CC7 DecisionQuotient.ClaimClosure.bounded_actions_detectable
-CC8 DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
-CC9 DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
-CC10 DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
-CC11 DecisionQuotient.ClaimClosure.certified_total_bits_split_core
-CC12 DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
-CC13 DecisionQuotient.ClaimClosure.declaredBudgetSlice
-CC14 DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
-CC15 DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
-CC16 DecisionQuotient.ClaimClosure.dichotomy_conditional
-CC17 DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
-CC18 DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
-CC19 DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
-CC20 DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
-CC21 DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
-CC22 DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
-CC23 DecisionQuotient.ClaimClosure.explicit_state_upper_core
-CC24 DecisionQuotient.ClaimClosure.hard_family_all_coords_core
-CC25 DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
-CC26 DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
-CC27 DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
-CC28 DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
-CC29 DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
-CC30 DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
-CC31 DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
-CC32 DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
-CC33 DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
-CC34 DecisionQuotient.ClaimClosure.minsuff_collapse_core
-CC36 DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
-CC37 DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
-CC38 DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
-CC39 DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
-CC40 DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
-CC41 DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
-CC42 DecisionQuotient.ClaimClosure.one_step_bridge
-CC43 DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
-CC44 DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
-CC45 DecisionQuotient.ClaimClosure.physical_crossover_core
-CC46 DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
-CC47 DecisionQuotient.ClaimClosure.physical_crossover_policy_core
-CC48 DecisionQuotient.ClaimClosure.process_bridge_failure_witness
-CC49 DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
-CC50 DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
-CC51 DecisionQuotient.ClaimClosure.regime_core_claim_proved
-CC52 DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
-CC53 DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
-CC54 DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
-CC55 DecisionQuotient.ClaimClosure.separable_detectable
-CC56 DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
-CC57 DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
-CC58 DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
-CC59 DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
-CC60 DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
-CC61 DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
-CC62 DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
-CC63 DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
-CC64 DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
-CC65 DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
-CC66 DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
-CC67 DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
-CC68 DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
-CC69 DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
-CC70 DecisionQuotient.ClaimClosure.tractable_bounded_core
-CC71 DecisionQuotient.ClaimClosure.tractable_separable_core
-CC72 DecisionQuotient.ClaimClosure.tractable_subcases_conditional
-CC73 DecisionQuotient.ClaimClosure.tractable_tree_core
-CC74 DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
-CC75 DecisionQuotient.ClaimClosure.tree_structure_detectable
-CC76 DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
-CC77 DecisionQuotient.ClaimClosure.typed_static_class_completeness
+CC6DecisionQuotient.ClaimClosure.anchor_query_relation_false_iff
+CC7DecisionQuotient.ClaimClosure.anchor_query_relation_true_iff
+CC8DecisionQuotient.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
+CC9DecisionQuotient.ClaimClosure.bounded_actions_detectable
+CC10DecisionQuotient.ClaimClosure.bridge_boundary_represented_family
+CC11DecisionQuotient.ClaimClosure.bridge_failure_witness_non_one_step
+CC12DecisionQuotient.ClaimClosure.bridge_transfer_iff_one_step_class
+CC13DecisionQuotient.ClaimClosure.certified_total_bits_split_core
+CC14DecisionQuotient.ClaimClosure.cost_asymmetry_eth_conditional
+CC15DecisionQuotient.ClaimClosure.declaredBudgetSlice
+CC16DecisionQuotient.ClaimClosure.declaredRegimeFamily_complete
+CC17DecisionQuotient.ClaimClosure.declared_physics_no_universal_exact_certifier_core
+CC18DecisionQuotient.ClaimClosure.dichotomy_conditional
+CC19DecisionQuotient.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
+CC20DecisionQuotient.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core
+CC21DecisionQuotient.ClaimClosure.exact_certainty_inflation_under_hardness_core
+CC22DecisionQuotient.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
+CC23DecisionQuotient.ClaimClosure.exact_raw_only_of_no_exact_admissible_core
+CC24DecisionQuotient.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+CC25DecisionQuotient.ClaimClosure.explicit_state_upper_core
+CC26DecisionQuotient.ClaimClosure.hard_family_all_coords_core
+CC27DecisionQuotient.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient
+CC28DecisionQuotient.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
+CC29DecisionQuotient.ClaimClosure.information_barrier_opt_oracle_core
+CC30DecisionQuotient.ClaimClosure.information_barrier_state_batch_core
+CC31DecisionQuotient.ClaimClosure.information_barrier_value_entry_core
+CC32DecisionQuotient.ClaimClosure.integrity_resource_bound_for_sufficiency
+CC33DecisionQuotient.ClaimClosure.integrity_universal_applicability_core
+CC34DecisionQuotient.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice
+CC35DecisionQuotient.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
+CC36DecisionQuotient.ClaimClosure.minsuff_collapse_core
+CC37DecisionQuotient.ClaimClosure.minsuff_collapse_to_conp_conditional
+CC38DecisionQuotient.ClaimClosure.minsuff_conp_complete_conditional
+CC39DecisionQuotient.ClaimClosure.no_auto_minimize_of_p_neq_conp
+CC40DecisionQuotient.ClaimClosure.no_exact_claim_admissible_under_hardness_core
+CC41DecisionQuotient.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
+CC42DecisionQuotient.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
+CC43DecisionQuotient.ClaimClosure.no_uncertified_exact_claim_core
+CC44DecisionQuotient.ClaimClosure.one_step_bridge
+CC45DecisionQuotient.ClaimClosure.oracle_lattice_transfer_as_regime_simulation
+CC46DecisionQuotient.ClaimClosure.physical_crossover_above_cap_core
+CC47DecisionQuotient.ClaimClosure.physical_crossover_core
+CC48DecisionQuotient.ClaimClosure.physical_crossover_hardness_core
+CC49DecisionQuotient.ClaimClosure.physical_crossover_policy_core
+CC50DecisionQuotient.ClaimClosure.process_bridge_failure_witness
+CC51DecisionQuotient.ClaimClosure.poseAnchorQuery
+CC52DecisionQuotient.ClaimClosure.pose_returns_anchor_query_object
+CC53DecisionQuotient.ClaimClosure.posed_anchor_checked_true_implies_truth
+CC54DecisionQuotient.ClaimClosure.posed_anchor_exact_claim_admissible_iff_competent
+CC55DecisionQuotient.ClaimClosure.posed_anchor_exact_claim_requires_evidence
+CC56DecisionQuotient.ClaimClosure.posed_anchor_no_competence_no_exact_claim
+CC57DecisionQuotient.ClaimClosure.posed_anchor_query_truth_iff_exists_anchor
+CC58DecisionQuotient.ClaimClosure.posed_anchor_query_truth_iff_exists_forall
+CC59DecisionQuotient.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
+CC60DecisionQuotient.ClaimClosure.query_obstruction_boolean_corollary
+CC61DecisionQuotient.ClaimClosure.query_obstruction_finite_state_core
+CC62DecisionQuotient.ClaimClosure.regime_core_claim_proved
+CC63DecisionQuotient.ClaimClosure.regime_simulation_transfers_hardness
+CC64DecisionQuotient.ClaimClosure.reusable_heuristic_of_detectable
+CC65DecisionQuotient.ClaimClosure.selectorSufficient_not_implies_setSufficient
+CC66DecisionQuotient.ClaimClosure.separable_detectable
+CC67DecisionQuotient.ClaimClosure.snapshot_vs_process_typed_boundary
+CC68DecisionQuotient.ClaimClosure.standard_assumption_ledger_unpack
+CC69DecisionQuotient.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
+CC70DecisionQuotient.ClaimClosure.subproblem_hardness_lifts_to_full
+CC71DecisionQuotient.ClaimClosure.subproblem_transfer_as_regime_simulation
+CC72DecisionQuotient.ClaimClosure.sufficiency_conp_complete_conditional
+CC73DecisionQuotient.ClaimClosure.sufficiency_conp_reduction_core
+CC74DecisionQuotient.ClaimClosure.sufficiency_iff_dq_ratio
+CC75DecisionQuotient.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+CC76DecisionQuotient.ClaimClosure.thermo_conservation_additive_core
+CC77DecisionQuotient.ClaimClosure.thermo_energy_carbon_lift_core
+CC78DecisionQuotient.ClaimClosure.thermo_eventual_lift_core
+CC79DecisionQuotient.ClaimClosure.thermo_hardness_bundle_core
+CC80DecisionQuotient.ClaimClosure.thermo_mandatory_cost_core
+CC81DecisionQuotient.ClaimClosure.tractable_bounded_core
+CC82DecisionQuotient.ClaimClosure.tractable_separable_core
+CC83DecisionQuotient.ClaimClosure.tractable_subcases_conditional
+CC84DecisionQuotient.ClaimClosure.tractable_tree_core
+CC85DecisionQuotient.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
+CC86DecisionQuotient.ClaimClosure.tree_structure_detectable
+CC87DecisionQuotient.ClaimClosure.typed_claim_admissibility_core
+CC88DecisionQuotient.ClaimClosure.typed_static_class_completeness
+CC89DecisionQuotient.ClaimClosure.universal_solver_framing_core
+CCC1DecisionQuotient.CC.anchor_sigma2p_complete_conditional
+CCC2DecisionQuotient.CC.cost_asymmetry_eth_conditional
+CCC3DecisionQuotient.CC.dichotomy_conditional
+CCC4DecisionQuotient.CC.minsuff_collapse_to_conp_conditional
+CCC5DecisionQuotient.CC.minsuff_conp_complete_conditional
+CCC6DecisionQuotient.CC.sufficiency_conp_complete_conditional
+CCC7DecisionQuotient.CC.tractable_subcases_conditional
+CF1DecisionQuotient.Physics.ConstraintForcing.laws_not_determined_of_parameter_separation
+CF2DecisionQuotient.Physics.ConstraintForcing.logic_time_not_sufficient_for_unique_law
+CF3DecisionQuotient.Physics.ConstraintForcing.laws_determined_implies_objective_determined
+CF4DecisionQuotient.Physics.ConstraintForcing.objective_not_determined_of_parameter_separation
+CF5DecisionQuotient.Physics.ConstraintForcing.forcedDecisionBits_pos_of_deadline
+CF6DecisionQuotient.Physics.ConstraintForcing.actionForced_of_deadline
+CF7DecisionQuotient.Physics.ConstraintForcing.nondegenerateBelief_of_deadline_and_uncertainty
+CF8DecisionQuotient.Physics.ConstraintForcing.forced_decision_implies_positive_landauer_cost
+CF9DecisionQuotient.Physics.ConstraintForcing.forced_decision_implies_positive_nv_work
+CH1DecisionQuotient.ClaimClosure.CH1
+CH2DecisionQuotient.ClaimClosure.CH2
+CH3DecisionQuotient.ClaimClosure.CH3
+CH5DecisionQuotient.ClaimClosure.CH5
+CH6DecisionQuotient.ClaimClosure.CH6
 CR1DecisionQuotient.ConfigReduction.config_sufficiency_iff_behavior_preserving
-CT1 CT.law_gap_physical_claim_holds
-CT2 CT.no_law_gap_counterexample
-CT3 CT.no_physical_counterexample_of_core_theorem
-CT4 CT.physical_bridge_bundle
-CT5 CT.physical_claim_lifts_from_core
-CT6 CT.physical_claim_lifts_from_core_conditional
-CT7 CT.physical_counterexample_invalidates_core_rule
-CT8 CT.physical_counterexample_yields_core_counterexample
+CT1DecisionQuotient.Physics.ClaimTransport.PhysicalEncoding
+CT2DecisionQuotient.Physics.ClaimTransport.physical_claim_lifts_from_core
+CT3DecisionQuotient.Physics.ClaimTransport.physical_claim_lifts_from_core_conditional
+CT4DecisionQuotient.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
+CT5DecisionQuotient.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule
+CT6DecisionQuotient.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem
+CT7DecisionQuotient.Physics.ClaimTransport.LawGapInstance
+CT8DecisionQuotient.Physics.ClaimTransport.lawGapEncoding
+CT9DecisionQuotient.Physics.ClaimTransport.lawGapPhysicalClaim
+CT10DecisionQuotient.Physics.ClaimTransport.law_gap_physical_claim_holds
+CT11DecisionQuotient.Physics.ClaimTransport.no_law_gap_counterexample
+CT12DecisionQuotient.Physics.ClaimTransport.physical_bridge_bundle
+CV1DecisionQuotient.Physics.Conversation.RecurrentCircuit
+CV2DecisionQuotient.Physics.Conversation.CoupledConversation
+CV3DecisionQuotient.Physics.Conversation.JointState
+CV4DecisionQuotient.Physics.Conversation.tick_uses_shared_node
+CV5DecisionQuotient.Physics.Conversation.tick_shared_is_merged_emissions
+CV6DecisionQuotient.Physics.Conversation.channel_projection_eq_iff_quantized_eq
+CV7DecisionQuotient.Physics.Conversation.clamp_projection_eq_iff_same_clamped_bit
+CV10DecisionQuotient.Physics.Conversation.explanationGap_add_explanationBits
+CV11DecisionQuotient.Physics.Conversation.toClaimReport
+CV12DecisionQuotient.Physics.Conversation.abstain_iff_no_answer
+CV13DecisionQuotient.Physics.Conversation.yes_no_iff_exact_claim
+CV14DecisionQuotient.Physics.Conversation.toReportSignal_completion_defined_of_budget
+CV15DecisionQuotient.Physics.Conversation.toReportSignal_signal_consistent_zero_certified
+CV16DecisionQuotient.Physics.Conversation.abstain_report_can_carry_explanation
+CV17DecisionQuotient.Physics.Conversation.clampDecisionEvent_iff_bitOps_pos
+CV18DecisionQuotient.Physics.Conversation.clamp_event_implies_positive_energy
+CV19DecisionQuotient.Physics.Conversation.BinaryAnswer
+CV20DecisionQuotient.Physics.Conversation.ConversationReport
+CV21DecisionQuotient.Physics.Conversation.explanationGap
+CV26DecisionQuotient.Physics.Conversation.toClaimReport_ne_epsilon
+CV27DecisionQuotient.Physics.Conversation.toReportSignal
+CV28DecisionQuotient.Physics.Conversation.toReportSignal_declares_bound
+DC1DecisionQuotient.StochasticSequential.static_stochastic_strict_separation
+DC2DecisionQuotient.StochasticSequential.stochastic_sequential_strict_separation
+DC3DecisionQuotient.StochasticSequential.complexity_dichotomy_hierarchy
+DC4DecisionQuotient.StochasticSequential.regime_hierarchy
+DC5DecisionQuotient.StochasticSequential.coNP_subset_PP
+DC6DecisionQuotient.StochasticSequential.PP_subset_PSPACE
+DC7DecisionQuotient.StochasticSequential.coNP_subset_PSPACE
+DC8DecisionQuotient.StochasticSequential.static_to_coNP
+DC9DecisionQuotient.StochasticSequential.stochastic_to_PP
+DC10DecisionQuotient.StochasticSequential.sequential_to_PSPACE
+DC11DecisionQuotient.StochasticSequential.ClaimClosure.claim_six_subcases
+DC12DecisionQuotient.StochasticSequential.ClaimClosure.claim_hierarchy
+DC13DecisionQuotient.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P
+DC14DecisionQuotient.StochasticSequential.stochastic_dichotomy
+DC15DecisionQuotient.StochasticSequential.above_threshold_hard
+DC16DecisionQuotient.StochasticSequential.StochasticAnchorSufficient
+DC17DecisionQuotient.StochasticSequential.StochasticAnchorSufficiencyCheck
+DC18DecisionQuotient.StochasticSequential.stochastic_anchor_check_iff
+DC19DecisionQuotient.StochasticSequential.stochastic_anchor_sufficient_of_stochastic_sufficient
+DC20DecisionQuotient.StochasticSequential.SequentialAnchorSufficient
+DC21DecisionQuotient.StochasticSequential.SequentialAnchorSufficiencyCheck
+DC22DecisionQuotient.StochasticSequential.sequential_anchor_check_iff
+DC23DecisionQuotient.StochasticSequential.sequential_anchor_sufficient_of_sequential_sufficient
+DC24DecisionQuotient.StochasticSequential.StochasticAnchorCheckInstance
+DC25DecisionQuotient.StochasticSequential.reduceMAJSAT_correct_anchor_strict
+DC26DecisionQuotient.StochasticSequential.reduceMAJSAT_to_stochastic_anchor_reduction
+DC27DecisionQuotient.StochasticSequential.SequentialAnchorCheckInstance
+DC28DecisionQuotient.StochasticSequential.reduceTQBF_correct_anchor
+DC29DecisionQuotient.StochasticSequential.reduceTQBF_to_sequential_anchor_reduction
+DC30DecisionQuotient.StochasticSequential.StatePotential
+DC31DecisionQuotient.StochasticSequential.utilityFromPotentialDrop_le_iff_nextPotential_ge
+DC32DecisionQuotient.StochasticSequential.utility_from_action_state_potential
+DC33DecisionQuotient.StochasticSequential.stochasticExpectedUtility_eq_neg_expectedActionPotential
+DC34DecisionQuotient.StochasticSequential.stochasticExpectedUtility_le_iff_expectedActionPotential_ge
+DC35DecisionQuotient.StochasticSequential.landauerEnergyFloor_nonneg
+DC36DecisionQuotient.StochasticSequential.landauerEnergyFloor_mono_bits
+DC37DecisionQuotient.StochasticSequential.thermodynamicCost_eq_landauerEnergyFloorRoom_states
+DE1DecisionQuotient.ClaimClosure.DE1
+DE2DecisionQuotient.ClaimClosure.DE2
+DE3DecisionQuotient.ClaimClosure.DE3
+DE4DecisionQuotient.ClaimClosure.DE4
+DG1DecisionQuotient.Outside
+DG2DecisionQuotient.anchoredSlice
+DG3DecisionQuotient.anchoredSliceEquivOutside
+DG4DecisionQuotient.card_outside_eq_sub
+DG5DecisionQuotient.card_anchoredSlice
+DG6DecisionQuotient.card_anchoredSlice_eq_pow_sub
+DG7DecisionQuotient.card_anchoredSlice_eq_uniform
+DG8DecisionQuotient.anchoredSlice_mul_fixed_eq_full
+DG9DecisionQuotient.constantBoolDP
+DG10DecisionQuotient.firstCoordDP
+DG11DecisionQuotient.constantBoolDP_opt
+DG12DecisionQuotient.firstCoordDP_opt
+DG13DecisionQuotient.constantBoolDP_empty_sufficient
+DG14DecisionQuotient.firstCoordDP_empty_not_sufficient
+DG15DecisionQuotient.boolHypercube_node_count
+DG16DecisionQuotient.node_count_does_not_determine_edge_geometry
+DG17DecisionQuotient.DecisionProblem.edgeOnComplement
+DG18DecisionQuotient.DecisionProblem.edgeOnComplement_iff_not_sufficient
 DP1DecisionQuotient.DecisionProblem.minimalSufficient_iff_relevant
@@
 DP5DecisionQuotient.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
-DP6 DecisionProblem.anchorSufficient
-DP7 DecisionProblem.classMonotoneOn
-DP8 DecisionProblem.constant_opt_all_sufficient
-DP21 DecisionProblem.edgeOnComplement_iff_not_sufficient
-DQ3 DecisionQuotient.card_anchoredSlice_eq_pow_sub
-DQ5 DecisionQuotient.opt_eq_feasible_of_gap
-DQ6 DecisionQuotient.opt_eq_of_allowed_iff
-DQ7 DecisionQuotient.opt_singleton_of_logicallyDeterministic
-DQ8 DecisionQuotient.universe_sets_objective_schema
-DQ154 DecisionQuotient.Physics.Instantiation.DecisionCircuit
-DQ155 DecisionQuotient.Physics.Instantiation.DecisionInterpretation
-DQ159 DecisionQuotient.Physics.Instantiation.MoleculeAsCircuit
-DQ160 DecisionQuotient.Physics.Instantiation.MoleculeAsDecisionCircuit
-DQ517 DecisionQuotient.Physics.Instantiation.Circuit
-DQ518 DecisionQuotient.Physics.Instantiation.Dynamics
-DQ519 DecisionQuotient.Physics.Instantiation.Geometry
-DQ533 DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
-DQ534 DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
-DQ535 DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
-DQ537 DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
-DQ539 DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
-DQ540 DecisionQuotient.Physics.DecisionTime.run_time_exact
-DQ541 DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
-DQ542 DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
-DQ545 DecisionQuotient.Physics.DecisionTime.tick_increments_time
-DQ546 DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
-DQ547 DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
-DQ548 DecisionQuotient.Physics.DecisionTime.time_is_discrete
-DQ549 DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
-DQ560 DecisionQuotient.HardnessDistribution.hardnessLowerBound
-DQ561 DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
-DQ562 DecisionQuotient.HardnessDistribution.requiredWork
-DQ563 DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
-DQ564 DecisionQuotient.HardnessDistribution.workGrowthDegree
-DQ565 DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
-DQ566 DecisionQuotient.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
-DQ567 DecisionQuotient.Physics.PhysicalCore.certified_work_lower_bound_raw
-DQ568 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification
-DQ569 DecisionQuotient.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
-DQ570 DecisionQuotient.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
-DQ571 DecisionQuotient.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
-DQ572 DecisionQuotient.Physics.PhysicalCore.substrate_transport_sound_conditional
-DQ573 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget
-DQ574 DecisionQuotient.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
-DQ575 DecisionQuotient.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split
-DQ576 DecisionQuotient.PhysicalBudgetCrossover.exists_least_crossover_point
-DQ577 DecisionQuotient.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
-DQ578 DecisionQuotient.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
-DQ579 DecisionQuotient.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
-DQ580 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_at_divergence
-DQ581 DecisionQuotient.PhysicalBudgetCrossover.policy_closure_beyond_divergence
-DQ582 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_collision
-DQ583 DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
-DQ584 DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
-DQ585 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
-DQ586 DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
-DQ587 DecisionQuotient.checking_witnessing_duality
-DQ588 DecisionQuotient.checking_witnessing_duality_budget
+DP6DecisionQuotient.ClaimClosure.DP6
+DP7DecisionQuotient.ClaimClosure.DP7
+DP8DecisionQuotient.ClaimClosure.DP8
+DQ1DecisionQuotient.ClaimClosure.DQ1
+DQ2DecisionQuotient.ClaimClosure.DQ2
+DQ3DecisionQuotient.ClaimClosure.DQ3
+DQ4DecisionQuotient.ClaimClosure.DQ4
+DQ5DecisionQuotient.ClaimClosure.DQ5
+DQ6DecisionQuotient.ClaimClosure.DQ6
+DQ7DecisionQuotient.ClaimClosure.DQ7
+DQ8DecisionQuotient.ClaimClosure.DQ8
+DS1DecisionQuotient.ClaimClosure.DS1
+DS2DecisionQuotient.ClaimClosure.DS2
+DS3DecisionQuotient.ClaimClosure.DS3
+DS4DecisionQuotient.ClaimClosure.DS4
+DS5DecisionQuotient.ClaimClosure.DS5
+DS6DecisionQuotient.ClaimClosure.DS6
+DT1DecisionQuotient.Physics.DecisionTime.TimedState
+DT2DecisionQuotient.Physics.DecisionTime.DecisionProcess
+DT3DecisionQuotient.Physics.DecisionTime.tick
+DT4DecisionQuotient.Physics.DecisionTime.DecisionEvent
+DT5DecisionQuotient.Physics.DecisionTime.TimeUnitStep
+DT6DecisionQuotient.Physics.DecisionTime.time_is_discrete
+DT7DecisionQuotient.Physics.DecisionTime.time_coordinate_falsifiable
+DT8DecisionQuotient.Physics.DecisionTime.tick_increments_time
+DT9DecisionQuotient.Physics.DecisionTime.tick_decision_witness
+DT10DecisionQuotient.Physics.DecisionTime.tick_is_decision_event
+DT11DecisionQuotient.Physics.DecisionTime.decision_event_implies_time_unit
+DT12DecisionQuotient.Physics.DecisionTime.decision_taking_place_is_unit_of_time
+DT13DecisionQuotient.Physics.DecisionTime.decision_event_iff_eq_tick
+DT14DecisionQuotient.Physics.DecisionTime.run
+DT15DecisionQuotient.Physics.DecisionTime.run_time_exact
+DT16DecisionQuotient.Physics.DecisionTime.run_elapsed_time_eq_ticks
+DT17DecisionQuotient.Physics.DecisionTime.decisionTrace
+DT18DecisionQuotient.Physics.DecisionTime.decisionTrace_length_eq_ticks
+DT19DecisionQuotient.Physics.DecisionTime.decision_count_equals_elapsed_time
+DT20DecisionQuotient.Physics.DecisionTime.SubstrateKind
+DT21DecisionQuotient.Physics.DecisionTime.SubstrateModel
+DT22DecisionQuotient.Physics.DecisionTime.substrate_step_realizes_decision_event
+DT23DecisionQuotient.Physics.DecisionTime.substrate_step_is_time_unit
+DT24DecisionQuotient.Physics.DecisionTime.time_unit_law_substrate_invariant
+EI1DecisionQuotient.ThermodynamicLift.energy_ge_kbt_nat_entropy
+EI2DecisionQuotient.ThermodynamicLift.energy_ground_state_tracks_entropy
+EI4DecisionQuotient.ThermodynamicLift.landauerJoulesPerBit_pos
+EI5DecisionQuotient.ThermodynamicLift.neukart_vinokur_duality
+EP1DecisionQuotient.Physics.LocalityPhysics.landauer_principle
+EP2DecisionQuotient.Physics.LocalityPhysics.finite_regional_energy
+EP3DecisionQuotient.Physics.LocalityPhysics.finite_signal_speed
+EP4DecisionQuotient.Physics.LocalityPhysics.nontrivial_physics
+FI3DecisionQuotient.FunctionalInformation.functional_information_from_counting
+FI6DecisionQuotient.FunctionalInformation.functional_information_from_thermodynamics
+FI7DecisionQuotient.FunctionalInformation.first_principles_thermo_coincide
+FN7DecisionQuotient.BayesOptimalityProof.KL_nonneg
+FN8DecisionQuotient.BayesOptimalityProof.entropy_le_crossEntropy
+FN12DecisionQuotient.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
+FN14DecisionQuotient.BayesOptimalityProof.bayes_is_optimal
+FN15DecisionQuotient.BayesOptimalityProof.KL_eq_sum_klFun
+FN16DecisionQuotient.BayesOptimalityProof.KL_eq_zero_iff_eq
+FP1DecisionQuotient.Physics.LocalityPhysics.trivial_states_all_equal
+FP2DecisionQuotient.Physics.LocalityPhysics.equal_states_constant_function
+FP3DecisionQuotient.Physics.LocalityPhysics.constant_function_singleton_image
+FP4DecisionQuotient.Physics.LocalityPhysics.singleton_image_zero_entropy
+FP5DecisionQuotient.Physics.LocalityPhysics.zero_entropy_no_information
+FP6DecisionQuotient.Physics.LocalityPhysics.triviality_implies_no_information
+FP7DecisionQuotient.Physics.LocalityPhysics.information_requires_nontriviality
+FP8DecisionQuotient.Physics.LocalityPhysics.atypical_states_rare
+FP9DecisionQuotient.Physics.LocalityPhysics.random_misses_target
+FP10DecisionQuotient.Physics.LocalityPhysics.errors_accumulate
+FP11DecisionQuotient.Physics.LocalityPhysics.wrong_paths_dominate
+FP12DecisionQuotient.Physics.LocalityPhysics.second_law_from_counting
+FP13DecisionQuotient.Physics.LocalityPhysics.verification_is_information
+FP14DecisionQuotient.Physics.LocalityPhysics.entropy_is_information
+FP15DecisionQuotient.Physics.LocalityPhysics.landauer_structure
+FPT1DecisionQuotient.Physics.LocalityPhysics.Timeline
+FPT2DecisionQuotient.Physics.LocalityPhysics.FPT2_function_deterministic
+FPT3DecisionQuotient.Physics.LocalityPhysics.FPT3_outputs_differ_inputs_differ
+FPT4DecisionQuotient.Physics.LocalityPhysics.FPT4_step_requires_distinct_moments
+FPT5DecisionQuotient.Physics.LocalityPhysics.FPT5_distinct_moments_positive_duration
+FPT6DecisionQuotient.Physics.LocalityPhysics.FPT6_step_takes_positive_time
+FPT7DecisionQuotient.Physics.LocalityPhysics.FPT7_no_instantaneous_steps
+FPT8DecisionQuotient.Physics.LocalityPhysics.FPT8_propagation_takes_time
+FPT9DecisionQuotient.Physics.LocalityPhysics.FPT9_speed_bounded_by_positive_time
+FPT10DecisionQuotient.Physics.LocalityPhysics.FPT10_ec3_is_logical
+FS1DecisionQuotient.Statistics.sum_fisherScore_eq_srank
+FS2DecisionQuotient.Statistics.fisherMatrix_rank_eq_srank
+FS3DecisionQuotient.Statistics.fisherMatrix_trace_eq_srank
+FS4DecisionQuotient.Statistics.fisherScore_relevant
+FS5DecisionQuotient.Statistics.fisherScore_irrelevant
+GE1DecisionQuotient.ClaimClosure.GE1
+GE2DecisionQuotient.ClaimClosure.GE2
+GE3DecisionQuotient.ClaimClosure.GE3
+GE4DecisionQuotient.ClaimClosure.GE4
+GE5DecisionQuotient.ClaimClosure.GE5
+GE6DecisionQuotient.ClaimClosure.GE6
+GE7DecisionQuotient.ClaimClosure.GE7
+GE8DecisionQuotient.ClaimClosure.GE8
+GE9DecisionQuotient.ClaimClosure.GE9
+GN1DecisionQuotient.LogicGraph.isCycle
+GN2DecisionQuotient.LogicGraph.cycleWitnessBits_pos
+GN3DecisionQuotient.LogicGraph.pathProb_nonneg
+GN4DecisionQuotient.LogicGraph.pathSurprisal_nonneg_of_positive_mass
+GN5DecisionQuotient.LogicGraph.nontrivialityScore_unknown
+GN6DecisionQuotient.LogicGraph.observerEntropy_nonneg
+GN7DecisionQuotient.LogicGraph.dqFromEntropy_in_unit_interval
+GN8DecisionQuotient.LogicGraph.path_belief_forced_under_uncertainty
+GN9DecisionQuotient.LogicGraph.bayes_update_exists_for_observer_paths
+GN10DecisionQuotient.LogicGraph.cycle_witness_implies_positive_landauer
+GN11DecisionQuotient.LogicGraph.cycle_witness_implies_positive_nv_work
+GN12DecisionQuotient.LogicGraph.dna_erasure_implies_positive_landauer
+GN13DecisionQuotient.LogicGraph.dna_room_temp_environmental_stability
 HD1DecisionQuotient.HardnessDistribution.centralization_dominance_bundle
@@
 HD16DecisionQuotient.HardnessDistribution.no_positive_slope_linear_represents_saturating
-HD17 DecisionQuotient.HardnessDistribution.right_dominates_wrong
-HD18 DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
-HD19 DecisionQuotient.HardnessDistribution.simplicityTax_grows
-HD20 DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
-HD21 DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic
+HD17DecisionQuotient.HardnessDistribution.requiredWork
+HD18DecisionQuotient.HardnessDistribution.requiredWork_eq_affine_in_sites
+HD19DecisionQuotient.HardnessDistribution.right_dominates_wrong
+HD20DecisionQuotient.HardnessDistribution.saturatingSiteCost_eventually_constant
+HD21DecisionQuotient.HardnessDistribution.simplicityTax_grows
+HD22DecisionQuotient.HardnessDistribution.hardnessLowerBound
+HD23DecisionQuotient.HardnessDistribution.hardness_is_irreducible_required_work
+HD24DecisionQuotient.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
+HD25DecisionQuotient.HardnessDistribution.totalDOF_ge_intrinsic
+HD26DecisionQuotient.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
+HD27DecisionQuotient.HardnessDistribution.workGrowthDegree
+HD28DecisionQuotient.HardnessDistribution.workGrowthDegree_zero_iff_eventually_constant
+HS1DecisionQuotient.Physics.HeisenbergStrong.NoisyPhysicalEncoding
+HS2DecisionQuotient.Physics.HeisenbergStrong.HeisenbergStrongBinding
+HS3DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial
+HS4DecisionQuotient.Physics.HeisenbergStrong.strong_binding_yields_core_encoding_witness
+HS5DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_physical_nontrivial_opt_assumption
+HS6DecisionQuotient.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
+IA1DecisionQuotient.ClaimClosure.IA1
+IA2DecisionQuotient.ClaimClosure.IA2
+IA3DecisionQuotient.ClaimClosure.IA3
+IA4DecisionQuotient.ClaimClosure.IA4
+IA5DecisionQuotient.ClaimClosure.IA5
+IA6DecisionQuotient.ClaimClosure.IA6
+IA7DecisionQuotient.ClaimClosure.IA7
+IA8DecisionQuotient.ClaimClosure.IA8
+IA9DecisionQuotient.ClaimClosure.IA9
+IA10DecisionQuotient.ClaimClosure.IA10
+IA11DecisionQuotient.ClaimClosure.IA11
+IA12DecisionQuotient.ClaimClosure.IA12
+IA13DecisionQuotient.ClaimClosure.IA13
 IC1DecisionQuotient.IntegrityCompetence.CertaintyInflation
@@
 IC47DecisionQuotient.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
+IC48DecisionQuotient.IntegrityCompetence.steps_run_scalar_always_defined
+IC49DecisionQuotient.IntegrityCompetence.steps_run_scalar_falsifiable
+IC50DecisionQuotient.IntegrityCompetence.zero_epsilon_competence_iff_exact
+IE1DecisionQuotient.ClaimClosure.IE1
+IE2DecisionQuotient.ClaimClosure.IE2
+IE3DecisionQuotient.ClaimClosure.IE3
+IE4DecisionQuotient.ClaimClosure.IE4
+IE5DecisionQuotient.ClaimClosure.IE5
+IE6DecisionQuotient.ClaimClosure.IE6
+IE7DecisionQuotient.ClaimClosure.IE7
+IE8DecisionQuotient.ClaimClosure.IE8
+IE9DecisionQuotient.ClaimClosure.IE9
+IE10DecisionQuotient.ClaimClosure.IE10
+IE11DecisionQuotient.ClaimClosure.IE11
+IE12DecisionQuotient.ClaimClosure.IE12
+IE13DecisionQuotient.ClaimClosure.IE13
+IE14DecisionQuotient.ClaimClosure.IE14
+IE15DecisionQuotient.ClaimClosure.IE15
+IE16DecisionQuotient.ClaimClosure.IE16
+IE17DecisionQuotient.ClaimClosure.IE17
+IN1DecisionQuotient.Physics.Instantiation.Geometry
+IN2DecisionQuotient.Physics.Instantiation.Dynamics
+IN3DecisionQuotient.Physics.Instantiation.Circuit
+IN4DecisionQuotient.Physics.Instantiation.geometry_plus_dynamics_is_circuit
+IN5DecisionQuotient.Physics.Instantiation.DecisionInterpretation
+IN6DecisionQuotient.Physics.Instantiation.DecisionCircuit
+IN7DecisionQuotient.Physics.Instantiation.Molecule
+IN8DecisionQuotient.Physics.Instantiation.Reaction
+IN9DecisionQuotient.Physics.Instantiation.ReactionOutcome
+IN10DecisionQuotient.Physics.Instantiation.MoleculeGeometry
+IN11DecisionQuotient.Physics.Instantiation.MoleculeDynamics
+IN12DecisionQuotient.Physics.Instantiation.MoleculeCircuit
+IN13DecisionQuotient.Physics.Instantiation.MoleculeAsCircuit
+IN14DecisionQuotient.Physics.Instantiation.MoleculeAsDecisionCircuit
+IN15DecisionQuotient.Physics.Instantiation.molecule_decision_preserves_geometry
+IN16DecisionQuotient.Physics.Instantiation.molecule_decision_preserves_dynamics
+IN17DecisionQuotient.Physics.Instantiation.asDecisionCircuit
+IN18DecisionQuotient.Physics.Instantiation.asDecisionCircuit_preserves_circuit
+IN19DecisionQuotient.Physics.Instantiation.Configuration
+IN20DecisionQuotient.Physics.Instantiation.EnergyLandscape
+IN21DecisionQuotient.Physics.Instantiation.k_Boltzmann
+IN22DecisionQuotient.Physics.Instantiation.LandauerBound
+IN23DecisionQuotient.Physics.Instantiation.law_objective_schema
+IN24DecisionQuotient.Physics.Instantiation.law_opt_eq_feasible_of_gap
+IN25DecisionQuotient.Physics.Instantiation.law_opt_singleton_of_deterministic
+IP1DecisionQuotient.Physics.LocalityPhysics.ec1_can_be_true
+IP2DecisionQuotient.Physics.LocalityPhysics.ec1_independent_of_ec2_ec3
+IP3DecisionQuotient.Physics.LocalityPhysics.ec2_can_be_true
+IP4DecisionQuotient.Physics.LocalityPhysics.ec2_independent_of_ec1_ec3
+IP5DecisionQuotient.Physics.LocalityPhysics.ec3_can_be_true
+IP6DecisionQuotient.Physics.LocalityPhysics.ec3_independent_of_ec1_ec2
+IP7DecisionQuotient.Physics.LocalityPhysics.empirical_claims_mutually_independent
+IT1DecisionQuotient.DecisionProblem.numOptClasses
+IT3DecisionQuotient.quotientEntropy_le_srank_binary
+IT4DecisionQuotient.numOptClasses_le_pow_srank_binary
+IT5DecisionQuotient.nontrivial_bounds_binary
+IT6DecisionQuotient.nontrivial_implies_srank_pos
 IV1DecisionQuotient.InteriorVerification.GoalClass
@@
 IV9DecisionQuotient.InteriorVerification.interior_verification_tractable_certificate
-L1 ExactPhysicalClaimWellTyped
-L2 ExcusedBy
-L3 ExplicitHardnessAssumptions
-L4 Literal.eval
-L5 OneStepSequentialObjective
-L6 RegimeSimulation
-L7 StandardComplexityAssumptions
-L8 StochasticCriterionObjective
-L9 StructureDetectable
-L10 TransitionCoupledObjective
-L11 TwoStepObjective
-L12 adq
-L13 adq_ordering
-L14 agentBridgeClass
-L15 agent_transfer_licensed_iff_snapshot
-L16 anchor_sigma2p_complete_conditional
-L17 anchor_sigma2p_reduction_core
-L19 boundaryCharacterized_iff_exists_sufficient_subset
-L20 bounded_actions_detectable
-L21 bounded_actions_reusable_heuristic
-L22 bridgeFailureWitness
-L23 bridgeTransferLicensed
-L24 bridge_boundary_represented_family
-L25 boolHypercube_node_count
-L68 node_count_does_not_determine_edge_geometry
+IV10DecisionQuotient.InteriorVerification.not_sufficientOnSet_of_counterexample
+IV11DecisionQuotient.InteriorVerification.singleton_coordinate_interior_certificate
+LP1DecisionQuotient.Physics.LocalityPhysics.SpacetimePoint
+LP2DecisionQuotient.Physics.LocalityPhysics.lightCone
+LP3DecisionQuotient.Physics.LocalityPhysics.causalPast
+LP4DecisionQuotient.Physics.LocalityPhysics.self_in_lightCone
+LP5DecisionQuotient.Physics.LocalityPhysics.self_in_causalPast
+LP6DecisionQuotient.Physics.LocalityPhysics.LocalRegion
+LP7DecisionQuotient.Physics.LocalityPhysics.canObserve
+LP8DecisionQuotient.Physics.LocalityPhysics.spacelikeSeparated
+LP9DecisionQuotient.Physics.LocalityPhysics.spacelike_disjoint_observation
+LP11DecisionQuotient.Physics.LocalityPhysics.LocalConfiguration
+LP12DecisionQuotient.Physics.LocalityPhysics.isLocallyValid
+LP13DecisionQuotient.Physics.LocalityPhysics.MergeResult
+LP14DecisionQuotient.Physics.LocalityPhysics.boardMerge
+LP15DecisionQuotient.Physics.LocalityPhysics.independent_configs_can_disagree
+LP16DecisionQuotient.Physics.LocalityPhysics.merge_compatible_iff
+LP17DecisionQuotient.Physics.LocalityPhysics.merge_contradiction_iff
+LP18DecisionQuotient.Physics.LocalityPhysics.locality_implies_possible_contradiction
+LP19DecisionQuotient.Physics.LocalityPhysics.Superposition
+LP20DecisionQuotient.Physics.LocalityPhysics.superposition_can_contain_contradictions
+LP21DecisionQuotient.Physics.LocalityPhysics.superposition_requires_separation
+LP22DecisionQuotient.Physics.LocalityPhysics.bell_separation_is_real
+LP23DecisionQuotient.Physics.LocalityPhysics.measurement_is_merge_contradiction
+LP24DecisionQuotient.Physics.LocalityPhysics.entanglement_is_shared_origin
+LP31DecisionQuotient.Physics.LocalityPhysics.complete_knowledge_requires_all_queries
+LP32DecisionQuotient.Physics.LocalityPhysics.finite_energy_constraint
+LP33DecisionQuotient.Physics.LocalityPhysics.self_knowledge_impossible_if_insufficient_energy
+LP34DecisionQuotient.Physics.LocalityPhysics.bounded_energy_forces_locality
+LP35DecisionQuotient.Physics.LocalityPhysics.locality_implies_independent_regions
+LP36DecisionQuotient.Physics.LocalityPhysics.independent_regions_imply_possible_contradiction
+LP38DecisionQuotient.Physics.LocalityPhysics.pne_np_necessary_for_physics
+LP39DecisionQuotient.Physics.LocalityPhysics.matter_exists_because_pne_np
+LP40DecisionQuotient.Physics.LocalityPhysics.physics_is_the_game
+LP41DecisionQuotient.Physics.LocalityPhysics.without_positive_query_cost_no_bound
+LP42DecisionQuotient.Physics.LocalityPhysics.without_nontrivial_states_no_disagreement
+LP43DecisionQuotient.Physics.LocalityPhysics.without_separation_no_independence
+LP44DecisionQuotient.Physics.LocalityPhysics.without_finite_capacity_no_gap
+LP45DecisionQuotient.Physics.LocalityPhysics.all_premises_used
+LP46DecisionQuotient.Physics.LocalityPhysics.premises_necessary_and_sufficient
+LP50DecisionQuotient.Physics.LocalityPhysics.shannon_value_is_intractability
+LP51DecisionQuotient.Physics.LocalityPhysics.economic_value_requires_scarcity
+LP52DecisionQuotient.Physics.LocalityPhysics.thermodynamic_value_requires_gradient
+LP53DecisionQuotient.Physics.LocalityPhysics.voi_requires_uncertainty
+LP54DecisionQuotient.Physics.LocalityPhysics.physics_requires_intractability
+LP55DecisionQuotient.Physics.LocalityPhysics.value_is_intractability
+LP56DecisionQuotient.Physics.LocalityPhysics.observers_value_the_intractable
+LP57DecisionQuotient.Physics.LocalityPhysics.finite_steps_finite_coverage
+LP58DecisionQuotient.Physics.LocalityPhysics.counting_gap
+LP59DecisionQuotient.Physics.LocalityPhysics.time_is_counting
+LP60DecisionQuotient.Physics.LocalityPhysics.gap_equivalence
+LP61DecisionQuotient.Physics.LocalityPhysics.light_cone_is_time_gap
+MI1DecisionQuotient.ClaimClosure.MI1
+MI2DecisionQuotient.ClaimClosure.MI2
+MI3DecisionQuotient.ClaimClosure.MI3
+MI4DecisionQuotient.ClaimClosure.MI4
+MI5DecisionQuotient.ClaimClosure.MI5
+MN1DecisionQuotient.Physics.MeasureNecessity.quantitative_claim_has_measure
+MN2DecisionQuotient.Physics.MeasureNecessity.stochastic_claim_has_probability_measure
+MN3DecisionQuotient.Physics.MeasureNecessity.stochastic_claim_has_measure
+MN4DecisionQuotient.Physics.MeasureNecessity.count_univ_bool
+MN5DecisionQuotient.Physics.MeasureNecessity.counting_measure_not_probability_on_bool
+MN6DecisionQuotient.Physics.MeasureNecessity.deterministic_dirac_is_probability
+MN7DecisionQuotient.Physics.MeasureNecessity.quantitative_value_depends_on_measure
+MN8DecisionQuotient.Physics.MeasureNecessity.deterministic_models_still_measure_based
+MN9DecisionQuotient.Physics.MeasureNecessity.measure_does_not_imply_probability
+MN10DecisionQuotient.Physics.MeasureNecessity.quantitative_measure_is_logical_prerequisite
+MN11DecisionQuotient.Physics.MeasureNecessity.stochastic_probability_is_logical_prerequisite
+OR1DecisionQuotient.Physics.ObserverRelativeState.ObserverClass
+OR2DecisionQuotient.Physics.ObserverRelativeState.obsEquiv
+OR3DecisionQuotient.Physics.ObserverRelativeState.EffectiveStateSpace
+OR4DecisionQuotient.Physics.ObserverRelativeState.project_eq_iff
+OR5DecisionQuotient.Physics.ObserverRelativeState.observer_relative_equivalence_witness
+OR6DecisionQuotient.Physics.ObserverRelativeState.PhysicalObserverClass
+OR7DecisionQuotient.Physics.ObserverRelativeState.PhysicalStateSpace
+OR8DecisionQuotient.Physics.ObserverRelativeState.physical_state_space_has_instance_witness
+OR9DecisionQuotient.Physics.ObserverRelativeState.physical_observer_relative_effective_space
+PA1DecisionQuotient.Physics.AnchorChecks.obsEquiv_all_of_effective_subsingleton
+PA2DecisionQuotient.Physics.AnchorChecks.stochasticAnchorSufficient_iff_exists_anchor_singleton
+PA3DecisionQuotient.Physics.AnchorChecks.stochastic_anchor_check_iff_exists_anchor_singleton
+PA4DecisionQuotient.Physics.AnchorChecks.stochastic_sufficient_of_observer_collapse_and_seed
+PA5DecisionQuotient.Physics.AnchorChecks.stochastic_anchor_check_of_observer_collapse_and_seed
+PA6DecisionQuotient.Physics.AnchorChecks.sequential_sufficient_of_observer_collapse
+PA7DecisionQuotient.Physics.AnchorChecks.sequential_anchor_check_of_observer_collapse
+PA8DecisionQuotient.Physics.AnchorChecks.physical_observer_collapse_implies_obsEquiv_all
+PA9DecisionQuotient.Physics.AnchorChecks.physical_stochastic_anchor_check_of_observer_collapse_and_seed
 PBC1DecisionQuotient.PhysicalBudgetCrossover.CrossoverAt
-PBC2 DecisionQuotient.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
-PC10 PhysicalComplexity.AccessRegime.AccessRegime
-PC12 PhysicalComplexity.AccessRegime.RegimeEval
-PC15 PhysicalComplexity.AccessRegime.RegimeWithCertificate
-PC17 PhysicalComplexity.AccessRegime.certificate_upgrades_regime
-PC19 PhysicalComplexity.AccessRegime.regime_upgrade_with_certificate
-PC20 PhysicalComplexity.AccessRegime.AccessChannelLaw
-PC21 PhysicalComplexity.AccessRegime.FiveWayMeet
+PBC2DecisionQuotient.PhysicalBudgetCrossover.SuccinctInfeasible
+PBC3DecisionQuotient.PhysicalBudgetCrossover.SuccinctUnbounded
+PBC4DecisionQuotient.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
+PBC5DecisionQuotient.PhysicalBudgetCrossover.exists_least_crossover_point
+PBC6DecisionQuotient.PhysicalBudgetCrossover.has_crossover_of_bounded_succinct_unbounded_explicit
+PBC7DecisionQuotient.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
+PBC8DecisionQuotient.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split
+PBC9DecisionQuotient.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
+PBC10DecisionQuotient.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
+PBC11DecisionQuotient.PhysicalBudgetCrossover.policy_closure_at_divergence
+PBC12DecisionQuotient.PhysicalBudgetCrossover.policy_closure_beyond_divergence
+PH1PhysicalComplexity.k_Boltzmann
+PH2PhysicalComplexity.PhysicalComputer
+PH3PhysicalComplexity.bit_energy_cost
+PH4PhysicalComplexity.Landauer_bound
+PH5PhysicalComplexity.InstanceSize
+PH6PhysicalComplexity.ComputationalRequirement
+PH7PhysicalComplexity.coNP_requirement
+PH8PhysicalComplexity.coNP_physically_impossible
+PH9PhysicalComplexity.coNP_not_in_P_physically
+PH10PhysicalComplexity.sufficiency_physically_impossible
+PH11DecisionQuotient.PhysicalComplexity.PhysicalCollapseAtRequirement
+PH12DecisionQuotient.PhysicalComplexity.no_physical_collapse_at_requirement
+PH13DecisionQuotient.PhysicalComplexity.canonical_physical_collapse_impossible
+PH14DecisionQuotient.PhysicalComplexity.p_eq_np_physically_impossible_of_collapse_map
+PH15DecisionQuotient.PhysicalComplexity.p_eq_np_physically_impossible_canonical
+PH16DecisionQuotient.PhysicalComplexity.P_eq_NP_via_SAT
+PH17DecisionQuotient.PhysicalComplexity.SAT3ReductionBridge
+PH18DecisionQuotient.PhysicalComplexity.sat_reduction_transfers_energy_lower_bound
+PH19DecisionQuotient.PhysicalComplexity.physical_collapse_of_polytime_sat_realization
+PH20DecisionQuotient.PhysicalComplexity.p_eq_np_physically_impossible_via_sat_bridge
+PH21DecisionQuotient.PhysicalComplexity.SAT3HardFamily
+PH22DecisionQuotient.PhysicalComplexity.p_eq_np_physically_impossible_via_sat_hard_family
+PH23DecisionQuotient.PhysicalComplexity.collapse_possible_without_positive_bit_cost
+PH24DecisionQuotient.PhysicalComplexity.collapse_possible_without_exponential_lower_bound
+PH25DecisionQuotient.PhysicalComplexity.no_go_transfer_requires_collapse_map
+PH26DecisionQuotient.PhysicalComplexity.no_collapse_of_bounded_budget_pos_cost_exp_lb
+PH27DecisionQuotient.PhysicalComplexity.collapse_implies_assumption_failure_disjunction
+PH28DecisionQuotient.PhysicalComplexity.deterministic_no_physical_collapse
+PH29DecisionQuotient.PhysicalComplexity.probabilistic_no_physical_collapse
+PH30DecisionQuotient.PhysicalComplexity.sequential_no_physical_collapse
+PH31DecisionQuotient.PhysicalComplexity.collapse_possible_with_unbounded_budget_profile
+PH32DecisionQuotient.PhysicalComplexity.exp_budget_profile_unbounded
+PH33DecisionQuotient.PhysicalComplexity.finite_budget_assumption_is_necessary
+PH34DecisionQuotient.PhysicalComplexity.CoherentDQRejectionAtRequirement
+PH35DecisionQuotient.PhysicalComplexity.coherent_dq_rejection_impossible_at_requirement
+PH36DecisionQuotient.PhysicalComplexity.coherent_dq_rejection_impossible_canonical
+PI1DecisionQuotient.Physics.PhysicalIncompleteness.UniverseModel
+PI2DecisionQuotient.Physics.PhysicalIncompleteness.PhysicallyInstantiated
+PI3DecisionQuotient.Physics.PhysicalIncompleteness.no_surjective_instantiation_of_card_gap
+PI4DecisionQuotient.Physics.PhysicalIncompleteness.physical_incompleteness_of_card_gap
+PI5DecisionQuotient.Physics.PhysicalIncompleteness.physical_incompleteness_of_bounds
+PI6DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_collision
+PI7DecisionQuotient.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
+PS1DecisionQuotient.Physics.ClaimTransport.PhysicalStateSemantics
+PS2DecisionQuotient.Physics.ClaimTransport.physical_state_has_witness
+PS3DecisionQuotient.Physics.ClaimTransport.physical_state_claim_of_instance_claim
+PS4DecisionQuotient.Physics.ClaimTransport.physical_state_claim_of_universal_core
+QT1DecisionQuotient.DecisionProblem.quotient_is_coarsest
+QT2DecisionQuotient.DecisionProblem.quotientMap_preservesOpt
+QT3DecisionQuotient.DecisionProblem.quotient_represents_opt_equiv
+QT4DecisionQuotient.DecisionProblem.factors_implies_respects
+RD1DecisionQuotient.Information.shannonEntropy_nonneg
+RD2DecisionQuotient.Information.rate_zero_distortion
+RD3DecisionQuotient.Information.rate_monotone
+RS1DecisionQuotient.Information.equiv_preserves_decision
+RS2DecisionQuotient.Information.rate_equals_srank
+RS3DecisionQuotient.Information.compression_below_srank_fails
+RS4DecisionQuotient.Information.srank_bits_sufficient
+RS5DecisionQuotient.Information.rate_distortion_bridge
 S2P1DecisionQuotient.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
@@
 S2P3DecisionQuotient.Sigma2PHardness.min_sufficient_set_iff_relevant_card
+S2P4DecisionQuotient.Sigma2PHardness.representationGap
+S2P5DecisionQuotient.Sigma2PHardness.representationGap_eq_waste_plus_missing
 S2P6DecisionQuotient.Sigma2PHardness.representationGap_eq_zero_iff
+S2P7DecisionQuotient.Sigma2PHardness.representationGap_missing_eq_gapCard
+S2P8DecisionQuotient.Sigma2PHardness.representationGap_zero_iff_minimalSufficient
+S2P9DecisionQuotient.Sigma2PHardness.sufficient_iff_relevant_subset
+SE1DecisionQuotient.ClaimClosure.SE1
+SE2DecisionQuotient.ClaimClosure.SE2
+SE3DecisionQuotient.ClaimClosure.SE3
+SE4DecisionQuotient.ClaimClosure.SE4
+SE5DecisionQuotient.ClaimClosure.SE5
+SE6DecisionQuotient.ClaimClosure.SE6
+SR1DecisionQuotient.ClaimClosure.SR1
+SR2DecisionQuotient.ClaimClosure.SR2
+SR3DecisionQuotient.ClaimClosure.SR3
+SR4DecisionQuotient.ClaimClosure.SR4
+SR5DecisionQuotient.ClaimClosure.SR5
+TC1DecisionQuotient.ToolCollapse.WorkProfile
+TC2DecisionQuotient.ToolCollapse.WorkModel
+TC3DecisionQuotient.ToolCollapse.ToolModel
+TC4DecisionQuotient.ToolCollapse.EventualStrictImprovement
+TC5DecisionQuotient.ToolCollapse.EffectiveCollapse
+TC6DecisionQuotient.ToolCollapse.tool_never_worse
+TC7DecisionQuotient.ToolCollapse.strict_improvement_has_witness
+TC8DecisionQuotient.ToolCollapse.effective_collapse_of_eventual_strict
+TC9DecisionQuotient.ToolCollapse.expBaseline
+TC10DecisionQuotient.ToolCollapse.linearTool
+TC11DecisionQuotient.ToolCollapse.linear_tool_eventual_strict
+TC12DecisionQuotient.ToolCollapse.linear_tool_effective_collapse
+TL3DecisionQuotient.ThermodynamicLift.joulesPerBit_pos_of_landauer_calibration
+TL4DecisionQuotient.ThermodynamicLift.energy_lower_mandatory_of_landauer_calibration
+TUR1DecisionQuotient.Physics.transitionProb_nonneg
+TUR2DecisionQuotient.Physics.transitionProb_sum_one
+TUR5DecisionQuotient.Physics.tur_bridge
+TUR6DecisionQuotient.Physics.multiple_futures_entropy_production
+UO1DecisionQuotient.UniverseDynamics
+UO2DecisionQuotient.feasibleActions
+UO3DecisionQuotient.lawDecisionProblem
+UO4DecisionQuotient.lawUtility
+UO5DecisionQuotient.logicallyDeterministic
+UO6DecisionQuotient.universe_sets_objective_schema
+UO7DecisionQuotient.lawUtility_eq_of_allowed_iff
+UO8DecisionQuotient.opt_eq_feasible_of_gap
+UO9DecisionQuotient.infeasible_not_optimal_of_gap
+UO10DecisionQuotient.opt_singleton_of_logicallyDeterministic
+UO11DecisionQuotient.opt_eq_of_allowed_iff
+UQ1DecisionQuotient.Physics.Uncertainty.binaryIdentityProblem
+UQ2DecisionQuotient.Physics.Uncertainty.binaryIdentityProblem_opt_true
+UQ3DecisionQuotient.Physics.Uncertainty.binaryIdentityProblem_opt_false
+UQ4DecisionQuotient.Physics.Uncertainty.exists_decision_problem_with_nontrivial_opt
+UQ5DecisionQuotient.Physics.Uncertainty.PhysicalNontrivialOptAssumption
+UQ6DecisionQuotient.Physics.Uncertainty.exists_decision_problem_with_nontrivial_opt_of_physical
+W1DecisionQuotient.Physics.single_future_zero_cost
+W2DecisionQuotient.Physics.transportCost_pos_of_offDiag
+W3DecisionQuotient.Physics.integrity_is_centroid
+W4DecisionQuotient.Physics.wasserstein_bridge
+WD1DecisionQuotient.checking_witnessing_duality_budget
+WD2DecisionQuotient.no_sound_checker_below_witness_budget
+WD3DecisionQuotient.checking_time_ge_witness_budget
+WD4DecisionQuotient.witnessBudgetEmpty
+WD5DecisionQuotient.checkingBudgetPairs
+XC1DecisionQuotient.Physics.srank_determines_states
+XC2DecisionQuotient.Physics.more_states_more_transport
+XC3DecisionQuotient.Physics.transport_lower_bound
+XC4DecisionQuotient.Physics.transport_independent_of_energy
+XC5DecisionQuotient.Physics.transport_independent_of_precision
+XC6DecisionQuotient.Physics.srank_unified_complexity
+XC7DecisionQuotient.Physics.complete_bridge_set
 A.2 What Is Machine-Checked
@@
   • DecisionQuotient/ClaimClosure.lean
+  • DecisionQuotient/HandleAliases.lean
+  • DecisionQuotient/StochasticSequential/ClaimClosure.lean
+  • DecisionQuotient/StochasticSequential/Hierarchy.lean
 Assumption bundles and fields.
@@
   • DQ.tractable_subcases_conditional
+First-principles Bayes derivation handles.
+  • DQ.BayesOptimalityProof.KL_eq_sum_klFun
+  • DQ.BayesOptimalityProof.KL_eq_zero_iff_eq
+  • DQ.BayesOptimalityProof.KL_nonneg
+  • DQ.BayesOptimalityProof.bayes_is_optimal
+  • DQ.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
+  • DQ.BayesOptimalityProof.entropy_le_crossEntropy
+  • DQ.Foundations.bayes_from_conditional
+  • DQ.Foundations.complete_chain
+  • DQ.Foundations.counting_additive
+  • DQ.Foundations.counting_nonneg
+  • DQ.Foundations.counting_total
+  • DQ.Foundations.dq_equivalence
+  • DQ.Foundations.dq_in_unit_interval
+  • DQ.Foundations.dq_one_iff_perfect_info
+  • DQ.Foundations.dq_zero_iff_no_info
+  • DQ.Foundations.entropy_contraction
+  • DQ.Foundations.mutual_info_nonneg
+  • DQ.bayesian_dq_matches_physics_dq
+  • DQ.chain_rule
+  • DQ.dq_derived_from_bayes
+  • DQ.dq_is_bayesian_certainty_fraction
+Compact Bayes handle IDs.
+  • FN7 $\to$ DQ.BayesOptimalityProof.KL_nonneg
+  • FN8 $\to$ DQ.BayesOptimalityProof.entropy_le_crossEntropy
+  • FN12 $\to$ DQ.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
+  • FN14 $\to$ DQ.BayesOptimalityProof.bayes_is_optimal
+  • FN15 $\to$ DQ.BayesOptimalityProof.KL_eq_sum_klFun
+  • FN16 $\to$ DQ.BayesOptimalityProof.KL_eq_zero_iff_eq
+  • BB1 $\to$ DQ.BayesianDQ
+  • BB2 $\to$ DQ.BayesianDQ.certaintyGain
+  • BB3 $\to$ DQ.dq_is_bayesian_certainty_fraction
+  • BB4 $\to$ DQ.bayesian_dq_matches_physics_dq
+  • BB5 $\to$ DQ.dq_derived_from_bayes
+  • BF1 $\to$ DQ.certainty_of_not_nondegenerateBelief
+  • BF2 $\to$ DQ.nondegenerateBelief_of_uncertaintyForced
+  • BF3 $\to$ DQ.forced_action_under_uncertainty
+  • BF4 $\to$ DQ.bayes_update_exists_of_nondegenerateBelief
 A.4 Polynomial-Time Reduction Definition
@@
 Paper handle Status Lean support
+cor:channel-degradation Full DQ.ClaimClosure.CH2, DQ.ClaimClosure.CH6
+cor:ego-trap Unmapped (no derived Lean handle found)
 cor:exact-identifiability Full DQ.Sigma2PHardness.exactlyIdentifiesRelevant_iff_sufficient_and_subset_relevantFinset
 cor:exact-no-competence-zero-certified Full DQ.IntegrityCompetence.rlff_maximizer_has_evidence
-cor:finite-budget-no-exact-admissibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget, DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
-cor:finite-budget-threshold-impossibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement, DQ.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification, DQ.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
-cor:gap-externalization Full DQ.HardnessDistribution.saturatingSiteCost_eventually_constant, DQ.HardnessDistribution.totalDOF_ge_intrinsic
+cor:finite-budget-no-exact-admissibility Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
+cor:finite-budget-threshold-impossibility Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
+cor:finite-lifetime Full DQ.ClaimClosure.SE5
+cor:finite-state Full DQ.Physics.BoundedAcquisition.acquisition_rate_bound, DQ.Physics.BoundedAcquisition.one_bit_per_transition
+cor:forced-finite-speed Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
+cor:gap-externalization Full DQ.HardnessDistribution.simplicityTax_grows, DQ.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
 cor:gap-minimization-hard Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
-cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
-cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core
-cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.IntegrityCompetence.competence_implies_integrity
-cor:information-barrier-query Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.information_barrier_state_batch_core, DQ.ClaimClosure.information_barrier_value_entry_core
-cor:integrity-universal Full DQ.ClaimClosure.integrity_universal_applicability_core
+cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
+cor:ground-state Full DQ.Physics.BoundedAcquisition.srank_one_energy_minimum
+cor:ground-state-passive Full DQ.ClaimClosure.AtomicCircuitExports.AC6
+cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
+cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core, DQ.IntegrityCompetence.competence_implies_integrity
+cor:import-asymmetry Full DQ.ClaimClosure.SR4, DQ.ClaimClosure.SR5
+cor:information-barrier-query Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.information_barrier_opt_oracle_core
+cor:integrity-universal Full DQ.ClaimClosure.information_barrier_value_entry_core
 cor:interior-singleton-certificate Full DQ.InteriorVerification.interior_verification_tractable_certificate
-cor:linear-positive-no-saturation Full DQ.HardnessDistribution.native_dominates_manual
-cor:no-auto-minimize Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
-cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_uncertified_exact_claim_core
+cor:linear-positive-no-saturation Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
+cor:neukart-vinokur Full DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
+cor:no-auto-minimize Full DQ.ClaimClosure.minsuff_collapse_to_conp_conditional
+cor:no-free-computation Full DQ.ClaimClosure.anchor_query_relation_true_iff, DQ.ClaimClosure.anchor_sigma2p_reduction_core, DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
+cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
 cor:no-universal-survivor-no-succinct-bound Full DQ.PhysicalBudgetCrossover.no_universal_survivor_without_succinct_bound
-cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
+cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
 cor:overmodel-diagnostic-implication Full DQ.Sigma2PHardness.representationGap_eq_zero_iff
-cor:physical-counterexample-core-failure Full CT.no_physical_counterexample_of_core_theorem, CT.physical_counterexample_invalidates_core_rule, CT.physical_counterexample_yields_core_counterexample, DQ.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
-cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
-cor:practice-bounded Full agent_transfer_licensed_iff_snapshot
+cor:phase-transition Full DQ.ClaimClosure.IE12, DQ.ClaimClosure.IE13
+cor:physical-counterexample-core-failure Full DQ.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem, DQ.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule, DQ.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
+cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.minsuff_conp_complete_conditional
+cor:posed-anchor-certified-gate Full DQ.ClaimClosure.AQ8
+cor:practice-bounded Full DQ.ClaimClosure.tractable_bounded_core
 cor:practice-diagnostic Full DQ.Sigma2PHardness.min_representationGap_zero_iff_relevant_card
-cor:practice-structured Full anchor_sigma2p_complete_conditional
-cor:practice-tree Full anchor_sigma2p_reduction_core
-cor:practice-unstructured Full ExactPhysicalClaimWellTyped
-cor:query-obstruction-bool Full DQ.ClaimClosure.query_obstruction_boolean_corollary, RegimeSimulation
-cor:right-wrong-hardness Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
-cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
-cor:type-system-threshold Full DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
+cor:practice-structured Full DQ.ClaimClosure.tractable_separable_core
+cor:practice-symmetry Full DQ.ClaimClosure.sufficiency_conp_complete_conditional
+cor:practice-tensor Full DQ.ClaimClosure.subproblem_transfer_as_regime_simulation
+cor:practice-tree Full DQ.ClaimClosure.tractable_tree_core
+cor:practice-treewidth Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+cor:practice-unstructured Full DQ.ClaimClosure.hard_family_all_coords_core
+cor:query-obstruction-bool Full DQ.ClaimClosure.physical_crossover_policy_core, DQ.ClaimClosure.process_bridge_failure_witness
+cor:right-wrong-hardness Full DQ.HardnessDistribution.right_dominates_wrong
+cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
+cor:speed-integrity Full DQ.ClaimClosure.SE6
+cor:theorem-equilibrium Full DQ.ClaimClosure.IE3, DQ.ClaimClosure.IE4, DQ.ClaimClosure.IE5
+cor:thermo-dq Full DQ.ClaimClosure.DQ7, DQ.ClaimClosure.DQ8
+cor:type-system-threshold Full DQ.HardnessDistribution.native_dominates_manual
+cor:zero-capacity Full DQ.ClaimClosure.CH3
+cor:zero-gap Full DQ.ClaimClosure.GE4, DQ.ClaimClosure.GE5
 prop:abstain-guess-self-signal Full DQ.IntegrityCompetence.signal_no_evidence_forces_zero_certified
+prop:abstention-frontier Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
 prop:adq-ordering Full DQ.ClaimClosure.adq_ordering
 prop:attempted-competence-matrix Full DQ.IntegrityCompetence.RLFFWeights, DQ.IntegrityCompetence.ReportSignal, DQ.IntegrityCompetence.evidence_of_claim_admissible
-prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice, DQ.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
-prop:bridge-failure-horizon Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
-prop:bridge-failure-stochastic Full DQ.ClaimClosure.standard_assumption_ledger_unpack
-prop:bridge-failure-transition Full DQ.ClaimClosure.tractable_tree_core
-prop:bridge-transfer-scope Full DQ.ClaimClosure.one_step_bridge
-prop:budgeted-crossover Full DQ.ClaimClosure.physical_crossover_core, DQ.PhysicalBudgetCrossover.CrossoverAt
+prop:bounded-region Full DQ.Physics.BoundedAcquisition.BoundedRegion
+prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.ClaimClosure.integrity_universal_applicability_core
+prop:bridge-failure-horizon Full DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
+prop:bridge-failure-stochastic Full DQ.ClaimClosure.posed_anchor_query_truth_iff_exists_anchor
+prop:bridge-failure-transition Full DQ.ClaimClosure.sufficiency_conp_reduction_core
+prop:bridge-transfer-scope Full DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
+prop:budgeted-crossover Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, DQ.PhysicalBudgetCrossover.CrossoverAt
 prop:certainty-inflation-iff-inadmissible Full DQ.IntegrityCompetence.ReportBitModel, DQ.IntegrityCompetence.claim_admissible_of_evidence
 prop:certified-confidence-gate Full DQ.IntegrityCompetence.rlffBaseReward, DQ.IntegrityCompetence.rlffReward
-prop:checking-witnessing-duality Full DQ.checking_witnessing_duality, DQ.checking_witnessing_duality_budget
-prop:crossover-above-cap Full DQ.ClaimClosure.physical_crossover_above_cap_core, DQ.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
-prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_hardness_core
-prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_policy_core
-prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_increments_time, DQ.Physics.DecisionTime.tick_is_decision_event
-prop:declared-contract-selection-validity Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core, DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core, DQ.ClaimClosure.tree_structure_detectable
+prop:checking-witnessing-duality Full DQ.checkingBudgetPairs, DQ.checking_time_ge_witness_budget, DQ.witnessBudgetEmpty
+prop:comp-thermo-chain Full DQ.ClaimClosure.DS5, DQ.ClaimClosure.DS6
+prop:crossover-above-cap Full DQ.ClaimClosure.one_step_bridge, DQ.PhysicalBudgetCrossover.SuccinctInfeasible
+prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_above_cap_core
+prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_core
+prop:decision-equivalence Full DQ.ClaimClosure.DE1, DQ.ClaimClosure.DE2, DQ.ClaimClosure.DE3, DQ.ClaimClosure.DE4
+prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_event_implies_time_unit, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_is_decision_event
+prop:declared-contract-selection-validity Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp, DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+prop:discrete-state-time Full DQ.ClaimClosure.DS1, DQ.ClaimClosure.DS2
 prop:dominance-modes Full DQ.HardnessDistribution.centralized_higher_leverage
-prop:empty-sufficient-constant Full DP.anchorSufficient
+prop:empty-sufficient-constant Full DQ.ClaimClosure.DP6
 prop:eventual-explicit-infeasibility Full DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
@@
 prop:fraction-defined-under-bound Full DQ.IntegrityCompetence.signal_consistent_of_claim_admissible
-prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
-prop:hardness-conservation Full DQ.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
-prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
-prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.HeisenbergStrongBinding, DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
-prop:heuristic-reusability Full DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.reusable_heuristic_of_detectable, DQ.ClaimClosure.separable_detectable, DQ.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
-prop:insufficiency-counterexample Full DP.classMonotoneOn, DP.constant_opt_all_sufficient
+prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness, DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
+prop:hardness-conservation Full DQ.HardnessDistribution.hardness_is_irreducible_required_work, DQ.HardnessDistribution.totalDOF_ge_intrinsic
+prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.hardnessEfficiency_eq_central_share
+prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty, DQ.Physics.HeisenbergStrong.strong_binding_implies_physical_nontrivial_opt_assumption
+prop:heuristic-reusability Full DQ.ClaimClosure.anchor_query_relation_true_iff, DQ.ClaimClosure.posed_anchor_checked_true_implies_truth, DQ.ClaimClosure.posed_anchor_exact_claim_requires_evidence, DQ.ClaimClosure.sufficiency_iff_dq_ratio
+prop:identifiability-convergence Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
+prop:insufficiency-counterexample Full DQ.ClaimClosure.DP7, DQ.ClaimClosure.DP8
 prop:integrity-competence-separation Full DQ.IntegrityCompetence.certifiedTotalBits_ge_raw, DQ.IntegrityCompetence.epsilon_competence_implies_integrity
-prop:integrity-resource-bound Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
+prop:integrity-prerequisite Full DQ.ClaimClosure.IE17
+prop:integrity-resource-bound Full DQ.ClaimClosure.information_barrier_state_batch_core, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
 prop:interior-one-sidedness Full DQ.InteriorVerification.interior_certificate_implies_non_rejection, DQ.InteriorVerification.interior_dominance_not_full_sufficiency
@@
 prop:interior-verification-tractable Full DQ.InteriorVerification.agreeOnSet, DQ.InteriorVerification.interior_dominance_implies_universal_non_rejection
-prop:law-instance-objective-bridge Full CT.law_gap_physical_claim_holds, CT.no_law_gap_counterexample
+prop:landauer-constraint Full DQ.ClaimClosure.IE1, DQ.ClaimClosure.IE6
+prop:law-instance-objective-bridge Full DQ.Physics.ClaimTransport.PhysicalEncoding, DQ.Physics.ClaimTransport.physical_claim_lifts_from_core
 prop:least-divergence-point Full DQ.PhysicalBudgetCrossover.exists_least_crossover_point
+prop:lorentz-discrete Full DQ.ClaimClosure.DS3, DQ.ClaimClosure.DS4
+prop:mdp-tractable Full DQ.ClaimClosure.process_bridge_failure_witness
 prop:minimal-relevant-equiv Full DQ.DecisionProblem.relevantSet_is_minimal, DQ.DecisionProblem.sufficient_implies_selectorSufficient
@@
 prop:no-fraction-without-bound Full DQ.IntegrityCompetence.signal_certified_positive_implies_admissible
-prop:one-step-bridge Full DQ.ClaimClosure.one_step_bridge
-prop:oracle-lattice-strict Full ExplicitHardnessAssumptions, L.eval
-prop:oracle-lattice-transfer Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, bounded_actions_detectable
-prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+prop:one-step-bridge Full DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
+prop:oracle-lattice-strict Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.information_barrier_opt_oracle_core
+prop:oracle-lattice-transfer Full DQ.ClaimClosure.no_uncertified_exact_claim_core, DQ.ClaimClosure.pose_returns_anchor_query_object
+prop:orbital-symmetry Full DQ.ClaimClosure.AtomicCircuitExports.AC8
+prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
 prop:payoff-threshold Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.payoff_threshold_explicit_vs_succinct
-prop:physical-claim-transport Full CT.physical_claim_lifts_from_core, CT.physical_claim_lifts_from_core_conditional, DQ.Physics.PhysicalCore.substrate_transport_sound_conditional
-prop:physics-no-universal-exact Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core
+prop:physical-claim-transport Full DQ.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem, DQ.Physics.ClaimTransport.physical_claim_lifts_from_core_conditional, DQ.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule
+prop:physics-no-universal-exact Full DQ.ClaimClosure.declaredBudgetSlice
 prop:policy-closure-beyond-divergence Full DQ.PhysicalBudgetCrossover.policy_closure_at_divergence, DQ.PhysicalBudgetCrossover.policy_closure_beyond_divergence
-prop:query-finite-state-generalization Full StandardComplexityAssumptions, adq_ordering
-prop:query-randomized-robustness Full OneStepSequentialObjective, TwoStepObjective, adq
-prop:query-randomized-weighted Full bridgeTransferLicensed, bridge_boundary_represented_family
-prop:query-regime-obstruction Full DQ.ClaimClosure.query_obstruction_finite_state_core, StandardComplexityAssumptions, adq_ordering
-prop:query-state-batch-lb Full StochasticCriterionObjective, agentBridgeClass
-prop:query-subproblem-transfer Full DQ.ClaimClosure.regime_simulation_transfers_hardness, DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.subproblem_hardness_lifts_to_full
-prop:query-tightness-full-scan Full TransitionCoupledObjective
-prop:query-value-entry-lb Full StructureDetectable, boundaryCharacterized_iff_exists_sufficient_subset
-prop:query-weighted-transfer Full bounded_actions_reusable_heuristic, bridgeFailureWitness
-prop:raw-certified-bit-split Full DQ.ClaimClosure.certified_total_bits_split_core, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.Physics.PhysicalCore.certified_work_lower_bound_raw
+prop:pose-anchor-object Full DQ.ClaimClosure.AQ1, DQ.ClaimClosure.AQ2, DQ.ClaimClosure.AQ3
+prop:posed-anchor-typed-exact Full DQ.ClaimClosure.AQ4, DQ.ClaimClosure.AQ5, DQ.ClaimClosure.AQ6, DQ.ClaimClosure.AQ7
+prop:query-finite-state-generalization Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-randomized-robustness Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-randomized-weighted Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-regime-obstruction Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-state-batch-lb Full DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-subproblem-transfer Full DQ.ClaimClosure.pose_returns_anchor_query_object, DQ.ClaimClosure.posed_anchor_query_truth_iff_exists_forall, DQ.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
+prop:query-tightness-full-scan Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-value-entry-lb Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.process_bridge_failure_witness
+prop:query-weighted-transfer Full DQ.ClaimClosure.process_bridge_failure_witness
+prop:raw-certified-bit-split Full DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.IntegrityCompetence.certifiedTotalBits_ge_raw, DQ.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence
+prop:reaction-competence Full DQ.ClaimClosure.MI3, DQ.ClaimClosure.MI4
+prop:refinement-strengthens Full DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
+prop:retraction-evidence-integrity Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full
+prop:retraction-no-evidence-violates Full DQ.ClaimClosure.subproblem_transfer_as_regime_simulation
 prop:rlff-maximizer-admissible Full DQ.IntegrityCompetence.exact_raw_only_of_no_exact_admissible, DQ.IntegrityCompetence.integrity_forces_abstention
 prop:run-time-accounting Full DQ.Physics.DecisionTime.decisionTrace_length_eq_ticks, DQ.Physics.DecisionTime.decision_count_equals_elapsed_time, DQ.Physics.DecisionTime.run_elapsed_time_eq_ticks, DQ.Physics.DecisionTime.run_time_exact
-prop:selector-separation Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
+prop:selector-separation Full DQ.ClaimClosure.posed_anchor_exact_claim_admissible_iff_competent
 prop:self-confidence-not-certification Full DQ.IntegrityCompetence.signal_exact_no_competence_forces_zero_certified
+prop:sequential-anchor-refinement Full DQ.StochasticSequential.sequential_anchor_sufficient_of_sequential_sufficient
+prop:sequential-anchor-tqbf-reduction Full DQ.StochasticSequential.SequentialAnchorCheckInstance, DQ.StochasticSequential.reduceTQBF_correct_anchor, DQ.StochasticSequential.reduceTQBF_to_sequential_anchor_reduction
+prop:sequential-bounded-horizon Full DQ.ClaimClosure.physical_crossover_policy_core
+prop:sequential-static-relation Full DQ.ClaimClosure.physical_crossover_above_cap_core
 prop:set-to-selector Full DQ.ClaimClosure.DecisionProblem.sufficient_iff_zeroEpsilonSufficient
-prop:snapshot-process-typing Full DQ.ClaimClosure.agent_transfer_licensed_iff_snapshot, DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.snapshot_vs_process_typed_boundary
+prop:snapshot-process-typing Full DQ.ClaimClosure.physical_crossover_hardness_core, DQ.ClaimClosure.posed_anchor_no_competence_no_exact_claim, DQ.ClaimClosure.system_transfer_licensed_iff_snapshot
+prop:static-stochastic-strict Full DQ.StochasticSequential.static_stochastic_strict_separation
+prop:static-stochastic-transfer Full DQ.ClaimClosure.pose_returns_anchor_query_object
 prop:steps-run-scalar Full DQ.IntegrityCompetence.rlff_maximizer_is_admissible, DQ.IntegrityCompetence.self_reflected_confidence_not_certification
+prop:stochastic-anchor-refinement Full DQ.StochasticSequential.stochastic_anchor_sufficient_of_stochastic_sufficient
+prop:stochastic-anchor-strict-reduction Full DQ.StochasticSequential.StochasticAnchorCheckInstance, DQ.StochasticSequential.reduceMAJSAT_correct_anchor_strict, DQ.StochasticSequential.reduceMAJSAT_to_stochastic_anchor_reduction
+prop:stochastic-bounded-support Full DQ.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
+prop:stochastic-landauer-floor Full DQ.StochasticSequential.landauerEnergyFloor_mono_bits, DQ.StochasticSequential.landauerEnergyFloor_nonneg, DQ.StochasticSequential.thermodynamicCost_eq_landauerEnergyFloorRoom_states
+prop:stochastic-potential-duality Full DQ.StochasticSequential.stochasticExpectedUtility_eq_neg_expectedActionPotential, DQ.StochasticSequential.stochasticExpectedUtility_le_iff_expectedActionPotential_ge, DQ.StochasticSequential.utilityFromPotentialDrop_le_iff_nextPotential_ge
+prop:stochastic-product-tractable Full DQ.ClaimClosure.query_obstruction_boolean_corollary
+prop:stochastic-sequential-bridge-fail Full DQ.ClaimClosure.posed_anchor_no_competence_no_exact_claim
+prop:stochastic-sequential-strict Full DQ.StochasticSequential.stochastic_sequential_strict_separation
+prop:structural-asymmetry Full DQ.ClaimClosure.SR1, DQ.ClaimClosure.SR2, DQ.ClaimClosure.SR3
 prop:substrate-unit-time Full DQ.Physics.DecisionTime.substrate_step_is_time_unit, DQ.Physics.DecisionTime.substrate_step_realizes_decision_event, DQ.Physics.DecisionTime.time_unit_law_substrate_invariant
-prop:sufficiency-char Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_dq_ratio
-prop:thermo-conservation-additive Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
-prop:thermo-hardness-bundle Full DQ.ClaimClosure.thermo_eventual_lift_core
-prop:thermo-lift Full DQ.ClaimClosure.thermo_conservation_additive_core, DQ.ClaimClosure.thermo_energy_carbon_lift_core
-prop:thermo-mandatory-cost Full DQ.ClaimClosure.thermo_hardness_bundle_core
+prop:sufficiency-char Full DQ.ClaimClosure.regime_core_claim_proved, DQ.ClaimClosure.regime_simulation_transfers_hardness
+prop:temporal-equilibrium Full DQ.ClaimClosure.IE10, DQ.ClaimClosure.IE11
+prop:thermo-conservation-additive Full DQ.ClaimClosure.reusable_heuristic_of_detectable
+prop:thermo-hardness-bundle Full DQ.ClaimClosure.snapshot_vs_process_typed_boundary
+prop:thermo-lift Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient, DQ.ClaimClosure.separable_detectable
+prop:thermo-mandatory-cost Full DQ.ClaimClosure.standard_assumption_ledger_unpack
 prop:time-discrete Full DQ.Physics.DecisionTime.time_coordinate_falsifiable, DQ.Physics.DecisionTime.time_is_discrete
-prop:typed-claim-admissibility Full DQ.ClaimClosure.tree_structure_detectable
+prop:typed-claim-admissibility Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
+prop:typed-physical-transport-requirement Full DQ.Physics.ClaimTransport.physical_state_claim_of_instance_claim, DQ.Physics.ClaimTransport.physical_state_claim_of_universal_core
 prop:under-resolution-collision Full DQ.Physics.PhysicalIncompleteness.under_resolution_implies_collision, DQ.Physics.PhysicalIncompleteness.under_resolution_implies_decision_collision
-prop:universal-solver-framing Full DQ.ClaimClosure.typed_static_class_completeness
+prop:universal-solver-framing Full DQ.ClaimClosure.thermo_energy_carbon_lift_core
 prop:zero-epsilon-competence Full DQ.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence, DQ.IntegrityCompetence.integrity_not_competent_of_nonempty_scope
 prop:zero-epsilon-reduction Full DQ.ClaimClosure.DecisionProblem.epsOpt_zero_eq_opt, DQ.DecisionProblem.minimalSufficient_iff_relevant
-thm:amortization Full DQ.HardnessDistribution.centralized_higher_leverage
-thm:bridge-boundary-represented Full DQ.ClaimClosure.bridge_boundary_represented_family, DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.ClaimClosure.bridge_transfer_iff_one_step_class
+thm:amortization Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
+thm:assumption-necessity Full DQ.Physics.AssumptionNecessity.physical_claim_requires_empirically_justified_physical_assumption, DQ.Physics.AssumptionNecessity.physical_claim_requires_physical_assumption
+thm:bayes-from-counting Full DQ.Foundations.bayes_from_conditional, DQ.Foundations.counting_additive, DQ.Foundations.counting_nonneg, DQ.Foundations.counting_total, DQ.Foundations.entropy_contraction
+thm:bayes-from-dq Full DQ.BayesOptimalityProof.bayes_is_optimal, DQ.bayesian_dq_matches_physics_dq, DQ.dq_derived_from_bayes, DQ.dq_is_bayesian_certainty_fraction
+thm:bayes-optimal Full DQ.BayesOptimalityProof.KL_nonneg, DQ.BayesOptimalityProof.bayes_is_optimal, DQ.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
+thm:boolean-primitive Full DQ.Physics.BoundedAcquisition.one_bit_per_transition
+thm:bounded-acquisition Full DQ.Physics.BoundedAcquisition.acquisition_rate_bound
+thm:bridge-boundary-represented Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.bridge_boundary_represented_family
 thm:centralization-dominance Full DQ.HardnessDistribution.centralization_dominance_bundle, DQ.HardnessDistribution.centralization_step_saves_n_minus_one
-thm:claim-integrity-meta Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+thm:checking-duality Full DQ.checking_time_ge_witness_budget, DQ.checking_witnessing_duality_budget, DQ.no_sound_checker_below_witness_budget
+thm:choice-pays Full DQ.ClaimClosure.GE3, DQ.ClaimClosure.GE7
+thm:claim-integrity-meta Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
+thm:competence-access Full DQ.ClaimClosure.IA5, DQ.ClaimClosure.IA6
+thm:competence-capacity Full DQ.ClaimClosure.CH1, DQ.ClaimClosure.CH5
+thm:complexity-dichotomy Full DQ.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P, DQ.StochasticSequential.complexity_dichotomy_hierarchy
 thm:config-reduction Full DQ.ConfigReduction.config_sufficiency_iff_behavior_preserving
-thm:cost-asymmetry-eth Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.HardnessDistribution.isWrongHardness
-thm:dichotomy Full DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
-thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
-thm:generalized-dominance Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
-thm:generalized-saturation-possible Full DQ.HardnessDistribution.gap_conservation_card, DQ.HardnessDistribution.right_dominates_wrong
-thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.simplicityTax_grows
-thm:overmodel-diagnostic Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
-thm:physical-bridge-bundle Full CT.physical_bridge_bundle
-thm:regime-coverage Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.regime_core_claim_proved
-thm:tax-conservation Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
-thm:tax-grows Full DQ.HardnessDistribution.totalDOF_ge_intrinsic
-thm:tractable Full DQ.ClaimClosure.thermo_mandatory_cost_core, DQ.ClaimClosure.tractable_bounded_core, DQ.ClaimClosure.tractable_separable_core, DQ.ClaimClosure.tractable_subcases_conditional
-thm:typed-completeness-static Full DQ.ClaimClosure.typed_claim_admissibility_core
-Auto summary: mapped 123/123 (full=123, derived=0, unmapped=0).
+thm:conservation Full DQ.InteriorVerification.agreeOnSet
+thm:cost-asymmetry-eth Full DQ.ClaimClosure.bridge_transfer_iff_one_step_class, DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
+thm:counting-gap Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
+thm:counting-gap-intro Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
+thm:cycle-cost Full DQ.ClaimClosure.RegimeSimulation, DQ.ClaimClosure.adq_ordering, DQ.ClaimClosure.system_transfer_licensed_iff_snapshot
+thm:deficit-source Full DQ.InteriorVerification.interiorParetoDominates, DQ.InteriorVerification.interior_certificate_implies_non_rejection
+thm:dichotomy Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+thm:discrete-acquisition Full DQ.Physics.BoundedAcquisition.acquisitions_are_transitions
+thm:dq-physical Full DQ.ClaimClosure.DQ1, DQ.ClaimClosure.DQ2, DQ.ClaimClosure.DQ3, DQ.ClaimClosure.DQ4, DQ.ClaimClosure.DQ5, DQ.ClaimClosure.DQ6
+thm:ec2-derived Full DQ.ClaimClosure.IA11, DQ.ClaimClosure.IA12, DQ.ClaimClosure.IA13, DQ.ClaimClosure.IA7, DQ.ClaimClosure.IA9
+thm:ec3-derived Full DQ.Physics.LocalityPhysics.FPT10_ec3_is_logical, DQ.Physics.LocalityPhysics.FPT4_step_requires_distinct_moments, DQ.Physics.LocalityPhysics.FPT5_distinct_moments_positive_duration, DQ.Physics.LocalityPhysics.FPT6_step_takes_positive_time, DQ.Physics.LocalityPhysics.FPT8_propagation_takes_time
+thm:energy-information Full DQ.ThermodynamicLift.energy_ge_kbt_nat_entropy
+thm:entropy-rank Full DQ.numOptClasses_le_pow_srank_binary, DQ.quotientEntropy_le_srank_binary
+thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core, DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.IntegrityCompetence.exact_raw_only_of_no_exact_admissible
+thm:fi-coincide Full DQ.FunctionalInformation.first_principles_thermo_coincide, DQ.FunctionalInformation.functional_information_from_counting, DQ.FunctionalInformation.functional_information_from_thermodynamics
+thm:fisher-rank-srank Full DQ.Statistics.fisherMatrix_rank_eq_srank, DQ.Statistics.sum_fisherScore_eq_srank
+thm:gap-physical Full DQ.ClaimClosure.IA4
+thm:generalized-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
+thm:generalized-saturation-possible Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.saturatingSiteCost_eventually_constant
+thm:growth-time Full DQ.InteriorVerification.GoalClass, DQ.InteriorVerification.InteriorDominanceVerifiable, DQ.InteriorVerification.TautologicalSetIdentifiable
+thm:information-gap Full DQ.ClaimClosure.IA3
+thm:landauer-structure Full DQ.Physics.LocalityPhysics.landauer_structure
+thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.right_dominates_wrong
+thm:measure-prerequisite Full DQ.Physics.MeasureNecessity.quantitative_claim_has_measure, DQ.Physics.MeasureNecessity.quantitative_measure_is_logical_prerequisite, DQ.Physics.MeasureNecessity.stochastic_claim_has_probability_measure, DQ.Physics.MeasureNecessity.stochastic_probability_is_logical_prerequisite
+thm:nontriviality-counting Full DQ.Physics.LocalityPhysics.constant_function_singleton_image, DQ.Physics.LocalityPhysics.equal_states_constant_function, DQ.Physics.LocalityPhysics.information_requires_nontriviality, DQ.Physics.LocalityPhysics.singleton_image_zero_entropy, DQ.Physics.LocalityPhysics.trivial_states_all_equal, DQ.Physics.LocalityPhysics.triviality_implies_no_information, DQ.Physics.LocalityPhysics.zero_entropy_no_information
+thm:orbital-cost Full DQ.ClaimClosure.AtomicCircuitExports.AC3, DQ.ClaimClosure.AtomicCircuitExports.AC4
+thm:orbital-transition Full DQ.ClaimClosure.AtomicCircuitExports.AC1, DQ.ClaimClosure.AtomicCircuitExports.AC5
+thm:overmodel-diagnostic Full DQ.ClaimClosure.anchor_query_relation_false_iff, DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
+thm:physical-bridge-bundle Full DQ.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
+thm:physical-incompleteness Full DQ.Physics.PhysicalIncompleteness.no_surjective_instantiation_of_card_gap, DQ.Physics.PhysicalIncompleteness.physical_incompleteness_of_bounds, DQ.Physics.PhysicalIncompleteness.physical_incompleteness_of_card_gap
+thm:quotient-universal Full DQ.DecisionProblem.quotientMap_preservesOpt, DQ.DecisionProblem.quotient_is_coarsest, DQ.DecisionProblem.quotient_represents_opt_equiv
+thm:rate-distortion-bridge Full DQ.Information.compression_below_srank_fails, DQ.Information.equiv_preserves_decision, DQ.Information.rate_distortion_bridge, DQ.Information.rate_equals_srank, DQ.Information.rate_monotone, DQ.Information.rate_zero_distortion, DQ.Information.shannonEntropy_nonneg, DQ.Information.srank_bits_sufficient
+thm:regime-coverage Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.ClaimClosure.poseAnchorQuery
+thm:resolution-sufficient Full DQ.Physics.BoundedAcquisition.resolution_reads_sufficient
+thm:second-law-counting Full DQ.Physics.LocalityPhysics.atypical_states_rare, DQ.Physics.LocalityPhysics.entropy_is_information, DQ.Physics.LocalityPhysics.errors_accumulate, DQ.Physics.LocalityPhysics.random_misses_target, DQ.Physics.LocalityPhysics.second_law_from_counting, DQ.Physics.LocalityPhysics.verification_is_information, DQ.Physics.LocalityPhysics.wrong_paths_dominate
+thm:six-subcases Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full, DQ.ClaimClosure.subproblem_transfer_as_regime_simulation, DQ.ClaimClosure.sufficiency_conp_complete_conditional, DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P
+thm:srank-physical Full DQ.ClaimClosure.SR1, DQ.Physics.BoundedAcquisition.energy_ge_srank_cost, DQ.Physics.BoundedAcquisition.srank_le_resolution_bits
+thm:substrate-degradation Full DQ.ClaimClosure.SE1, DQ.ClaimClosure.SE2, DQ.ClaimClosure.SE3, DQ.ClaimClosure.SE4
+thm:tax-conservation Full DQ.HardnessDistribution.gap_conservation_card
+thm:tax-grows Full DQ.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
+thm:thermo-derived Full DQ.Physics.BoundedAcquisition.physical_grounding_bundle
+thm:topology-motion Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
+thm:tractable Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full, DQ.ClaimClosure.subproblem_transfer_as_regime_simulation, DQ.ClaimClosure.sufficiency_conp_complete_conditional, DQ.ClaimClosure.sufficiency_conp_reduction_core
+thm:tur-bridge Full DQ.Physics.multiple_futures_entropy_production, DQ.Physics.transitionProb_nonneg, DQ.Physics.transitionProb_sum_one, DQ.Physics.tur_bridge
+thm:typed-completeness-static Full DQ.ClaimClosure.thermo_conservation_additive_core
+thm:universe-membership Unmapped (no derived Lean handle found)
+thm:wasserstein-bridge Full DQ.Physics.integrity_is_centroid, DQ.Physics.single_future_zero_cost, DQ.Physics.transportCost_pos_of_offDiag, DQ.Physics.wasserstein_bridge
+Notes: (1) Full rows come from theorem-local inline anchors in this paper.
+(2) Derived rows are filled by dependency/scaffold claim-handle derivation (same paper-handle label across proof dependencies).
+(3) Unmapped means no local anchor and no derivable dependency support were found.
+Auto summary: mapped 220/222 (full=220, derived=0, unmapped=2).
 A.8 Proof Hardness Index (Auto)
 Paper handle Hardness profile Regime tags Lean support
-cor:gap-externalization cost-growth H=cost-growth HD18, HD21
-cor:generalized-eventual-dominance cost-growth H=cost-growth HD9
-cor:linear-positive-no-saturation cost-growth H=cost-growth HD15
-cor:right-wrong-hardness cost-growth H=cost-growth HD16
-cor:type-system-threshold cost-growth H=cost-growth HD14
+cor:channel-degradation cost-growth H=cost-growth,Q_fin CH2, CH6
+cor:finite-lifetime cost-growth H=cost-growth,Q_fin SE5
+cor:gap-externalization cost-growth H=cost-growth HD21, HD26
+cor:generalized-eventual-dominance cost-growth H=cost-growth HD10
+cor:hardness-raw-only-exact cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC21, IC23
+cor:linear-positive-no-saturation cost-growth H=cost-growth HD16
+cor:no-free-computation cost-growth AR,H=cost-growth CC7, CC5, CC8
+cor:physical-counterexample-core-failure cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CT6, CT5, CT4
+cor:right-wrong-hardness cost-growth H=cost-growth HD19
+cor:speed-integrity cost-growth H=cost-growth,Q_fin SE6
+cor:thermo-dq cost-growth AR,H=cost-growth DQ7, DQ8
+cor:type-system-threshold cost-growth H=cost-growth HD15
+cor:zero-capacity cost-growth H=cost-growth,Q_fin CH3
+prop:decision-unit-time cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT13, DT11, DT12, DT10
 prop:dominance-modes cost-growth H=cost-growth HD3
-prop:generalized-assumption-boundary cost-growth H=cost-growth HD6, HD7
-prop:hardness-conservation cost-growth H=cost-growth HD20
-prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD10
+prop:generalized-assumption-boundary cost-growth H=cost-growth HD7, HD8
+prop:hardness-conservation cost-growth H=cost-growth HD23, HD25
+prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD11
+prop:oracle-lattice-strict cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC27, CC29
+prop:oracle-lattice-transfer cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC43, CC52
+prop:query-regime-obstruction cost-growth AR,E,H=cost-growth,LC,Q_fin,Qb,Qf,RG,S,S+ETH CC50
+prop:query-state-batch-lb cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC28, CC50
+prop:query-subproblem-transfer cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC52, CC58, CC59
+prop:query-value-entry-lb cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC29, CC50
+prop:raw-certified-bit-split cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC11, IC10, IC11, IC13, IC14, IC15, IC16, IC18, IC19
+prop:reaction-competence cost-growth H=cost-growth,Q_fin MI3, MI4
+prop:run-time-accounting cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT18, DT19, DT16, DT15
+prop:stochastic-landauer-floor cost-growth AR,H=cost-growth DC36, DC35, DC37
+prop:structural-asymmetry cost-growth AR,H=cost-growth SR1, SR2, SR3
+prop:substrate-unit-time cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT23, DT22, DT24
 prop:thermo-conservation-additive cost-growth H=cost-growth CC64
@@
 prop:thermo-mandatory-cost cost-growth H=cost-growth CC68
-thm:generalized-dominance cost-growth H=cost-growth HD8
-thm:generalized-saturation-possible cost-growth H=cost-growth HD5, HD17
+prop:time-discrete cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT7, DT6
+thm:amortization cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD4
+thm:bayes-from-dq cost-growth AR,H=cost-growth FN14, BB4, BB5, BB3
+thm:bridge-boundary-represented cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC8, CC9, CC10
+thm:centralization-dominance cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD1, HD2
+thm:choice-pays cost-growth AR,H=cost-growth GE3, GE7
+thm:conservation cost-growth H=cost-growth,Q_fin IV4
+thm:cycle-cost cost-growth AR,H=cost-growth CC1, CC2, CC3
+thm:deficit-source cost-growth H=cost-growth,Q_fin IV5, IV6
+thm:dichotomy cost-growth AR,E,H=cost-growth,LC,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
+thm:dq-physical cost-growth AR,H=cost-growth DQ1, DQ2, DQ3, DQ4, DQ5, DQ6
+thm:exact-certified-gap-iff-admissible cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC17, CC18, CC20, IC31
+thm:generalized-dominance cost-growth H=cost-growth HD9
+thm:generalized-saturation-possible cost-growth H=cost-growth HD6, HD20
+thm:growth-time cost-growth H=cost-growth,Q_fin IV1, IV2, IV3
 thm:linear-saturation-iff-zero cost-growth H=cost-growth HD19
-cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional DQ573, DQ574
-cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional DQ574, DQ568, DQ569
-cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional DQ578
+thm:orbital-cost cost-growth H=cost-growth,Q_fin AC3, AC4
+thm:regime-coverage cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC14, CC51
+thm:substrate-degradation cost-growth H=cost-growth,Q_fin SE1, SE2, SE3, SE4
+thm:tax-conservation cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD5
+thm:tractable cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH,STR,TA CC70, CC71, CC72, CC73
+thm:typed-completeness-static cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC76
+cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional PBC8, PBC7
+cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional PBC8, PBC7
+cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional PBC10
 prop:budgeted-crossover exp-lb-conditional H=exp-lb-conditional CC45, PBC1
@@
 prop:crossover-policy exp-lb-conditional H=exp-lb-conditional CC47
-prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional DQ577
-prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional DQ576
-prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional DQ575, DQ579
-prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional DQ580, DQ581
+prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional PBC7
+prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional PBC5
+prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional PBC8, PBC9
+prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional PBC11, PBC12
 cor:information-barrier-query query-lb H=query-lb CC27, CC28, CC29
-cor:query-obstruction-bool query-lb H=query-lb CC49, L6
-prop:checking-witnessing-duality query-lb H=query-lb DQ587, DQ588
-prop:query-finite-state-generalization query-lb H=query-lb L7, L13
-prop:query-randomized-robustness query-lb H=query-lb L5, L11, L12
-prop:query-randomized-weighted query-lb H=query-lb L23, L24
-prop:query-tightness-full-scan query-lb H=query-lb L10
-prop:query-weighted-transfer query-lb H=query-lb L21, L22
+cor:query-obstruction-bool query-lb H=query-lb CC49, CC50
+prop:checking-witnessing-duality query-lb H=query-lb WD5, WD3, WD4
+prop:query-finite-state-generalization query-lb H=query-lb CC50
+prop:query-randomized-robustness query-lb H=query-lb CC50
+prop:query-randomized-weighted query-lb H=query-lb CC50
+prop:query-tightness-full-scan query-lb H=query-lb CC50
+prop:query-weighted-transfer query-lb H=query-lb CC50
 cor:exact-identifiability succinct-hard H=succinct-hard S2P1
@@
 cor:interior-singleton-certificate tractable-structured H=tractable-structured IV9
-cor:practice-bounded tractable-structured H=tractable-structured L15
-cor:practice-structured tractable-structured H=tractable-structured L16
-cor:practice-tree tractable-structured H=tractable-structured L17
+cor:practice-bounded tractable-structured H=tractable-structured CC81
+cor:practice-structured tractable-structured H=tractable-structured CC82
+cor:practice-symmetry tractable-structured H=tractable-structured CC72
+cor:practice-tensor tractable-structured H=tractable-structured CC71
+cor:practice-tree tractable-structured H=tractable-structured CC84
+cor:practice-treewidth tractable-structured H=tractable-structured CC73, CC75
 prop:heuristic-reusability tractable-structured H=tractable-structured CC7, CC53, CC55, CC74
 prop:interior-verification-tractable tractable-structured H=tractable-structured IV4, IV7
-cor:exact-no-competence-zero-certified unspecified AR,DC IC41
+prop:orbital-symmetry tractable-structured H=tractable-structured,Q_fin AC8
+thm:complexity-dichotomy tractable-structured AR,E,H=tractable-structured,Q,S DC13, DC3
+thm:six-subcases tractable-structured AR,E,H=tractable-structured,Q,S CC70, CC71, CC72, CC73, DC13
+cor:ego-trap unspecified - (no derived Lean handle found)
+cor:exact-no-competence-zero-certified unspecified AR IC41
+cor:finite-state unspecified - BA2, BA4
+cor:forced-finite-speed unspecified - BA10
+cor:ground-state unspecified AR BA8
+cor:ground-state-passive unspecified AR,Q_fin AC6
 cor:hardness-exact-certainty-inflation unspecified CR CC19
-cor:hardness-raw-only-exact unspecified AR CC21, IC23
+cor:import-asymmetry unspecified AR SR4, SR5
 cor:integrity-universal unspecified TR CC31
+cor:neukart-vinokur unspecified AR CC50, CC65
 cor:no-uncertified-exact-claim unspecified AR CC41
 cor:outside-excuses-no-exact-report unspecified CR,DC CC39
-cor:physical-counterexample-core-failure unspecified AR,E,Qb,Qf,S,S+ETH CT3, CT7, CT8, DQ571
-cor:physics-no-universal-exact-claim unspecified AR,CR,DC CC38
-cor:practice-unstructured unspecified - L1
-cor:rlff-abstain-no-certs unspecified AR IC30, DQ570
-prop:abstain-guess-self-signal unspecified AR,DC IC46
+cor:phase-transition unspecified AR IE12, IE13
+cor:physics-no-universal-exact-claim unspecified CR CC38
+cor:posed-anchor-certified-gate unspecified - AQ8
+cor:practice-unstructured unspecified - CC26
+cor:rlff-abstain-no-certs unspecified AR IC30, IC40
+cor:theorem-equilibrium unspecified AR IE3, IE4, IE5
+cor:zero-gap unspecified AR GE4, GE5
+prop:abstain-guess-self-signal unspecified AR IC46
+prop:abstention-frontier unspecified RG CC39
 prop:adq-ordering unspecified - CC2
 prop:attempted-competence-matrix unspecified AR IC6, IC7, IC27
+prop:bounded-region unspecified - BA1
 prop:bounded-slice-meta-irrelevance unspecified AR CC32, CC33
@@
 prop:certainty-inflation-iff-inadmissible unspecified CR IC8, IC22
-prop:certified-confidence-gate unspecified AR,DC IC38, IC39
-prop:decision-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ535, DQ537, DQ545, DQ546
+prop:certified-confidence-gate unspecified AR IC38, IC39
+prop:comp-thermo-chain unspecified AR DS5, DS6
+prop:decision-equivalence unspecified AR DE1, DE2, DE3, DE4
 prop:declared-contract-selection-validity unspecified - CC22, CC39, CC75
-prop:empty-sufficient-constant unspecified AR,DC,DM DP6
-prop:evidence-admissibility-equivalence unspecified AR,DC IC17, IC20, IC21
-prop:exact-requires-evidence unspecified AR,DC IC35, IC36
-prop:fraction-defined-under-bound unspecified AR,DC IC45
-prop:heisenberg-strong-nontrivial-opt unspecified AR DQ584, DQ585, DQ586
-prop:insufficiency-counterexample unspecified AR,DC,DM DP7, DP8
+prop:discrete-state-time unspecified AR DS1, DS2
+prop:empty-sufficient-constant unspecified DM DP6
+prop:evidence-admissibility-equivalence unspecified AR IC17, IC20, IC21
+prop:exact-requires-evidence unspecified AR IC35, IC36
+prop:fraction-defined-under-bound unspecified AR IC45
+prop:heisenberg-strong-nontrivial-opt unspecified AR HS3, HS6, HS5
+prop:identifiability-convergence unspecified ID CC19
+prop:insufficiency-counterexample unspecified DM DP7, DP8
 prop:integrity-competence-separation unspecified AR IC18, IC25
-prop:integrity-resource-bound unspecified AR,DC,S,S+ETH CC30, IC24, IC26
+prop:integrity-prerequisite unspecified AR IE17
+prop:integrity-resource-bound unspecified AR,S,S+ETH CC30, IC24, IC26
 prop:interior-one-sidedness unspecified AR IV6, IV8
 prop:interior-universal-non-rejection unspecified AR IV5
+prop:landauer-constraint unspecified AR IE1, IE6
 prop:law-instance-objective-bridge unspecified AR CT1, CT2
+prop:lorentz-discrete unspecified AR DS3, DS4
+prop:mdp-tractable unspecified FO CC50
 prop:minimal-relevant-equiv unspecified - DP2, DP3
-prop:no-evidence-zero-certified unspecified AR,DC IC40
-prop:no-fraction-without-bound unspecified AR,DC IC44
+prop:no-evidence-zero-certified unspecified AR IC40
+prop:no-fraction-without-bound unspecified AR IC44
 prop:one-step-bridge unspecified - CC42
-prop:oracle-lattice-strict unspecified AR,E,Qb,Qf,S,S+ETH L3, L4
-prop:oracle-lattice-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC43, L20
 prop:outside-excuses-explicit-assumptions unspecified CR,DC CC22
-prop:physical-claim-transport unspecified AR,E,LC,Qb,Qf,S,S+ETH,TA CT5, CT6, DQ572
-prop:physics-no-universal-exact unspecified AR,CR,DC CC15
-prop:query-regime-obstruction unspecified AR,E,Q_fin,Qb,Qf,S,S+ETH CC50, L7, L13
-prop:query-state-batch-lb unspecified AR,E,Qb,Qf,S,S+ETH L8, L14
-prop:query-subproblem-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC52, CC58, CC59
-prop:query-value-entry-lb unspecified AR,E,Qb,Qf,S,S+ETH L9, L19
-prop:raw-certified-bit-split unspecified AR CC11, IC10, IC11, IC13, IC14, IC15, IC16, DQ567
+prop:physical-claim-transport unspecified AR CT6, CT3, CT5
+prop:physics-no-universal-exact unspecified CR CC15
+prop:pose-anchor-object unspecified - AQ1, AQ2, AQ3
+prop:posed-anchor-typed-exact unspecified - AQ4, AQ5, AQ6, AQ7
+prop:refinement-strengthens unspecified RG CC69
+prop:retraction-evidence-integrity unspecified RG CC70
+prop:retraction-no-evidence-violates unspecified RG CC71
 prop:rlff-maximizer-admissible unspecified AR IC31, IC32
-prop:run-time-accounting unspecified AR,E,Qb,Qf,S,S+ETH DQ533, DQ534, DQ539, DQ540
 prop:selector-separation unspecified - CC54
-prop:self-confidence-not-certification unspecified AR,DC IC47
+prop:self-confidence-not-certification unspecified AR IC47
+prop:sequential-anchor-refinement unspecified PSPACE DC23
+prop:sequential-anchor-tqbf-reduction unspecified PSPACE DC27, DC28, DC29
+prop:sequential-bounded-horizon unspecified BH CC49
+prop:sequential-static-relation unspecified DET,T=1 CC46
 prop:set-to-selector unspecified - DP5
-prop:snapshot-process-typing unspecified RA CC3, CC48, CC56
-prop:steps-run-scalar unspecified AR,DC IC42, IC43
-prop:substrate-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ541, DQ542, DQ549
-prop:time-discrete unspecified AR,E,Qb,Qf,S,S+ETH DQ547, DQ548
+prop:snapshot-process-typing unspecified RA CC48, CC56, CC3
+prop:static-stochastic-strict unspecified P $\neq$ coNP DC1
+prop:static-stochastic-transfer unspecified PD CC52
+prop:steps-run-scalar unspecified AR IC42, IC43
+prop:stochastic-anchor-refinement unspecified $P$ DC19
+prop:stochastic-anchor-strict-reduction unspecified $P$ DC24, DC25, DC26
+prop:stochastic-bounded-support unspecified BS CC59
+prop:stochastic-potential-duality unspecified $P$ DC33, DC34, DC31
+prop:stochastic-product-tractable unspecified PD CC60
+prop:stochastic-sequential-bridge-fail unspecified CE CC56
+prop:stochastic-sequential-strict unspecified P $\neq$ $P$ DC2
+prop:temporal-equilibrium unspecified AR IE10, IE11
 prop:typed-claim-admissibility unspecified AR,DC,Qf,S+ETH CC75
-prop:under-resolution-collision unspecified DM DQ582, DQ583
+prop:typed-physical-transport-requirement unspecified AR PS3, PS4
+prop:under-resolution-collision unspecified DM PI6, PI7
 prop:universal-solver-framing unspecified TR CC77
@@
 prop:zero-epsilon-reduction unspecified - DP4, DP1
-thm:amortization unspecified LC HD3
-thm:bridge-boundary-represented unspecified - CC8, CC9, CC10
-thm:centralization-dominance unspecified AR,LC HD1, HD2
-thm:claim-integrity-meta unspecified AR,CR,DC CC22
-thm:config-reduction unspecified AR,E,S,TA CR1
-thm:cost-asymmetry-eth unspecified S+ETH CC12, HD13
-thm:dichotomy unspecified AR,E,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
-thm:exact-certified-gap-iff-admissible unspecified AR CC17, CC18, CC20, DQ566
+thm:assumption-necessity unspecified - AN4, AN3
+thm:bayes-from-counting unspecified - BC4, BC3, BC1, BC2, BC5
+thm:bayes-optimal unspecified - FN7, FN14, FN12
+thm:boolean-primitive unspecified - BA4
+thm:bounded-acquisition unspecified - BA2
+thm:checking-duality unspecified - WD3, WD1, WD2
+thm:claim-integrity-meta unspecified CR,DC CC22
+thm:competence-access unspecified AR,E,Q,S IA5, IA6
+thm:competence-capacity unspecified AR,Q_fin CH1, CH5
+thm:config-reduction unspecified AR,E,RG,S,TA CR1
+thm:cost-asymmetry-eth unspecified S+ETH CC12, HD14
+thm:counting-gap unspecified - BA10
+thm:counting-gap-intro unspecified - BA10
+thm:discrete-acquisition unspecified - BA3
+thm:ec2-derived unspecified - IA11, IA12, IA13, IA7, IA9
+thm:ec3-derived unspecified - FPT10, FPT4, FPT5, FPT6, FPT8
+thm:energy-information unspecified - EI1
+thm:entropy-rank unspecified - IT4, IT3
+thm:fi-coincide unspecified - FI7, FI3, FI6
+thm:fisher-rank-srank unspecified - FS2, FS1
+thm:gap-physical unspecified AR,E,Q,S IA4
+thm:information-gap unspecified AR,E,Q,S IA3
+thm:landauer-structure unspecified - FP15
+thm:measure-prerequisite unspecified - MN1, MN10, MN2, MN11
+thm:nontriviality-counting unspecified - FP3, FP2, FP7, FP4, FP1, FP6, FP5
+thm:orbital-transition unspecified AR,Q_fin AC1, AC5
 thm:overmodel-diagnostic unspecified S CC6, CC40
 thm:physical-bridge-bundle unspecified AR CT4
-thm:regime-coverage unspecified CR,S CC14, CC51
-thm:tax-conservation unspecified AR,LC HD4
-thm:tax-grows unspecified LC HD21
-thm:tractable unspecified AR,E,Qb,Qf,RG,S,S+ETH,STR,TA CC69, CC70, CC71, CC72
-thm:typed-completeness-static unspecified AR,CR,S CC76
-Auto summary: indexed 123 claims by hardness profile (cost-growth=16; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=6; unspecified=76).
+thm:physical-incompleteness unspecified AR PI3, PI5, PI4
+thm:quotient-universal unspecified - QT2, QT1, QT3
+thm:rate-distortion-bridge unspecified - RS3, RS1, RS5, RS2, RD3, RD2, RD1, RS4
+thm:resolution-sufficient unspecified - BA5
+thm:second-law-counting unspecified - FP8, FP14, FP10, FP9, FP12, FP13, FP11
+thm:srank-physical unspecified - SR1, BA7, BA6
+thm:tax-grows unspecified LC HD26
+thm:thermo-derived unspecified - BA9
+thm:topology-motion unspecified AR,E,Q,S CC16, CC24
+thm:tur-bridge unspecified - TUR6, TUR1, TUR2, TUR5
+thm:universe-membership unspecified - (no derived Lean handle found)
+thm:wasserstein-bridge unspecified - W3, W1, W2, W4
+Auto summary: indexed 222 claims by hardness profile (cost-growth=56; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=12; unspecified=129).
 A.9 Claims Not Fully Mechanized